### v1

# A Calculus of Overlays

Abstract.
Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) as the foundation of functional programming, Overlay-Calculus uses three primitives (record, definition, inheritance) as the foundation of declarative programming. It trivially embeds the $\lambda$-calculus, although the entire semantics builds on only naive set theory; as a consequence, all constructs including inheritance are inherently commutative, idempotent, and associative; the linearization problem of multiple inheritance simply does not arise. This induces a fully abstract semantics of the lazy $\lambda$-calculus with respect to Böhm tree equivalence. Overlay-Calculus is distilled from the Overlay language, a practical implementation in which we observed further emergent phenomena: the Expression Problem dissolves, programs are CPS-agnostic, records natively encode random-access memory, and self-reference resolves to multiple targets. These properties suggest applications to configuration languages, dependency injection, object-oriented programming, composable effect systems, modular software architectures, file-system-as-compiler, general-purpose programming, and no-code development.
declarative programming, inheritance, denotational semantics, self-referential records, expression problem, $\lambda$-calculus, Böhm trees, configuration languages
This work was conducted independently prior to the author’s employment at Figure AI.
$c$s: Theory of computation Denotational semanticsccs: Theory of computation Object oriented constructsccs: Theory of computation Functional constructsccs: Theory of computation Program schemesccs: Software and its engineering Formal language definitionsccs: Software and its engineering Extensible languagesccs: Software and its engineering Object oriented languages
1. Introduction
Declarative and configuration languages are ubiquitous in modern software engineering. Systems such as NixOS modules (Dolstra et al., 2010; NixOS Contributors, 2024), Jsonnet (Cunningham, 2014), Hydra (Yadan and Droppo, 2023), CUE (van Lohuizen, 2019), Dhall (Gonzalez, 2017), Kustomize (Kubernetes SIG CLI, 2018), and JSON Patch (Bryan et al., 2013) all provide mechanisms for composing structured data through inheritance or overlay. Among these, the NixOS module system stands out: its recursive attribute set merging with fixed-point semantics and deferred modules (NixOS Contributors, 2024) achieves remarkable expressiveness in practice, and the mechanism has been adopted well beyond NixOS itself—Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), disko (Lassulus and nix-community Contributors, 2022), flake-parts (Hensing, 2022), dream2nix (Hauer, 2021), devenv (Kožar, 2022), KubeNix (Xtruder and Hall, 2022), and nixidy (Ingólfsson, 2024) use it to manage user environments, macOS configuration, disk partitioning, flake structure, multi-language packaging, developer environments, and Kubernetes clusters respectively. Yet no computational theory explains why this mechanism is so powerful.
The $\lambda$-calculus serves as the foundational computational model for functional programming. No analogous calculus exists for declarative programming. This gap matters because the two paradigms differ in fundamental ways. Configuration languages are inherently declarative: their values are immutable and they have no first-class functions. At first glance, Turing completeness appears incompatible with these constraints. The three classical models of computation each violate at least one of them:
Model Turing complete Immutable No $\lambda$
Turing Machine $\checkmark$ $\times$ $\checkmark$
$\lambda$-calculus $\checkmark$ $\checkmark$ $\times$
RAM Machine $\checkmark$ $\times$ $\checkmark$
The Turing Machine and RAM Machine require mutable state; the $\lambda$-calculus requires first-class functions. A computational model that is Turing complete, immutable, and function-free would sit in the empty cell of this table. The NixOS module system already suggests that such a model exists: its inheritance-based composition over recursive records, without explicit functions, is expressive enough to configure entire operating systems.
Conventionally, the value domain of configuration languages is assumed to consist of finite, well-founded structures—initial algebras in the sense of universal algebra. We challenge this assumption: configuration values are better understood as lazily observable, possibly infinite structures—in the spirit of F-coalgebras (Rutten, 2000)—where semantics is determined by the observer and a finite prefix suffices for any finite observation. This is not merely a theoretical distinction: the nixpkgs package collection (Dolstra, 2006)—over 100,000 packages—behaves as a lazily observed structure in practice, allowing any single package to be evaluated without materializing the entire set. Guided by this observation and by the NixOS module system’s recursive merging mechanism, we built the Overlay language (Yang, 2025b)—a declarative programming language with only three constructs (record literals, properties, and inheritance) and no functions or let-bindings.
In developing real programs in the Overlay language—Church-encoded data types (Appendix B), an Expression Problem solution (Section 5), trie-based random-access memory (Appendix A), and continuation patterns (Appendix G)—we discovered that several long-standing problems in programming language theory do not arise. The linearization problem that has plagued mixin-based systems (Bracha and Cook, 1990; Barrett et al., 1996) does not exist—not because it is resolved, but because the three primitives make it structurally impossible (Section 3). Self-reference naturally resolves to multiple objects rather than one; prior systems assume single-object resolution and reject the multi-path situation (Scala, NixOS module system). Self-referential records require no fixed-point combinator—the inheritance tree is the call stack, and qualified this resolution is a first-order set-theoretic query, not a fixed-point construction. These are not properties we engineered; they are phenomena we observed in working programs.
This paper distills these observations into Overlay-Calculus, a minimal computational model for declarative programming. It turns out that function application has a straightforward translation into inheritance: a $\lambda$-abstraction corresponds to a record with an $\mathrm{argument}$ slot and a $\mathrm{result}$ slot, and function application corresponds to inheriting from the function record and a record that supplies the argument. Since the $\lambda$-calculus embeds into Overlay-Calculus, Overlay-Calculus is Turing complete. The translation reveals an asymmetry: while embedding the $\lambda$-calculus into Overlay-Calculus is trivial, the reverse is non-trivial because Overlay-Calculus provides primitive operations—self-referential records, open recursion, symmetric inheritance—that have no direct counterparts in the $\lambda$-calculus. Moreover, Overlay-Calculus naturally encodes random-access memory: records are tries, and inheritance is trie union, yielding an immutable analogue of the RAM machine (Appendix A). This asymmetry suggests that Overlay-Calculus is not merely a syntactic variant of the $\lambda$-calculus but captures different computational patterns.
Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) to serve as the foundation of functional programming, Overlay-Calculus uses three primitives (record, definition, inheritance) to serve as the foundation of declarative programming.
All constructions in this paper have been implemented and tested in the Overlay language, available as open source (Yang, 2025b), with a test suite covering every example. All tests are mechanically verified against expected output snapshots.
Contributions
  • We present Overlay-Calculus, a minimal computational model for declarative programming that contains only overlay constructs (record, definition, inheritance) and no functions or scalar types.
  • Inheritance operates at the tree level: a record is a set of elements—definitions and inheritance sources—so inheriting from two overlays that share a label recursively merges their subtrees (union mount semantics), rather than overriding at the method level as in mixin and trait calculi. Since records are sets, inheritance is inherently commutative, idempotent, and associative. The linearization problem inherent in prior mixin and trait calculi simply does not arise.
  • The semantics is purely observational: there are no reduction rules, no evaluation order, and no notion of a result that a term reduces to. The observer drives computation by querying paths in a lazily constructed tree.
  • Self-referential records require no fixed-point combinator. The $\lambda$-calculus translation maps each application to a inheritance; the resulting inheritance tree is the call stack. The $\mathtt{this}$ function of Section 3 resolves self-references by walking this tree—a first-order set-theoretic query, not a fixed-point construction.
  • Self-reference resolves to multiple objects, not one. Path-dependent types with inheritance naturally produce this situation; prior systems assume single-object resolution and reject it (Scala, Appendix C; NixOS module system, Appendix D). Since inheritance is idempotent, all objects contribute equally—no disambiguation is needed.
  • The $\lambda$-calculus (in A-normal form) translates into Overlay-Calculus in five rules, so Overlay-Calculus is Turing complete. Composing this translation with the semantic functions induces a semantics on the $\lambda$-calculus that requires no function spaces, no substitution, and no reduction rules—only set membership in powerset lattices (Section 4). The translation is adequate and fully abstract with respect to Böhm tree equivalence (Section 4.2): two terms have the same overlay-convergence behavior in all contexts if and only if they have the same Böhm tree. This yields a fully abstract semantics of the lazy $\lambda$-calculus that lives entirely in naive set theory—powerset lattices and the Knaster–Tarski fixed-point theorem—without domain theory, inverse limits, or Scott continuity. Standard data types (booleans, natural numbers) can be Church-encoded without scalars (Appendix B).
  • We observe several emergent properties in practice: the Expression Problem (Wadler, 1998) does not arise (Section 5), programs are CPS-agnostic (Appendix G), and records natively encode random-access memory (Appendix A).
Section 5 demonstrates these properties through a case study of the Expression Problem.
2. Syntax
Let $e$ denote an expression. Let $\ell$ denote a label (property name). Let $k$ denote a non-negative integer (path length).
$e\quad::=\quad$ $\{c_{1},\;\ldots,\;c_{m}\}$ $\text{(}m\geq 0\text{)}$
$\mid\quad$ $c$ $\text{(sugar for }\{c\}\text{)}$
$c\quad::=\quad$ $\ell\mapsto e$ (definition)
$\mid\quad$ $[\ell_{\mathrm{up}}.\mathbf{this}.\mid\uparrow^{\!n}.]\;\ell_{\mathrm{down},1}.\ell_{\mathrm{down},2}\ldots\ell_{\mathrm{down},k}$ $\text{(inheritance, }k\geq 0\text{ with qualifier,}$
$k\geq 1\text{ without)}$
An expression is a set of elements enclosed in braces. Each element is either a definition $\ell\mapsto e$ or an inheritance source. Because expressions are sets, element order is irrelevant and duplicates have no effect—inheritance is inherently commutative, idempotent, and associative.
The $\mapsto$ in $\ell\mapsto e$ defines a property, not a let-binding. There is no variable binding in Overlay-Calculus. The Overlay language also provides scalar types through a foreign-function interface (FFI); Overlay-Calculus is what remains after removing the FFI. None of the examples in this paper uses the FFI—scalars simply did not arise.
Inheritance
An inheritance is written as a reference that navigates the scope hierarchy upward and then projects properties downward. The upward navigation admits two equivalent notations:
  • Named form: $\ell_{\mathrm{up}}.\mathbf{this}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$, analogous to Java’s Outer.this.field. Each $\{\ldots\}$ record has a label in its enclosing scope; $\ell_{\mathrm{up}}$ names the target enclosing scope.
  • Indexed form: $\uparrow^{\!n}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$, where $n$ is a de Bruijn index (de Bruijn, 1972) counting the number of enclosing scope levels to traverse.
Both forms retrieve the target scope—the fully inherited record after all inheritance has been applied—and then project properties $\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$ from it. The named form desugars to the indexed form during parsing (Section 3).
A new scope level is created at each $\{\ldots\}$ (record literal). The scope contains all properties of that overlay, including those inherited via inheritance sources. For example, $\{a\mapsto\{\},\;b\mapsto\{\}\}$ is a single overlay whose scope contains both $a$ and $b$; so is $\{r,\;a\mapsto\{\}\}$ where $r$ is a reference to a record containing $b$. Inheritance sources do not create additional scope levels.
  • $\mathrm{Foo}.\mathbf{this}$ (equivalently, $\uparrow^{\!n}$ where $\mathrm{Foo}$ is $n$ scope levels above) retrieves the enclosing record named $\mathrm{Foo}$.
  • $\mathrm{Foo}.\mathbf{this}.\ell$ (equivalently, $\uparrow^{\!n}.\ell$) projects property $\ell$ from $\mathrm{Foo}$.
  • When $k=0$, the reference inherits the entire enclosing scope, which may produce an infinitely deep tree.
When the qualifier is unambiguous, it may be omitted: the shorthand $\ell_{\mathrm{down},1}.\ell_{\mathrm{down},2}\ldots\ell_{\mathrm{down},k}$ ($k\geq 1$) resolves $\ell_{\mathrm{down},1}$ against the innermost enclosing scope whose $\mathtt{defines}$ contains it, and desugars to the corresponding indexed form. Since this uses $\mathtt{defines}$, the shorthand can only access properties defined in record literals, not those inherited through inheritance; one of the two qualified forms is required for inherited properties or to bypass variable shadowing.
3. Overlay Trees
Given an AST, what are the properties visible at each position in the tree? This section answers the question by defining a collection of mutually recursive functions over paths.
[…]
Most of these functions have direct analogues in object-oriented languages: $\mathtt{Path}$ identifies a position in the tree, $\mathtt{init}$ and $\mathtt{last}$ navigate paths, $\mathtt{defines}$ extracts the locally defined members from the AST, $\mathtt{inherits}$ extracts the inheritance declarations from the AST, $\mathtt{properties}$ computes all visible members at a path including both locally defined and inherited ones, $\mathtt{supers}$ computes the transitive inheritance closure, and $\mathtt{bases}$ collects the direct base classes. One function has no object-oriented analogue: $\mathtt{overlays}$ collects the identity-sharing paths that arise when inheritance introduces multiple definitions of the same label at the same scope level. Two functions, the reference resolution function $\mathtt{resolve}$ and the qualified this resolution function $\mathtt{this}$, have object-oriented analogues whose semantics differ significantly; they are discussed in detail below.
All functions are pure and may be cached (memoised) without changing semantics.
Path
A $\mathtt{Path}$ is a sequence of labels $(\ell_{1},\ell_{2},\ldots,\ell_{n})$ that identifies a position in the tree. [^1] Since paths are referentially transparent, an implementation may intern paths so that structural equality reduces to pointer equality. We write $p$ for a path. The root path is the empty sequence $()$. Given a path $p$ and a label $\ell$, $p\mathbin{\triangleright}\ell$ is the sequence $p$ extended with $\ell$. For a non-root path, the parent path $\mathtt{init}(p)$ is $p$ with its last element removed, and the final label $\mathtt{last}(p)$ is the last element of $p$. For convenience, we sometimes write $\ell$ for $\mathtt{last}(p)$ when the path is clear from context. With paths in hand, we can state what the AST provides at each path.
AST
An expression $e$ (Section 2) is parsed into an AST. The AST provides two primitive functions at each path $p$:
  • $\mathtt{defines}(p)$: the set of labels that have subtrees at $p$.
  • $\mathtt{inherits}(p)$: the set of reference pairs $(n,\;\ell_{\mathrm{down},*})$, where $n$ is the de Bruijn index (de Bruijn, 1972) (zero-based) and $\ell_{\mathrm{down},*}$ is the list of downward projections.
During parsing, all three syntactic forms are resolved to de Bruijn index pairs $(n,\;\ell_{\mathrm{down},*})$. The indexed form $\uparrow^{\!n}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$ already carries the de Bruijn index $n$ directly; no resolution is needed. A named reference $\ell_{\mathrm{up}}.\mathbf{this}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$ at path $p$ is resolved by finding the last occurrence of $\ell_{\mathrm{up}}$ among the labels of $p$. Let $p_{\mathrm{def}}$ be the prefix of $p$ up to and including that occurrence. The de Bruijn index is $n=|p|-|p_{\mathrm{def}}|$ and the projections are $\ell_{\mathrm{down},*}=(\ell_{\mathrm{down},1},\ldots,\ell_{\mathrm{down},k})$. A lexical reference $\ell_{\mathrm{down},1}.\ell_{\mathrm{down},2}\ldots\ell_{\mathrm{down},k}$ at path $p$ is resolved by finding the nearest prefix $p^{\prime}$ of $p$ such that $\ell_{\mathrm{down},1}\in\mathtt{defines}(p^{\prime})$. The de Bruijn index is $n=|p|-|p^{\prime}|$ and the projections are $\ell_{\mathrm{down},*}=(\ell_{\mathrm{down},1},\ldots,\ell_{\mathrm{down},k})$. All three forms produce the same representation; the semantic functions below operate only on de Bruijn index pairs.
Parsing populates $\mathtt{defines}$ and $\mathtt{inherits}$ from each element of a record. A definition $\ell\mapsto e$ contributes $\ell$ to $\mathtt{defines}(p)$. A reference contributes $(n,\;\ell_{\mathrm{down},*})$ to $\mathtt{inherits}(p)$. Since records are sets, the order of elements is irrelevant and duplicate elements have no effect. Both $\mathtt{defines}$ and $\mathtt{inherits}$ are pure data; they are not functions of runtime state. Given these two primitives, we can now answer the question posed at the beginning of this section.
Properties
The properties of a path are its own properties together with those inherited from all supers:
(1) $\mathtt{properties}(p)=\bigl\{\;\ell\;\big|\;(\_,\;p_{\mathrm{overlay}})\in\mathtt{supers}(p),\;\ell\in\mathtt{defines}(p_{\mathrm{overlay}})\;\bigr\}$
The remainder of this section defines $\mathtt{supers}$, the transitive inheritance closure, and its dependencies.
Supers
Intuitively, $\mathtt{supers}(p)$ collects every path that $p$ inherits from: the identity-sharing paths $\mathtt{overlays}$ of $p$ itself, plus the $\mathtt{overlays}$ of each direct base $\mathtt{bases}$ of $p$, and so on transitively. Each result is paired with the inheritance-site context $\mathtt{init}(p_{\mathrm{base}})$ through which it is reached. This provenance is needed by qualified this resolution, i.e. the $\mathtt{this}$ function defined below, to map a definition-site overlay back to the inheritance-site paths that incorporate it:
(2) $\mathtt{supers}(p)=\bigl\{\;(\mathtt{init}(p_{\mathrm{base}}),\;p_{\mathrm{overlay}})\;\big|\;p_{\mathrm{base}}\in\mathtt{bases}^{*}(p),\;p_{\mathrm{overlay}}\in\mathtt{overlays}(p_{\mathrm{base}})\;\bigr\}$
The $\mathtt{supers}$ formula depends on two functions: the identity-sharing paths $\mathtt{overlays}$ and the one-hop reference targets $\mathtt{bases}$. We define $\mathtt{overlays}$ first.
Overlays
Intuitively, inheritance can introduce multiple definitions of the same label at the same scope level; $\mathtt{overlays}(p)$ collects all such paths that share the same identity as $p$. Concretely, the overlays of $p$ include $p$ itself and $p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)$ for every branch $p_{\mathrm{branch}}$ of $\mathtt{init}(p)$ that also defines $\mathtt{last}(p)$:
(3) $\mathtt{overlays}(p)=\begin{cases}\{p\}&\text{if }p=()\\[6.0pt] \{p\}\;\cup\;\left\{\;p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)\;\left|\;\begin{aligned} &(\_,\;p_{\mathrm{branch}})\in\mathtt{supers}(\mathtt{init}(p)),\\ &\text{s.t.}\;\mathtt{last}(p)\in\mathtt{defines}(p_{\mathrm{branch}})\end{aligned}\right.\right\}&\text{if }p\neq()\end{cases}$
It remains to define $\mathtt{bases}$, the other dependency of $\mathtt{supers}$.
Bases
Intuitively, $\mathtt{bases}(p)$ are the paths that $p$ directly inherits from via references—analogous to the direct base classes in object-oriented languages. Concretely, $\mathtt{bases}$ resolves every reference in $p$’s $\mathtt{overlays}$ one step:
[…]
The $\mathtt{bases}$ formula calls the reference resolution function $\mathtt{resolve}$, which we define next.
Reference resolution
Intuitively, $\mathtt{resolve}$ turns a syntactic reference into the set of paths it points to in the fully inherited tree. It takes an inheritance-site path $p_{\mathrm{site}}$, a definition-site path $p_{\mathrm{def}}$, a de Bruijn index $n$, and downward projections $\ell_{\mathrm{down},*}$. Resolution proceeds in two phases: first, the qualified this resolution function $\mathtt{this}$ performs $n$ upward steps starting from the enclosing scope $\mathtt{init}(p_{\mathrm{def}})$, mapping the definition-site path to inheritance-site paths; then the downward projections $\ell_{\mathrm{down},*}$ are appended:
(5) $\mathtt{resolve}(p_{\mathrm{site}},\;p_{\mathrm{def}},\;n,\;\ell_{\mathrm{down},*})=\\ \bigl\{\;p_{\mathrm{current}}\mathbin{\triangleright}\ell_{\mathrm{down},1}\mathbin{\triangleright}\cdots\mathbin{\triangleright}\ell_{\mathrm{down},k}\;\big|\;p_{\mathrm{current}}\in\mathtt{this}(\{p_{\mathrm{site}}\},\;\mathtt{init}(p_{\mathrm{def}}),\;n)\;\bigr\}$
When multiple routes exist (due to multi-path inheritance), they may yield different target paths; all are collected. We now define $\mathtt{this}$.
Qualified this resolution
Intuitively, $\mathtt{this}$ answers the question: “in the fully inherited tree, where does the definition-site scope $p_{\mathrm{def}}$ actually live?” The design of $\mathtt{this}$ reconciles two historically opposing approaches to scope resolution. Early Lisp implementations used dynamic scope (McCarthy, 1978), where variable references are resolved against the run-time call stack; this is flexible but unpredictable. Scheme (Sussman and Steele Jr., 1975) corrected this with lexical scope, where a de Bruijn index resolves in a single step by indexing into the statically determined environment; this is predictable but rigid. Scala (Odersky et al., 2004) and the NixOS module system (NixOS Contributors, 2024) follow the lexical approach: each this reference is bound to a single statically known path. When inheritance introduces multiple inheritance routes to the same scope, Scala rejects the ambiguity at compile time (Appendix C) and the NixOS module system rejects the duplicate option declaration (Appendix D)—neither can express that all routes are equally valid.
Overlay-Calculus combines the strengths of both approaches. The reference itself is lexically determined at the definition site as a de Bruijn index, so what $\mathtt{this}$ looks up is predictable. But because inheritance interleaves scopes from different overlays, the result depends on the inherited tree structure: $\mathtt{this}$ walks upward one scope level at a time, consulting $\mathtt{supers}$ at each level, and naturally tracks a set of inheritance-site paths rather than a single one. All routes contribute equally to the inherited result, and since inheritance is idempotent, duplicated contributions are harmless. The only observable is property presence, making multi-path resolution well-defined without disambiguation.
Concretely, each step finds, among the supers of every path in the frontier $S$, those whose overlay component matches $p_{\mathrm{def}}$, and collects the corresponding inheritance-site paths as the new frontier. After $n$ steps the frontier contains the answer:
(6) $\small\mathtt{this}(S,\;p_{\mathrm{def}},\;n)=\begin{cases}S&\text{if }n=0\\[6.0pt] \mathtt{this}\!\left(\left\{\;p_{\mathrm{site}}\;\left|\;\begin{aligned} &p_{\mathrm{current}}\in S,\\ &(p_{\mathrm{site}},\;p_{\mathrm{overlay}})\in\mathtt{supers}(p_{\mathrm{current}}),\\ &\text{s.t.}\;p_{\mathrm{overlay}}=p_{\mathrm{def}}\end{aligned}\right.\right\},\;\mathtt{init}(p_{\mathrm{def}}),\;n-1\right)&\text{if }n>0\end{cases}$
At each step $\mathtt{init}$ shortens $p_{\mathrm{def}}$ by one label and $n$ decreases by one. Since $n$ is a non-negative integer, the recursion terminates.
Alternative approaches considered
Our first three implementations of the Overlay language used standard techniques for self-reference resolution: closure-based fixed points following Cook (Cook, 1989), and stack-based environment lookup using de Bruijn indices (de Bruijn, 1972). All three produced subtle bugs when inheritance introduced multiple inheritance routes to the same enclosing scope.
In retrospect, the difficulty was inevitable. A closure captures a single environment, so a fixed-point combinator solves for a single self-reference; when multiple routes exist, it attempts to unify them into one value. A stack-based environment is a linear chain—each scope level has one parent—so looking up a de Bruijn index returns a single value. Both data structures embed a single-valued assumption that is incompatible with the multi-path situation that arises when independent overlays inherit from the same enclosing scope (Appendix C). Equation (6) emerged from recognizing this: it operates on paths as first-order data and tracks a frontier set $S$ rather than a single current scope.
This completes the chain of definitions needed to compute $\mathtt{properties}(p)$.
Observational semantics without reduction
The definitions above constitute the complete semantics of Overlay-Calculus. There are no reduction rules, no evaluation contexts, and no small-step or big-step judgments. This is a deliberate design choice, not an omission. In the $\lambda$-calculus, a term such as $(\lambda x.\,x)\;y$ must be reduced to $y$ before its meaning is apparent; the operational semantics specifies how and in what order reductions occur. In Overlay-Calculus, an inheritance tree such as $\{a\mapsto\{\},\;b\mapsto\{\}\}$ does not reduce to anything—it is the value. The functions $\mathtt{properties}$, $\mathtt{supers}$, etc. determine what is observable at each path, and that is all. Computation arises not from rewriting terms but from querying an inductively deeper path in a lazily constructed tree: the observer drives the computation by choosing which path to inspect, and the mutually recursive definitions unfold on demand. Formally, the mutually recursive equations (1)–(6) define set membership inductively: $\ell\in\mathtt{properties}(p)$ holds if and only if it can be established by a finite chain of applications of these equations. The corresponding operator on the product lattice of path sets is monotone, and its least fixed point exists by the Knaster–Tarski theorem (Tarski, 1955). The detailed well-definedness argument—identifying the lattice, verifying positivity of each equation, and handling the transitive closure $\mathtt{bases}^{*}$—is given in Appendix E. When no finite derivation exists for a given query, the membership is not in this least fixed point and the semantics is undefined at that point—analogous to divergence in the $\lambda$-calculus, and the expected consequence of Turing completeness. This observation model is inspired by F-coalgebras (Rutten, 2000), where identity is determined by observable behavior rather than by internal structure.
4. Translation from $\lambda$-Calculus
The following translation maps the $\lambda$-calculus in A-normal form (ANF) (Flanagan et al., 1993) to Overlay-Calculus. In ANF, every intermediate result is bound to a name and each expression contains at most one application:
[…]
Every $\lambda$-term can be mechanically converted to ANF by naming intermediate results, and every ANF term is a $\lambda$-term, so ANF is neither a restriction nor an extension of the $\lambda$-calculus. The ANF structure matches Overlay-Calculus naturally: each $\mathbf{let}$-binding becomes a named record property, and each application is wrapped and projected via $\mathrm{result}.$
Let $\mathcal{T}$ denote the translation function. Each $\lambda$-abstraction introduces one scope level in the translated overlay tree.
$x$ ($\lambda$-bound, index $n$)   $\longrightarrow$ $\uparrow^{\!n}.\mathrm{argument}$
$x$ (let-bound)   $\longrightarrow$ $x.\mathrm{result}$
$\lambda x.\,M$   $\longrightarrow$ $\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
$\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M$   $\longrightarrow$ $\{x\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
$V_{1}\;V_{2}$ (tail call)   $\longrightarrow$ $\{\_\mathrm{tail}\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\_\mathrm{tail}.\mathrm{result}\}$
A $\lambda$-bound variable $x$ with de Bruijn index $n$ (the number of enclosing $\lambda$-abstractions between the reference and its binder) becomes $\uparrow^{\!n}.\mathrm{argument}$. The indexed reference $\uparrow^{\!n}$ traverses $n$ scope levels upward to reach the abstraction that binds $x$, giving access to the $\mathrm{argument}$ slot after all inheritance. A let-bound variable $x$ becomes the lexical reference $x.\mathrm{result}$, projecting the application result from the sibling property $x$ in the enclosing record.
An abstraction $\lambda x.\,M$ translates to a record with own properties $\{\mathrm{argument},\mathrm{result}\}$; we call this the abstraction shape. In the $\mathbf{let}$-binding rule, the name $x$ binds to the inherited record $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$, and $\mathcal{T}(M)$ may reference $x.\mathrm{result}$ to obtain the value of the application. In the tail-call rule, the fresh label $\mathrm{tail}$ serves the same purpose, and $\mathrm{result}$ projects the answer. In both cases the inheritance that performs the application is encapsulated behind a named property, so its internal structure (the $\mathrm{argument}$ and $\mathrm{result}$ labels of the callee) does not leak to the enclosing scope.
The five rules above are a complete translation from the $\lambda$-calculus (via ANF) to Overlay-Calculus: every ANF term has an image, and the construction is compositional. Since the $\lambda$-calculus is Turing complete, so is Overlay-Calculus; Appendix A further shows that Overlay-Calculus primitives directly encode a random-access machine.
An elementary semantics of the $\lambda$-calculus
Composing the translation $\mathcal{T}$ with the semantic functions of Section 3 induces a semantics on the $\lambda$-calculus itself. Given a $\lambda$-term $M$, one translates it to $\mathcal{T}(M)$ and then determines convergence behavior by querying $\ell\in\mathtt{properties}(p)$ via equations (1)–(6), following the $\mathrm{result}$ chain until the abstraction shape appears or the fixed point yields the empty set.
This induced semantics is more elementary than traditional approaches in the mathematical machinery it requires:
  • No function spaces. Scott’s denotational semantics (Scott, 1976) constructs a reflexive domain $D$ satisfying $D\cong[D\to D]$ via an inverse limit in a category of continuous lattices. Here, the semantic domain is a product of powerset lattices—no function spaces, no continuity conditions, no inverse limits. The only fixed-point construction is Knaster–Tarski on powerset lattices (Appendix E), which requires only set-theoretic monotonicity—strictly weaker than Scott continuity.
  • No substitution. Operational semantics defines $\beta$-reduction via capture-avoiding substitution, a mechanism whose subtleties motivated de Bruijn indices (de Bruijn, 1972) and explicit substitution calculi. Here, variable references are resolved by the $\mathtt{this}$ function walking the inheritance tree—no substitution occurs at any point.
  • No reduction rules. There are no $\beta$-reduction steps, no evaluation contexts, and no choice between call-by-name and call-by-value (Plotkin, 1975). The semantics is observational: one queries path membership, and the mutually recursive equations unfold on demand.
That a semantics of the $\lambda$-calculus can be given without function spaces is perhaps unexpected: function application is encoded as record inheritance, and the resulting record is observed through path queries rather than applied as a function. Section 4.2 shows that the convergence behavior of the translated overlay tree corresponds exactly to the Böhm tree (Barendregt, 1984) of the original term: the translation is adequate (a term converges iff it has a head normal form) and fully abstract (two terms are overlay-contextually equivalent iff they have the same Böhm tree).
Combining these two facts—elementary machinery and full abstraction—yields a noteworthy consequence: the lazy $\lambda$-calculus admits a fully abstract semantics that lives entirely in naive set theory. The only mathematical prerequisites are powerset lattices, set-theoretic monotonicity, and the Knaster–Tarski fixed-point theorem. No continuous lattices, no inverse limits, no Scott topology, no game-theoretic machinery. Scott’s $D_{\infty}$ model (Scott, 1976) uses heavier machinery yet is only adequate, not fully abstract (Plotkin, 1975); Hyland and Ong’s game semantics (Hyland and Ong, 2000) achieves full abstraction but at the cost of even greater mathematical complexity. The overlay tree semantics achieves full abstraction with strictly simpler foundations than either.
The reverse direction is non-trivial, as we discuss next.
4.1. Expressive Asymmetry
The translation above consists of five rules; the full computational power of the $\lambda$-calculus embeds into Overlay-Calculus with no additional machinery.
The reverse direction—encoding Overlay-Calculus in pure $\lambda$-calculus—is non-trivial. Defining the meaning of Overlay-Calculus requires machinery for lazy allocation, recursive data structures, and set-valued resolution that have no direct counterparts in the $\lambda$-calculus. This sophistication is not accidental—it reflects three features:
  1. (1)
    Self-referential records with lazy evaluation. An overlay’s properties can reference other properties and enclosing scopes through qualified this without explicit binding. In the $\lambda$-calculus, encoding such self-referential structures requires explicit allocation of mutable or lazy references (e.g., Haskell-style thunks or ML-style ref cells), obscuring the declarative intent.
  2. (2)
    Open recursion with symmetric inheritance. Inheritance merges definitions from independent sources, all sharing the same inherited overlay tree. Encoding this in the $\lambda$-calculus requires maintaining an extensible dictionary of methods that can be merged from multiple directions. Cook’s denotational semantics of inheritance (Cook, 1989) devoted an entire dissertation to formalizing this pattern in a functional setting.
  3. (3)
    Commutative, idempotent merge. Inheritance in Overlay-Calculus is symmetric: a record’s elements form a set, so reordering them produces the same observable result. In the $\lambda$-calculus, function composition $(f\circ g)$ is neither commutative nor idempotent, and simulating these properties requires additional machinery.
Overlay-Calculus is not the $\lambda$-calculus in disguise: it provides primitive operations for computational patterns that are derivable but complex in the $\lambda$-calculus.
The reverse asymmetry also exists: nested expressions such as f(gx)f\;(g\;x) are primitive in the $\lambda$-calculus but require manual A-normal form conversion in Overlay-Calculus—the intermediate result gxg\;x must be given an explicit name. However, ANF conversion is a trivial, mechanical transformation (Flanagan et al., 1993); a surface language built on Overlay-Calculus can support nested expressions as syntactic sugar without extending the calculus itself.
4.2. Böhm Tree Correspondence
The translation $\mathcal{T}$ induces a semantics on the $\lambda$-calculus via the overlay tree equations (1)–(6). We show that the induced semantics corresponds precisely to the Böhm tree (Barendregt, 1984) of the original $\lambda$-term.
An overlay tree $T$ converges, written $T{\Downarrow}$, if following finitely many $\mathrm{result}$ projections from the root reaches a node with both $\mathrm{argument}$ and $\mathrm{result}$ among its properties—the abstraction shape produced by $\mathcal{T}(\lambda x.\,M)$. We write $T{\Uparrow}$ if no such depth exists.
Definition 4.1 (Overlay-convergence).
An overlay tree $T$ converges, written $T{\Downarrow}$, if there exists $n\geq 0$ such that:
$\{\mathrm{argument},\,\mathrm{result}\}\;\subseteq\;\mathtt{properties}\!\bigl(\,\underbrace{\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}\cdots\mathbin{\triangleright}\mathrm{result}}_{n}\,\bigr)$
in the least fixed point of Section 3.
Theorem 4.2 (Adequacy).
A closed ANF $\lambda$-term $M$ has a head normal form if and only if $\mathcal{T}(M){\Downarrow}$.
Theorem 4.3 (Full Abstraction).
For closed $\lambda$-terms $M$ and $N$, define $\mathcal{T}(M)\approx\mathcal{T}(N)$ iff for every closing context $C[\cdot]$, $\mathcal{T}(C[M]){\Downarrow}\Leftrightarrow\mathcal{T}(C[N]){\Downarrow}$. Then:
$\mathcal{T}(M)\approx\mathcal{T}(N)\quad\Longleftrightarrow\quad\mathrm{BT}(M)=\mathrm{BT}(N)$
Together, Adequacy and Full Abstraction establish that the overlay tree semantics is a computation mechanism for Böhm trees (Barendregt, 1984)—the canonical fully abstract model of the lazy $\lambda$-calculus—using only powerset lattices and the Knaster–Tarski fixed-point theorem, without the topological infrastructure of domain theory. This is stronger than Scott’s denotational semantics (Scott, 1976; Wadsworth, 1976), which is adequate but not fully abstract—it identifies terms that observational equivalence distinguishes (Plotkin, 1975). The full proof, including the Single-Path Lemma, Substitution Lemma, and Convergence Preservation, is given in Appendix F.
5. Case Study: The Expression Problem
The Expression Problem (Wadler, 1998) asks how to extend a data type with both new cases and new operations, without modifying existing code. Existing solutions require object algebras (d. S. Oliveira and Cook, 2012; Wang and d. S. Oliveira, 2016), finally tagless interpreters (Carette et al., 2009), or type-class machinery. In a typed setting the problem additionally requires static type safety; Overlay-Calculus is untyped, so we focus on the extensibility aspect. Suppose a standard library provides a $\mathrm{binnat}$ module for binary natural arithmetic (see Appendix B). We define an expression language with an evaluation operation:
$\{\quad\mathrm{expression}$ $\mapsto\{\mathrm{Constant}\mapsto\{\mathrm{magnitude}\mapsto\{\}\},\;\mathrm{Addition}\mapsto\{\mathrm{left}\mapsto\{\},\;\mathrm{right}\mapsto\{\}\}\},$
$\mathrm{evaluation}$ $\mapsto\{\mathrm{Constant}\mapsto\{\mathrm{expression}.\mathrm{Constant},\;\mathrm{outcome}\mapsto\mathrm{Constant}.\mathbf{this}.\mathrm{magnitude}\},$
$\qquad\mathrm{Addition}\mapsto\{\mathrm{expression}.\mathrm{Addition},\;$
$\qquad\qquad\mathrm{left\_outcome}\mapsto\mathrm{Addition}.\mathbf{this}.\mathrm{left}.\mathrm{outcome},$
$\qquad\qquad\mathrm{right\_outcome}\mapsto\mathrm{Addition}.\mathbf{this}.\mathrm{right}.\mathrm{outcome},$
$\qquad\qquad\mathrm{sum}\mapsto\{\mathrm{binnat}.\mathrm{Addition},\;\mathrm{augend}\mapsto\mathrm{sum}.\mathbf{this}.\mathrm{left\_outcome},\;\mathrm{addend}\mapsto\mathrm{sum}.\mathbf{this}.\mathrm{right\_outcome}\},$
$\qquad\qquad\mathrm{outcome}\mapsto\mathrm{sum}.\mathrm{outcome}\}\},$
$\mathrm{binnat}$ $\mapsto\{\mathrm{Addition}\mapsto\{\mathrm{augend}\mapsto\{\},\;$
$\qquad\qquad\mathrm{addend}\mapsto\{\},\;\mathrm{outcome}\mapsto\{\}\}\}\quad\}$
A new operation (e.g., display) is added by defining a new module that inherits from the same $\mathrm{expression}$ schemas:
$\mathrm{display}\mapsto\{\mathrm{Constant}$ $\mapsto\{\mathrm{expression}.\mathrm{Constant},\;\mathrm{representation}\mapsto\mathrm{Constant}.\mathbf{this}.\mathrm{magnitude}\},$
$\mathrm{Addition}$ $\mapsto\{\mathrm{expression}.\mathrm{Addition},\;$
$\qquad\mathrm{representation}\mapsto\{$
$\qquad\qquad\mathrm{left}\mapsto\mathrm{Addition}.\mathbf{this}.\mathrm{left}.\mathrm{representation},\;$
$\qquad\qquad\mathrm{right}\mapsto\mathrm{Addition}.\mathbf{this}.\mathrm{right}.\mathrm{representation}\}\}\}$
A new case (e.g., negation) is added with its operation handlers and an empty API for the arithmetic primitive:
$\mathrm{negation}$ $\mapsto\{\mathrm{Negation}\mapsto\{\mathrm{operand}\mapsto\{\}\}\},$
$\mathrm{negation\_evaluation}$ $\mapsto\{\mathrm{Negation}\mapsto\{\mathrm{negation}.\mathrm{Negation},\;$
$\qquad\mathrm{operand\_outcome}\mapsto\mathrm{Negation}.\mathbf{this}.\mathrm{operand}.\mathrm{outcome},$
$\qquad\mathrm{negated}\mapsto\{\mathrm{binnat}.\mathrm{Negation},\;\mathrm{operand}\mapsto\mathrm{operand\_outcome}\},$
$\qquad\mathrm{outcome}\mapsto\mathrm{negated}.\mathrm{outcome}\}\},$
$\mathrm{binnat}$ $\mapsto\{\mathrm{Negation}\mapsto\{\mathrm{operand}\mapsto\{\},\;\mathrm{outcome}\mapsto\{\}\}\}$
All modules compose freely; none requires modification of existing definitions. A complete executable version of this example—including both the evaluation and display operations, a negation extension, and their free inheritance—is available in the implementation (Yang, 2025b), along with test cases verifying evaluation and display for nested expressions mixing old and new cases.
Note that $\mathrm{magnitude}\mapsto\{\}$ in $\mathrm{Constant}$ resembles a type annotation but is purely structural in Overlay-Calculus: the evaluator works as long as the required properties appear somewhere in the inheritance chain, regardless of any schema declarations.
6. Discussion
6.1. Relationship to $\lambda$-Calculus
Three facts establish the relationship between Overlay-Calculus and the $\lambda$-calculus.
More elementary and stronger than Scott domains.
The semantics induced by the translation (Section 4) requires only powerset lattices and the Knaster–Tarski fixed-point theorem—no function spaces, no inverse limits, no Scott continuity. Yet this semantics is fully abstract (Theorem 4.3), whereas Scott’s $D_{\infty}$ model (Scott, 1976; Wadsworth, 1976) is adequate but not fully abstract—it identifies terms that observational equivalence distinguishes (Plotkin, 1975).
$\lambda$-Calculus embeds into Overlay-Calculus.
The five-rule translation of Section 4 maps every ANF $\lambda$-term to an overlay tree. The translation uses only single-target qualified this—the full multi-target semantics is not needed to embed $\lambda$-calculus.
Multi-target qualified this is the natural primitive.
If the semantic functions of Section 3 were restricted to single-target resolution (as in DOT’s single-valued self variable (Amin et al., 2016)), the formulas would need to reject various structurally legitimate cases—adding complexity, not removing it. Multi-target qualified this is not an artificial extension; it is what falls out when no restrictions are imposed. The $\lambda$-calculus translation uses only single-target qualified this, suggesting that $\lambda$-calculus itself represents an artificially restricted fragment of the natural model. Appendix H presents a detailed analysis of the asymmetry in problem difficulty between the two systems.
6.2. Observed Phenomena
Several properties emerge in practice without having been designed:
  • The Expression Problem does not arise (Section 5, Appendix H).
  • Programs are CPS-agnostic: empty record slots serve as dependency injection points, making the same program text CPS or direct-style depending on assembly (Appendix G).
  • Records are tries, and inheritance is trie union—an immutable analogue of random-access memory (Appendix A).
  • Unlike SKI combinators (Schönfinkel, 1924; Curry and Feys, 1958), which are isomorphic to $\lambda$-calculus in problem difficulty, Overlay-Calculus exhibits genuine asymmetry (Section 4.1).
6.3. Properties of Inheritance
Because records are sets, inheritance is inherently commutative and idempotent, making Overlay-Calculus transparent to linearization order. Deduplication of inherited overlays is purely an optimization.
6.4. Relationship to Practical Systems
Several classes of declarative systems that have emerged independently in practice turn out to be instances of Overlay-Calculus.
Configuration and object-oriented languages
The NixOS module system (Dolstra et al., 2010; NixOS Contributors, 2024)—whose recursive attribute merging motivated this work—can be understood as an implementation of Overlay-Calculus with additional features (type checking, error reporting, FFI to the Nix language). Detailed comparisons with CUE, Jsonnet, Dhall, with mixin and trait calculi, and with the NixOS module system’s lazy evaluation limitations under multiple inheritance routes are given in Section 7.
Effect systems
Algebraic effects with handlers (Plotkin and Power, 2003; Pretnar, 2015) and monad transformers structure computational effects through composition. The CPS-agnostic property (Appendix G) shows how Overlay-Calculus naturally supports effect handler composition: empty API slots can be filled with effect handlers (CPS interpretation) or direct implementations (direct-style interpretation). The same pattern appears in algebraic effect systems, without explicit monad machinery.
Modular software architectures
Plugin systems, component frameworks, and dependency injection frameworks enable software extensibility through declarative inheritance. The symmetric, associative nature of inheritance (Section 3) explains why such systems work: components can be inherited in any order without changing behavior. These systems implement Overlay-Calculus patterns ad hoc.
Union file systems
Union file systems (UnionFS, OverlayFS, AUFS) layer multiple directory hierarchies to present a unified view. Their semantics resembles Overlay-Calculus—later layers override earlier layers, files from all layers remain accessible—but without a formal foundation these systems resort to ad-hoc choices: asymmetric inheritance, non-commutative conflict resolution for scalar files, and no late-binding or dynamic dispatch across layers. We implemented Overlay-Calculus directly as a union file system (Yang, 2025a), obtaining late-binding semantics and dynamic dispatch for file lookups with no additional mechanism beyond the three primitives. In package managers, build systems, and OS distributions, external toolchains transform configuration into file trees; here, the file system serves as configuration to itself, and the transformation is inheritance—a file-system-as-compiler in which the directory tree is both the source program and the compilation target.
7. Related Work
Denotational semantics of inheritance
Cook (Cook, 1989) gave the first denotational semantics of inheritance, modeling objects as recursive records and inheritance as composition of generators (functions from self to complete object) and wrappers (functions that modify generators), solved by a fixed-point construction. Cook’s key insight—that inheritance is a general mechanism applicable to any form of recursive definition, not only object-oriented methods—is one of the starting points of the present work. The model has four primitives (records, functions, generators, wrappers) and requires a fixed-point combinator; composition is asymmetric (a wrapper modifies a generator, not vice versa). Whether the same insight can be captured with fewer primitives and symmetric inheritance was a natural question.
Mixin-based inheritance
Bracha and Cook (Bracha and Cook, 1990) formalized mixins as abstract subclasses—functions from a superclass parameter to a subclass—unifying the inheritance mechanisms of Smalltalk, Beta, and CLOS under a single model. Because mixin application is function composition, it is neither commutative nor idempotent: applying the same mixin twice may produce different results, and applying two mixins in different orders may yield different linearizations. The C3 linearization algorithm (Barrett et al., 1996) was later developed to impose a deterministic order; whether linearization is inherent to inheritance or an artifact of non-commutativity was worth examining. Mixin composition operates at the method level: applying a mixin replaces or extends individual methods in a flat namespace. There is no recursive merging of nested structure—composing two mixins that define the same method name results in an override, not a union mount. Finally, mixins presuppose the $\lambda$-calculus as their computational foundation: method bodies are functions, and mixin application is function composition. Mixins organize and compose functions but cannot replace them—without the underlying $\lambda$-calculus, there is nothing to compose. The mixin mechanism itself is not Turing complete.
Traits
Schärli et al. (Schärli et al., 2003) introduced traits as composable units of behavior. Trait composition is symmetric and commutative: the sum of two traits is order-independent, which is a significant improvement over mixins. However, when two traits provide methods with the same name, a conflict arises and must be explicitly resolved by the programmer through exclusion or aliasing. Ducasse et al. (Ducasse et al., 2006) extended the model with a flattening property: the semantics of a class is independent of whether its methods originate from traits or are defined directly. Like mixins, trait composition is flat: it operates on individual method names, not on nested structures, so same-name definitions are conflicts rather than opportunities for deeper merging. And like mixins, traits presuppose the $\lambda$-calculus: trait methods are functions, and the computational power resides in those functions, not in the trait mechanism itself.
Family polymorphism and virtual classes
Ernst (Ernst, 2001) introduced family polymorphism, where families of mutually dependent classes can be refined together in subclasses. Ernst, Ostermann, and Cook (Ernst et al., 2006) formalized this in the virtual class calculus, where classes are class-valued attributes of objects, accessed via path expressions of the form this.out.C. The path this.out navigates to an enclosing scope, analogous to qualified this ($\ell_{\mathrm{up}}.\mathbf{this}$) in this paper. A notable restriction is that each such path resolves to a single class in a single enclosing object. When multiple inheritance routes lead to the same scope—a common situation under inheritance—single-valued resolution must choose one route and discard the others, losing information. Whether qualified this can resolve to a set of paths, letting all routes contribute equally, had not been explored.
Object calculi
Abadi and Cardelli (Abadi and Cardelli, 1996) developed a comprehensive theory of objects, treating objects—rather than classes—as the primitive notion. Their calculi include method update and self-referential objects, but methods are fundamentally functions (taking self as a parameter), and object extension is asymmetric (method override replaces the previous definition). Boudol (Boudol, 2004) studied the recursive record semantics of objects in a call-by-value setting, showing that self-referential records require an unsafe fixed-point operator. Whether self-reference can be handled without a fixed-point operator in the object language—for instance, through coalgebraic observation rather than reduction—was a question these calculi left open.
Record calculi
Harper and Pierce (Harper and Pierce, 1991) presented a record calculus based on symmetric concatenation with a type system that tracks the absence of fields. When two records define the same label, concatenation is undefined (a type error); the calculus avoids same-label composition rather than giving it meaning. Cardelli (Cardelli, 1994) studied extensible records with subtyping; Rémy (Rémy, 1989) gave a type system for records and variants in ML. These calculi operate on flat records (no nesting) and treat records as data (initial algebras): a record is a finite map from labels to values, fully determined at construction time. Nested records, self-reference, and lazy observation—the properties needed for records to serve as a computational substrate—are outside their scope.
DOT calculus
Amin et al. (Amin et al., 2016) formalized path-dependent types in the DOT calculus, the theoretical foundation of Scala’s type system. DOT paths navigate through objects to reach nested type members, providing a rich dependent type structure. DOT’s self variable in the object constructor $\nu(x:T)\,d$ binds $x$ to a single object; when multiple inheritance routes lead to the same enclosing scope, the resulting intersection type is unrealizable and Scala rejects it at compile time (Appendix C). DOT includes the $\lambda$-calculus as a sub-language; removing it while retaining path-dependent types would yield a calculus closer to pure overlay inheritance (see Section 8).
Configuration languages
CUE (van Lohuizen, 2019) unifies types and values in a single lattice where unification (&) is commutative, associative, and idempotent. CUE’s design was motivated by Google’s Borg Configuration Language (GCL), which used graph unification. CUE’s lattice includes scalar types with a conflict semantics (unifying incompatible scalars yields $\bot$); whether scalars are necessary for a configuration language, or whether tree structure alone suffices, is an interesting design question. Jsonnet (Cunningham, 2014) provides object inheritance via the + operator with mixin semantics: composition is not commutative (the right-hand side wins on scalar conflicts), and deep merging requires explicit opt-in (+: syntax) on a per-field basis—so union mount semantics is available but not the default. Dhall (Gonzalez, 2017) takes a functional approach: it is a typed $\lambda$-calculus with records, where composition is a record merge operator, not an inheritance mechanism.
Module systems and union mount semantics
The NixOS module system (Dolstra et al., 2010; NixOS Contributors, 2024) composes modules by recursively merging nested attribute sets, not by linearization or method-level override—the same union mount semantics that distinguishes tree-level inheritance from mixin and trait composition. When two modules define the same nested path, their subtrees are merged rather than overridden. The NixOS module system got nearly all of the high-level design right: union mount inheritance, recursive self-reference, deferred evaluation of modules, and a rich type-checking layer on top. Its practical success is remarkable—the same mechanism powers NixOS system configuration, Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), flake-parts (Hensing, 2022), and dozens of other ecosystems, collectively managing configurations of arbitrary complexity. The one aspect that appears not to have been fully addressed is the semantics of self-reference under multiple inheritance. When two independent modules that inherit from the same deferred module are composed, each import introduces its own declaration of the shared options; the module system’s mergeOptionDecls rejects duplicate declarations with a static error (Appendix D). The module system cannot express that both routes contribute equally to the same option—the multi-path situation that recursive merging naturally gives rise to—which motivated the coalgebraic semantics of Section 3.
SKI combinator calculus
Schönfinkel (Schönfinkel, 1924) and Curry (Curry and Feys, 1958) showed that the $\lambda$-calculus can be reformulated without bound variables using three combinators (S, K, I). SKI is function-free at the object level and Turing complete. However, bracket abstraction translates between SKI and the $\lambda$-calculus preserving computational structure, so the two systems are isomorphic in terms of problem difficulty (Section 4): what is hard in one remains hard in the other. Whether a function-free calculus can exhibit a genuine shift in problem difficulty—making some problems structurally easier than in the $\lambda$-calculus, rather than merely encoding them differently—is the question that motivates the comparison in Section 4.
8. Future Work
The Overlay language (Yang, 2025b) is an executable implementation of Overlay-Calculus that already goes beyond the untyped calculus presented here: it includes compile-time checking that all references resolve to valid paths in the inheritance, and a foreign-function interface (FFI) that introduces scalar values from the host language. The future work described below concerns the distance between this implementation and a fully practical language—spanning both theoretical foundations and library-level encodings.
Type system
The Overlay language’s existing compile-time checks verify that every reference path resolves to a property that exists in the inherited result. Formalizing the soundness of these checks—proving that well-typed programs do not produce dangling references at runtime—is one direction of future work. Such a formalization would relate to the DOT calculus (Amin et al., 2016) in a manner analogous to how Overlay-Calculus relates to the $\lambda$-calculus: a Typed Overlay-Calculus can be viewed as DOT without $\lambda$, retaining path-dependent types while replacing functions with overlay inheritance. A formal type system would be amenable to mechanization in proof assistants (Coq, Agda) for verifying type safety, progress, and preservation properties.
Beyond soundness of the existing checks, additional type system features would strengthen the language. A key example is totality checking: verifying that a inheritance provides implementations for all required slots, not merely that references resolve. In untyped Overlay-Calculus, writing $\mathrm{magnitude}\mapsto\{\}$ is purely documentary—it defines a structural slot but imposes no constraint on what must fill it. Totality checking would enforce such declarations statically, rejecting inheritances that leave required slots unfilled.
Standard library
The remaining items concern the standard library: they require no extensions to the calculus or type system, but involve nontrivial encodings within the existing framework.
Overlay-Calculus is open by default: inheritance can freely merge any two overlays, and a value may simultaneously inhabit multiple constructors (e.g., $\{\mathrm{Zero},\;\mathrm{Odd},\;\mathrm{half}\mapsto\mathrm{Zero}\}$). This is not a defect—it is the natural trie semantics of the calculus, and the basis for solving the expression problem (Section 4.1). However, many operations (e.g., equality testing) assume that values are linear—inhabiting exactly one constructor. The observer pattern (Appendix B) does not enforce this: applied to a multi-constructor value, all callbacks fire and their results are inherited.
Schema validation can be implemented at the library level using existing primitives. An observer can test whether a particular constructor is present, returning a Boolean; a second Boolean dispatch then branches on the result. By chaining such tests, a factory can validate that a value matches exactly one constructor and carries the required fields. This does not restrict inheritance itself—it remains unconstrained—but provides a way to detect invariant violations before they propagate.
The same technique—observer returning Boolean, followed by Boolean dispatch—yields closed pattern matching. The observer pattern of Appendix B is inherently open: new constructor cases can be added through inheritance. Closed matching, where exactly one branch is taken, can be encoded as a visitor chain: a linked list of if-then-else nodes, each testing one constructor and falling through on mismatch. This is analogous to GHC.Generics’ $\mathrm{(:+:)}$ sum representation (Magalhães et al., 2010), where each link corresponds to one summand. An important consequence is that such chains do not commute—they have a fixed order—and therefore cannot be extended through open inheritance, which is the correct semantics for closed dispatch.
The trie operations in Appendix A are hardcoded for binary natural number keys. Since every overlay is a trie (Appendix A) and insertion is simply inheritance, only operations that cannot be expressed as inheritance need explicit implementation: deletion, dynamic lookup, and prefix lookup. Generalizing to arbitrary key types requires each key factory to provide two interfaces—SelectTrieChild (select a subtrie by constructor) and ReplaceTrieChild (replace a subtrie by constructor while preserving the rest)—corresponding to the getter and setter of a lens (Foster et al., 2007) focused on the constructor-determined child. Prefix lookup is structurally analogous to equality testing: both synchronously traverse a key (a single-path trie) alongside another structure, dispatching at each level. The difference is the return type—a Boolean for equality, a subtrie for prefix lookup.
Appendix A Trie Operations
[…]
$\mathrm{Zero}$ $\;\longrightarrow\;\mathrm{value}$
$\mathrm{Odd}(n)$ $\;\longrightarrow\;\mathrm{at\_odd},\;\text{then path for }n$
$\mathrm{Even}(n)$ $\;\longrightarrow\;\mathrm{at\_even},\;\text{then path for }n$
Write (insertion)
Inserting an entry with value $v$ at key $k$ is inheriting a singleton trie whose nesting mirrors the bit structure of $k$. For example, inserting at key $5=\mathrm{Odd}(\mathrm{Even}(\mathrm{Zero}))$:
$\mathrm{trie}^{\prime}=\{\mathrm{trie},\;\mathrm{at\_odd}\mapsto\{\mathrm{at\_even}\mapsto\{\mathrm{value}\mapsto v\}\}\}$
Each inheritance creates a new immutable trie; the original is unchanged.
Read (static lookup)
When the key is known statically, retrieving the entry at key $k$ is reference-path navigation: the reference follows the same label sequence that was used for insertion. For the same key $5$:
$\mathrm{enclosing\_scope}.\mathbf{this}.\mathrm{trie}.\mathrm{at\_odd}.\mathrm{at\_even}.\mathrm{value}$
where $\mathrm{enclosing\_scope}$ is the qualified-this scope name containing the trie definition. The reference mechanism traverses the nested record structure—which is the trie—returning the value at the final node.
Read (dynamic lookup)
When the key is a BinNat value constructed at runtime via $\mathrm{Zero}$, $\mathrm{Odd}$, and $\mathrm{Even}$, the lookup path must be computed dynamically. This is achieved by folding the BinNat key into a Selector—an overlay that, given a trie, navigates to the appropriate node and returns the stored entry. A Selector declares the trie shape and a result slot:
$\mathrm{Selector}\mapsto\{\mathrm{trie}\mapsto\{\mathrm{value}\mapsto\{\},\;\mathrm{at\_odd}\mapsto\{\},\;\mathrm{at\_even}\mapsto\{\}\},\;\mathrm{result}\mapsto\{\}\}$
Three observer callbacks build Selectors bottom-up. These callbacks ($\mathrm{on\_zero}$, $\mathrm{on\_odd}$, $\mathrm{on\_even}$) are interpretation points in the sense of Appendix G: they can be filled externally by an orchestrator (CPS interpretation) or inherited locally (direct interpretation), demonstrating the CPS-agnostic property. The $\mathrm{on\_zero}$ callback returns a Selector that reads $\mathrm{value}$ from the trie:
$\mathrm{lookup\_on\_zero}\mapsto\{\mathrm{Selector},\;\mathrm{result}\mapsto\mathrm{lookup\_on\_zero}.\mathbf{this}.\mathrm{trie}.\mathrm{value}\}$
The $\mathrm{on\_odd}$ callback takes a Selector (via the $\mathrm{argument}/\mathrm{result}$ encoding of Section 4) and returns a new Selector that navigates to $\mathrm{at\_odd}$, then delegates to the argument Selector:
$\mathrm{lookup\_on\_odd}$ $\mapsto\{\mathrm{argument}\mapsto\mathrm{Selector},\;\mathrm{result}\mapsto\{\mathrm{Selector},\;$
$\qquad\mathrm{applied}\mapsto\{\mathrm{argument},\;\mathrm{trie}\mapsto\mathrm{lookup\_on\_odd}.\mathbf{this}.\mathrm{trie}.\mathrm{at\_odd}\},$
$\qquad\mathrm{result}\mapsto\mathrm{applied}.\mathrm{result}\}\}$
The $\mathrm{on\_even}$ callback is symmetric, navigating to $\mathrm{at\_even}$ instead.
The $\mathrm{Lookup}$ combinator folds a BinNat key with these callbacks to obtain a Selector, then applies it to the trie:
$\mathrm{Lookup}$ $\mapsto\{$
$\quad\mathrm{key}\mapsto\mathrm{trie\_module}.\mathbf{this}.\mathrm{BinNat},$
$\quad\mathrm{trie\_input}\mapsto\{\},$
$\quad\mathrm{applied\_key}\mapsto\{\mathrm{key},\;$
$\qquad\quad\mathrm{on\_zero}\mapsto\mathrm{lookup\_on\_zero},\;$
$\qquad\quad\mathrm{on\_odd}\mapsto\mathrm{lookup\_on\_odd},\;$
$\qquad\quad\mathrm{on\_even}\mapsto\mathrm{lookup\_on\_even}\},$
$\quad\mathrm{applied\_selector}\mapsto\{\mathrm{applied\_key}.\mathrm{result},\;$
$\qquad\quad\mathrm{trie}\mapsto\mathrm{trie\_input}\},$
$\quad\mathrm{result}\mapsto\mathrm{applied\_selector}.\mathrm{result}\}$
For example, looking up key $5=\mathrm{Odd}(\mathrm{Even}(\mathrm{Zero}))$ folds into a Selector that navigates $\mathrm{at\_odd}.\mathrm{at\_even}.\mathrm{value}$—the same path as the static reference, but computed from the key value. The entire construction uses only records (with inheritance sources) and references; no additional primitives are needed.
Delete
Since inheritance is additive—it can only merge properties, never remove them—deletion cannot be expressed as a single inheritance. Instead, deletion reconstructs the trie, copying every subtrie except at the targeted key.
We fold the BinNat key into a Rebuilder—an overlay that, given an input trie, produces a new trie as its result. A Rebuilder declares the trie shape on both its input and output:
$\mathrm{Rebuilder}\mapsto\{$ $\mathrm{trie}\mapsto\{\mathrm{value}\mapsto\{\},\;\mathrm{at\_odd}\mapsto\{\},\;\mathrm{at\_even}\mapsto\{\}\},\;$
$\mathrm{result}\mapsto\{\mathrm{value}\mapsto\{\},\;\mathrm{at\_odd}\mapsto\{\},\;\mathrm{at\_even}\mapsto\{\}\}\}$
Three observer callbacks build Rebuilders bottom-up. The $\mathrm{on\_zero}$ callback returns a Rebuilder that copies the subtries but omits the value:
$\mathrm{delete\_on\_zero}\mapsto\{$ $\mathrm{Rebuilder},\;\mathrm{result}\mapsto\{$
$\quad\mathrm{at\_odd}\mapsto\mathrm{delete\_on\_zero}.\mathbf{this}.\mathrm{trie}.\mathrm{at\_odd},\;$
$\quad\mathrm{at\_even}\mapsto\mathrm{delete\_on\_zero}.\mathbf{this}.\mathrm{trie}.\mathrm{at\_even}\}\}\}$
The $\mathrm{result}.\mathrm{value}$ slot inherits only the empty declaration from $\mathrm{Rebuilder}$, effectively deleting the entry.
The $\mathrm{on\_odd}$ callback takes a Rebuilder (the recursive result for the subtrie) and returns a new Rebuilder that preserves $\mathrm{value}$ and $\mathrm{at\_even}$, but replaces $\mathrm{at\_odd}$ with the recursively rebuilt subtrie:
$\mathrm{delete\_on\_odd}$ $\mapsto\{\mathrm{argument}\mapsto\mathrm{Rebuilder},\;$
$\quad\mathrm{result}\mapsto\{\mathrm{Rebuilder},\;$
$\qquad\mathrm{applied}\mapsto\{\mathrm{argument},\;$
$\qquad\quad\mathrm{trie}\mapsto\mathrm{delete\_on\_odd}.\mathbf{this}.\mathrm{trie}.\mathrm{at\_odd}\},$
$\qquad\mathrm{result}\mapsto\{$
$\qquad\quad\mathrm{value}\mapsto\mathrm{delete\_on\_odd}.\mathbf{this}.\mathrm{trie}.\mathrm{value},\;$
$\qquad\quad\mathrm{at\_odd}\mapsto\mathrm{applied}.\mathrm{result},\;$
$\qquad\quad\mathrm{at\_even}\mapsto\mathrm{delete\_on\_odd}.\mathbf{this}.\mathrm{trie}.\mathrm{at\_even}\}\}\}$
The $\mathrm{on\_even}$ callback is symmetric.
The $\mathrm{Delete}$ combinator folds a BinNat key with these callbacks to obtain a Rebuilder, then applies it to the trie:
$\mathrm{Delete}$ $\mapsto\{$
$\quad\mathrm{key}\mapsto\mathrm{trie\_module}.\mathbf{this}.\mathrm{BinNat},$
$\quad\mathrm{trie\_input}\mapsto\{\},$
$\quad\mathrm{applied\_key}\mapsto\{\mathrm{key},\;$
$\qquad\quad\mathrm{on\_zero}\mapsto\mathrm{delete\_on\_zero},\;$
$\qquad\quad\mathrm{on\_odd}\mapsto\mathrm{delete\_on\_odd},\;$
$\qquad\quad\mathrm{on\_even}\mapsto\mathrm{delete\_on\_even}\},$
$\quad\mathrm{applied\_rebuilder}\mapsto\{\mathrm{applied\_key}.\mathrm{result},\;$
$\qquad\quad\mathrm{trie}\mapsto\mathrm{trie\_input}\},$
$\quad\mathrm{result}\mapsto\mathrm{applied\_rebuilder}.\mathrm{result}\}$
Since inheritance is idempotent, the empty scaffolding left by the Rebuilder schema merges harmlessly with any subsequent insertions.
Update
Updating the entry at key $k$ to a new value $v$ combines deletion and insertion. Because $\mathrm{Delete}$ is a combinator whose output lives in the result property, one must first project out the rebuilt trie before inheriting the new singleton. In ANF style:
$\mathrm{deleted}$ $\mapsto\{\mathrm{Delete},\;\mathrm{key}\mapsto k,\;\mathrm{trie\_input}\mapsto\mathrm{trie}\}$
$\mathrm{result}$ $\mapsto\{\mathrm{deleted}.\mathrm{result},\;\mathrm{at\_odd}\mapsto\{\mathrm{at\_even}\mapsto\{\mathrm{value}\mapsto v\}\}\}$
where the second line uses key $5=\mathrm{Odd}(\mathrm{Even}(\mathrm{Zero}))$ as a concrete example. The projection $\mathrm{result}$ is essential: $\mathrm{Delete}$ carries internal scaffolding ($\mathrm{key}$, $\mathrm{trie\_input}$, $\mathrm{applied\_key}$, etc.), and only its $\mathrm{result}$ property holds the rebuilt trie. This intermediate projection prevents Update from being a single symmetric inheritance.
Appendix B Church Encoding of Standard Types
This appendix demonstrates that standard data types can be encoded in Overlay-Calculus without scalar types, using Church encoding. These encodings serve as a proof of concept that a standard library can be built entirely within Overlay-Calculus. Each encoding is presented as a module overlay. In practical implementations, more efficient representations (e.g., binary naturals) or FFI to host language primitives would typically be used. The implementation (Yang, 2025b) includes executable implementations of these encodings with test cases for $\mathrm{Not}(\mathrm{True})$, $\mathrm{And}(\mathrm{True},\mathrm{True})$, $\mathrm{Or}(\mathrm{False},\mathrm{True})$, and natural number addition, all verified against expected output snapshots.
B.1. Boolean Module
The boolean module contains a schema ($\mathrm{Boolean}$), constructors ($\mathrm{True}$, $\mathrm{False}$), and operations ($\mathrm{Not}$, $\mathrm{And}$, $\mathrm{Or}$). All definitions are siblings within a single module overlay.
$\{\quad\mathrm{boolean}\mapsto\{\quad\mathrm{Boolean}$ $\mapsto\{\mathrm{on\_true}\mapsto\{\},\;\mathrm{on\_false}\mapsto\{\},\;\mathrm{result}\mapsto\{\}\},$
$\mathrm{True}$ $\mapsto\{\mathrm{Boolean},\;\mathrm{result}\mapsto\mathrm{True}.\mathbf{this}.\mathrm{on\_true}\},$
$\mathrm{False}$ $\mapsto\{\mathrm{Boolean},\;\mathrm{result}\mapsto\mathrm{False}.\mathbf{this}.\mathrm{on\_false}\},$
$\mathrm{Not}$ $\mapsto\{\mathrm{Boolean},$
$\qquad\mathrm{operand}\mapsto\mathrm{Boolean},$
$\qquad\mathrm{applied\_operand}\mapsto\{\mathrm{operand},$
$\qquad\qquad\mathrm{on\_true}\mapsto\mathrm{Not}.\mathbf{this}.\mathrm{on\_false},$
$\qquad\qquad\mathrm{on\_false}\mapsto\mathrm{Not}.\mathbf{this}.\mathrm{on\_true}\},$
$\qquad\mathrm{result}\mapsto\mathrm{applied\_operand}.\mathrm{result}\},$
$\mathrm{And}$ $\mapsto\{\mathrm{Boolean},$
$\qquad\mathrm{left}\mapsto\mathrm{Boolean},$
$\qquad\mathrm{right}\mapsto\mathrm{Boolean},$
$\qquad\mathrm{applied\_left}\mapsto\{\mathrm{left},$
$\qquad\qquad\mathrm{on\_true}\mapsto\mathrm{right},$
$\qquad\qquad\mathrm{on\_false}\mapsto\mathrm{False}\},$
$\qquad\mathrm{result}\mapsto\mathrm{applied\_left}.\mathrm{result}\},$
$\mathrm{Or}$ $\mapsto\{\mathrm{Boolean},$
$\qquad\mathrm{left}\mapsto\mathrm{Boolean},$
$\qquad\mathrm{right}\mapsto\mathrm{Boolean},$
$\qquad\mathrm{applied\_left}\mapsto\{\mathrm{left},$
$\qquad\qquad\mathrm{on\_true}\mapsto\mathrm{True},$
$\qquad\qquad\mathrm{on\_false}\mapsto\mathrm{right}\},$
$\qquad\mathrm{result}\mapsto\mathrm{applied\_left}.\mathrm{result}\}\quad\}\quad\}$
Reference explanation
Within the $\mathrm{boolean}$ module, own properties such as $\mathrm{Boolean}$, $\mathrm{True}$, $\mathrm{False}$, $\mathrm{Not}$, $\mathrm{And}$, and $\mathrm{Or}$ are accessed via lexical references (e.g., the inheritance source $\mathrm{Boolean}$ in each definition). Qualified this is required in two situations:
  • Inherited properties. Inside True’s definition, $\mathrm{True}.\mathbf{this}.\mathrm{on\_true}$ requires qualified this because $\mathrm{on\_true}$ is inherited from Boolean via inheritance, not defined as an own property. Similarly, $\mathrm{Not}.\mathbf{this}.\mathrm{on\_false}$ and $\mathrm{Not}.\mathbf{this}.\mathrm{on\_true}$ access inherited callbacks.
  • Variable shadowing. Inside Not’s $\mathrm{applied\_operand}$, the labels $\mathrm{on\_true}$ and $\mathrm{on\_false}$ are own properties of $\mathrm{applied\_operand}$ (with swapped values), so a lexical reference $\mathrm{on\_false}$ would resolve to the local definition rather than Not’s inherited callback. Qualified this $\mathrm{Not}.\mathbf{this}.\mathrm{on\_false}$ bypasses the shadowing.
Worked example
Consider the expression $\{\mathrm{boolean}.\mathrm{Not},\;\mathrm{operand}\mapsto\mathrm{boolean}.\mathrm{True},\;\mathrm{on\_true}\mapsto A,\;\mathrm{on\_false}\mapsto B\}$ at the root scope level, where $A$ and $B$ are arbitrary expressions. The result is an object whose $\mathrm{result}$ label contains the same observable tree as $B$, because $\mathrm{Not}$ swaps $\mathrm{on\_true}$ and $\mathrm{on\_false}$, and $\mathrm{True}$ selects $\mathrm{on\_true}$, which after swapping becomes $B$.
B.2. Natural Number Module
The natural number module contains a schema ($\mathrm{Nat}$), constructors ($\mathrm{Zero}$, $\mathrm{Succ}$), and an operation ($\mathrm{Add}$).
$\{\quad\mathrm{natural}\mapsto\{\quad\mathrm{Nat}$ $\mapsto\{\mathrm{successor}\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\{\}\},\;\mathrm{zero}\mapsto\{\},\;\mathrm{result}\mapsto\{\}\},$
$\mathrm{Zero}$ $\mapsto\{\mathrm{Nat},\;\mathrm{result}\mapsto\mathrm{Zero}.\mathbf{this}.\mathrm{zero}\},$
$\mathrm{Succ}$ $\mapsto\{\mathrm{Nat},$
$\qquad\mathrm{predecessor}\mapsto\mathrm{Nat},$
$\qquad\mathrm{applied\_predecessor}\mapsto\{\mathrm{predecessor},$
$\qquad\qquad\mathrm{successor}\mapsto\mathrm{Succ}.\mathbf{this}.\mathrm{successor},$
$\qquad\qquad\mathrm{zero}\mapsto\mathrm{Succ}.\mathbf{this}.\mathrm{zero}\},$
$\qquad\mathrm{applied\_successor}\mapsto\{\mathrm{Succ}.\mathbf{this}.\mathrm{successor},$
$\qquad\qquad\mathrm{argument}\mapsto\mathrm{applied\_predecessor}.\mathrm{result}\},$
$\qquad\mathrm{result}\mapsto\mathrm{applied\_successor}.\mathrm{result}\},$
$\mathrm{Add}$ $\mapsto\{\mathrm{Nat},$
$\qquad\mathrm{augend}\mapsto\mathrm{Nat},$
$\qquad\mathrm{addend}\mapsto\mathrm{Nat},$
$\qquad\mathrm{applied\_addend}\mapsto\{\mathrm{addend},$
$\qquad\qquad\mathrm{successor}\mapsto\mathrm{Add}.\mathbf{this}.\mathrm{successor},$
$\qquad\qquad\mathrm{zero}\mapsto\mathrm{Add}.\mathbf{this}.\mathrm{zero}\},$
$\qquad\mathrm{applied\_augend}\mapsto\{\mathrm{augend},$
$\qquad\qquad\mathrm{successor}\mapsto\mathrm{Add}.\mathbf{this}.\mathrm{successor},$
$\qquad\qquad\mathrm{zero}\mapsto\mathrm{applied\_addend}.\mathrm{result}\},$
$\qquad\mathrm{result}\mapsto\mathrm{applied\_augend}.\mathrm{result}\}\quad\}\quad\}$
Concrete example
At the root scope level,
$\mathrm{one}\mapsto\{\mathrm{natural}.\mathrm{Succ},\;\mathrm{predecessor}\mapsto\mathrm{natural}.\mathrm{Zero}\}.$
Applying this with a Peano-style successor and zero: $\{\mathrm{one},\;\mathrm{successor}\mapsto S,\;\mathrm{zero}\mapsto Z\}$ produces an object whose $\mathrm{result}$ label contains the same observable tree as $\{S,\;\mathrm{argument}\mapsto Z.\mathrm{result}\}.\mathrm{result}$, i.e., successor applied once to zero. Here $S$ and $Z$ are arbitrary expressions representing a successor operation and a zero value respectively.
Appendix C Multi-Path Self-Reference in Scala
This appendix demonstrates that Scala 3 rejects the multi-path self-reference pattern that Overlay-Calculus handles naturally. Consider two objects that independently extend an outer class, each providing its own copy of an inner trait:
class MyOuter:
  trait MyInner:
    def outer = MyOuter.this
object Object1 extends MyOuter
object Object2 extends MyOuter
object MyObjectA extends Object1.MyInner
                     with Object2.MyInner
Scala 3 rejects MyObjectA with the error:
trait MyInner is extended twice
object MyObjectA cannot be instantiated since
  it has conflicting base types
  Object1.MyInner and Object2.MyInner
The rejection is not a surface-level restriction but a consequence of DOT’s type-theoretic foundations (Amin et al., 2016). In DOT, an object is constructed as $\nu(x:T)\,d$, where the self variable $x$ binds to a single object. Object1.MyInner and Object2.MyInner are distinct path-dependent types, each constraining the outer self-reference to a different object. Their intersection requires outer to simultaneously return Object1 and Object2—but DOT’s self variable is single-valued, making this intersection unrealizable.
The equivalent Overlay-Calculus definition is:
$\{\quad\mathrm{MyOuter}$ $\mapsto\{\mathrm{MyInner}\mapsto\{\mathrm{outer}\mapsto\mathrm{MyOuter}.\mathbf{this}\}\},$
$\mathrm{Object1}$ $\mapsto\{\mathrm{MyOuter}\},$
$\mathrm{Object2}$ $\mapsto\{\mathrm{MyOuter}\},$
$\mathrm{MyObjectA}$ $\mapsto\{\mathrm{Object1}.\mathrm{MyInner},\;\mathrm{Object2}.\mathrm{MyInner}\}\quad\}$
This is well-defined in Overlay-Calculus. Inside $\mathrm{MyInner}$, the reference $\mathrm{MyOuter}.\mathbf{this}$ has de Bruijn index $n=1$ (one scope level up from $\mathrm{MyInner}$ to $\mathrm{MyOuter}$). When $\mathrm{MyObjectA}$ inherits from both $\mathrm{Object1}.\mathrm{MyInner}$ and $\mathrm{Object2}.\mathrm{MyInner}$, the $\mathtt{this}$ function (equation 6) resolves $\mathrm{MyOuter}.\mathbf{this}$ by searching through $\mathtt{supers}(\mathrm{MyObjectA})$ for overlay paths matching $\mathrm{MyInner}$’s definition site. It finds two inheritance-site paths—one through $\mathrm{Object1}$ and one through $\mathrm{Object2}$—and returns both. Both paths lead to records that inherit from the same $\mathrm{MyOuter}$, so querying $\mathrm{MyObjectA}.\mathrm{outer}$ yields the $\mathtt{properties}$ of $\mathrm{MyOuter}$ regardless of which route is taken. Since inheritance is idempotent, the two routes contribute the same properties and no disambiguation is needed.
Appendix D Multi-Path Self-Reference in the NixOS Module System
This appendix demonstrates that the NixOS module system rejects the multi-path self-reference pattern with a static error. The translation uses the module system’s own abstractions: deferredModule for traits (unevaluated module values imported into other fixpoints) and submoduleWith for objects (evaluated in their own fixpoints).
let
  lib = (import <nixpkgs> { }).lib;
  result = lib.evalModules {
    modules = [
      (toplevel@{ config, ... }: {
        # class MyOuter { trait MyInner {
        #   def outer = MyOuter.this } }
        options.MyOuter = lib.mkOption {
          default = { };
          type = lib.types.deferredModuleWith {
            staticModules = [
              (MyOuter: {
                options.MyInner = lib.mkOption {
                  default = { };
                  type = lib.types.deferredModuleWith {
                    staticModules = [
                      (MyInner: {
                        options.outer = lib.mkOption {
                          default = { };
                          type =
                            lib.types.deferredModuleWith {
                              staticModules =
                                [ toplevel.config.MyOuter ];
        # object Object1 extends MyOuter
        options.Object1 = lib.mkOption {
          default = { };
          type = lib.types.submoduleWith {
            modules = [ toplevel.config.MyOuter ];
        # object Object2 extends MyOuter
        options.Object2 = lib.mkOption {
          default = { };
          type = lib.types.submoduleWith {
            modules = [ toplevel.config.MyOuter ];
        # object MyObjectA extends
        #   Object1.MyInner with Object2.MyInner
        options.MyObjectA = lib.mkOption {
          default = { };
          type = lib.types.submoduleWith {
            modules = [
              toplevel.config.Object1.MyInner
              toplevel.config.Object2.MyInner
in
  builtins.attrNames result.config.MyObjectA.outer
The NixOS module system rejects this with:
error: The option ‘MyObjectA.outer’
  in ‘<unknown-file>’
  is already declared
  in ‘<unknown-file>’.
The rejection occurs because Object1.MyInner and Object2.MyInner each carry the full staticModules of the MyInner deferred module, including the declaration options.outer. When MyObjectA imports both, the module system’s mergeOptionDecls encounters two declarations of the same option and raises a static error—it cannot express that both declarations originate from the same trait and should be unified.
The rejection is analogous to Scala’s “conflicting base types” error (Appendix C): both systems assume that each option (or type member) has a single declaration site, and reject the multi-path situation where two inheritance routes introduce the same declaration independently. In Overlay-Calculus, the same pattern is well-defined because $\mathtt{overlays}$ (equation 3) recognizes that both routes lead to the same definition, and $\mathtt{supers}$ (equation 2) collects both inheritance-site contexts without duplication.
Appendix E Well-Definedness of the Semantic Functions
This appendix proves that the mutually recursive equations (1)–(6) have a unique least fixed point, and therefore the semantics of Section 3 is well-defined.
E.1. The Product Lattice
Fix an AST with its primitive functions $\mathtt{defines}$ and $\mathtt{inherits}$. Let $\mathcal{L}$ denote the set of all labels and $\mathcal{P}$ the set of all paths (finite sequences of labels). Define the following domains, one for each semantic function:
$D_{\mathtt{properties}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{L})$ (sets of labels)
$D_{\mathtt{supers}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P}\times\mathcal{P})$ (sets of site–overlay pairs)
$D_{\mathtt{overlays}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P})$ (sets of paths)
$D_{\mathtt{bases}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P})$ (sets of paths)
$D_{\mathtt{resolve}}$ $=(\mathcal{P}\times\mathcal{P}\times\mathbb{N}\times\mathcal{L}^{*})\to\mathcal{P}(\mathcal{P})$ (sets of paths)
$D_{\mathtt{this}}$ $=(\mathcal{P}(\mathcal{P})\times\mathcal{P}\times\mathbb{N})\to\mathcal{P}(\mathcal{P})$ (sets of paths)
Each $D_{i}$ is a complete lattice under the pointwise subset ordering: $f\sqsubseteq g$ iff $f(x)\subseteq g(x)$ for all $x$. The product $D=D_{\mathtt{properties}}\times D_{\mathtt{supers}}\times D_{\mathtt{overlays}}\times D_{\mathtt{bases}}\times D_{\mathtt{resolve}}\times D_{\mathtt{this}}$ is a complete lattice under the componentwise ordering, with bottom element $\bot=(\lambda x.\,\varnothing,\ldots,\lambda x.\,\varnothing)$.
E.2. The Operator
Equations (1)–(6) define a function $F:D\to D$ that takes a tuple of candidate interpretations $(\mathtt{properties},\mathtt{supers},\mathtt{overlays},\mathtt{bases},\mathtt{resolve},\mathtt{this})$ and produces a new tuple by evaluating the right-hand side of each equation. The transitive closure $\mathtt{bases}^{*}$ appearing in equation (2) is derived from $\mathtt{bases}$: for any fixed $\mathtt{bases}$, the set $\mathtt{bases}^{*}(p)$ is defined as the least fixed point of the monotone operator $X\mapsto\{p\}\cup\bigcup_{q\in X}\mathtt{bases}(q)$ on $\mathcal{P}(\mathcal{P})$. If $\mathtt{bases}\sqsubseteq\mathtt{bases}^{\prime}$ (pointwise), then $\mathtt{bases}^{*}\sqsubseteq\mathtt{bases}^{\prime*}$ (pointwise), since enlarging the one-step relation can only enlarge the transitive closure.
The $\mathtt{this}$ function involves a recursion on the de Bruijn index $n$. Since $n$ decreases by one at each step and is a non-negative integer, this recursion terminates for any fixed interpretation of $\mathtt{supers}$. The result of $\mathtt{this}(S,p_{\mathrm{def}},n)$ for all valid $(S,p_{\mathrm{def}},n)$ is therefore well-defined as a function of $\mathtt{supers}$, and it remains to show that it is monotone in $\mathtt{supers}$.
E.3. Monotonicity
Proposition E.1.
The operator $F$ is monotone: if $\vec{d}\sqsubseteq\vec{d}^{\prime}$ in $D$, then $F(\vec{d})\sqsubseteq F(\vec{d}^{\prime})$.
Proof.
It suffices to show that each component of $F$ is monotone in every argument, i.e., that enlarging any of the input interpretations (in the $\subseteq$ ordering) can only enlarge the output. We verify this by inspecting the right-hand side of each equation and confirming that every occurrence of a mutually defined function appears in a positive position—as the domain of an existential quantifier in a set comprehension (Aczel, 1977).
Equation (1): $\mathtt{properties}$
$\mathtt{properties}(p)=\bigl\{\,\ell\bigm|(\_,p_{\mathrm{overlay}})\in\mathtt{supers}(p),\;\ell\in\mathtt{defines}(p_{\mathrm{overlay}})\,\bigr\}$. The function $\mathtt{supers}$ appears as the set being iterated over ($\in\mathtt{supers}(p)$). The predicate $\ell\in\mathtt{defines}(\cdot)$ involves only the AST primitive $\mathtt{defines}$, not any mutually defined function. Enlarging $\mathtt{supers}(p)$ can only add pairs, hence can only add labels to the result. Positive in $\mathtt{supers}$.
Equation (2): $\mathtt{supers}$
$\mathtt{supers}(p)=\bigl\{\,(\mathtt{init}(p_{\mathrm{base}}),p_{\mathrm{overlay}})\bigm|p_{\mathrm{base}}\in\mathtt{bases}^{*}(p),\;p_{\mathrm{overlay}}\in\mathtt{overlays}(p_{\mathrm{base}})\,\bigr\}$. Both $\mathtt{bases}^{*}$ and $\mathtt{overlays}$ appear as sets being iterated over. As noted above, $\mathtt{bases}^{*}$ is monotone in $\mathtt{bases}$. Enlarging either $\mathtt{bases}$ or $\mathtt{overlays}$ can only add elements to the result. Positive in $\mathtt{bases}$ and $\mathtt{overlays}$.
Equation (3): $\mathtt{overlays}$
For $p\neq()$:
$\mathtt{overlays}(p)=\{p\}\cup\bigl\{\,p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)\bigm|(\_,p_{\mathrm{branch}})\in\mathtt{supers}(\mathtt{init}(p)),\;\mathtt{last}(p)\in\mathtt{defines}(p_{\mathrm{branch}})\,\bigr\}.$
The function $\mathtt{supers}$ appears as the set being iterated over. The filtering condition $\mathtt{last}(p)\in\mathtt{defines}(\cdot)$ involves only the AST primitive. Enlarging $\mathtt{supers}$ can only add branches, hence can only add paths to the result. Positive in $\mathtt{supers}$.
Equation (4): $\mathtt{bases}$
$\mathtt{bases}(p)=\bigl\{\,p_{\mathrm{target}}\bigm|p_{\mathrm{overlay}}\in\mathtt{overlays}(p),\;(n,\ell_{\mathrm{down},*})\in\mathtt{inherits}(p_{\mathrm{overlay}}),\;p_{\mathrm{target}}\in\mathtt{resolve}(\ldots)\,\bigr\}.$
Both $\mathtt{overlays}$ and $\mathtt{resolve}$ appear as sets being iterated over. The function $\mathtt{inherits}$ is an AST primitive. Enlarging either $\mathtt{overlays}$ or $\mathtt{resolve}$ can only add targets to the result. Positive in $\mathtt{overlays}$ and $\mathtt{resolve}$.
Equation (5): $\mathtt{resolve}$
$\mathtt{resolve}(\ldots)=\bigl\{\,p_{\mathrm{current}}\mathbin{\triangleright}\ell_{\mathrm{down},1}\mathbin{\triangleright}\cdots\bigm|p_{\mathrm{current}}\in\mathtt{this}(\{p_{\mathrm{site}}\},\mathtt{init}(p_{\mathrm{def}}),n)\,\bigr\}.$
The function $\mathtt{this}$ appears as the set being iterated over. Enlarging $\mathtt{this}$ can only add paths to the result. Positive in $\mathtt{this}$.
Equation (6): $\mathtt{this}$
For $n>0$: $\mathtt{this}(S,p_{\mathrm{def}},n)=\mathtt{this}(S^{\prime},\mathtt{init}(p_{\mathrm{def}}),n-1)$ where $S^{\prime}=\bigl\{\,p_{\mathrm{site}}\bigm|p_{\mathrm{current}}\in S,\;(p_{\mathrm{site}},p_{\mathrm{overlay}})\in\mathtt{supers}(p_{\mathrm{current}}),\;p_{\mathrm{overlay}}=p_{\mathrm{def}}\,\bigr\}$. The function $\mathtt{supers}$ appears as the set being iterated over. The filtering condition $p_{\mathrm{overlay}}=p_{\mathrm{def}}$ is an equality test on path data, not involving any mutually defined function. Enlarging $\mathtt{supers}$ can only enlarge $S^{\prime}$, and by induction on $n$ (the base case $n=0$ returns $S$ unchanged), the recursive call is monotone in $S^{\prime}$. Positive in $\mathtt{supers}$.
In summary, every mutually defined function appears on the right-hand side of every equation only inside the pattern $x\in f(\ldots)$ within a set comprehension—i.e., as the domain of an existential quantifier. No complement, set difference, or negated membership test appears in any equation. Therefore $F$ is a monotone operator on the complete lattice $D$. ∎
E.4. Existence of the Least Fixed Point
Theorem E.2.
The semantic functions $\mathtt{properties}$, $\mathtt{supers}$, $\mathtt{overlays}$, $\mathtt{bases}$, $\mathtt{resolve}$, $\mathtt{this}$ are well-defined as the least fixed point of $F$.
Proof.
By Proposition E.1, $F$ is monotone on the complete lattice $D$. By the Knaster–Tarski theorem (Tarski, 1955), $F$ has a least fixed point $\vec{d}^{*}=\bigsqcup_{n\geq 0}F^{n}(\bot)$, where $\bot$ assigns the empty set to every input. The components of $\vec{d}^{*}$ are the semantic functions of Section 3.
Concretely, the least fixed point is the union of the finite approximation chain: $\bot\sqsubseteq F(\bot)\sqsubseteq F^{2}(\bot)\sqsubseteq\cdots$. A membership $\ell\in\mathtt{properties}(p)$ holds in $\vec{d}^{*}$ if and only if it appears at some finite stage $F^{k}(\bot)$, which corresponds to a finite derivation tree using the equations. When no finite derivation exists, the membership is absent from the least fixed point and the query is undefined—the analogue of divergence. ∎
Appendix F Böhm Tree Correspondence: Proofs
This appendix contains the full proofs of Adequacy (Theorem 4.2) and Full Abstraction (Theorem 4.3) stated in Section 4.2.
F.1. Böhm Trees
We briefly recall the definition of Böhm trees (Barendregt, 1984). A head reduction $M\to_{h}M^{\prime}$ contracts the outermost $\beta$-redex only: if $M$ has the form $(\lambda x.\,e)\;v\;M_{1}\cdots M_{k}$, then $M\to_{h}e[v/x]\;M_{1}\cdots M_{k}$. A term $M$ is in head normal form (HNF) if it has no head redex, i.e., $M=\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}$ where $y$ is a variable.
The Böhm tree $\mathrm{BT}(M)$ of a $\lambda$-term $M$ is an infinite labeled tree defined by:
$\mathrm{BT}(M)=\begin{cases}\bot&\text{if }M\text{ has no HNF}\\[6.0pt] \lambda x_{1}\ldots x_{m}.\;y\bigl(\mathrm{BT}(M_{1}),\;\ldots,\;\mathrm{BT}(M_{k})\bigr)&\text{if }M\to^{*}_{h}\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}\end{cases}$
Böhm tree equivalence $\mathrm{BT}(M)=\mathrm{BT}(N)$ is the standard observational equivalence of the lazy $\lambda$-calculus (Barendregt, 1984; Hyland and Ong, 2000).
F.2. Path Encoding
We define a correspondence between positions in the Böhm tree and paths in the overlay tree. When the translation $\mathcal{T}$ is applied to an ANF $\lambda$-term, the resulting overlay tree has a specific shape:
  • An abstraction $\lambda x.\,M$ translates to a record with own properties $\{\mathrm{argument},\mathrm{result}\}$ (the abstraction shape).
  • A $\mathbf{let}$-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M$ translates to a record with own properties $\{x,\mathrm{result}\}$, where $x$ holds the encapsulated application $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$.
  • A tail call $V_{1}\;V_{2}$ translates to a record with own properties $\{\_\mathrm{tail},\mathrm{result}\}$, where $\mathrm{tail}$ holds the encapsulated application and $\mathrm{result}$ projects $\mathrm{tail}.\mathrm{result}$.
  • A variable $x$ with de Bruijn index $n$ translates to an indexed reference $\uparrow^{\!n}.\mathrm{argument}$.
Each Böhm tree position is a sequence of navigation steps from the root. In the Böhm tree, the possible steps at a node $\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}$ are: entering the body (peeling off one abstraction layer) and entering the $i$-th argument $M_{i}$ of the head variable. In the overlay tree, these correspond to following the $\mathrm{result}$ label (for entering the body) and the $\mathrm{argument}$ label after $\mathtt{this}$ resolution (for entering an argument position).
Rather than comparing absolute paths across different overlay trees (which would be sensitive to internal structure), we observe only the convergence behavior accessible via the $\mathrm{result}$ projection. Recall the definition of overlay-convergence (Definition 4.1). Each $\mathrm{result}$-following step corresponds to one head reduction step in the $\lambda$-calculus: a tail call’s $\mathrm{result}$ projects through the encapsulated application, and a $\mathbf{let}$-binding’s $\mathrm{result}$ enters the continuation. The abstraction shape at the end signals that a weak head normal form has been reached.
F.3. Single-Path Lemma
The $\mathtt{this}$ function (equation 6) is designed for the general case of multi-path overlay inheritance. For translated $\lambda$-terms, we show it degenerates to single-path resolution.
Lemma F.1 (Single Path).
For any closed ANF $\lambda$-term $M$, at every invocation of $\mathtt{this}(S,\;p_{\mathrm{def}},\;n)$ during evaluation of $\mathcal{T}(M)$, the frontier set $S$ contains exactly one path.
Proof.
By induction on the structure of the ANF term.
Case $\mathcal{T}(\lambda x.\,M)=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
This is a record literal with $\mathtt{defines}=\{\mathrm{argument},\mathrm{result}\}$ and no inheritance sources, so $\mathtt{inherits}=\varnothing$. At the root path $p$, we have $\mathtt{overlays}(p)=\{p\}$ (a single path) since there are no inheritance sources to introduce additional branches. Inside $\mathcal{T}(M)$, any reference to $x$ is $\uparrow^{\!n}.\mathrm{argument}$ where $n$ is the de Bruijn index pointing to this scope level. Since $\mathcal{T}(\lambda x.\,M)$ introduces exactly one scope level, there is exactly one overlay at that level, hence $\mathtt{this}$ finds exactly one matching pair $(p_{\mathrm{site}},\;p_{\mathrm{overlay}})$ in $\mathtt{supers}$.
Case $\mathcal{T}(\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M)=\{x\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
The outer record has own properties $\{x,\mathrm{result}\}$ and no inheritance sources at the outer level ($\mathtt{inherits}=\varnothing$), so $\mathtt{overlays}(\text{root})=\{\text{root}\}$. Inside the $x$ subtree, the inheritance $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$ has exactly one inheritance source ($\mathcal{T}(V_{1})$), creating exactly one $\mathtt{bases}$ entry. By induction on $V_{1}$, the single-path property holds inside $\mathcal{T}(V_{1})$. Inside $\mathcal{T}(M)$, the induction hypothesis applies directly.
Case $\mathcal{T}(V_{1}\;V_{2})=\{\_\mathrm{tail}\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\_\mathrm{tail}.\mathrm{result}\}$
Identical to the $\mathbf{let}$-binding case: the outer record has own properties $\{\_\mathrm{tail},\mathrm{result}\}$ with no inheritance sources at the outer level, and the $\mathrm{tail}$ subtree contains exactly one inheritance source.
Case $\mathcal{T}(x)=\uparrow^{\!n}.\mathrm{argument}$
This is a reference, not a record. The $\mathtt{this}$ resolution walks up $n$ scope levels. By the inductive hypothesis on the enclosing translations, each step of $\mathtt{this}$ encounters exactly one matching pair in $\mathtt{supers}$.
The key structural invariant is that the translation $\mathcal{T}$ never produces two independent inheritance sources at the same scope level that both define the same label. In every case, application inheritances are encapsulated inside a named property ($x$ or $\mathrm{tail}$), and each such encapsulated inheritance has exactly one inheritance source. This prevents diamond inheritance and ensures single-path resolution throughout. ∎
F.4. Substitution Lemma
The core mechanism of the translation is that inheritance with $\mathrm{argument}\mapsto\mathcal{T}(V)$ plays the role of substitution. We make this precise.
Lemma F.2 (Substitution).
Let $M$ be an ANF term in which $x$ may occur free, and let $V$ be a closed value. Define the inherited tree $C=\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$. Then for every path $p$ under $C\mathbin{\triangleright}\mathrm{result}$ and every label $\ell$:
$\ell\in\mathtt{properties}(p)\text{ in }C\mathbin{\triangleright}\mathrm{result}\quad\Longleftrightarrow\quad\ell\in\mathtt{properties}(p)\text{ in }\mathcal{T}(M[V/x])$
where $M[V/x]$ is the usual capture-avoiding substitution.
Proof.
By structural induction on $M$.
Base case: $M=x$
Then $\mathcal{T}(x)=\uparrow^{\!n}.\mathrm{argument}$ (where $n$ is the de Bruijn index of $x$), and $C\mathbin{\triangleright}\mathrm{result}$ contains this reference. In the inherited tree $C$, $\mathtt{this}$ resolves the reference by walking up $n$ scope levels to the binding $\lambda$, where the inheritance $\{\mathcal{T}(\lambda x.\,x),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$ overlays the $\mathrm{argument}$ slot with $\mathcal{T}(V)$. Since $C$ is the $\mathrm{tail}$ subtree of $\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)$, which is a closed ANF term, Lemma F.1 applies and $\mathtt{this}$ finds exactly one path, so $\mathtt{resolve}$ returns $\mathcal{T}(V)$’s subtree. Since $M[V/x]=V$, we have $\mathcal{T}(M[V/x])=\mathcal{T}(V)$, and the properties coincide.
Base case: $M=y$ where $y\neq x$
Then $\uparrow^{\!m}.\mathrm{argument}$ (where $m$ is the de Bruijn index of $y$, $m\neq n$) walks up to a different scope level, so the inheritance with $\mathrm{argument}\mapsto\mathcal{T}(V)$ at the binding scope of $x$ has no effect. Since $M[V/x]=y$, the properties are identical.
Inductive case: $M=\lambda y.\,M^{\prime}$
Then $\mathcal{T}(M)=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$. The subtree at $C\mathbin{\triangleright}\mathrm{result}$ is another abstraction-shaped record. References to $x$ inside $\mathcal{T}(M^{\prime})$ resolve through $\mathtt{this}$ in exactly the same way (with one additional scope level from the inner $\lambda$), and by induction on $M^{\prime}$, the properties under $C\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}\mathrm{result}$ coincide with those of $\mathcal{T}(M^{\prime}[V/x])$. Since $(\lambda y.\,M^{\prime})[V/x]=\lambda y.\,(M^{\prime}[V/x])$ (assuming $y$ is fresh), the result follows.
Inductive case: $M=\mathbf{let}\;z=V_{1}\;V_{2}$ $\mathbf{in}\;M^{\prime}$
Then $\mathcal{T}(M)=\{z\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$. The substitution distributes: $M[V/x]=\mathbf{let}\;z=V_{1}[V/x]\;V_{2}[V/x]\;\mathbf{in}\;M^{\prime}[V/x]$. By induction on $V_{1}$, $V_{2}$, and $M^{\prime}$, the properties under each subtree ($z$ and $\mathrm{result}$) coincide, since references to $x$ inside each subtree resolve to $\mathcal{T}(V)$ via the same $\mathtt{this}$ mechanism.
Inductive case: $M=V_{1}\;V_{2}$ (tail call)
Analogous to the $\mathbf{let}$-binding case with $\mathrm{tail}$ in place of $z$. ∎
F.5. Convergence Preservation
The convergence criterion (Definition 4.1) follows the $\mathrm{result}$ chain from the root. Each step in this chain corresponds to one head reduction step in the $\lambda$-calculus. We make this precise.
Lemma F.3 (Result-step).
Let $(\lambda x.\,M)\;V$ be a tail call (a $\beta$-redex). Then for every path $p$ and label $\ell$:
$\ell\in\mathtt{properties}\!\bigl(\,\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p\,\bigr)\text{ in }\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)\;\Longleftrightarrow\;\ell\in\mathtt{properties}\!\bigl(\,\text{root}\mathbin{\triangleright}p\,\bigr)\text{ in }\mathcal{T}(M[V/x])$
That is, following one $\mathrm{result}$ projection in the tail call’s overlay tree yields the same properties as the root of the reduct’s overlay tree.
Proof.
The translation gives:
$\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)=\{\_\mathrm{tail}\mapsto\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\},\;\mathrm{result}\mapsto\_\mathrm{tail}.\mathrm{result}\}$
The $\mathrm{result}$ label at the root is defined as $\mathrm{tail}.\mathrm{result}$, so $\text{root}\mathbin{\triangleright}\mathrm{result}$ resolves to the $\mathrm{result}$ path inside the inheritance $\mathrm{tail}=\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$. This inheritance is exactly the inherited tree $C$ of Lemma F.2, and $\mathrm{tail}\mathbin{\triangleright}\mathrm{result}$ is $C\mathbin{\triangleright}\mathrm{result}$. By Lemma F.2, the properties under $C\mathbin{\triangleright}\mathrm{result}$ coincide with those of $\mathcal{T}(M[V/x])$.
Let-binding variant. An analogous result holds for $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}$ where $V_{1}=\lambda y.\,M^{\prime\prime}$. The translation gives:
$\mathcal{T}\bigl(\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}\bigr)=\{x\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$
Since $\mathrm{result}\mapsto\mathcal{T}(M^{\prime})$ is a property definition, the subtree at $\text{root}\mathbin{\triangleright}\mathrm{result}$ is $\mathcal{T}(M^{\prime})$. Inside $\mathcal{T}(M^{\prime})$, each reference to $x$ is $x.\mathrm{result}$, which resolves to the $\mathrm{result}$ path inside the $x$ subtree $\{\mathcal{T}(\lambda y.\,M^{\prime\prime}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$. This is the inherited tree $C$ of Lemma F.2, so $x.\mathrm{result}$ has the same properties as $\mathcal{T}(M^{\prime\prime}[V_{2}/y])$. In the $\lambda$-calculus, $M=(\lambda x.\,M^{\prime})(V_{1}\;V_{2})$, and the head reduct is $N=M^{\prime}[(V_{1}\;V_{2})/x]$. In $\mathcal{T}(N)$, each occurrence of $x$ is replaced by $(V_{1}\;V_{2})$, whose result, by the tail-call variant above, also yields $\mathcal{T}(M^{\prime\prime}[V_{2}/y])$. Therefore the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p$ in $\mathcal{T}(M)$ coincide with those at $\text{root}\mathbin{\triangleright}p$ in $\mathcal{T}(N)$ for all $p$. ∎
Theorem F.4 (Convergence preservation).
If $M\to_{h}N$ (head reduction), then $\mathcal{T}(M){\Downarrow}$ if and only if $\mathcal{T}(N){\Downarrow}$.
Proof.
A head reduction step in ANF takes one of two forms: a tail call $(\lambda x.\,M^{\prime})\;V\to_{h}M^{\prime}[V/x]$, or a $\mathbf{let}$-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}\to_{h}M^{\prime}[(V_{1}\;V_{2})/x]$ (since $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}=(\lambda x.\,M^{\prime})(V_{1}\;V_{2})$ in the $\lambda$-calculus). In both cases, Lemma F.3 (tail-call and let-binding variants respectively) shows that the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}^{n}$ in $\mathcal{T}(M)$ coincide with the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}^{n-1}$ in $\mathcal{T}(N)$ for all $n\geq 1$, where $N$ is the head reduct. Hence the abstraction shape appears at depth $n$ in the pre-reduct if and only if it appears at depth n1n-1 in the post-reduct. ∎
F.6. Adequacy
A closed ANF term at the top level is either an abstraction (a value) or a computation (a $\mathbf{let}$-binding or tail call). An abstraction translates to a record whose root immediately has the abstraction shape; a computation’s root has the form $\{x,\mathrm{result}\}$ or $\{\_\mathrm{tail},\mathrm{result}\}$, and one must follow the $\mathrm{result}$ chain to find the eventual value.
Proof of Theorem 4.2.
($\Rightarrow$) Suppose $M\to^{*}_{h}\lambda x.\,M^{\prime}$ in $k$ head reduction steps. We show $\mathcal{T}(M)$ converges at depth $\leq k$ by induction on $k$.
Base case ($k=0$): $M$ is already an abstraction $\lambda x.\,M^{\prime}$. Then $\mathcal{T}(M)=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$, whose root has $\mathtt{defines}=\{\mathrm{argument},\mathrm{result}\}$. These labels appear at stage $F^{1}(\bot)$ of the approximation chain, so $\mathcal{T}(M)$ converges at depth $n=0$.
Inductive step ($k\geq 1$): $M$ is not an abstraction, so $M$ is either a tail call or a $\mathbf{let}$-binding, both of which are $\beta$-redexes. Let $N$ be the head reduct, so $M\to_{h}N\to^{k-1}_{h}\lambda x.\,M^{\prime}$. By Lemma F.3 (tail-call or let-binding variant), the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p$ in $\mathcal{T}(M)$ coincide with those at $\text{root}\mathbin{\triangleright}p$ in $\mathcal{T}(N)$. By the induction hypothesis, $\mathcal{T}(N)$ converges at depth $\leq k-1$, so $\mathcal{T}(M)$ converges at depth $\leq k$.
($\Leftarrow$) Suppose $\mathcal{T}(M){\Downarrow}$ at depth $n$. We show $M$ has a head normal form by induction on $n$.
Base case ($n=0$): the root of $\mathcal{T}(M)$ has abstraction shape $\{\mathrm{argument},\mathrm{result}\}$. Only the abstraction rule $\mathcal{T}(\lambda x.\,M^{\prime})$ produces a root with own property $\mathrm{argument}$. (For tail calls and $\mathbf{let}$-bindings, the root has $\mathtt{defines}=\{\_\mathrm{tail},\mathrm{result}\}$ or $\{x,\mathrm{result}\}$, and $\mathrm{argument}$ cannot appear via $\mathtt{supers}$ because $\mathtt{inherits}=\varnothing$ at the root, giving $\mathtt{bases}(\text{root})=\varnothing$.) Therefore $M$ is an abstraction, hence already in head normal form.
Inductive step ($n\geq 1$): the abstraction shape appears at depth $n$ but not at depth 0. By the base case argument, $M$ is not an abstraction, so $M$ is either a tail call $V_{1}\;V_{2}$ or a $\mathbf{let}$-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}$. In both cases, $M$ is a $\beta$-redex (since $V_{1}=\lambda x.\,M^{\prime}$ or $V_{1}=\lambda y.\,M^{\prime\prime}$ respectively, the only closed ANF value). By Lemma F.3 (tail-call or let-binding variant respectively), the properties at depth n1n-1 in $\mathcal{T}(N)$ match those at depth $n$ in $\mathcal{T}(M)$, where $N$ is the head reduct. So $\mathcal{T}(N)$ converges at depth n1n-1. By the induction hypothesis, $N$ has a head normal form, and since $M\to_{h}N$, so does $M$. ∎
F.7. Full Abstraction
Adequacy relates a single term’s head normal form to its overlay-convergence. Full abstraction lifts this to an equivalence between terms: the translation $\mathcal{T}$ preserves and reflects observational equivalence.
Definition F.5 (Overlay-contextual equivalence).
For closed $\lambda$-terms $M$ and $N$, write $\mathcal{T}(M)\approx\mathcal{T}(N)$ if for every closing $\lambda$-calculus context $C[\cdot]$:
$\mathcal{T}(C[M]){\Downarrow}\;\Longleftrightarrow\;\mathcal{T}(C[N]){\Downarrow}$
This definition observes only convergence (a single bit), but quantifying over all contexts makes it a fine-grained equivalence: different contexts can supply different arguments and project different results, probing every aspect of the term’s behavior.
Proof of Theorem 4.3.
The translation $\mathcal{T}$ is compositional: for every $\lambda$-calculus context $C[\cdot]$, the overlay tree $\mathcal{T}(C[M])$ is determined by the overlay tree $\mathcal{T}(M)$ and the translation of the context. This means $\mathcal{T}$ preserves the context structure needed for the argument.
($\Leftarrow$) Suppose $\mathrm{BT}(M)=\mathrm{BT}(N)$. Böhm tree equivalence is a congruence (Barendregt, 1984), so for every context $C[\cdot]$, $\mathrm{BT}(C[M])=\mathrm{BT}(C[N])$. In particular, $C[M]$ has a head normal form iff $C[N]$ does. By Theorem 4.2, $\mathcal{T}(C[M]){\Downarrow}$ iff $\mathcal{T}(C[N]){\Downarrow}$. Hence $\mathcal{T}(M)\approx\mathcal{T}(N)$.
($\Rightarrow$) Suppose $\mathrm{BT}(M)\neq\mathrm{BT}(N)$. Böhm tree equivalence coincides with observational equivalence for the lazy $\lambda$-calculus (Barendregt, 1984; Hyland and Ong, 2000): there exists a context $C[\cdot]$ such that $C[M]$ converges and $C[N]$ diverges (or vice versa). By Theorem 4.2, $\mathcal{T}(C[M]){\Downarrow}$ and $\mathcal{T}(C[N]){\Uparrow}$. Hence $\mathcal{T}(M)\not\approx\mathcal{T}(N)$. ∎
The proof rests on two pillars: Adequacy (Theorem 4.2), which is our contribution, and the classical result that Böhm tree equivalence equals observational equivalence for the lazy $\lambda$-calculus (Barendregt, 1984; Hyland and Ong, 2000). The overlay tree semantics inherits full abstraction from the $\lambda$-calculus without needing to define or compare absolute paths inside overlay trees.
Appendix G CPS-Agnostic Programs
In the $\lambda$-calculus, a program is either in continuation-passing style (CPS) or direct style. Converting between these representations requires explicit transformation—the CPS conversion (Plotkin, 1975; Sabry and Felleisen, 1993). In Overlay-Calculus, the distinction does not exist: the same program text is CPS or direct-style depending on how dependencies are assembled.
Consider a generic computation step:
$\mathrm{Step}\mapsto\{\mathrm{input}\mapsto\{\},\;\mathrm{process}\mapsto\{\},\;\mathrm{result}\mapsto\{\}\}$
The empty $\mathrm{process}$ slot is an interpretation point. If filled externally by an orchestrator, it acts as a continuation parameter (CPS interpretation). If inherited locally with an implementation, it acts as a direct dependency (direct-style interpretation). The $\mathrm{Step}$ code remains unchanged.
CPS interpretation via external orchestration
When $\mathrm{process}$ is injected from outside, the computation passes its intermediate result to an external handler:
$\{\mathrm{Step}.\mathbf{this},\;$
$\quad\mathrm{input}\mapsto\{\mathrm{data}\mapsto\{\}\},\;$
$\quad\mathrm{process}\mapsto\{$
$\qquad\mathrm{result}\mapsto\{\mathrm{handled}\mapsto\{\}\}$
$\quad\}$
The $\mathrm{process}$ slot acts as a continuation: it receives the result and determines what happens next. This is CPS-style programming.
Direct-style interpretation via local inheritance
Alternatively, $\mathrm{process}$ can be filled by local inheritance:
$\{\mathrm{Step}.\mathbf{this},\;$
$\quad\mathrm{input}\mapsto\{\mathrm{data}\mapsto\{\}\},\;$
$\quad\mathrm{process.result}\mapsto\{\mathrm{computed}\mapsto\{\}\}$
Here $\mathrm{process}$ is implemented directly within the inheritance. This is direct-style programming. The $\mathrm{Step}$ code is identical in both cases—only the assembly strategy differs.
Representation independence
This property is a form of representation independence (Reynolds, 1983) applied to control flow: client code using $\mathrm{Step}$ need not know whether $\mathrm{process}$ will be filled continuation-style or directly. Library authors write code once; users choose the control flow style at assembly time.
The $\lambda$-calculus lacks this flexibility. A function written in direct style must be explicitly transformed (CPS conversion) before it can be used continuation-style. Overlay-Calculus programs are inherently in A-normal form (Flanagan et al., 1993), where every intermediate result has a name.
However, Overlay-Calculus ANF differs critically from ANF in functional programming. In functional language compilers, ANF conversion requires specifying a target monad—either a concrete monad or an abstract one (Moggi, 1991). Compiler-internal ANF transformations (the ANF form used in intermediate representations) typically target the identity monad, which cannot express CPS transformations. This is analogous to how imperative compilers use SSA (Static Single Assignment) form, which similarly lacks control flow flexibility. Overlay-Calculus ANF, by contrast, has built-in dependency injection: empty API slots can be filled by external orchestrators. This capability enables the same ANF program to serve both CPS and direct-style interpretations without monad specification or transformation.
Multiple symmetric continuations
When interpreted continuation-style, Overlay-Calculus naturally supports multiple exit points:
$\mathrm{Choice}\mapsto\{\mathrm{on\_success}$ $\mapsto\{\},\;$
$\mathrm{on\_failure}$ $\mapsto\{\},\;$
$\mathrm{on\_timeout}$ $\mapsto\{\},\;$
$\mathrm{result}$ $\mapsto\{\}\}$
All continuation slots are symmetric elements of the same record. Traditional CPS in the $\lambda$-calculus distinguishes a single continuation parameter; additional continuations must be encoded (e.g., as pairs or disjoint sums). Overlay-Calculus treats all continuation slots uniformly, eliminating the need for specialized encodings.
Abstract and concrete are syntactically identical
From a category-theoretic perspective, abstract algebra and concrete implementation have the same syntactic structure in Overlay-Calculus:
Abstract: $\mathrm{API}\mapsto\{\}$
Concrete: $\mathrm{API}\mapsto\{\text{implementation}\}$
This is the abstract/concrete correspondence: the same functor applies to both. In traditional languages, abstract interfaces (e.g., Java interfaces, Haskell type classes) have different syntax from concrete implementations. In Overlay-Calculus, they are syntactically identical—only the interpretation differs. An empty API is simultaneously an abstract requirement (when viewed as a dependency) and a concrete extension point (when viewed as an injection site).
The Selector and Rebuilder combinators (Appendix A) demonstrate this flexibility: their callback slots ($\mathrm{on\_zero}$, $\mathrm{on\_odd}$, $\mathrm{on\_even}$) can be filled externally (CPS interpretation) or inherited locally (direct interpretation).
Practical implications
A library component with empty API slots can be used in multiple contexts without modification:
  • In synchronous code, fill slots with direct implementations
  • In asynchronous code, fill slots with continuation handlers
  • In effect systems, compose with algebraic effect handlers (Plotkin and Power, 2003)
  • In test code, fill slots with mock implementations
No dual APIs are needed. The same component serves all use cases. This eliminates a common source of code duplication in traditional languages, where async and sync versions of the same functionality must be maintained separately.
The implementation (Yang, 2025b) includes executable examples of all three patterns described above: escape continuations analogous to call/cc, success/failure continuations analogous to exception handling, and protected computations with multiple symmetric exit points. Each example includes test cases verifying that the correct continuation path is taken.
CPS-agnosticism is not a designed feature of Overlay-Calculus; it is an emergent property of records and inheritance. The calculus defines no notion of continuation—the three syntax constructs and six semantic functions of Section 3 are the entire language.
Appendix H Asymmetry in Problem Difficulty
A skeptical reader might ask whether Overlay-Calculus is simply the $\lambda$-calculus with different syntax. The test of whether two computational models are “the same” is whether they make the same problems easy and the same problems hard. They do not:
System Expression Problem Random-access memory
$\lambda$-calculus Object algebras (d. S. Oliveira and Cook, 2012), Explicit trie encoding
finally tagless (Carette et al., 2009)
SKI combinators Same difficulty as $\lambda$ Same difficulty as $\lambda$
Overlay-Calculus Direct inheritance Records are tries
Overlay-Calculus requires no additional machinery for either problem class; the $\lambda$-calculus requires elaborate encodings for both.
The semantic domain of Overlay-Calculus (Section 3) is entirely first-order: paths are label sequences, overlays are the record trees produced by evaluation, AST nodes are pure data, and the resolution functions ($\mathtt{overlays}$, $\mathtt{supers}$, $\mathtt{resolve}$, etc.) are ordinary set-valued functions over these structures. No function types appear in the object language or its semantics. The crucial question is not what notation is used, but what computational patterns are primitive in the language itself.
In Overlay-Calculus, the primitive operations are record construction, qualified-this reference, and inheritance. Function application is derived from these primitives (Section 4), not primitive itself. The asymmetry of Section 4.1 makes this concrete: embedding the $\lambda$-calculus into Overlay-Calculus requires five rules; the reverse requires sophisticated encoding.
Why tree-level inheritance dissolves the Expression Problem.
Each operation is an independent overlay that extends a factory’s subtree. Inheritance recursively merges subtrees that share a label, so every entity in the factory acquires all operations from all inherited overlays. Crucially, this applies to return types: if operation $A$ returns values from factory $F$, and an independently defined overlay $B$ adds a property to $F$, then inheriting from $A$ and $B$ causes the values returned by $A$ to carry that property automatically—without modifying either $A$ or $B$. In object algebras (d. S. Oliveira and Cook, 2012) and finally tagless interpreters (Carette et al., 2009), independently defined operations do not automatically enrich each other’s return types; additional boilerplate or type-class machinery is required.
Multi-path qualified this (equation 6) is essential to this dissolution—not merely to solving the Expression Problem (for which many solutions exist (d. S. Oliveira and Cook, 2012; Carette et al., 2009)), but to making it not arise in the first place. In the case study of Section 5, $\mathrm{evaluation}.\mathrm{Constant}$ and $\mathrm{display}.\mathrm{Constant}$ both inherit $\mathrm{expression}.\mathrm{Constant}$. When the two operations are inherited, the resulting $\mathrm{Constant}$ inherits the $\mathrm{expression}$ schema through two independent routes. The reference $\mathrm{Constant}.\mathbf{this}.\mathrm{magnitude}$ inside each operation resolves through both routes, and since inheritance is idempotent, all routes contribute the same properties. In Scala, the same pattern—two independently defined inner classes extending the same outer class’s trait—triggers a “conflicting base types” rejection (Appendix C), because DOT’s self variable is single-valued and cannot resolve through multiple inheritance routes simultaneously. The Expression Problem can still be solved in Scala using object algebras or similar patterns, but it cannot be made to not arise: free inheritance of independently defined operations inherently produces multi-path inheritance to shared schemas, which DOT’s single-valued self variable cannot accommodate.