### v2

# A Calculus of Overlays

Abstract.
Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) as the foundation of functional programming, overlay-calculus uses three primitives (record, definition, inheritance) as the foundation of declarative programming. It trivially embeds the $\lambda$-calculus, although the entire semantics rests solely on naive set theory; as a consequence, all constructs including inheritance are inherently commutative, idempotent, and associative; the linearization problem of multiple inheritance does not arise. This induces a fully abstract semantics of the lazy $\lambda$-calculus with respect to Böhm tree equivalence (Barendregt, 1984). Overlay-calculus is distilled from the Overlay language, a practical implementation in which we observed further emergent phenomena: the Expression Problem dissolves, programs are function color blind (Nystrom, 2015), ordinary arithmetic yields the relational semantics of logic programming, and self-reference resolves to multiple targets, making overlay-calculus strictly more expressive than the $\lambda$-calculus in Felleisen’s sense (Felleisen, 1991). These properties suggest applications to configuration languages, dependency injection, object-oriented programming, composable effect systems, modular software architectures, file-system-as-compiler, general-purpose programming, and no-code development.
declarative programming, inheritance, denotational semantics, self-referential records, expression problem, $\lambda$-calculus, Böhm trees, configuration languages
This work was conducted independently prior to the author’s employment at Figure AI.
$c$s: Theory of computation Denotational semanticsccs: Theory of computation Object oriented constructsccs: Theory of computation Functional constructsccs: Theory of computation Program schemesccs: Software and its engineering Formal language definitionsccs: Software and its engineering Extensible languagesccs: Software and its engineering Object oriented languages
1. Introduction
Declarative and configuration languages are ubiquitous in modern software engineering. Systems such as NixOS modules (Dolstra et al., 2010; NixOS Contributors, 2024), Jsonnet (Cunningham, 2014), Hydra (Yadan and Droppo, 2023), CUE (van Lohuizen, 2019), Dhall (Gonzalez, 2017), Kustomize (Kubernetes SIG CLI, 2018), and JSON Patch (Bryan et al., 2013) all provide mechanisms for composing structured data through inheritance or overlay. Among these, the NixOS module system stands out: its recursive attribute set merging with fixed-point semantics and deferred modules (NixOS Contributors, 2024) achieves considerable expressiveness in practice, and the mechanism has been adopted well beyond NixOS itself: Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), disko (Lassulus and nix-community Contributors, 2022), flake-parts (Hensing, 2022), dream2nix (Hauer, 2021), devenv (Kožar, 2022), KubeNix (Xtruder and Hall, 2022), and nixidy (Ingólfsson, 2024) use it to manage user environments, macOS configuration, disk partitioning, flake structure, multi-language packaging, developer environments, and Kubernetes clusters respectively. Yet no computational theory explains why this mechanism is so expressive.
The $\lambda$-calculus serves as the foundational computational model for functional programming. No analogous calculus exists for declarative programming. This gap matters because the two paradigms differ in fundamental ways. Configuration languages are inherently declarative: their values are immutable and they have no first-class functions. At first glance, Turing completeness appears incompatible with these constraints. The three classical models of computation each violate at least one of them:
Model Turing complete Immutable No $\lambda$
Turing Machine $\checkmark$ $\times$ $\checkmark$
$\lambda$-calculus $\checkmark$ $\checkmark$ $\times$
RAM Machine $\checkmark$ $\times$ $\checkmark$
The Turing Machine and RAM Machine require mutable state; the $\lambda$-calculus requires first-class functions. A computational model that is Turing complete, immutable, and function-free would sit in the empty cell of this table. The NixOS module system already suggests that such a model exists: its inheritance-based composition over recursive records, without explicit functions, is expressive enough to configure entire operating systems.
Conventionally, the value domain of configuration languages is assumed to consist of finite, well-founded structures, that is, initial algebras in the sense of universal algebra. We challenge this assumption: configuration values are better understood as lazily observable, possibly infinite structures, in the spirit of F-coalgebras (Rutten, 2000), where semantics is determined by the observer and a finite prefix suffices for any finite observation. This is not merely a theoretical distinction: the nixpkgs package collection (Dolstra, 2006) (over 100,000 packages) behaves as a lazily observed structure in practice, allowing any single package to be evaluated without materializing the entire set. Guided by this observation, we set out to reduce the NixOS module system to a minimal set of primitives. Through iterative refinement, beginning with an early version implemented in Nix, the system evolved into the Overlay language (Yang, 2025b), now implemented in Python: a declarative programming language with only three constructs (record literals, properties, and inheritance) and no functions or let-bindings.
In developing real programs in the Overlay language, including boolean logic and natural number arithmetic (Section 5) and a web application (Section 5), we found that the linearization problem of mixin-based systems (Bracha and Cook, 1990; Barrett et al., 1996) is structurally absent: because records are sets, inheritance is inherently commutative, idempotent, and associative (Section 3). Self-reference naturally resolves to multiple objects rather than one; prior systems assume single-object resolution and reject the multi-path situation (Scala, NixOS module system). Self-referential records require no fixed-point combinator—the inheritance tree is the call stack, and qualified this resolution is a first-order set-theoretic query, not a fixed-point construction. We also found that boolean logic and arithmetic, implemented as separate overlay files, extend in both dimensions of the Expression Problem (Wadler, 1998) without any dedicated mechanism; that Church-encoded values automatically form trie structures; and that operations on tries automatically become Cartesian products, so that adding $\{1,2\}$ to $\{3,4\}$ yields $\{4,5,6\}$, which is the relational semantics of logic programming (Section 5); and that the same web application overlay file runs under both Python asyncio and synchronous execution without modification, an instance of function color blindness (Nystrom, 2015), equivalently CPS-agnostic behavior (Section 5).
This paper distills these observations into overlay-calculus, a minimal computational model for declarative programming. Function application translates directly into inheritance: a $\lambda$-abstraction corresponds to a record with an $\mathrm{argument}$ slot and a $\mathrm{result}$ slot, and function application corresponds to inheriting from the function record and a record that supplies the argument. Since the $\lambda$-calculus embeds into overlay-calculus, overlay-calculus is Turing complete. The translation reveals an asymmetry: embedding the $\lambda$-calculus into overlay-calculus requires five local, syntax-directed rules (Section 4), but the reverse is impossible—overlay enables open extension that changes operational equivalence on $\lambda$-terms, establishing a strict separation in Felleisen’s sense of expressiveness (Felleisen, 1991) (Section 4.2).
Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) to serve as the foundation of functional programming, overlay-calculus uses three primitives (record, definition, inheritance) to serve as the foundation of declarative programming.
All constructions in this paper have been implemented and tested in the Overlay language, available as open source (Yang, 2025b), with a test suite covering every example. [^1] The Overlay language is statically typed and therefore stricter than the untyped overlay-calculus. All examples in this paper are ported from the Overlay language and conform to both semantics. Elements added to satisfy the type checker are harmless for the algorithms presented here and may aid readability. All tests are mechanically verified against expected output snapshots.
Contributions
  • We present overlay-calculus, a minimal computational model for declarative programming that contains only overlay constructs (record, definition, inheritance) and no functions or scalar types.
  • Inheritance operates at the tree level: a record is a set of elements, both definitions and inheritance sources, so inheriting from two overlays that share a label recursively merges their subtrees (deep merge), rather than overriding at the method level as in mixin and trait calculi. Since records are sets, inheritance is inherently commutative, idempotent, and associative. The linearization problem inherent in prior mixin and trait calculi does not arise.
  • The semantics is purely observational: there are no reduction rules, no evaluation order, and no notion of a result that a term reduces to. The observer drives computation by querying paths in a lazily constructed tree.
  • Self-referential records require no fixed-point combinator. The $\lambda$-calculus translation maps each application to a inheritance; the resulting inheritance tree is the call stack. The $\mathtt{this}$ function of Section 3 resolves self-references by walking this tree: a first-order set-theoretic query, not a fixed-point construction.
  • Self-reference resolves to multiple objects, not one. Path-dependent types with inheritance naturally produce this situation; prior systems assume single-object resolution and reject it (Scala and the NixOS module system, Appendix B). Since inheritance is idempotent, all objects contribute equally; no disambiguation is needed.
  • The $\lambda$-calculus (in A-normal form) translates into overlay-calculus in five rules (Section 4). The translation is fully abstract with respect to Böhm tree equivalence (Section 4.1), yielding a semantics of the lazy $\lambda$-calculus that requires no domain theory.
  • We observe further properties: extensibility in the sense of the Expression Problem (Wadler, 1998) (Section 5), relational semantics of logic programming from ordinary arithmetic (Section 5), and function color blindness (Nystrom, 2015) (Section 5).
  • Open extension via overlay changes the operational equivalence of $\lambda$-terms, establishing via Felleisen’s Theorem 3.14 (Felleisen, 1991) that the $\lambda$-calculus cannot macro-express overlay (Section 4.2).
Section 5 demonstrates these properties through a case study of Nat arithmetic.
2. Syntax
Let $e$ denote an expression. Let $\ell$ denote a label (property name). Let $k$ denote a non-negative integer (path length).
$e\quad::=\quad$ $\{c_{1},\;\ldots,\;c_{m}\}$ $\text{(}m\geq 0\text{)}$
$\mid\quad$ $c$ $\text{(sugar for }\{c\}\text{)}$
$c\quad::=\quad$ $\ell\mapsto e$ (definition)
$\mid\quad$ $[\ell_{\mathrm{up}}.\mathbf{this}.\mid\uparrow^{\!n}.]\;\ell_{\mathrm{down},1}.\ell_{\mathrm{down},2}\ldots\ell_{\mathrm{down},k}$ $\text{(inheritance, }k\geq 0\text{ with qualifier,}$
$k\geq 1\text{ without)}$
An expression is a set of elements enclosed in braces. Each element is either a definition $\ell\mapsto e$ or an inheritance source. Because expressions are sets, element order is irrelevant and duplicates have no effect; inheritance is inherently commutative, idempotent, and associative.
The $\mapsto$ in $\ell\mapsto e$ defines a property, not a let-binding. There is no variable binding in overlay-calculus. The Overlay language also provides scalar types through a foreign-function interface (FFI); overlay-calculus is what remains after removing the FFI. None of the examples in this paper uses the FFI; scalars did not arise.
Inheritance
An inheritance is written as a reference that navigates the scope hierarchy upward and then projects properties downward. The upward navigation admits two equivalent notations:
  • Named form: $\ell_{\mathrm{up}}.\mathbf{this}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$, analogous to Java’s Outer.this.field. Each $\{\ldots\}$ record has a label in its enclosing scope; $\ell_{\mathrm{up}}$ names the target enclosing scope.
  • Indexed form: $\uparrow^{\!n}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$, where $n\geq 0$ is a de Bruijn index (de Bruijn, 1972): $n=0$ refers to the reference’s enclosing scope, and each increment of $n$ moves one scope level further out.
Both forms retrieve the target scope, that is, the fully inherited record after all inheritance has been applied, and then project properties $\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$ from it. The named form desugars to the indexed form during parsing (Section 3).
A new scope level is created at each $\{\ldots\}$ record literal. The scope contains all properties of that overlay, including those inherited via inheritance sources. For example, $\{a\mapsto\{\},\;b\mapsto\{\}\}$ is a single overlay whose scope contains both $a$ and $b$; so is $\{r,\;a\mapsto\{\}\}$ where $r$ is a reference to a record containing $b$. Inheritance sources do not create additional scope levels.
  • $\mathrm{Foo}.\mathbf{this}$ (equivalently, $\uparrow^{\!n}$ where $\mathrm{Foo}$ is $n$ steps above the enclosing scope) retrieves the record named $\mathrm{Foo}$.
  • $\mathrm{Foo}.\mathbf{this}.\ell$ (equivalently, $\uparrow^{\!n}.\ell$) projects property $\ell$ from $\mathrm{Foo}$.
  • When $k=0$, the reference inherits the entire enclosing scope, which may produce an infinitely deep tree.
When the qualifier is unambiguous, it may be omitted: the shorthand $\ell_{\mathrm{down},1}.\ell_{\mathrm{down},2}\ldots\ell_{\mathrm{down},k}$ ($k\geq 1$) resolves $\ell_{\mathrm{down},1}$ against the innermost enclosing scope whose $\mathtt{defines}$ contains it, and desugars to the corresponding indexed form. Since this uses $\mathtt{defines}$, the shorthand can only access properties defined in record literals, not those inherited through inheritance; one of the two qualified forms is required for inherited properties or to bypass variable shadowing.
3. Overlay Trees
Given an AST, what are the properties visible at each position in the tree? This section answers the question by defining a collection of mutually recursive functions over paths.
Unlike the $\lambda$-calculus, which requires reduction rules ($\beta$-reduction) to specify how terms compute, overlay-calculus has no reduction steps. There is no rewriting of terms, no evaluation order, and no notion of a “result” that a term reduces to. Instead, the semantics is purely observational: one queries whether a label $\ell$ is a property at a path $p$ in the inheritance tree. The mutually recursive functions defined below, namely $\mathtt{properties}$, $\mathtt{supers}$, $\mathtt{overlays}$, $\mathtt{bases}$, $\mathtt{resolve}$, and $\mathtt{this}$, constitute the entire semantics. This is the observational perspective introduced in Section 1: like an F-coalgebra (Rutten, 2000), an overlay tree’s behavior is determined by observation (property presence), not by reduction.
Most of these functions have direct analogues in object-oriented languages: $\mathtt{Path}$ identifies a position in the tree, $\mathtt{init}$ and $\mathtt{last}$ navigate paths, $\mathtt{defines}$ extracts the locally defined members from the AST, $\mathtt{inherits}$ extracts the inheritance declarations from the AST, $\mathtt{properties}$ computes all visible members at a path including both locally defined and inherited ones, $\mathtt{supers}$ computes the transitive inheritance closure, and $\mathtt{bases}$ collects the direct base classes. One function has no object-oriented analogue: $\mathtt{overlays}$ collects the identity-sharing paths that arise when inheritance introduces multiple definitions of the same label at the same scope level. Two functions, the reference resolution function $\mathtt{resolve}$ and the qualified this resolution function $\mathtt{this}$, have object-oriented analogues whose semantics differ significantly; they are discussed in detail below.
All functions are pure and may be cached (memoised) without changing semantics.
Path
A $\mathtt{Path}$ is a sequence of labels $(\ell_{1},\ell_{2},\ldots,\ell_{n})$ that identifies a position in the tree. [^2] Since paths are referentially transparent, an implementation may intern paths so that structural equality reduces to pointer equality. We write $p$ for a path. The root path is the empty sequence $()$. Given a path $p$ and a label $\ell$, $p\mathbin{\triangleright}\ell$ is the sequence $p$ extended with $\ell$. For a non-root path, the parent path $\mathtt{init}(p)$ is $p$ with its last element removed, and the final label $\mathtt{last}(p)$ is the last element of $p$. For convenience, we sometimes write $\ell$ for $\mathtt{last}(p)$ when the path is clear from context. With paths in hand, we can state what the AST provides at each path.
AST
An expression $e$ (Section 2) is parsed into an AST. The AST provides two primitive functions at each path $p$:
  • $\mathtt{defines}(p)$: the set of labels that have subtrees at $p$.
  • $\mathtt{inherits}(p)$: the set of reference pairs $(n,\;\ell_{\mathrm{down},*})$, where $n$ is the de Bruijn index and $\ell_{\mathrm{down},*}$ is the list of downward projections.
During parsing, all three syntactic forms are resolved to de Bruijn index pairs $(n,\;\ell_{\mathrm{down},*})$. The indexed form $\uparrow^{\!n}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$ already carries the de Bruijn index $n$ directly; no resolution is needed. A named reference $\ell_{\mathrm{up}}.\mathbf{this}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$ at path $p$ is resolved by finding the last occurrence of $\ell_{\mathrm{up}}$ among the labels of $p$. Let $p_{\mathrm{target}}$ be the prefix of $p$ up to and including that occurrence. The de Bruijn index is $n=|p|-|p_{\mathrm{target}}|-1$ and the projections are $\ell_{\mathrm{down},*}=(\ell_{\mathrm{down},1},\ldots,\ell_{\mathrm{down},k})$. A lexical reference $\ell_{\mathrm{down},1}.\ell_{\mathrm{down},2}\ldots\ell_{\mathrm{down},k}$ at path $p$ is resolved by finding the nearest prefix $p^{\prime}$ of $p$ such that $\ell_{\mathrm{down},1}\in\mathtt{defines}(p^{\prime})$. The de Bruijn index is $n=|p|-|p^{\prime}|-1$ and the projections are $\ell_{\mathrm{down},*}=(\ell_{\mathrm{down},1},\ldots,\ell_{\mathrm{down},k})$. All three forms produce the same representation; the semantic functions below operate only on de Bruijn index pairs.
Parsing populates $\mathtt{defines}$ and $\mathtt{inherits}$ from each element of a record. A definition $\ell\mapsto e$ contributes $\ell$ to $\mathtt{defines}(p)$. A reference contributes $(n,\;\ell_{\mathrm{down},*})$ to $\mathtt{inherits}(p)$. Since records are sets, the order of elements is irrelevant and duplicate elements have no effect. Both $\mathtt{defines}$ and $\mathtt{inherits}$ are pure data; they are not functions of runtime state. Given these two primitives, we can now answer the question posed at the beginning of this section.
Properties
The properties of a path are its own properties together with those inherited from all supers:
(1) $\mathtt{properties}(p)=\bigl\{\;\ell\;\big|\;(\_,\;p_{\mathrm{overlay}})\in\mathtt{supers}(p),\;\ell\in\mathtt{defines}(p_{\mathrm{overlay}})\;\bigr\}$
The remainder of this section defines $\mathtt{supers}$, the transitive inheritance closure, and its dependencies.
Supers
Intuitively, $\mathtt{supers}(p)$ collects every path that $p$ inherits from: the identity-sharing paths $\mathtt{overlays}$ of $p$ itself, plus the $\mathtt{overlays}$ of each direct base $\mathtt{bases}$ of $p$, and so on transitively. Each result is paired with the inheritance-site context $\mathtt{init}(p_{\mathrm{base}})$ through which it is reached. This provenance is needed by qualified this resolution, i.e. the $\mathtt{this}$ function defined below, to map a definition-site overlay back to the inheritance-site paths that incorporate it:
(2) $\mathtt{supers}(p)=\bigl\{\;(\mathtt{init}(p_{\mathrm{base}}),\;p_{\mathrm{overlay}})\;\big|\;p_{\mathrm{base}}\in\mathtt{bases}^{*}(p),\;p_{\mathrm{overlay}}\in\mathtt{overlays}(p_{\mathrm{base}})\;\bigr\}$
Here $\mathtt{bases}^{*}$ denotes the reflexive-transitive closure of $\mathtt{bases}$. The $\mathtt{supers}$ formula depends on two functions: the identity-sharing paths $\mathtt{overlays}$ and the one-hop reference targets $\mathtt{bases}$. We define $\mathtt{overlays}$ first.
Overlays
Intuitively, inheritance can introduce multiple definitions of the same label at the same scope level; $\mathtt{overlays}(p)$ collects all such paths that share the same identity as $p$, so that their definitions merge rather than shadow each other, enabling deep merge. Concretely, the overlays of $p$ include $p$ itself and $p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)$ for every branch $p_{\mathrm{branch}}$ of $\mathtt{init}(p)$ that also defines $\mathtt{last}(p)$:
(3) $\mathtt{overlays}(p)=\begin{cases}\{p\}&\text{if }p=()\\[6.0pt] \{p\}\;\cup\;\left\{\;p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)\;\left|\;\begin{aligned} &(\_,\;p_{\mathrm{branch}})\in\mathtt{supers}(\mathtt{init}(p)),\\ &\text{s.t.}\;\mathtt{last}(p)\in\mathtt{defines}(p_{\mathrm{branch}})\end{aligned}\right.\right\}&\text{if }p\neq()\end{cases}$
It remains to define $\mathtt{bases}$, the other dependency of $\mathtt{supers}$.
Bases
Intuitively, $\mathtt{bases}(p)$ are the paths that $p$ directly inherits from via references, analogous to the direct base classes in object-oriented languages. Concretely, $\mathtt{bases}$ resolves every reference in $p$’s $\mathtt{overlays}$ one step:
[…]
The $\mathtt{bases}$ formula calls the reference resolution function $\mathtt{resolve}$, which we define next.
Reference resolution
Intuitively, $\mathtt{resolve}$ turns a syntactic reference into the set of paths it points to in the fully inherited tree. It takes an inheritance-site path $p_{\mathrm{site}}$, a definition-site path $p_{\mathrm{def}}$, a de Bruijn index $n$, and downward projections $\ell_{\mathrm{down},*}$. Resolution proceeds in two phases: $\mathtt{this}$ performs $n$ upward steps starting from the enclosing scope $\mathtt{init}(p_{\mathrm{def}})$, mapping the definition-site path to inheritance-site paths; then the downward projections $\ell_{\mathrm{down},*}$ are appended:
(5) $\mathtt{resolve}(p_{\mathrm{site}},\;p_{\mathrm{def}},\;n,\;\ell_{\mathrm{down},*})=\\ \bigl\{\;p_{\mathrm{current}}\mathbin{\triangleright}\ell_{\mathrm{down},1}\mathbin{\triangleright}\cdots\mathbin{\triangleright}\ell_{\mathrm{down},k}\;\big|\;p_{\mathrm{current}}\in\mathtt{this}(\{p_{\mathrm{site}}\},\;\mathtt{init}(p_{\mathrm{def}}),\;n)\;\bigr\}$
When multiple routes exist due to multi-path inheritance, they may yield different target paths; all are collected. We now define $\mathtt{this}$.
Qualified this resolution
$\mathtt{this}$ answers the question: “in the fully inherited tree, where does the definition-site scope $p_{\mathrm{def}}$ actually live?” Each step finds, among the supers of every path in the frontier $S$, those whose overlay component matches $p_{\mathrm{def}}$, and collects the corresponding inheritance-site paths as the new frontier. After $n$ steps the frontier contains the answer:
(6) $\mathtt{this}(S,\;p_{\mathrm{def}},\;n)=\begin{cases}S&\text{if }n=0\\[6.0pt] \mathtt{this}\!\left(\left\{\;p_{\mathrm{site}}\;\left|\;\begin{aligned} &p_{\mathrm{current}}\in S,\\ &(p_{\mathrm{site}},\;p_{\mathrm{overlay}})\in\mathtt{supers}(p_{\mathrm{current}}),\\ &\text{s.t.}\;p_{\mathrm{overlay}}=p_{\mathrm{def}}\end{aligned}\right.\right\},\;\mathtt{init}(p_{\mathrm{def}}),\;n-1\right)&\text{if }n>0\end{cases}$
At each step $\mathtt{init}$ shortens $p_{\mathrm{def}}$ by one label and $n$ decreases by one. Since $n$ is a non-negative integer, the recursion terminates.
This completes the chain of definitions needed to compute $\mathtt{properties}(p)$.
Observational semantics without reduction
The definitions above constitute the complete semantics of overlay-calculus. There are no reduction rules, no evaluation contexts, and no small-step or big-step judgments. In the $\lambda$-calculus, a term such as $(\lambda x.\,x)\;y$ must be reduced to $y$ before its meaning is apparent; the operational semantics specifies how and in what order reductions occur. In overlay-calculus, an inheritance tree such as $\{a\mapsto\{\},\;b\mapsto\{\}\}$ does not reduce to anything: it is the value. The functions $\mathtt{properties}$, $\mathtt{supers}$, etc. determine what is observable at each path, and that is all. Computation arises not from rewriting terms but from querying an inductively deeper path in a lazily constructed tree: the observer drives the computation by choosing which path to inspect, and the mutually recursive definitions unfold on demand. Formally, the mutually recursive equations (1)–(6) define set membership inductively: $\ell\in\mathtt{properties}(p)$ holds if and only if it can be established by a finite chain of applications of these equations. The corresponding operator on the product lattice of path sets is monotone, and its least fixed point exists by the Knaster–Tarski theorem (Tarski, 1955). The detailed well-definedness argument, which identifies the lattice, verifies positivity of each equation, and handles the transitive closure $\mathtt{bases}^{*}$, is given in Appendix C. When no finite derivation exists for a given query, the membership is not in this least fixed point and the semantics is undefined at that point, analogous to divergence in the $\lambda$-calculus, and the expected consequence of Turing completeness. This observation model is inspired by F-coalgebras (Rutten, 2000), where identity is determined by observable behavior rather than by internal structure.
4. Embedding the $\lambda$-Calculus
The following translation maps the $\lambda$-calculus in A-normal form (ANF) (Flanagan et al., 1993) to overlay-calculus. In ANF, every intermediate result is bound to a name and each expression contains at most one application:
[…]
Every $\lambda$-term can be mechanically converted to ANF by naming intermediate results, and every ANF term is a $\lambda$-term, so ANF is neither a restriction nor an extension of the $\lambda$-calculus. The ANF structure matches overlay-calculus naturally: each $\mathbf{let}$-binding becomes a named record property, and each application is wrapped and projected via $\mathrm{result}.$
Let $\mathcal{T}$ denote the translation function. Each $\lambda$-abstraction introduces one scope level in the translated overlay tree.
$x$ ($\lambda$-bound, index $n$)   $\longrightarrow$ $\uparrow^{\!n}.\mathrm{argument}$
$x$ (let-bound)   $\longrightarrow$ $x.\mathrm{result}$
$\lambda x.\,M$   $\longrightarrow$ $\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
$\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M$   $\longrightarrow$ $\{x\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
$V_{1}\;V_{2}$ (tail call)   $\longrightarrow$ $\{\_\mathrm{tail}\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\_\mathrm{tail}.\mathrm{result}\}$
[…]
An abstraction $\lambda x.\,M$ translates to a record with own properties $\{\mathrm{argument},\mathrm{result}\}$; we call this the abstraction shape. In the $\mathbf{let}$-binding rule, the name $x$ binds to the inherited record $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$, and $\mathcal{T}(M)$ may reference $x.\mathrm{result}$ to obtain the value of the application. In the tail-call rule, the fresh label $\mathrm{tail}$ serves the same purpose, and $\mathrm{result}$ projects the answer. In both cases the inheritance that performs the application is encapsulated behind a named property, so its internal structure (the $\mathrm{argument}$ and $\mathrm{result}$ labels of the callee) does not leak to the enclosing scope.
The five rules above are a complete translation from the $\lambda$-calculus (via ANF) to overlay-calculus: every ANF term has an image, and the construction is compositional. Since the $\lambda$-calculus is Turing complete, so is overlay-calculus.
An elementary semantics of the $\lambda$-calculus
Composing the translation $\mathcal{T}$ with the semantic functions of Section 3 induces a semantics on the $\lambda$-calculus itself. Given a $\lambda$-term $M$, one translates it to $\mathcal{T}(M)$ and then determines convergence behavior by querying $\ell\in\mathtt{properties}(p)$ via equations (1)–(6), following the $\mathrm{result}$ chain until the abstraction shape appears or the fixed point yields the empty set.
This induced semantics is more elementary than traditional approaches in the mathematical machinery it requires:
  • No function spaces. Scott’s denotational semantics (Scott, 1976) constructs a reflexive domain $D$ satisfying $D\cong[D\to D]$ via an inverse limit in a category of continuous lattices. Here, the semantic domain is a product of powerset lattices, with no function spaces, no continuity conditions, and no inverse limits. The only fixed-point construction is Knaster–Tarski on powerset lattices (Appendix C), which requires only set-theoretic monotonicity, which is strictly weaker than Scott continuity.
  • No substitution. Operational semantics defines $\beta$-reduction via capture-avoiding substitution, a mechanism whose subtleties motivated de Bruijn indices (de Bruijn, 1972) and explicit substitution calculi. Here, variable references are resolved by the $\mathtt{this}$ function walking the inheritance tree; no substitution occurs at any point.
  • No reduction rules. There are no $\beta$-reduction steps, no evaluation contexts, and no choice between call-by-name and call-by-value (Plotkin, 1975). The semantics is observational: one queries path membership, and the mutually recursive equations unfold on demand.
The key step is that function application is encoded as record inheritance, and the resulting record is observed through path queries rather than applied as a function; no function spaces are required. Section 4.1 shows that the convergence behavior of the translated overlay tree corresponds exactly to the Böhm tree (Barendregt, 1984) of the original term: the translation is adequate, meaning a term converges iff it has a head normal form, and fully abstract, meaning two terms are overlay-contextually equivalent iff they have the same Böhm tree.
Combining these two facts, elementary machinery and full abstraction, yields a noteworthy consequence: the lazy $\lambda$-calculus admits a fully abstract semantics that lives entirely in naive set theory. The only mathematical prerequisites are powerset lattices, set-theoretic monotonicity, and the Knaster–Tarski fixed-point theorem. No continuous lattices, no inverse limits, no Scott topology, no game-theoretic machinery. Scott’s $D_{\infty}$ model (Scott, 1976) uses heavier machinery yet is only adequate, not fully abstract (Plotkin, 1975); Abramsky and McCusker’s game semantics (Abramsky and McCusker, 1995) achieves full abstraction but at the cost of even greater mathematical complexity. The overlay tree semantics achieves full abstraction with strictly simpler foundations than either.
4.1. Böhm Tree Correspondence
The translation $\mathcal{T}$ induces a semantics on the $\lambda$-calculus via the overlay tree equations (1)–(6). We show that the induced semantics corresponds precisely to the Böhm tree (Barendregt, 1984) of the original $\lambda$-term.
An overlay tree $T$ converges, written $T{\Downarrow}$, if following finitely many $\mathrm{result}$ projections from the root reaches a node with both $\mathrm{argument}$ and $\mathrm{result}$ among its properties, that is, the abstraction shape produced by $\mathcal{T}(\lambda x.\,M)$. We write $T{\Uparrow}$ if no such depth exists.
Definition 4.1 (Overlay-convergence).
An overlay tree $T$ converges, written $T{\Downarrow}$, if there exists $n\geq 0$ such that:
$\{\mathrm{argument},\,\mathrm{result}\}\;\subseteq\;\mathtt{properties}\!\bigl(\,\underbrace{\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}\cdots\mathbin{\triangleright}\mathrm{result}}_{n}\,\bigr)$
in the least fixed point of Section 3.
Theorem 4.2 (Adequacy).
A closed ANF $\lambda$-term $M$ has a head normal form if and only if $\mathcal{T}(M){\Downarrow}$.
Theorem 4.3 (Full Abstraction).
For closed $\lambda$-terms $M$ and $N$, define $\mathcal{T}(M)\approx\mathcal{T}(N)$ iff for every closing context $C[\cdot]$, $\mathcal{T}(C[M]){\Downarrow}\Leftrightarrow\mathcal{T}(C[N]){\Downarrow}$. Then:
$\mathcal{T}(M)\approx\mathcal{T}(N)\quad\Longleftrightarrow\quad\mathrm{BT}(M)=\mathrm{BT}(N)$
Together, Adequacy and Full Abstraction establish that the overlay tree semantics is a computation mechanism for Böhm trees (Barendregt, 1984), the canonical fully abstract model of the lazy $\lambda$-calculus, using only powerset lattices and the Knaster–Tarski fixed-point theorem, without the topological infrastructure of domain theory. This is stronger than Scott’s denotational semantics (Scott, 1976; Wadsworth, 1976), which is adequate but not fully abstract, as it identifies terms that observational equivalence distinguishes (Plotkin, 1975). The full proof, including the Single-Path Lemma, Substitution Lemma, and Convergence Preservation, is given in Appendix D.
4.2. Expressive Asymmetry
Two $\lambda$-terms that are Böhm-tree equivalent—and therefore indistinguishable in the $\lambda$-calculus—can be separated in overlay-calculus by overlaying new observable projections onto an existing definition via open extension. By Felleisen’s expressiveness criterion (Felleisen, 1991), overlay-calculus is strictly more expressive than the $\lambda$-calculus: the forward embedding is macro-expressible (Theorem A.2), but the reverse is not (Theorem A.3). The formal definitions, constructions, and proofs are given in Appendix A; the separating construction is itself a miniature instance of the Expression Problem, adding a new observable projection to an existing definition without modifying it.
The underlying mechanism is open extension: overlay definitions compose via open recursion, so existing definitions gain new observable properties without modification. In the $\lambda$-calculus, a function’s result is closed—there is no way to add new observable projections to an already-computed value without modifying the function’s definition.
5. Case Study: Nat Arithmetic
We implemented boolean logic and natural number arithmetic, both unary and binary, as overlay files (Yang, 2025b). This section traces the unary Nat case in detail. The implementation follows a declarative object-oriented style: each concern is one file; UpperCamelCase names are constructors and operations; lowerCamelCase names are fields and parameters; and since there are no methods, all names are nouns or adjectives: $\mathrm{Plus}$ is a structure, not an action. Operations expose their result through a projection field (e.g., $\mathrm{sum}$).
NatData
Following the Church-encoding convention (Yang, 2025b), we define a factory containing the abstract product type and its two constructors. A $\mathrm{Zero}$ value inherits $\mathrm{Product}$; a $\mathrm{Successor}$ value inherits $\mathrm{Product}$ and exposes a $\mathrm{predecessor}$ field. The file name $\mathrm{NatData}$ is itself the top-level scope:
$\mathrm{NatData}\mapsto\{\;$ $\mathrm{NatFactory}\mapsto\{\mathrm{Product}\mapsto\{\},$
$\quad\mathrm{Zero}\mapsto\{\mathrm{Product}\},$
$\quad\mathrm{Successor}\mapsto\{\mathrm{Product},\;\mathrm{predecessor}\mapsto\{\mathrm{Product}\}\}\},$
$\mathrm{Nat}\mapsto\{\mathrm{NatFactory},\;\mathrm{NatFactory}.\mathrm{Product}\}\}$
NatPlus
Addition is a separate overlay file that inherits $\mathrm{NatData}$ and extends each constructor with a $\mathrm{Plus}$ operation exposing a $\mathrm{sum}$ field. The base case is $0+m=m$; the recursive case reduces $\mathrm{S}(n_{0})+m$ to $n_{0}+\mathrm{S}(m)$ by delegating to $n_{0}$’s $\mathrm{Plus}$ with an incremented addend:
$\mathrm{NatPlus}\mapsto\{\;$ $\mathrm{NatData},$
$\mathrm{NatFactory}\mapsto\{$
$\quad\mathrm{Product}\mapsto\{\mathrm{Plus}\mapsto\{\mathrm{sum}\mapsto\{\mathrm{Product}\}\}\},$
$\quad\mathrm{Zero}\mapsto\{\mathrm{Plus}\mapsto\{\mathrm{addend}\mapsto\{\mathrm{Product}\},\;\mathrm{sum}\mapsto\{\mathrm{addend}\}\}\},$
$\quad\mathrm{Successor}\mapsto\{\mathrm{Plus}\mapsto\{\mathrm{addend}\mapsto\{\mathrm{Product}\},$
$\qquad\mathrm{\_increasedAddend}\mapsto\{\mathrm{Successor},\;\mathrm{predecessor}\mapsto\{\mathrm{addend}\}\},$
$\qquad\mathrm{\_recursiveAddition}\mapsto\{\mathrm{Successor}.\mathbf{this}.\mathrm{predecessor}.\mathrm{Plus},$
$\qquad\quad\mathrm{addend}\mapsto\{\mathrm{\_increasedAddend}\}\},$
$\qquad\mathrm{sum}\mapsto\{\mathrm{\_recursiveAddition}.\mathrm{sum}\}\}\}\}\}$
NatEquality
To test that $2+3=5$, we needed equality on Nat values. Equality requires case analysis on the constructor; we use the Visitor pattern. $\mathrm{NatVisitor}$ adds a $\mathrm{Visitor}$ operation to each constructor, with a $\mathrm{Visit}$ field that delegates to the appropriate $\mathrm{VisitZero}$ or $\mathrm{VisitSuccessor}$ branch:
$\mathrm{NatVisitor}\mapsto\{\;$ $\mathrm{NatData},$
$\mathrm{NatFactory}\mapsto\{$
$\quad\mathrm{Product}\mapsto\{\mathrm{Visitor}\mapsto\{\mathrm{Visit}\mapsto\{\}\}\},$
$\quad\mathrm{Zero}\mapsto\{\mathrm{Visitor}\mapsto\{\mathrm{VisitZero}\mapsto\{\},\;\mathrm{Visit}\mapsto\{\mathrm{VisitZero}\}\}\},$
$\quad\mathrm{Successor}\mapsto\{\mathrm{Visitor}\mapsto\{\mathrm{VisitSuccessor}\mapsto\{\},\;\mathrm{Visit}\mapsto\{\mathrm{VisitSuccessor}\}\}\}\}\}$
$\mathrm{BooleanData}$ defines the output sort:
$\mathrm{BooleanData}\mapsto\{\;$ $\mathrm{BooleanFactory}\mapsto\{\mathrm{Product}\mapsto\{\},\;\mathrm{True}\mapsto\{\mathrm{Product}\},\;\mathrm{False}\mapsto\{\mathrm{Product}\}\},$
$\mathrm{Boolean}\mapsto\{\mathrm{BooleanFactory},\;\mathrm{BooleanFactory}.\mathrm{Product}\}\}$
$\mathrm{NatEquality}$ inherits $\mathrm{NatVisitor}$ and $\mathrm{BooleanData}$, and adds an $\mathrm{Equal}$ operation to each constructor. $\mathrm{Zero}$ is equal only to another $\mathrm{Zero}$; $\mathrm{Successor}$ is equal only to another $\mathrm{Successor}$ with an equal predecessor. Case analysis on $\mathrm{other}$ is done via its $\mathrm{Visitor}$:
$\mathrm{NatEquality}\mapsto\{\;$ $\mathrm{NatVisitor},\;\mathrm{BooleanData},$
$\mathrm{NatFactory}\mapsto\{$
$\qquad\mathrm{Product}\mapsto\{\mathrm{Equal}\mapsto\{\mathrm{other}\mapsto\{\mathrm{Product}\},$
$\hskip 18.49988pt\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{Boolean}\}\}\},$
$\qquad\mathrm{Zero}\mapsto\{\mathrm{Equal}\mapsto\{\mathrm{other}\mapsto\{\mathrm{Product}\},$
$\hskip 18.49988pt\mathrm{OtherVisitor}\mapsto\{\mathrm{other}.\mathrm{Visitor},$
$\hskip 18.49988pt\;\mathrm{VisitZero}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{True}\}\},$
$\hskip 18.49988pt\;\mathrm{VisitSuccessor}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{False}\}\},$
$\hskip 18.49988pt\;\mathrm{Visit}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{Boolean}\}\}\},$
$\hskip 18.49988pt\mathrm{equal}\mapsto\{\mathrm{OtherVisitor}.\mathrm{Visit}.\mathrm{equal}\}\}\},$
$\qquad\mathrm{Successor}\mapsto\{\mathrm{Equal}\mapsto\{\mathrm{other}\mapsto\{\mathrm{Product},\;\mathrm{predecessor}\mapsto\{\mathrm{Product}\}\},$
$\hskip 18.49988pt\mathrm{RecursiveEquality}\mapsto\{\mathrm{Successor}.\mathbf{this}.\mathrm{predecessor}.\mathrm{Equal},$
$\hskip 18.49988pt\;\mathrm{other}\mapsto\{\mathrm{Equal}.\mathbf{this}.\mathrm{other}.\mathrm{predecessor}\}\},$
$\hskip 18.49988pt\mathrm{OtherVisitor}\mapsto\{\mathrm{other}.\mathrm{Visitor},$
$\hskip 18.49988pt\;\mathrm{VisitZero}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{False}\}\},$
$\hskip 18.49988pt\;\mathrm{VisitSuccessor}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{RecursiveEquality}.\mathrm{equal}\}\},$
$\hskip 18.49988pt\;\mathrm{Visit}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{Boolean}\}\}\},$
$\hskip 18.49988pt\mathrm{equal}\mapsto\{\mathrm{OtherVisitor}.\mathrm{Visit}.\mathrm{equal}\}\}\}\}\}$
Composing $\mathrm{NatPlus}$ with $\mathrm{NatEquality}$ by inheriting both, with no modifications to either file, gives the resulting Nat values both $\mathrm{Plus}$ and $\mathrm{Equal}$ automatically. The concrete numerals live in a shared file that inherits only $\mathrm{NatData}$:
$\mathrm{NatConstants}\mapsto\{\;$ $\mathrm{NatData},$
$\mathrm{One}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Successor},$
$\quad\quad\mathrm{predecessor}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Zero}\}\},$
$\mathrm{Two}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Successor},$
$\quad\quad\mathrm{predecessor}\mapsto\{\mathrm{One}\}\},$
$\mathrm{Three}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Successor},$
$\quad\quad\mathrm{predecessor}\mapsto\{\mathrm{Two}\}\},$
$\mathrm{Four}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Successor},$
$\quad\quad\mathrm{predecessor}\mapsto\{\mathrm{Three}\}\},$
$\mathrm{Five}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Successor},$
$\quad\quad\mathrm{predecessor}\mapsto\{\mathrm{Four}\}\}\}$
The arithmetic test inherits $\mathrm{NatConstants}$ together with the operations it needs:
$\mathrm{Test}\mapsto\{\;$ $\mathrm{NatConstants},\;\mathrm{NatPlus},\;\mathrm{NatEquality},$
$\mathrm{Addition}\mapsto\{\mathrm{Test}.\mathbf{this}.\mathrm{Two}.\mathrm{Plus},\;\mathrm{addend}\mapsto\{\mathrm{Test}.\mathbf{this}.\mathrm{Three}\}\},$
$\mathrm{Test2plus3}\mapsto\{\mathrm{Test}.\mathbf{this}.\mathrm{Five}.\mathrm{Equal},\;\mathrm{other}\mapsto\{\mathrm{Addition}.\mathrm{sum}\}\}\}$
Querying $\mathrm{Test}.\mathrm{Test2plus3}.\mathrm{equal}$ via the host-language API returns $\mathrm{True}$. No existing file was modified. $\mathrm{NatEquality}$, a new operation, and $\mathrm{Boolean}$, a new data type, each required only a new file: extensibility in both dimensions of the Expression Problem (Wadler, 1998).
Tree-level inheritance and return types
Each operation is an independent overlay that extends a factory’s subtree. Inheritance recursively merges subtrees that share a label, so every entity in the factory acquires all operations from all inherited overlays. Crucially, this applies to return types: $\mathrm{NatPlus}$ returns values from $\mathrm{NatFactory}$, and $\mathrm{NatEquality}$ independently adds $\mathrm{Equal}$ to the same factory. Inheriting both causes the values returned by $\mathrm{Plus}$ to carry $\mathrm{Equal}$ automatically, without modifying either file. In object algebras (d. S. Oliveira and Cook, 2012) and finally tagless interpreters (Carette et al., 2009), independently defined operations do not automatically enrich each other’s return types; additional boilerplate or type-class machinery is required.
Multi-path qualified this (equation 6) is essential. In the test file above, $\mathrm{NatPlus}.\mathrm{Successor}$ and $\mathrm{NatEquality}.\mathrm{Successor}$ both inherit $\mathrm{NatData}.\mathrm{Successor}$. When the two operations are composed, the resulting $\mathrm{Successor}$ inherits the $\mathrm{NatData}$ schema through two independent routes. The reference $\mathrm{Successor}.\mathbf{this}.\mathrm{predecessor}$ inside each operation resolves through both routes, and since inheritance is idempotent, all routes contribute the same properties. In Scala, the same pattern, where two independently defined inner classes extend the same outer class’s trait, triggers a “conflicting base types” rejection (Appendix B), because DOT’s self variable is single-valued and cannot resolve through multiple inheritance routes simultaneously. The Expression Problem can still be solved in Scala using object algebras or similar patterns, but composing independently defined operations inherently produces multi-path inheritance to shared schemas, which DOT’s single-valued self variable cannot accommodate.
Church-encoded Nats are tries
A Church-encoded Nat is a record whose structure mirrors the constructor path used to build it: $\mathrm{Zero}$ is a flat record; $\mathrm{S}(\mathrm{Zero})$ is a record with a $\mathrm{predecessor}$ field pointing to a $\mathrm{Zero}$ record; and so on. This is exactly the shape of a trie. A scope that simultaneously inherits two Nat values is a trie union. The following test applies $\mathrm{Plus}$ to a trie union of $\{1,2\}$ with addend a trie union of $\{3,4\}$, then compares the result to itself with $\mathrm{Equal}$:
$\mathrm{CartesianTest}\mapsto\{\;$ $\mathrm{NatConstants},\;\mathrm{NatPlus},\;\mathrm{NatEquality},$
$\mathrm{OneOrTwo}\mapsto\{\mathrm{CartesianTest}.\mathbf{this}.\mathrm{One},\;\mathrm{CartesianTest}.\mathbf{this}.\mathrm{Two}\},$
$\mathrm{ThreeOrFour}\mapsto\{\mathrm{CartesianTest}.\mathbf{this}.\mathrm{Three},\;\mathrm{CartesianTest}.\mathbf{this}.\mathrm{Four}\},$
$\mathrm{Result}\mapsto\{\mathrm{OneOrTwo}.\mathrm{Plus},\;\mathrm{addend}\mapsto\{\mathrm{ThreeOrFour}\}\},$
$\mathrm{Check}\mapsto\{\mathrm{Result}.\mathrm{sum}.\mathrm{Equal},\;\mathrm{other}\mapsto\{\mathrm{Result}.\mathrm{sum}\}\}\}$
$\mathrm{Result}.\mathrm{sum}$ evaluates to the trie $\{4,5,6\}$, the set of all pairwise sums. $\mathrm{Equal}$ then applies to every pair drawn from $\{4,5,6\}\times\{4,5,6\}$; querying $\mathrm{CartesianTest}.\mathrm{Check}.\mathrm{equal}$ via the host-language API returns $\{\mathrm{True},\mathrm{False}\}$. The Cartesian product arises directly from the semantic equations. $\mathrm{OneOrTwo}$ inherits both $\mathrm{One}$ and $\mathrm{Two}$, so $\mathtt{bases}(\mathrm{OneOrTwo})$ contains two paths. When $\mathrm{Result}$ inherits $\mathrm{OneOrTwo}.\mathrm{Plus}$, the $\mathtt{bases}$ of $\mathrm{Result}$ include the $\mathrm{Plus}$ of both $\mathrm{One}$ and $\mathrm{Two}$. Each $\mathrm{Plus}$ resolves its $\mathrm{addend}$ via $\mathtt{this}$, reaching $\mathrm{ThreeOrFour}$, which is itself a trie union of $\mathrm{Three}$ and $\mathrm{Four}$: the $\mathtt{supers}$ of the $\mathrm{addend}$ path contain both values. Since $\mathrm{Successor}.\mathrm{Plus}$ recursively delegates to $\mathrm{predecessor}.\mathrm{Plus}$ with an incremented addend, and $\mathrm{predecessor}$ also resolves through the trie union, the recursion branches at every trie node. The result $\mathrm{sum}$ therefore collects all paths reachable through any combination of the two operands, which is the Cartesian product.
The same $\mathrm{Plus}$ and $\mathrm{Equal}$ files that operate on single values thus produce, without modification, the relational semantics of logic programming: $\mathtt{overlays}$ and $\mathtt{supers}$ collect all inheritance paths, and every operation distributes over them.
6. Discussion
6.1. Emergent Phenomena
Several features of overlay-calculus are primitive yet have no direct counterpart in the $\lambda$-calculus. Encoding any one of them requires pervasive, non-local changes to program structure.
  1. (1)
    Self-referential records with lazy evaluation. An overlay’s properties can reference other properties and enclosing scopes through qualified this without explicit binding. In the $\lambda$-calculus, encoding such self-referential structures requires explicit allocation of mutable or lazy references (Boudol, 2004) (e.g., Haskell-style thunks or ML-style ref cells), obscuring the declarative intent.
  2. (2)
    Open recursion with symmetric inheritance. Inheritance merges definitions from independent sources, all sharing the same inherited overlay tree. Encoding this in the $\lambda$-calculus requires maintaining an extensible dictionary of methods that can be merged from multiple directions. Cook’s denotational semantics of inheritance (Cook, 1989) devoted an entire dissertation to formalizing this pattern in a functional setting.
  3. (3)
    Commutative, idempotent merge. Inheritance in overlay-calculus is symmetric: a record’s elements form a set, so reordering them produces the same observable result. In the $\lambda$-calculus, function composition $(f\circ g)$ is neither commutative nor idempotent, and simulating these properties requires additional machinery (Harper and Pierce, 1991; Cardelli, 1994).
  4. (4)
    Relational semantics. Every overlay-calculus operation is defined by recursion on overlay trees, not on individual values. When an overlay tree is viewed as a trie, the correspondence with relational algebra becomes precise: a path from the root to a leaf encodes a single value (e.g., $\mathrm{Successor}\to\mathrm{predecessor}\to\mathrm{Zero}$ encodes the natural number $1$); a trie union encodes a set of such values, i.e., a unary relation. An operation that takes two trie-union operands traverses every combination of their paths through $\mathtt{bases}$ (equation 4), which is a cross-join; the transitive closure $\mathtt{bases}^{*}$ in $\mathtt{supers}$ (equation 2) computes the union of all reachable paths; and $\mathtt{properties}$ (equation 1) projects the observable labels. The carrier is a trie rather than a flat relation, and observation is coinductive—the observer drives computation by querying progressively deeper paths—rather than Datalog’s bottom-up materialization, but the relational structure of the computation is the same. In the Case Study (Section 5), the $\mathrm{Plus}$ and $\mathrm{Equal}$ operations are defined on single Nat values. When applied to set-valued overlays $\{1,2\}$ and $\{3,4\}$, the same definitions, without any modification, produce the set of all pairwise sums $\{4,5,6\}$. This is relational semantics in action: each operation distributes over every combination of its inputs. The behavior is native: the semantic equations of Section 3 resolve every property to a set of paths, so single-valued and set-valued computation are the same mechanism. No code change is needed to move from one to the other.
  5. (5)
    Function color blindness. We implemented a web application as overlay files (Yang, 2025b), initially targeting Python’s synchronous sqlite3 and http.server modules. The business logic overlay file contained no async-specific code. Later, we replaced the FFI layer with one built on aiosqlite and starlette; the business logic overlay file did not change. This is an instance of function color blindness (Nystrom, 2015): async and sync functions are distinct types in the host language, yet the overlay file is agnostic to the distinction because its empty property slots are filled at assembly time, with coroutine handlers under asyncio, or with direct implementations under synchronous execution. More generally, empty slots can be filled with effect handlers (Plotkin and Power, 2003; Pretnar, 2015) under a CPS interpretation, or with direct implementations under a direct-style interpretation, without explicit monad machinery.
Each of these features is primitive in overlay-calculus. In the $\lambda$-calculus, encoding them requires pervasive changes to program structure: every scope must thread a mutable store for lazy allocation, every method invocation must consult a dynamically constructed dictionary for open recursion, and every composition must be wrapped in order-insensitive infrastructure for commutativity. These are global, cascading transformations, not local rewrites of individual terms.
6.2. Classification of Semantic Variants
The goals of declarative programming rule out several properties of the model: separate primitives for methods, classes, types, and objects; mutable reference cells; domain-theoretic approximation; and asymmetric override. Ruling these out leaves a minimal syntax of three constructs: records, definitions, and inheritance.
Two degrees of freedom remain. These dimensions are independent, and each combination yields a coherent semantics. They arise in any system whose definitions form a tree of named scopes, where each node is a path and references are path-dependent, even in systems ruled out above, such as Scala, which has separate primitives for methods, classes, objects, and types.
Existing designs occupy several of the resulting cells:
No deep merge Deep merge
Early binding CUE (van Lohuizen, 2019)
Dynamic scope union file systems, Jsonnet (Cunningham, 2014)
Lexical scope Scala (Odersky et al., 2004) NixOS modules (NixOS Contributors, 2024)
Multi-target $\mathtt{this}$ overlay-calculus
Early binding
With early binding, references resolve at definition time to a fixed path, independently of how the record is later inherited. CUE (van Lohuizen, 2019) exemplifies this: field references are statically resolved, and a field in a nested record cannot refer back to properties of its enclosing context contributed by later composition. Early binding combined with deep merge is coherent and practically useful, but cross-overlay cooperation requires explicit wiring rather than inheritance: a definition cannot observe properties contributed by overlays composed after it.
Dynamic scope
With dynamic scope, references resolve against the context at the point of use rather than the definition site. Early Lisp implementations used dynamic scope (McCarthy, 1978), where variable references are resolved against the run-time call stack. Union file systems exhibit the same property at the file-system level: a symbolic link’s relative path resolves against the mount context at dereference time, not against the layer that defined the link. Jsonnet (Cunningham, 2014) follows this approach: self always refers to the final merged object, not the object at the definition site. When the Scala multi-path example (Appendix B) is translated to Jsonnet, it does not error, but outer resolves to the merged MyObjectA rather than to Object1 or Object2 separately, silently discarding the distinction between the two inheritance routes. Dynamic scope is flexible but loses the ability to track which inheritance route a reference came from.
Deep merge
Without the second case of equation (3), $\mathtt{overlays}(p)$ would always return $\{p\}$, a single element. When two independently authored overlays both define a label with the same name at the same scope level, only the definition at the literal path $p$ would contribute to $\mathtt{supers}$ and hence to $\mathtt{properties}$; the other definitions would be silently invisible. This is the situation in most composition systems: pytest permits at most one fixture per name per scope; trait systems (Schärli et al., 2003) flag same-name methods as conflicts. Under any of these rules, two independently authored overlays that both extend an entity with the same name, which is the situation that arises naturally when solving the Expression Problem (Wadler, 1998), cannot be composed. The non-trivial case of equation (3) removes this limitation: it collects all paths that share the same label and merges their definitions rather than shadowing each other. This is the record-level counterpart of method override in object-oriented languages, and the mechanism by which the calculus admits extensibility in both dimensions of the Expression Problem (Wadler, 1998) (Section 5).
Multi-target qualified this
The qualified this reference $\mathtt{this}$ answers the question: “in the fully inherited tree, where does the definition-site scope $p_{\mathrm{def}}$ actually live?” Scheme (Sussman and Steele Jr., 1975) introduced lexical scope, where a de Bruijn index resolves in a single step by indexing into the statically determined environment; this is predictable but rigid. Scala (Odersky et al., 2004) and the NixOS module system (NixOS Contributors, 2024) follow the lexical approach: each this reference is bound to a single statically known path. When inheritance introduces multiple inheritance routes to the same scope, Scala rejects the ambiguity at compile time (Appendix B) and the NixOS module system rejects the duplicate option declaration (Appendix B); neither can express that all routes are equally valid.
Overlay-calculus combines the predictability of lexical scope with the flexibility of dynamic scope. The reference itself is lexically determined at the definition site as a de Bruijn index, so what $\mathtt{this}$ looks up is predictable. But because inheritance interleaves scopes from different overlays, the result depends on the inherited tree structure: $\mathtt{this}$ walks upward one scope level at a time, consulting $\mathtt{supers}$ at each level, and naturally tracks a set of inheritance-site paths rather than a single one. All routes contribute equally to the inherited result, and since inheritance is idempotent, duplicated contributions are harmless. The only observable is property presence, making multi-path resolution well-defined without disambiguation.
Our first three implementations of the Overlay language used standard techniques for self-reference resolution: closure-based fixed points following Cook (Cook, 1989), and stack-based environment lookup using de Bruijn indices (de Bruijn, 1972). All three produced subtle bugs when inheritance introduced multiple routes to the same enclosing scope. The difficulty was structural, not incidental: a closure captures a single environment, so a fixed-point combinator solves for a single self-reference; a stack-based environment is a linear chain where each scope level has one parent. Both data structures embed a single-valued assumption incompatible with the multi-path situation that multiple inheritance naturally produces (Appendix B). Equation (6) emerged from abandoning this assumption: it tracks a frontier set $S$ rather than a single current scope. If the equations were restricted to single-target resolution, as in DOT’s single-valued self variable (Amin et al., 2016), they would need to reject structurally legitimate cases, adding complexity rather than removing it.
Consequences of the alternative choices
Choosing differently in any dimension requires introducing machinery that is absent from the equations of Section 3. With early binding, the equations must freeze reference targets at definition time, so a definition cannot observe properties contributed by later-composed overlays. With dynamic scope, the equations must resolve references against a runtime context rather than the definition-site structure, losing the ability to distinguish inheritance routes. Restricting $\mathtt{overlays}(p)$ to $\{p\}$ means same-label definitions from different inheritance sources no longer share identity, requiring either a method resolution order (MRO) (Barrett et al., 1996) to linearize them or a shadowing rule to select one and discard the rest, which is why systems without deep merge rarely support path-dependent multiple inheritance in practice. Restricting $\mathtt{this}$ to a single path requires a mechanism to select among inheritance routes (Scala rejects the ambiguity at compile time; the NixOS module system raises a duplicate-declaration error), and loses contributions from all non-selected routes. These mechanisms, including early-binding freezing, runtime context threading, MRO, shadowing, route selection, and error reporting, are more artificial than their absence. The equations of Section 3 contain none of them.
7. Related Work
Minimal Turing-complete models
Schönfinkel (Schönfinkel, 1924) and Curry (Curry and Feys, 1958) showed that the $\lambda$-calculus reduces to three combinators $S$, $K$, II (II is redundant: $I=SKK$). The bracket abstraction that compiles $\lambda$-terms to SKI is mechanical but incurs exponential blow-up in term size; no new computational patterns emerge beyond those already present in the $\lambda$-calculus. Dolan (Dolan, 2013) showed that the x86 mov instruction alone is Turing complete by exploiting addressing modes as implicit arithmetic on mutable memory; the resulting programs are correct but provide no abstraction mechanisms beyond those of the RAM Machine. The Turing Machine itself, while universal, lacks random access: reading a cell at distance $d$ requires $d$ sequential tape movements. Each of these models achieves Turing completeness by removing structure from an existing model: fewer abstractions, worse ergonomics, no new expressiveness. None of these reductions exhibits emergent phenomena absent from the model it reduces.
Denotational semantics of inheritance
Cook (Cook, 1989) gave the first denotational semantics of inheritance, modeling objects as recursive records and inheritance as composition of generators, which are functions from self to complete object, and wrappers, which are functions that modify generators, solved by a fixed-point construction. Cook’s key insight, that inheritance is a general mechanism applicable to any form of recursive definition and not only object-oriented methods, is one of the starting points of the present work. The model has four primitives—records, functions, generators, and wrappers—and requires a fixed-point combinator; composition is asymmetric, as a wrapper modifies a generator but not vice versa.
Mixin-based inheritance
Bracha and Cook (Bracha and Cook, 1990) formalized mixins as abstract subclasses, that is, functions from a superclass parameter to a subclass, unifying the inheritance mechanisms of Smalltalk, Beta, and CLOS under a single model. Because mixin application is function composition, it is neither commutative nor idempotent: applying the same mixin twice may produce different results, and applying two mixins in different orders may yield different linearizations. The C3 linearization algorithm (Barrett et al., 1996) was later developed to impose a deterministic order; whether linearization is inherent to inheritance or an artifact of non-commutativity had not been established. Mixin composition operates at the method level: applying a mixin replaces or extends individual methods in a flat namespace. There is no recursive merging of nested structure: composing two mixins that define the same method name results in an override, not deep merge. Finally, mixins presuppose the $\lambda$-calculus as their computational foundation: method bodies are functions, and mixin application is function composition. Mixins organize and compose functions but cannot replace them; without the underlying $\lambda$-calculus, there is nothing to compose. The mixin mechanism itself is not Turing complete.
Traits
Schärli et al. (Schärli et al., 2003) introduced traits as composable units of behavior. Trait composition is symmetric and commutative: the sum of two traits is order-independent, which is a departure from mixin ordering. However, when two traits provide methods with the same name, a conflict arises and must be explicitly resolved by the programmer through exclusion or aliasing. Ducasse et al. (Ducasse et al., 2006) extended the model with a flattening property: the semantics of a class is independent of whether its methods originate from traits or are defined directly. Like mixins, trait composition is flat: it operates on individual method names, not on nested structures, so same-name definitions are conflicts rather than recursive merges. And like mixins, traits presuppose the $\lambda$-calculus: trait methods are functions, and the computational power resides in those functions, not in the trait mechanism itself.
Family polymorphism and virtual classes
Ernst (Ernst, 2001) introduced family polymorphism, where families of mutually dependent classes can be refined together in subclasses. Ernst, Ostermann, and Cook (Ernst et al., 2006) formalized this in the virtual class calculus, where classes are class-valued attributes of objects, accessed via path expressions of the form this.out.C. The path this.out navigates to an enclosing scope, analogous to qualified this ($\ell_{\mathrm{up}}.\mathbf{this}$) in this paper. A notable restriction is that each such path resolves to a single class in a single enclosing object. When multiple inheritance routes lead to the same scope, a common situation under inheritance, single-valued resolution must choose one route and discard the others, losing information. The framework provides no mechanism for set-valued resolution.
Object calculi
Abadi and Cardelli (Abadi and Cardelli, 1996) developed a comprehensive theory of objects, treating objects rather than classes as the primitive notion. Their calculi include method update and self-referential objects, but methods are fundamentally functions that take self as a parameter, and object extension is asymmetric since method override replaces the previous definition. Boudol (Boudol, 2004) studied the recursive record semantics of objects in a call-by-value setting, showing that self-referential records require an unsafe fixed-point operator. Whether self-reference can be handled without a fixed-point operator in the object language, for instance through coalgebraic observation rather than reduction, was not explored in these calculi.
Record calculi
Harper and Pierce (Harper and Pierce, 1991) presented a record calculus based on symmetric concatenation with a type system that tracks the absence of fields. When two records define the same label, concatenation is undefined and raises a type error; the calculus leaves same-label composition undefined. Cardelli (Cardelli, 1994) studied extensible records with subtyping; Rémy (Rémy, 1989) gave a type system for records and variants in ML. These calculi operate on flat records with no nesting and treat records as data, that is, initial algebras: a record is a finite map from labels to values, fully determined at construction time. Nested records, self-reference, and lazy observation, which are the properties needed for records to serve as a computational substrate, are outside their scope.
DOT calculus
Amin et al. (Amin et al., 2016) formalized path-dependent types in the DOT calculus, the theoretical foundation of Scala’s type system. DOT paths navigate through objects to reach nested type members, providing a rich dependent type structure. DOT’s self variable in the object constructor $\nu(x:T)\,d$ binds $x$ to a single object; when multiple inheritance routes lead to the same enclosing scope, the resulting intersection type is unrealizable and Scala rejects it at compile time (Appendix B). DOT stratifies its syntax into terms and types, with types further divided into abstract type members bounded by $L..U$ and concrete definitions; each stratum requires its own subtyping rules. Whether these three roles could be unified under a single construct remains an open question in the DOT line of work. DOT includes the $\lambda$-calculus as a sub-language; removing it while retaining path-dependent types would yield a calculus closer to pure overlay inheritance (see Section 8).
Configuration languages
CUE (van Lohuizen, 2019) unifies types and values in a single lattice where unification (&) is commutative, associative, and idempotent. CUE’s design was motivated by Google’s Borg Configuration Language (GCL), which used graph unification. CUE’s lattice includes scalar types with a conflict semantics where unifying incompatible scalars yields $\bot$; whether scalars are necessary for a configuration language, or whether tree structure alone suffices, is an interesting design question. Jsonnet (Cunningham, 2014) provides object inheritance via the + operator with mixin semantics: composition is not commutative since the right-hand side wins on scalar conflicts, and deep merging requires explicit opt-in via the +: syntax) on a per-field basis, so deep merge is available but not the default. Dhall (Gonzalez, 2017) takes a functional approach: it is a typed $\lambda$-calculus with records, where composition is a record merge operator, not an inheritance mechanism.
Module systems and deep merge
The NixOS module system (Dolstra et al., 2010; NixOS Contributors, 2024) composes modules by recursively merging nested attribute sets, not by linearization or method-level override, which is the same overlay semantics that distinguishes tree-level inheritance from mixin and trait composition. When two modules define the same nested path, their subtrees are merged rather than overridden. The NixOS module system incorporates deep merge, recursive self-reference, deferred evaluation of modules, and a rich type-checking layer on top. The same mechanism powers NixOS system configuration, Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), flake-parts (Hensing, 2022), and dozens of other ecosystems, collectively managing configurations of arbitrary complexity. The one aspect that appears not to have been fully addressed is the semantics of self-reference under multiple inheritance. When two independent modules that inherit from the same deferred module are composed, each import introduces its own declaration of the shared options; the module system’s mergeOptionDecls rejects duplicate declarations with a static error (Appendix B). The module system cannot express that both routes contribute equally to the same option, that is, the multi-path situation that recursive merging naturally gives rise to, which motivated the coalgebraic semantics of Section 3.
Modular software architectures
Plugin systems, component frameworks, and dependency injection frameworks enable software extensibility through declarative inheritance. The symmetric, associative nature of inheritance (Section 3) explains why such systems work: components can be inherited in any order without changing behavior. These systems implement overlay-calculus patterns ad hoc.
Union file systems
Union file systems (UnionFS, OverlayFS, AUFS) layer multiple directory hierarchies to present a unified view. Their semantics resembles overlay-calculus in that later layers override earlier layers and files from all layers remain accessible, but without a formal foundation these systems resort to ad-hoc choices: asymmetric inheritance, non-commutative conflict resolution for scalar files, and no late-binding or dynamic dispatch across layers. We implemented overlay-calculus directly as a union file system (Yang, 2025a), obtaining late-binding semantics and dynamic dispatch for file lookups with no additional mechanism beyond the three primitives. In package managers, build systems, and OS distributions, external toolchains transform configuration into file trees; here, the file system serves as configuration to itself, and the transformation is inheritance: a file-system-as-compiler in which the directory tree is both the source program and the compilation target.
8. Future Work
The Overlay language (Yang, 2025b) is an executable implementation of overlay-calculus that already goes beyond the untyped calculus presented here: it includes compile-time checking that all references resolve to valid paths in the inheritance, and a foreign-function interface (FFI) that introduces scalar values from the host language. The future work described below concerns the distance between this implementation and a fully practical language, spanning both theoretical foundations and library-level encodings.
Type system
The Overlay language’s existing compile-time checks verify that every reference path resolves to a property that exists in the inherited result. Formalizing the soundness of these checks, that is, proving that well-typed programs do not produce dangling references at runtime, is one direction of future work. Such a formalization would relate to the DOT calculus (Amin et al., 2016) in a manner analogous to how overlay-calculus relates to the $\lambda$-calculus: a Typed overlay-calculus can be viewed as DOT without $\lambda$, retaining path-dependent types while replacing functions with overlay inheritance. A formal type system would be amenable to mechanization in proof assistants (Coq, Agda) for verifying type safety, progress, and preservation properties.
Beyond soundness of the existing checks, additional type system features would strengthen the language. A key example is totality checking: verifying that a inheritance provides implementations for all required slots, not merely that references resolve. In untyped overlay-calculus, writing $\mathrm{magnitude}\mapsto\{\}$ is purely documentary: it defines a structural slot but imposes no constraint on what must fill it. Totality checking would enforce such declarations statically, rejecting inheritances that leave required slots unfilled.
Standard library
The remaining items concern the standard library: they require no extensions to the calculus or type system, but involve nontrivial encodings within the existing framework.
Overlay-calculus is open by default: inheritance can freely merge any two overlays, and a value may simultaneously inhabit multiple constructors (e.g., $\{\mathrm{Zero},\;\mathrm{Odd},\;\mathrm{half}\mapsto\mathrm{Zero}\}$). This is not a defect; it is the natural trie semantics of the calculus, and the basis for solving the expression problem (Section 5). However, many operations (e.g., equality testing) assume that values are linear, inhabiting exactly one constructor. The observer pattern (Section 5) does not enforce this: applied to a multi-constructor value, all callbacks fire and their results are inherited.
Schema validation can be implemented at the library level using existing primitives. An observer can test whether a particular constructor is present, returning a Boolean; a second Boolean dispatch then branches on the result. By chaining such tests, a factory can validate that a value matches exactly one constructor and carries the required fields. This does not restrict inheritance itself, which remains unconstrained, but provides a way to detect invariant violations before they propagate.
The same technique, an observer returning Boolean followed by Boolean dispatch, yields closed pattern matching. The observer pattern of Section 5 is inherently open: new constructor cases can be added through inheritance. Closed matching, where exactly one branch is taken, can be encoded as a visitor chain: a linked list of if-then-else nodes, each testing one constructor and falling through on mismatch. This is analogous to GHC.Generics’ $\mathrm{(:+:)}$ sum representation (Magalhães et al., 2010), where each link corresponds to one summand. An important consequence is that such chains do not commute, as they have a fixed order, and therefore cannot be extended through open inheritance, which is the correct semantics for closed dispatch.
9. Conclusion
The goals of declarative programming forced the semantics of overlay-calculus upon us. The emergent phenomena exhibit an expressive asymmetry between overlay-calculus and the $\lambda$-calculus. Inheritance is not a mechanism for organizing functions; functions are syntactic sugar for inheritance.
Appendix A Expressive Asymmetry: Proofs
Definition A.1 (Macro-expressibility, after Felleisen (Felleisen, 1991)).
A construct $C$ of language $\mathscr{L}_{1}$ is macro-expressible in language $\mathscr{L}_{0}$ if there exists a translation $\mathcal{E}$ such that for every program $P$ containing occurrences of $C$, the translation $\mathcal{E}(P)$ is obtained by replacing each occurrence of $C$ with an $\mathscr{L}_{0}$ expression that depends only on $C$’s subexpressions, leaving all other constructs of $P$ unchanged.
Theorem A.2 (Forward macro-expressibility).
The translation $\mathcal{T}$ (Section 4) is a macro-expressible embedding of the $\lambda$-calculus into overlay-calculus.
Proof.
Each of the five rules of $\mathcal{T}$ replaces one $\lambda$-calculus construct with an overlay-calculus expression that depends only on the construct’s immediate subexpressions. No rule inspects the surrounding context. ∎
Theorem A.3 (Non-expressibility of overlay).
There is no macro-expressible translation from overlay-calculus to the $\lambda$-calculus.
Proof.
Let $\mathscr{L}_{1}$ be the overlay-calculus and let $\mathscr{L}_{0}$ be the sublanguage consisting of overlay trees in the image of the translation $\mathcal{T}$ (Section 4), so that $\mathscr{L}_{0}\subset\mathscr{L}_{1}$. Define the operational equivalence $\cong_{0}$ on $\mathscr{L}_{0}$ via Böhm trees: $\mathcal{T}(M)\cong_{0}\mathcal{T}(N)$ iff $\mathrm{BT}(M)=\mathrm{BT}(N)$. Define $\cong_{1}$ on $\mathscr{L}_{1}$ as overlay-contextual equivalence: $S\cong_{1}S^{\prime}$ iff for every overlay context $C[\cdot]$, $C[S]{\Downarrow}\Leftrightarrow C[S^{\prime}]{\Downarrow}$.
By the Böhm Tree Correspondence (Theorems 4.2 and 4.3), $\mathscr{L}_{1}$ conservatively extends $\mathscr{L}_{0}$: for closed $\lambda$-terms $M$, $\mathcal{T}(M){\Downarrow}$ if and only if $M$ has a head normal form.
Define Church booleans via the translation $\mathcal{T}$ (Section 4): $\mathrm{true}=\lambda t.\,\lambda f.\,t$ and $\mathrm{false}=\lambda t.\,\lambda f.\,f$. In de Bruijn notation:
$\mathrm{true}$ $\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\uparrow^{\!1}.\mathrm{argument}\}\}$
$\mathrm{false}$ $\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\uparrow^{\!0}.\mathrm{argument}\}\}$
Define Church boolean equality $\mathrm{eq}=\lambda a.\,\lambda b.\,(a\;b)\;(b\;\mathrm{false}\;\mathrm{true})$. In ANF:
$\mathrm{eq}=\lambda a.\,\lambda b.\,\mathbf{let}\;x_{1}=b\;\mathrm{false}\;\mathbf{in}\;\mathbf{let}\;x_{2}=x_{1}\;\mathrm{true}\;\mathbf{in}\;\mathbf{let}\;x_{3}=a\;b\;\mathbf{in}\;x_{3}\;x_{2}$
Applying $\mathcal{T}$ (Section 4):
$\mathrm{eq}$ $\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\{\mathrm{argument}\mapsto\{\},$
$\quad x_{1}\mapsto\{\uparrow^{\!0}.\mathrm{argument},\;\mathrm{argument}\mapsto\mathrm{false}\},$
$\quad x_{2}\mapsto\{x_{1}.\mathrm{result},\;\mathrm{argument}\mapsto\mathrm{true}\},$
$\quad x_{3}\mapsto\{\uparrow^{\!1}.\mathrm{argument},\;\mathrm{argument}\mapsto\uparrow^{\!0}.\mathrm{argument}\},$
$\quad\_\mathrm{tail}\mapsto\{x_{3}.\mathrm{result},\;\mathrm{argument}\mapsto x_{2}.\mathrm{result}\},$
$\quad\mathrm{result}\mapsto\_\mathrm{tail}.\mathrm{result}\}\}$
Here $\uparrow^{\!1}.\mathrm{argument}$ in $x_{3}$ reaches the outer $\lambda a$ scope (one $\lambda$ up from $\lambda b$), and $\uparrow^{\!0}.\mathrm{argument}$ reaches $\lambda b$ itself.
In overlay-calculus, applying a function record to an argument is: inherit the function, override its $\mathrm{argument}$ slot, and project $\mathrm{result}$. Define overlay trees $e_{1}$ and $e_{2}$ that apply $\mathrm{eq}$ to its two arguments via this mechanism:
$e_{1}\;=\;\{$ $\mathrm{eqFalse}\mapsto\{\mathrm{eq},\;\mathrm{argument}\mapsto\mathrm{false}\},$
$\mathrm{applied}\mapsto\{\mathrm{eqFalse}.\mathrm{result},\;\mathrm{argument}\mapsto\mathrm{false}\},$
$\mathrm{result}\mapsto\mathrm{applied}.\mathrm{result}\;\}$
$e_{2}\;=\;\{$ $\mathrm{eqTrue}\mapsto\{\mathrm{eq},\;\mathrm{argument}\mapsto\mathrm{true}\},$
$\mathrm{applied}\mapsto\{\mathrm{eqTrue}.\mathrm{result},\;\mathrm{argument}\mapsto\mathrm{true}\},$
$\mathrm{result}\mapsto\mathrm{applied}.\mathrm{result}\;\}$
That is, $e_{1}=\mathcal{T}(\mathrm{ANF}(\mathrm{eq}\;\mathrm{false}\;\mathrm{false}))$ and $e_{2}=\mathcal{T}(\mathrm{ANF}(\mathrm{eq}\;\mathrm{true}\;\mathrm{true}))$, where $\mathrm{ANF}$ converts a $\lambda$-term to A-normal form by naming every intermediate application (Section 4). Since both $\mathrm{ANF}$ and $\mathcal{T}$ are macro-expressible, their composition is too, so $e_{1},e_{2}\in\mathscr{L}_{0}$. Both $e_{1}.\mathrm{result}$ and $e_{2}.\mathrm{result}$ evaluate to Church $\mathrm{true}$. By Full Abstraction (Theorem 4.3), $\mathrm{BT}(e_{1})=\mathrm{BT}(e_{2})$, so $e_{1}\cong_{0}e_{2}$.
[…]
$C={}$ $\{\mathrm{eq}\mapsto\{\mathrm{result}\mapsto\{\mathrm{repr}\mapsto\{\mathrm{firstOperand}\mapsto\uparrow^{\!2}.\mathrm{argument},\;\mathrm{secondOperand}\mapsto\uparrow^{\!1}.\mathrm{argument}\}\}\}\}$
The de Bruijn indices count scope levels from $\mathrm{repr}$: $\uparrow^{\!0}$ is $\mathrm{repr}$ itself, $\uparrow^{\!1}$ is the $\lambda b$ scope, and $\uparrow^{\!2}$ is the $\lambda a$ scope. Thus $\mathrm{firstOperand}$ projects $\lambda a$’s argument and $\mathrm{secondOperand}$ projects $\lambda b$’s argument.
In the composite scope $\{S,\,C\}$, the definition of $\mathrm{eq}$ gains $\mathrm{repr}$ as a new sibling of $x_{1},x_{2},x_{3}$ inside the $\lambda b$ scope. By open recursion, the de Bruijn references $\uparrow^{\!2}.\mathrm{argument}$ and $\uparrow^{\!1}.\mathrm{argument}$ resolve through the same scope chain as the existing let-bindings. Since $e_{1}$ and $e_{2}$ inherit $\mathrm{eq}$, they also inherit the newly added $\mathrm{repr}$: $e_{1}.\mathrm{repr}.\mathrm{firstOperand}=\mathrm{false}$ and $e_{2}.\mathrm{repr}.\mathrm{firstOperand}=\mathrm{true}$. These produce structurally distinct overlay trees (one converges to an abstraction selecting its first argument, the other to an abstraction selecting its second), so $e_{1}\not\cong_{1}e_{2}$.
Since $e_{1}\cong_{0}e_{2}$ but $e_{1}\not\cong_{1}e_{2}$, we have ${\cong_{0}}\neq{(\cong_{1}|_{\mathscr{L}_{0}})}$. By Theorem 3.14(i) of Felleisen (Felleisen, 1991), $\mathscr{L}_{0}$ cannot macro-express the overlay facilities of $\mathscr{L}_{1}$. ∎
Corollary A.4.
Overlay-calculus is strictly more expressive than the $\lambda$-calculus in the sense of Felleisen (Felleisen, 1991): the forward embedding is macro-expressible (Theorem A.2), but the reverse is not (Theorem A.3).
Appendix B Multi-Path Self-Reference in Other Systems
This appendix demonstrates that two representative systems, Scala 3 and the NixOS module system, reject the multi-path self-reference pattern that overlay-calculus handles naturally.
Scala 3
Consider two objects that independently extend an outer class, each providing its own copy of an inner trait:
class MyOuter:
  trait MyInner:
    def outer = MyOuter.this
object Object1 extends MyOuter
object Object2 extends MyOuter
object MyObjectA extends Object1.MyInner
                     with Object2.MyInner
Scala 3 rejects MyObjectA with the error:
trait MyInner is extended twice
object MyObjectA cannot be instantiated since
  it has conflicting base types
  Object1.MyInner and Object2.MyInner
The rejection is not a surface-level restriction but a consequence of DOT’s type-theoretic foundations (Amin et al., 2016). In DOT, an object is constructed as $\nu(x:T)\,d$, where the self variable $x$ binds to a single object. Object1.MyInner and Object2.MyInner are distinct path-dependent types, each constraining the outer self-reference to a different object. Their intersection requires outer to simultaneously return Object1 and Object2, but DOT’s self variable is single-valued, making this intersection unrealizable.
The equivalent overlay-calculus definition is:
$\{\quad\mathrm{MyOuter}$ $\mapsto\{\mathrm{MyInner}\mapsto\{\mathrm{outer}\mapsto\mathrm{MyOuter}.\mathbf{this}\}\},$
$\mathrm{Object1}$ $\mapsto\{\mathrm{MyOuter}\},$
$\mathrm{Object2}$ $\mapsto\{\mathrm{MyOuter}\},$
$\mathrm{MyObjectA}$ $\mapsto\{\mathrm{Object1}.\mathrm{MyInner},\;\mathrm{Object2}.\mathrm{MyInner}\}\quad\}$
This is well-defined in overlay-calculus. The reference $\mathrm{MyOuter}.\mathbf{this}$ inside $\mathrm{outer}$ has de Bruijn index $n=1$: starting from $\mathrm{outer}$’s enclosing scope $\mathrm{MyInner}$, one $\mathtt{this}$ step reaches $\mathrm{MyOuter}$. When $\mathrm{MyObjectA}$ inherits from both $\mathrm{Object1}.\mathrm{MyInner}$ and $\mathrm{Object2}.\mathrm{MyInner}$, the $\mathtt{this}$ function (equation 6) resolves $\mathrm{MyOuter}.\mathbf{this}$ by searching through $\mathtt{supers}(\mathrm{MyObjectA})$ for overlay paths matching $\mathrm{MyInner}$’s definition site. It finds two inheritance-site paths, one through $\mathrm{Object1}$ and one through $\mathrm{Object2}$, and returns both. Both paths lead to records that inherit from the same $\mathrm{MyOuter}$, so querying $\mathrm{MyObjectA}.\mathrm{outer}$ yields the $\mathtt{properties}$ of $\mathrm{MyOuter}$ regardless of which route is taken. Since inheritance is idempotent, the two routes contribute the same properties and no disambiguation is needed.
NixOS Module System
The NixOS module system rejects the same pattern with a static error. The translation uses the module system’s own abstractions: deferredModule for traits, which are unevaluated module values imported into other fixpoints, and submoduleWith for objects, which are evaluated in their own fixpoints.
let
  lib = (import <nixpkgs> { }).lib;
  result = lib.evalModules {
    modules = [
      (toplevel@{ config, ... }: {
        # class MyOuter { trait MyInner {
        #   def outer = MyOuter.this } }
        options.MyOuter = lib.mkOption {
          default = { };
          type = lib.types.deferredModuleWith {
            staticModules = [
              (MyOuter: {
                options.MyInner = lib.mkOption {
                  default = { };
                  type = lib.types.deferredModuleWith {
                    staticModules = [
                      (MyInner: {
                        options.outer = lib.mkOption {
                          default = { };
                          type =
                            lib.types.deferredModuleWith {
                              staticModules =
                                [ toplevel.config.MyOuter ];
        # object Object1 extends MyOuter
        options.Object1 = lib.mkOption {
          default = { };
          type = lib.types.submoduleWith {
            modules = [ toplevel.config.MyOuter ];
        # object Object2 extends MyOuter
        options.Object2 = lib.mkOption {
          default = { };
          type = lib.types.submoduleWith {
            modules = [ toplevel.config.MyOuter ];
        # object MyObjectA extends
        #   Object1.MyInner with Object2.MyInner
        options.MyObjectA = lib.mkOption {
          default = { };
          type = lib.types.submoduleWith {
            modules = [
            toplevel.config.Object1.MyInner
            toplevel.config.Object2.MyInner
in
  builtins.attrNames result.config.MyObjectA.outer
The NixOS module system rejects this with:
error: The option ‘MyObjectA.outer’
  in ‘<unknown-file>’
  is already declared
  in ‘<unknown-file>’.
The rejection occurs because Object1.MyInner and Object2.MyInner each carry the full staticModules of the MyInner deferred module, including the declaration options.outer. When MyObjectA imports both, the module system’s mergeOptionDecls encounters two declarations of the same option and raises a static error; it cannot express that both declarations originate from the same trait and should be unified.
The rejection is analogous to Scala’s “conflicting base types” error above: both systems assume that each option or type member has a single declaration site, and reject the multi-path situation where two inheritance routes introduce the same declaration independently. In overlay-calculus, the same pattern is well-defined because $\mathtt{overlays}$ (equation 3) recognizes that both routes lead to the same definition, and $\mathtt{supers}$ (equation 2) collects both inheritance-site contexts without duplication.
Appendix C Well-Definedness of the Semantic Functions
This appendix proves that the mutually recursive equations (1)–(6) have a unique least fixed point, and therefore the semantics of Section 3 is well-defined.
C.1. The Product Lattice
Fix an AST with its primitive functions $\mathtt{defines}$ and $\mathtt{inherits}$. Let $\mathcal{L}$ denote the set of all labels and $\mathcal{P}$ the set of all paths (finite sequences of labels). Define the following domains, one for each semantic function:
$D_{\mathtt{properties}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{L})$ (sets of labels)
$D_{\mathtt{supers}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P}\times\mathcal{P})$ (sets of site–overlay pairs)
$D_{\mathtt{overlays}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P})$ (sets of paths)
$D_{\mathtt{bases}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P})$ (sets of paths)
$D_{\mathtt{resolve}}$ $=(\mathcal{P}\times\mathcal{P}\times\mathbb{N}\times\mathcal{L}^{*})\to\mathcal{P}(\mathcal{P})$ (sets of paths)
$D_{\mathtt{this}}$ $=(\mathcal{P}(\mathcal{P})\times\mathcal{P}\times\mathbb{N})\to\mathcal{P}(\mathcal{P})$ (sets of paths)
Each $D_{i}$ is a complete lattice under the pointwise subset ordering: $f\sqsubseteq g$ iff $f(x)\subseteq g(x)$ for all $x$. The product $D=D_{\mathtt{properties}}\times D_{\mathtt{supers}}\times D_{\mathtt{overlays}}\times D_{\mathtt{bases}}\times D_{\mathtt{resolve}}\times D_{\mathtt{this}}$ is a complete lattice under the componentwise ordering, with bottom element $\bot=(\lambda x.\,\varnothing,\ldots,\lambda x.\,\varnothing)$.
C.2. The Operator
Equations (1)–(6) define a function $F:D\to D$ that takes a tuple of candidate interpretations $(\mathtt{properties},\mathtt{supers},\mathtt{overlays},\mathtt{bases},\mathtt{resolve},\mathtt{this})$ and produces a new tuple by evaluating the right-hand side of each equation. The transitive closure $\mathtt{bases}^{*}$ appearing in equation (2) is derived from $\mathtt{bases}$: for any fixed $\mathtt{bases}$, the set $\mathtt{bases}^{*}(p)$ is defined as the least fixed point of the monotone operator $X\mapsto\{p\}\cup\bigcup_{q\in X}\mathtt{bases}(q)$ on $\mathcal{P}(\mathcal{P})$. If $\mathtt{bases}\sqsubseteq\mathtt{bases}^{\prime}$ (pointwise), then $\mathtt{bases}^{*}\sqsubseteq\mathtt{bases}^{\prime*}$ (pointwise), since enlarging the one-step relation can only enlarge the transitive closure.
The $\mathtt{this}$ function involves a recursion on the de Bruijn index $n$. Since $n$ decreases by one at each step and is a non-negative integer, this recursion terminates for any fixed interpretation of $\mathtt{supers}$. The result of $\mathtt{this}(S,p_{\mathrm{def}},n)$ for all valid $(S,p_{\mathrm{def}},n)$ is therefore well-defined as a function of $\mathtt{supers}$, and it remains to show that it is monotone in $\mathtt{supers}$.
C.3. Monotonicity
Proposition C.1.
The operator $F$ is monotone: if $\vec{d}\sqsubseteq\vec{d}^{\prime}$ in $D$, then $F(\vec{d})\sqsubseteq F(\vec{d}^{\prime})$.
Proof.
It suffices to show that each component of $F$ is monotone in every argument, i.e., that enlarging any of the input interpretations (in the $\subseteq$ ordering) can only enlarge the output. We verify this by inspecting the right-hand side of each equation and confirming that every occurrence of a mutually defined function appears in a positive position—as the domain of an existential quantifier in a set comprehension (Aczel, 1977).
Equation (1): $\mathtt{properties}$
$\mathtt{properties}(p)=\bigl\{\,\ell\bigm|(\_,p_{\mathrm{overlay}})\in\mathtt{supers}(p),\;\ell\in\mathtt{defines}(p_{\mathrm{overlay}})\,\bigr\}.$
The function $\mathtt{supers}$ appears as the set being iterated over ($\in\mathtt{supers}(p)$). The predicate $\ell\in\mathtt{defines}(\cdot)$ involves only the AST primitive $\mathtt{defines}$, not any mutually defined function. Enlarging $\mathtt{supers}(p)$ can only add pairs, hence can only add labels to the result. Positive in $\mathtt{supers}$.
Equation (2): $\mathtt{supers}$
$\mathtt{supers}(p)=\bigl\{\,(\mathtt{init}(p_{\mathrm{base}}),p_{\mathrm{overlay}})\bigm|p_{\mathrm{base}}\in\mathtt{bases}^{*}(p),\;p_{\mathrm{overlay}}\in\mathtt{overlays}(p_{\mathrm{base}})\,\bigr\}$. Both $\mathtt{bases}^{*}$ and $\mathtt{overlays}$ appear as sets being iterated over. As noted above, $\mathtt{bases}^{*}$ is monotone in $\mathtt{bases}$. Enlarging either $\mathtt{bases}$ or $\mathtt{overlays}$ can only add elements to the result. Positive in $\mathtt{bases}$ and $\mathtt{overlays}$.
Equation (3): $\mathtt{overlays}$
For $p\neq()$:
$\mathtt{overlays}(p)=\{p\}\cup\bigl\{\,p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)\bigm|{}\\ (\_,p_{\mathrm{branch}})\in\mathtt{supers}(\mathtt{init}(p)),\;\mathtt{last}(p)\in\mathtt{defines}(p_{\mathrm{branch}})\,\bigr\}.$
The function $\mathtt{supers}$ appears as the set being iterated over. The filtering condition $\mathtt{last}(p)\in\mathtt{defines}(\cdot)$ involves only the AST primitive. Enlarging $\mathtt{supers}$ can only add branches, hence can only add paths to the result. Positive in $\mathtt{supers}$.
Equation (4): $\mathtt{bases}$
$\mathtt{bases}(p)=\bigl\{\,p_{\mathrm{target}}\bigm|p_{\mathrm{overlay}}\in\mathtt{overlays}(p),\\ (n,\ell_{\mathrm{down},*})\in\mathtt{inherits}(p_{\mathrm{overlay}}),\;p_{\mathrm{target}}\in\mathtt{resolve}(\ldots)\,\bigr\}.$
Both $\mathtt{overlays}$ and $\mathtt{resolve}$ appear as sets being iterated over. The function $\mathtt{inherits}$ is an AST primitive. Enlarging either $\mathtt{overlays}$ or $\mathtt{resolve}$ can only add targets to the result. Positive in $\mathtt{overlays}$ and $\mathtt{resolve}$.
Equation (5): $\mathtt{resolve}$
$\mathtt{resolve}(\ldots)=\bigl\{\,p_{\mathrm{current}}\mathbin{\triangleright}\ell_{\mathrm{down},1}\mathbin{\triangleright}\cdots\bigm|p_{\mathrm{current}}\in\mathtt{this}(\{p_{\mathrm{site}}\},\mathtt{init}(p_{\mathrm{def}}),n)\,\bigr\}.$
The function $\mathtt{this}$ appears as the set being iterated over. Enlarging $\mathtt{this}$ can only add paths to the result. Positive in $\mathtt{this}$.
Equation (6): $\mathtt{this}$
For $n>0$: $\mathtt{this}(S,p_{\mathrm{def}},n)=\mathtt{this}(S^{\prime},\mathtt{init}(p_{\mathrm{def}}),n-1)$ where $S^{\prime}=\bigl\{\,p_{\mathrm{site}}\bigm|p_{\mathrm{current}}\in S,\;(p_{\mathrm{site}},p_{\mathrm{overlay}})\in\mathtt{supers}(p_{\mathrm{current}}),\;p_{\mathrm{overlay}}=p_{\mathrm{def}}\,\bigr\}$. The function $\mathtt{supers}$ appears as the set being iterated over. The filtering condition $p_{\mathrm{overlay}}=p_{\mathrm{def}}$ is an equality test on path data, not involving any mutually defined function. Enlarging $\mathtt{supers}$ can only enlarge $S^{\prime}$, and by induction on $n$ (the base case $n=0$ returns $S$ unchanged), the recursive call is monotone in $S^{\prime}$. Positive in $\mathtt{supers}$.
In summary, every mutually defined function appears on the right-hand side of every equation only inside the pattern $x\in f(\ldots)$ within a set comprehension, i.e., as the domain of an existential quantifier. No complement, set difference, or negated membership test appears in any equation. Therefore $F$ is a monotone operator on the complete lattice $D$. ∎
C.4. Existence of the Least Fixed Point
Theorem C.2.
The semantic functions $\mathtt{properties}$, $\mathtt{supers}$, $\mathtt{overlays}$, $\mathtt{bases}$, $\mathtt{resolve}$, $\mathtt{this}$ are well-defined as the least fixed point of $F$.
Proof.
By Proposition C.1, $F$ is monotone on the complete lattice $D$. By the Knaster–Tarski theorem (Tarski, 1955), $F$ has a least fixed point $\vec{d}^{*}=\bigsqcup_{n\geq 0}F^{n}(\bot)$, where $\bot$ assigns the empty set to every input. The components of $\vec{d}^{*}$ are the semantic functions of Section 3.
Concretely, the least fixed point is the union of the finite approximation chain: $\bot\sqsubseteq F(\bot)\sqsubseteq F^{2}(\bot)\sqsubseteq\cdots$. A membership $\ell\in\mathtt{properties}(p)$ holds in $\vec{d}^{*}$ if and only if it appears at some finite stage $F^{k}(\bot)$, which corresponds to a finite derivation tree using the equations. When no finite derivation exists, the membership is absent from the least fixed point and the query is undefined, the analogue of divergence. ∎
Appendix D Böhm Tree Correspondence: Proofs
This appendix contains the full proofs of Adequacy (Theorem 4.2) and Full Abstraction (Theorem 4.3) stated in Section 4.1.
D.1. Böhm Trees
We briefly recall the definition of Böhm trees (Barendregt, 1984). A head reduction $M\to_{h}M^{\prime}$ contracts the outermost $\beta$-redex only: if $M$ has the form $(\lambda x.\,e)\;v\;M_{1}\cdots M_{k}$, then $M\to_{h}e[v/x]\;M_{1}\cdots M_{k}$. A term $M$ is in head normal form (HNF) if it has no head redex, i.e., $M=\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}$ where $y$ is a variable.
The Böhm tree $\mathrm{BT}(M)$ of a $\lambda$-term $M$ is an infinite labeled tree defined by:
$\mathrm{BT}(M)=\begin{cases}\bot&\text{if }M\text{ has no HNF}\\[6.0pt] \lambda x_{1}\ldots x_{m}.\;y\bigl(\mathrm{BT}(M_{1}),\;\ldots,\;\mathrm{BT}(M_{k})\bigr)&\begin{aligned} &\text{if }M\to^{*}_{h}\\[-2.0pt] &\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}\end{aligned}\end{cases}$
Böhm tree equivalence $\mathrm{BT}(M)=\mathrm{BT}(N)$ is the standard observational equivalence of the lazy $\lambda$-calculus (Barendregt, 1984; Abramsky and Ong, 1993).
D.2. Path Encoding
We define a correspondence between positions in the Böhm tree and paths in the overlay tree. When the translation $\mathcal{T}$ is applied to an ANF $\lambda$-term, the resulting overlay tree has a specific shape:
  • An abstraction $\lambda x.\,M$ translates to a record with own properties $\{\mathrm{argument},\mathrm{result}\}$ (the abstraction shape).
  • A $\mathbf{let}$-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M$ translates to a record with own properties $\{x,\mathrm{result}\}$, where $x$ holds the encapsulated application $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$.
  • A tail call $V_{1}\;V_{2}$ translates to a record with own properties $\{\_\mathrm{tail},\mathrm{result}\}$, where $\mathrm{tail}$ holds the encapsulated application and $\mathrm{result}$ projects $\mathrm{tail}.\mathrm{result}$.
  • A variable $x$ with de Bruijn index $n$ translates to an indexed reference $\uparrow^{\!n}.\mathrm{argument}$.
Each Böhm tree position is a sequence of navigation steps from the root. In the Böhm tree, the possible steps at a node $\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}$ are: entering the body (peeling off one abstraction layer) and entering the $i$-th argument $M_{i}$ of the head variable. In the overlay tree, these correspond to following the $\mathrm{result}$ label (for entering the body) and the $\mathrm{argument}$ label after $\mathtt{this}$ resolution (for entering an argument position).
Rather than comparing absolute paths across different overlay trees (which would be sensitive to internal structure), we observe only the convergence behavior accessible via the $\mathrm{result}$ projection. Recall the definition of overlay-convergence (Definition 4.1). Each $\mathrm{result}$-following step corresponds to one head reduction step in the $\lambda$-calculus: a tail call’s $\mathrm{result}$ projects through the encapsulated application, and a $\mathbf{let}$-binding’s $\mathrm{result}$ enters the continuation. The abstraction shape at the end signals that a weak head normal form has been reached.
D.3. Single-Path Lemma
The $\mathtt{this}$ function (equation 6) is designed for the general case of multi-path overlay inheritance. For translated $\lambda$-terms, we show it degenerates to single-path resolution.
Lemma D.1 (Single Path).
For any closed ANF $\lambda$-term $M$, at every invocation of $\mathtt{this}(S,\;p_{\mathrm{def}},\;n)$ during evaluation of $\mathcal{T}(M)$, the frontier set $S$ contains exactly one path.
Proof.
By induction on the structure of the ANF term.
Case $\mathcal{T}(\lambda x.\,M)=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
This is a record literal with $\mathtt{defines}=\{\mathrm{argument},\mathrm{result}\}$ and no inheritance sources, so $\mathtt{inherits}=\varnothing$. At the root path $p$, we have $\mathtt{overlays}(p)=\{p\}$ (a single path) since there are no inheritance sources to introduce additional branches. Inside $\mathcal{T}(M)$, any reference to $x$ is $\uparrow^{\!n}.\mathrm{argument}$ where $n$ is the de Bruijn index pointing to this scope level. Since $\mathcal{T}(\lambda x.\,M)$ introduces exactly one scope level, there is exactly one overlay at that level, hence $\mathtt{this}$ finds exactly one matching pair $(p_{\mathrm{site}},\;p_{\mathrm{overlay}})$ in $\mathtt{supers}$.
Case $\mathcal{T}(\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M)=\{x\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
The outer record has own properties $\{x,\mathrm{result}\}$ and no inheritance sources at the outer level ($\mathtt{inherits}=\varnothing$), so $\mathtt{overlays}(\text{root})=\{\text{root}\}$. Inside the $x$ subtree, the inheritance $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$ has exactly one inheritance source ($\mathcal{T}(V_{1})$), creating exactly one $\mathtt{bases}$ entry. By induction on $V_{1}$, the single-path property holds inside $\mathcal{T}(V_{1})$. Inside $\mathcal{T}(M)$, the induction hypothesis applies directly.
Case $\mathcal{T}(V_{1}\;V_{2})=\{\_\mathrm{tail}\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\_\mathrm{tail}.\mathrm{result}\}$
Identical to the $\mathbf{let}$-binding case: the outer record has own properties $\{\_\mathrm{tail},\mathrm{result}\}$ with no inheritance sources at the outer level, and the $\mathrm{tail}$ subtree contains exactly one inheritance source.
Case $\mathcal{T}(x)=\uparrow^{\!n}.\mathrm{argument}$
This is a reference, not a record. The $\mathtt{this}$ function walks up $n$ steps from the enclosing scope. By the inductive hypothesis on the enclosing translations, each step of $\mathtt{this}$ encounters exactly one matching pair in $\mathtt{supers}$.
The key structural invariant is that the translation $\mathcal{T}$ never produces two independent inheritance sources at the same scope level that both define the same label. In every case, application inheritances are encapsulated inside a named property ($x$ or $\mathrm{tail}$), and each such encapsulated inheritance has exactly one inheritance source. This prevents diamond inheritance and ensures single-path resolution throughout. ∎
D.4. Substitution Lemma
The core mechanism of the translation is that inheritance with $\mathrm{argument}\mapsto\mathcal{T}(V)$ plays the role of substitution. We make this precise.
Lemma D.2 (Substitution).
Let $M$ be an ANF term in which $x$ may occur free, and let $V$ be a closed value. Define the inherited tree $C=\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$. Then for every path $p$ under $C\mathbin{\triangleright}\mathrm{result}$ and every label $\ell$:
$\ell\in\mathtt{properties}(p)\text{ in }C\mathbin{\triangleright}\mathrm{result}\quad\Longleftrightarrow\quad\ell\in\mathtt{properties}(p)\text{ in }\mathcal{T}(M[V/x])$
where $M[V/x]$ is the usual capture-avoiding substitution.
Proof.
By structural induction on $M$.
Base case: $M=x$
Then $\mathcal{T}(x)=\uparrow^{\!n}.\mathrm{argument}$ (where $n$ is the de Bruijn index of $x$), and $C\mathbin{\triangleright}\mathrm{result}$ contains this reference. In the inherited tree $C$, $\mathtt{resolve}$ and $\mathtt{this}$ resolve the reference by navigating from the reference’s enclosing scope up $n$ steps to the binding $\lambda$, where the inheritance $\{\mathcal{T}(\lambda x.\,x),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$ overlays the $\mathrm{argument}$ slot with $\mathcal{T}(V)$. Since $C$ is the $\mathrm{tail}$ subtree of $\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)$, which is a closed ANF term, Lemma D.1 applies and $\mathtt{this}$ finds exactly one path, so $\mathtt{resolve}$ returns $\mathcal{T}(V)$’s subtree. Since $M[V/x]=V$, we have $\mathcal{T}(M[V/x])=\mathcal{T}(V)$, and the properties coincide.
Base case: $M=y$ where $y\neq x$
Then $\uparrow^{\!m}.\mathrm{argument}$ (where $m$ is the de Bruijn index of $y$, $m\neq n$) walks up to a different scope level, so the inheritance with $\mathrm{argument}\mapsto\mathcal{T}(V)$ at the binding scope of $x$ has no effect. Since $M[V/x]=y$, the properties are identical.
Inductive case: $M=\lambda y.\,M^{\prime}$
Then $\mathcal{T}(M)=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$. The subtree at $C\mathbin{\triangleright}\mathrm{result}$ is another abstraction-shaped record. References to $x$ inside $\mathcal{T}(M^{\prime})$ resolve through $\mathtt{this}$ in exactly the same way (with one additional scope level from the inner $\lambda$), and by induction on $M^{\prime}$, the properties under $C\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}\mathrm{result}$ coincide with those of $\mathcal{T}(M^{\prime}[V/x])$. Since $(\lambda y.\,M^{\prime})[V/x]=\lambda y.\,(M^{\prime}[V/x])$ (assuming $y$ is fresh), the result follows.
Inductive case: $M=\mathbf{let}\;z=V_{1}\;V_{2}$ $\mathbf{in}\;M^{\prime}$
Then
$\mathcal{T}(M)=\{z\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}.$
The substitution distributes: $M[V/x]=\mathbf{let}\;z=V_{1}[V/x]\;V_{2}[V/x]\;\mathbf{in}\;M^{\prime}[V/x]$. By induction on $V_{1}$, $V_{2}$, and $M^{\prime}$, the properties under each subtree ($z$ and $\mathrm{result}$) coincide, since references to $x$ inside each subtree resolve to $\mathcal{T}(V)$ via the same $\mathtt{this}$ mechanism.
Inductive case: $M=V_{1}\;V_{2}$ (tail call)
Analogous to the $\mathbf{let}$-binding case with $\mathrm{tail}$ in place of $z$. ∎
D.5. Convergence Preservation
The convergence criterion (Definition 4.1) follows the $\mathrm{result}$ chain from the root. Each step in this chain corresponds to one head reduction step in the $\lambda$-calculus. We make this precise.
Lemma D.3 (Result-step).
Let $(\lambda x.\,M)\;V$ be a tail call (a $\beta$-redex). Then for every path $p$ and label $\ell$:
$\ell\in\mathtt{properties}\!\bigl(\,\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p\,\bigr)\text{ in }\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)\;\Longleftrightarrow\;\ell\in\mathtt{properties}\!\bigl(\,\text{root}\mathbin{\triangleright}p\,\bigr)\text{ in }\mathcal{T}(M[V/x])$
That is, following one $\mathrm{result}$ projection in the tail call’s overlay tree yields the same properties as the root of the reduct’s overlay tree.
Proof.
The translation gives:
$\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)=\{\_\mathrm{tail}\mapsto\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\},\;\mathrm{result}\mapsto\_\mathrm{tail}.\mathrm{result}\}$
The $\mathrm{result}$ label at the root is defined as $\mathrm{tail}.\mathrm{result}$, so $\text{root}\mathbin{\triangleright}\mathrm{result}$ resolves to the $\mathrm{result}$ path inside the inheritance $\mathrm{tail}=\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$. This inheritance is exactly the inherited tree $C$ of Lemma D.2, and $\mathrm{tail}\mathbin{\triangleright}\mathrm{result}$ is $C\mathbin{\triangleright}\mathrm{result}$. By Lemma D.2, the properties under $C\mathbin{\triangleright}\mathrm{result}$ coincide with those of $\mathcal{T}(M[V/x])$.
Let-binding variant. An analogous result holds for $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}$ where $V_{1}=\lambda y.\,M^{\prime\prime}$. The translation gives:
$\mathcal{T}\bigl(\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}\bigr)=\{x\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$
Since $\mathrm{result}\mapsto\mathcal{T}(M^{\prime})$ is a property definition, the subtree at $\text{root}\mathbin{\triangleright}\mathrm{result}$ is $\mathcal{T}(M^{\prime})$. Inside $\mathcal{T}(M^{\prime})$, each reference to $x$ is $x.\mathrm{result}$, which resolves to the $\mathrm{result}$ path inside the $x$ subtree $\{\mathcal{T}(\lambda y.\,M^{\prime\prime}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$. This is the inherited tree $C$ of Lemma D.2, so $x.\mathrm{result}$ has the same properties as $\mathcal{T}(M^{\prime\prime}[V_{2}/y])$. In the $\lambda$-calculus, $M=(\lambda x.\,M^{\prime})(V_{1}\;V_{2})$, and the head reduct is $N=M^{\prime}[(V_{1}\;V_{2})/x]$. In $\mathcal{T}(N)$, each occurrence of $x$ is replaced by $(V_{1}\;V_{2})$, whose result, by the tail-call variant above, also yields $\mathcal{T}(M^{\prime\prime}[V_{2}/y])$. Therefore the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p$ in $\mathcal{T}(M)$ coincide with those at $\text{root}\mathbin{\triangleright}p$ in $\mathcal{T}(N)$ for all $p$. ∎
Theorem D.4 (Convergence preservation).
If $M\to_{h}N$ (head reduction), then $\mathcal{T}(M){\Downarrow}$ if and only if $\mathcal{T}(N){\Downarrow}$.
Proof.
A head reduction step in ANF takes one of two forms: a tail call $(\lambda x.\,M^{\prime})\;V\to_{h}M^{\prime}[V/x]$, or a $\mathbf{let}$-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}\to_{h}M^{\prime}[(V_{1}\;V_{2})/x]$ (since $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}=(\lambda x.\,M^{\prime})(V_{1}\;V_{2})$ in the $\lambda$-calculus). In both cases, Lemma D.3 (tail-call and let-binding variants respectively) shows that the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}^{n}$ in $\mathcal{T}(M)$ coincide with the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}^{n-1}$ in $\mathcal{T}(N)$ for all $n\geq 1$, where $N$ is the head reduct. Hence the abstraction shape appears at depth $n$ in the pre-reduct if and only if it appears at depth n1n-1 in the post-reduct. ∎
D.6. Adequacy
A closed ANF term at the top level is either an abstraction (a value) or a computation (a $\mathbf{let}$-binding or tail call). An abstraction translates to a record whose root immediately has the abstraction shape; a computation’s root has the form $\{x,\mathrm{result}\}$ or $\{\_\mathrm{tail},\mathrm{result}\}$, and one must follow the $\mathrm{result}$ chain to find the eventual value.
Proof of Theorem 4.2.
($\Rightarrow$) Suppose $M\to^{*}_{h}\lambda x.\,M^{\prime}$ in $k$ head reduction steps. We show $\mathcal{T}(M)$ converges at depth $\leq k$ by induction on $k$.
Base case ($k=0$): $M$ is already an abstraction $\lambda x.\,M^{\prime}$. Then $\mathcal{T}(M)=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$, whose root has $\mathtt{defines}=\{\mathrm{argument},\mathrm{result}\}$. These labels appear at stage $F^{1}(\bot)$ of the approximation chain, so $\mathcal{T}(M)$ converges at depth $n=0$.
Inductive step ($k\geq 1$): $M$ is not an abstraction, so $M$ is either a tail call or a $\mathbf{let}$-binding, both of which are $\beta$-redexes. Let $N$ be the head reduct, so $M\to_{h}N\to^{k-1}_{h}\lambda x.\,M^{\prime}$. By Lemma D.3 (tail-call or let-binding variant), the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p$ in $\mathcal{T}(M)$ coincide with those at $\text{root}\mathbin{\triangleright}p$ in $\mathcal{T}(N)$. By the induction hypothesis, $\mathcal{T}(N)$ converges at depth $\leq k-1$, so $\mathcal{T}(M)$ converges at depth $\leq k$.
($\Leftarrow$) Suppose $\mathcal{T}(M){\Downarrow}$ at depth $n$. We show $M$ has a head normal form by induction on $n$.
Base case ($n=0$): the root of $\mathcal{T}(M)$ has abstraction shape $\{\mathrm{argument},\mathrm{result}\}$. Only the abstraction rule $\mathcal{T}(\lambda x.\,M^{\prime})$ produces a root with own property $\mathrm{argument}$. (For tail calls and $\mathbf{let}$-bindings, the root has $\mathtt{defines}=\{\_\mathrm{tail},\mathrm{result}\}$ or $\{x,\mathrm{result}\}$, and $\mathrm{argument}$ cannot appear via $\mathtt{supers}$ because $\mathtt{inherits}=\varnothing$ at the root, giving $\mathtt{bases}(\text{root})=\varnothing$.) Therefore $M$ is an abstraction, hence already in head normal form.
Inductive step ($n\geq 1$): the abstraction shape appears at depth $n$ but not at depth 0. By the base case argument, $M$ is not an abstraction, so $M$ is either a tail call $V_{1}\;V_{2}$ or a $\mathbf{let}$-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}$. In both cases, $M$ is a $\beta$-redex (since $V_{1}=\lambda x.\,M^{\prime}$ or $V_{1}=\lambda y.\,M^{\prime\prime}$ respectively, the only closed ANF value). By Lemma D.3 (tail-call or let-binding variant respectively), the properties at depth n1n-1 in $\mathcal{T}(N)$ match those at depth $n$ in $\mathcal{T}(M)$, where $N$ is the head reduct. So $\mathcal{T}(N)$ converges at depth n1n-1. By the induction hypothesis, $N$ has a head normal form, and since $M\to_{h}N$, so does $M$. ∎
D.7. Full Abstraction
Adequacy relates a single term’s head normal form to its overlay-convergence. Full abstraction lifts this to an equivalence between terms: the translation $\mathcal{T}$ preserves and reflects observational equivalence.
Definition D.5 (Overlay-contextual equivalence).
For closed $\lambda$-terms $M$ and $N$, write $\mathcal{T}(M)\approx\mathcal{T}(N)$ if for every closing $\lambda$-calculus context $C[\cdot]$:
$\mathcal{T}(C[M]){\Downarrow}\;\Longleftrightarrow\;\mathcal{T}(C[N]){\Downarrow}$
This definition observes only convergence (a single bit), but quantifying over all contexts makes it a fine-grained equivalence: different contexts can supply different arguments and project different results, probing every aspect of the term’s behavior.
Proof of Theorem 4.3.
The translation $\mathcal{T}$ is compositional: for every $\lambda$-calculus context $C[\cdot]$, the overlay tree $\mathcal{T}(C[M])$ is determined by the overlay tree $\mathcal{T}(M)$ and the translation of the context. This means $\mathcal{T}$ preserves the context structure needed for the argument.
($\Leftarrow$) Suppose $\mathrm{BT}(M)=\mathrm{BT}(N)$. Böhm tree equivalence is a congruence (Barendregt, 1984), so for every context $C[\cdot]$, $\mathrm{BT}(C[M])=\mathrm{BT}(C[N])$. In particular, $C[M]$ has a head normal form iff $C[N]$ does. By Theorem 4.2, $\mathcal{T}(C[M]){\Downarrow}$ iff $\mathcal{T}(C[N]){\Downarrow}$. Hence $\mathcal{T}(M)\approx\mathcal{T}(N)$.
($\Rightarrow$) Suppose $\mathrm{BT}(M)\neq\mathrm{BT}(N)$. Böhm tree equivalence coincides with observational equivalence for the lazy $\lambda$-calculus (Barendregt, 1984; Abramsky and Ong, 1993): there exists a context $C[\cdot]$ such that $C[M]$ converges and $C[N]$ diverges (or vice versa). By Theorem 4.2, $\mathcal{T}(C[M]){\Downarrow}$ and $\mathcal{T}(C[N]){\Uparrow}$. Hence $\mathcal{T}(M)\not\approx\mathcal{T}(N)$. ∎
The proof rests on two pillars: Adequacy (Theorem 4.2), which is our contribution, and the classical result that Böhm tree equivalence equals observational equivalence for the lazy $\lambda$-calculus (Barendregt, 1984; Abramsky and Ong, 1993). The overlay tree semantics inherits full abstraction from the $\lambda$-calculus without needing to define or compare absolute paths inside overlay trees.