### v4

# A Calculus of Inheritance

Abstract.
Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) as the foundation of functional programming, inheritance-calculus uses three primitives (record, definition, inheritance) as the foundation of declarative programming. It trivially embeds the $\lambda$-calculus, although the entire semantics rests solely on naive set theory; as a consequence, all constructs including inheritance are inherently commutative, idempotent, and associative; the linearization problem of multiple inheritance does not arise. This induces a fully abstract semantics of the lazy $\lambda$-calculus with respect to Böhm tree equivalence (Barendregt, 1984). Inheritance-calculus is distilled from MIXINv2, a practical implementation in which we observed further emergent phenomena: the same code acts as different function colors (Nystrom, 2015); ordinary arithmetic yields the relational semantics of logic programming (van Emden and Kowalski, 1976); self-reference resolves to multiple targets; and programs are immune to the Expression Problem (Wadler, 1998). This makes inheritance-calculus strictly more expressive than the $\lambda$-calculus in both common sense and Felleisen’s sense (Felleisen, 1991). These properties suggest applications to configuration languages, dependency injection, object-oriented programming, composable effect systems, modular software architectures, file-system-as-compiler, general-purpose programming, and no-code development.
declarative programming, inheritance, denotational semantics, self-referential records, expression problem, $\lambda$-calculus, Böhm trees, configuration languages
This work was conducted independently prior to the author’s employment at Figure AI.
$c$s: Theory of computation Denotational semanticsccs: Theory of computation Object oriented constructsccs: Theory of computation Functional constructsccs: Theory of computation Program schemesccs: Software and its engineering Formal language definitionsccs: Software and its engineering Extensible languagesccs: Software and its engineering Object oriented languages
1. Introduction
Declarative and configuration languages are ubiquitous in modern software engineering. Systems such as NixOS modules (Dolstra et al., 2010; NixOS Contributors, 2024), Jsonnet (Cunningham, 2014), Hydra (Yadan and Droppo, 2023), CUE (van Lohuizen, 2019), Dhall (Gonzalez, 2017), Kustomize (Kubernetes SIG CLI, 2018), and JSON Patch (Bryan et al., 2013) all provide mechanisms for composing structured data through inheritance or merging. Among these, the NixOS module system stands out: its recursive attribute set merging with fixed-point semantics and deferred modules (NixOS Contributors, 2024) achieves considerable expressiveness in practice, and the mechanism has been adopted well beyond NixOS itself: Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), disko (Lassulus and nix-community Contributors, 2022), flake-parts (Hensing, 2022), dream2nix (Hauer, 2021), devenv (Kožar, 2022), KubeNix (Xtruder and Hall, 2022), and nixidy (Ingólfsson, 2024) use it to manage user environments, macOS configuration, disk partitioning, flake structure, multi-language packaging, developer environments, and Kubernetes clusters respectively. Yet no computational theory explains why this mechanism is so expressive.
The $\lambda$-calculus serves as the foundational computational model for functional programming. No analogous calculus exists for declarative programming. This gap matters because the two paradigms differ in fundamental ways. Configuration languages are inherently declarative: their values are immutable and they have no first-class functions. At first glance, Turing completeness appears incompatible with these constraints. The three classical models of computation each violate at least one of them:
Model Turing complete Immutable No $\lambda$
Turing Machine $\checkmark$ $\times$ $\checkmark$
$\lambda$-calculus $\checkmark$ $\checkmark$ $\times$
RAM Machine $\checkmark$ $\times$ $\checkmark$
The Turing Machine and RAM Machine require mutable state; the $\lambda$-calculus requires first-class functions. A computational model that is Turing complete, immutable, and function-free would sit in the empty cell of this table. The NixOS module system already suggests that such a model exists: its inheritance-based composition over recursive records, without explicit functions, is expressive enough to configure entire operating systems.
Conventionally, the value domain of configuration languages is assumed to consist of finite, well-founded structures, that is, initial algebras in the sense of universal algebra. We challenge this assumption: configuration values are better understood as lazily observable, possibly infinite structures, where semantics is determined by the observer and a finite prefix suffices for any finite observation. This is not merely a theoretical distinction: the nixpkgs package collection (Dolstra, 2006) (over 100,000 packages) behaves as a lazily observed structure in practice, allowing any single package to be evaluated without materializing the entire set. Guided by this observation, we set out to reduce the NixOS module system to a minimal set of primitives. The reduction produced MIXIN, an early implementation in Nix, and subsequently MIXINv2 (Yang, 2026), now implemented in Python: a declarative programming language with only three constructs (record literals, properties, and inheritance) and no functions or let-bindings.
In developing real programs in MIXINv2, including boolean logic and natural number arithmetic (Section 5) and a web application (Section Function color blindness: ), we found that the linearization problem of mixin-based systems (Bracha and Cook, 1990; Barrett et al., 1996) is structurally absent: because records are sets, inheritance is inherently commutative, idempotent, and associative (Section 3). Self-reference naturally resolves to multiple objects rather than one; prior systems assume single-object resolution and reject the multi-path situation (Scala, NixOS module system). Self-referential records require no fixed-point combinator—the inheritance tree is the call stack, and qualified this resolution is a first-order set-theoretic query, not a fixed-point construction. We also found that boolean logic and arithmetic, implemented as separate MIXINv2 files, extend in both dimensions of the Expression Problem (Wadler, 1998) without any dedicated mechanism; that Church-encoded values automatically form trie structures; and that operations on tries automatically become Cartesian products, so that adding $\{1,2\}$ to $\{3,4\}$ yields $\{4,5,6\}$, which is the relational semantics of logic programming (van Emden and Kowalski, 1976) (Section 5); and that the same web application MIXINv2 file runs under both Python asyncio and synchronous execution without modification, an instance of function color blindness (Nystrom, 2015), equivalently CPS-agnostic behavior (Section Function color blindness: ).
This paper distills these observations into inheritance-calculus, a minimal computational model for declarative programming. Function application translates directly into inheritance: a $\lambda$-abstraction corresponds to a record with an $\mathrm{argument}$ slot and a $\mathrm{result}$ slot, and function application corresponds to inheriting from the function record and a record that supplies the argument. Since the $\lambda$-calculus embeds into inheritance-calculus, inheritance-calculus is Turing complete. The translation reveals an asymmetry: embedding the $\lambda$-calculus into inheritance-calculus requires five local, syntax-directed rules (Section 4), but the reverse is not macro-expressible under the equivalence notions fixed in Section 4.2 and Appendix A— inheritance-calculus enables open extension that changes operational equivalence on $\lambda$-terms, establishing a strict separation in Felleisen’s sense of expressiveness (Felleisen, 1991) (Section 4.2).
Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) to serve as the foundation of functional programming, inheritance-calculus uses three primitives (record, definition, inheritance) to serve as the foundation of declarative programming.
All constructions in this paper have been implemented and tested in MIXINv2, available as open source (Yang, 2026), with a test suite covering every example. [^1] MIXINv2 is statically typed and therefore stricter than the untyped inheritance-calculus. All examples in this paper are ported from MIXINv2 and conform to both semantics. Elements added to satisfy the type checker are harmless for the algorithms presented here and may aid readability. All tests are mechanically verified against expected output snapshots.
Contributions
  • We present inheritance-calculus, a minimal computational model for declarative programming that contains only three constructs (record, definition, inheritance) and no functions or scalar types.
  • Inheritance operates at the tree level: a record is a set of elements, both definitions and inheritance sources, so inheriting from two mixins that share a label recursively merges their subtrees (deep merge), rather than overriding at the method level as in prior mixin and trait calculi. Since records are sets, inheritance is inherently commutative, idempotent, and associative. The linearization problem inherent in prior mixin and trait calculi does not arise.
  • The semantics is purely observational: there are no reduction rules, no evaluation order, and no notion of a result that a term reduces to. The observer drives computation by querying paths in a lazily constructed tree.
  • Self-referential records require no fixed-point combinator. The $\lambda$-calculus translation maps each application to an inheritance; the resulting inheritance tree is the call stack. The $\mathtt{this}$ function of Section 3 resolves self-references by walking this tree: a first-order set-theoretic query, not a fixed-point construction.
  • Self-reference resolves to multiple objects, not one. Path-dependent types with inheritance naturally produce this situation; prior systems assume single-object resolution and reject it (Scala and the NixOS module system, Appendix B). Since inheritance is idempotent, all objects contribute equally; this semantics does not introduce a route-priority disambiguation rule.
  • The $\lambda$-calculus (in A-normal form) translates into inheritance-calculus in five rules (Section 4). The translation is fully abstract with respect to Böhm tree equivalence under $\lambda$-definable contexts (Section 4.1), yielding a semantics of the lazy $\lambda$-calculus that requires no domain theory.
  • We observe further properties: extensibility in the sense of the Expression Problem (Wadler, 1998) (Section 5), relational semantics of logic programming from ordinary arithmetic (Section 5), and function color blindness (Nystrom, 2015) (Section Function color blindness: ).
  • Open extension via inheritance changes the operational equivalence of $\lambda$-terms, establishing via Felleisen’s Theorem 3.14 (Felleisen, 1991) that the $\lambda$-calculus cannot macro-express inheritance-calculus (Section 4.2).
Section 5 demonstrates these properties through a case study of Nat arithmetic.
2. Syntax
Let $e$ denote an expression. Let $\ell$ denote a label (property name). Let $k$ denote a non-negative integer (path length).
$e\quad::=\quad$ $\{c_{1},\;\ldots,\;c_{m}\}$ $\text{(}m\geq 0\text{)}$
$\mid\quad$ $c$ $\text{(sugar for }\{c\}\text{)}$
$c\quad::=\quad$ $\ell\mapsto e$ (definition)
$\mid\quad$ $[\ell_{\mathrm{up}}.\mathbf{this}.\mid\uparrow^{\!n}.]\;\ell_{\mathrm{down},1}.\ell_{\mathrm{down},2}\ldots\ell_{\mathrm{down},k}$ $\text{(inheritance, }k\geq 0\text{ with qualifier,}$
$k\geq 1\text{ without)}$
An expression is a set of elements enclosed in braces. Each element is either a definition $\ell\mapsto e$ or an inheritance source. Because expressions are sets, element order is irrelevant and duplicates have no effect; inheritance is inherently commutative, idempotent, and associative.
The $\mapsto$ in $\ell\mapsto e$ defines a property, not a let-binding. There is no variable binding in inheritance-calculus. MIXINv2 also provides scalar types through a foreign-function interface (FFI); inheritance-calculus is what remains after removing the FFI. None of the examples in this paper uses the FFI; scalars did not arise.
Inheritance
An inheritance is written as a reference that navigates the scope hierarchy upward and then projects properties downward. The upward navigation admits two equivalent notations:
Named form:
$\ell_{\mathrm{up}}.\mathbf{this}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$, analogous to Java’s Outer.this.field. Each $\{\ldots\}$ record has a label in its enclosing scope; $\ell_{\mathrm{up}}$ names the target enclosing scope.
Indexed form:
$\uparrow^{\!n}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$, where $n\geq 0$ is a de Bruijn index (de Bruijn, 1972): $n=0$ refers to the reference’s enclosing scope, and each increment of $n$ moves one scope level further out.
Both forms retrieve the target scope, that is, the fully inherited record after all inheritance has been applied, and then project properties $\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$ from it. The named form desugars to the indexed form during parsing (Section 3).
A new scope level is created at each $\{\ldots\}$ record literal. The scope contains all properties of that mixin, including those inherited via inheritance sources. For example, $\{a\mapsto\{\},\;b\mapsto\{\}\}$ is a single mixin whose scope contains both $a$ and $b$; so is $\{r,\;a\mapsto\{\}\}$ where $r$ is a reference to a record containing $b$. Inheritance sources do not create additional scope levels.
  • $\mathrm{Foo}.\mathbf{this}$ (equivalently, $\uparrow^{\!n}$ where $\mathrm{Foo}$ is $n$ steps above the enclosing scope) retrieves the record named $\mathrm{Foo}$.
  • $\mathrm{Foo}.\mathbf{this}.\ell$ (equivalently, $\uparrow^{\!n}.\ell$) projects property $\ell$ from $\mathrm{Foo}$.
  • When $k=0$, the reference inherits the entire enclosing scope, which may produce an infinitely deep tree.
When the qualifier is unambiguous, it may be omitted: the shorthand $\ell_{\mathrm{down},1}.\ell_{\mathrm{down},2}\ldots\ell_{\mathrm{down},k}$ ($k\geq 1$) resolves $\ell_{\mathrm{down},1}$ against the innermost enclosing scope whose $\mathtt{defines}$ contains it, and desugars to the corresponding indexed form. Since this uses $\mathtt{defines}$, the shorthand can only access properties defined in record literals, not those inherited through inheritance; one of the two qualified forms is required for inherited properties or to bypass variable shadowing.
3. Mixin Trees
Given an AST, what are the properties visible at each position in the tree? This section answers the question by defining a collection of mutually recursive functions over paths.
A mixin tree has a single abstraction, the mixin: every mixin can serve as field, method, class, and object at the same time in object-oriented terms.
Unlike the $\lambda$-calculus, which requires $\beta$-reduction to specify how terms compute, inheritance-calculus has no reduction steps: no rewriting of terms, no evaluation order, and no notion of a result that a term reduces to. Instead, the semantics is purely observational: one queries whether a label $\ell$ is a property at a path $p$ in the inheritance tree. The mutually recursive functions defined below, namely $\mathtt{properties}$, $\mathtt{supers}$, $\mathtt{overrides}$, $\mathtt{bases}$, $\mathtt{resolve}$, and $\mathtt{this}$, constitute the entire semantics.
Each function has an analogue in object-oriented languages: $\mathtt{Path}$ identifies a position in the tree, $\mathtt{init}$ and $\mathtt{last}$ navigate paths, $\mathtt{defines}$ extracts the locally defined members from the AST, $\mathtt{inherits}$ extracts the inheritance declarations from the AST, $\mathtt{properties}$ computes all visible members at a path including both locally defined and inherited ones, $\mathtt{supers}$ computes the transitive inheritance closure, $\mathtt{bases}$ collects the direct base classes, and $\mathtt{overrides}$ generalizes method override to the tree setting, collecting the identity-sharing paths that arise when inheritance introduces multiple definitions of the same label at the same scope level. Two functions, the reference resolution function $\mathtt{resolve}$ and the qualified this resolution function $\mathtt{this}$, have object-oriented analogues whose semantics differ significantly; they are discussed in detail below.
All functions are pure and may be cached (memoised) without changing semantics.
Path
A $\mathtt{Path}$ is a sequence of labels $(\ell_{1},\ell_{2},\ldots,\ell_{n})$ that identifies a position in the tree. [^2] Since paths are referentially transparent, an implementation may intern paths so that structural equality reduces to pointer equality. We write $p$ for a path. The root path is the empty sequence $()$. Given a path $p$ and a label $\ell$, $p\mathbin{\triangleright}\ell$ is the sequence $p$ extended with $\ell$. For a non-root path, the parent path $\mathtt{init}(p)$ is $p$ with its last element removed, and the final label $\mathtt{last}(p)$ is the last element of $p$. For convenience, we sometimes write $\ell$ for $\mathtt{last}(p)$ when the path is clear from context. With paths in hand, we can state what the AST provides at each path.
AST
An expression $e$ (Section 2) is parsed into an AST. The AST provides two primitive functions at each path $p$:
  • $\mathtt{defines}(p)$: the set of labels that have subtrees at $p$.
  • $\mathtt{inherits}(p)$: the set of reference pairs $(n,\;\ell_{\mathrm{down},*})$, where $n$ is the de Bruijn index and $\ell_{\mathrm{down},*}$ is the list of downward projections.
During parsing, all three syntactic forms are resolved to de Bruijn index pairs $(n,\;\ell_{\mathrm{down},*})$. The indexed form $\uparrow^{\!n}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$ already carries the de Bruijn index $n$ directly; no resolution is needed. A named reference $\ell_{\mathrm{up}}.\mathbf{this}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$ at path $p$ is resolved by finding the last occurrence of $\ell_{\mathrm{up}}$ among the labels of $p$. Let $p_{\mathrm{target}}$ be the prefix of $p$ up to and including that occurrence. The de Bruijn index is $n=|p|-|p_{\mathrm{target}}|-1$ and the projections are $\ell_{\mathrm{down},*}=(\ell_{\mathrm{down},1},\ldots,\ell_{\mathrm{down},k})$. A lexical reference $\ell_{\mathrm{down},1}.\ell_{\mathrm{down},2}\ldots\ell_{\mathrm{down},k}$ at path $p$ is resolved by finding the nearest prefix $p^{\prime}$ of $p$ such that $\ell_{\mathrm{down},1}\in\mathtt{defines}(p^{\prime})$. The de Bruijn index is $n=|p|-|p^{\prime}|-1$ and the projections are $\ell_{\mathrm{down},*}=(\ell_{\mathrm{down},1},\ldots,\ell_{\mathrm{down},k})$. All three forms produce the same representation; the semantic functions below operate only on de Bruijn index pairs.
Parsing populates $\mathtt{defines}$ and $\mathtt{inherits}$ from each element of a record. A definition $\ell\mapsto e$ contributes $\ell$ to $\mathtt{defines}(p)$. A reference contributes $(n,\;\ell_{\mathrm{down},*})$ to $\mathtt{inherits}(p)$. Since records are sets, the order of elements is irrelevant and duplicate elements have no effect. Both $\mathtt{defines}$ and $\mathtt{inherits}$ are pure data; they are not functions of runtime state. Given these two primitives, we can now answer the question posed at the beginning of this section.
Properties
The properties of a path are its own properties together with those inherited from all supers:
(1) $\mathtt{properties}(p)=\bigl\{\;\ell\;\big|\;(\_,\;p_{\mathrm{override}})\in\mathtt{supers}(p),\;\ell\in\mathtt{defines}(p_{\mathrm{override}})\;\bigr\}$
The remainder of this section defines $\mathtt{supers}$, the transitive inheritance closure, and its dependencies.
Supers
Intuitively, $\mathtt{supers}(p)$ collects every path that $p$ inherits from: the $\mathtt{overrides}$ of $p$ itself, plus the $\mathtt{overrides}$ of each direct base $\mathtt{bases}$ of $p$, and so on transitively. Each result is paired with the inheritance-site context $\mathtt{init}(p_{\mathrm{base}})$ through which it is reached. This provenance is needed by qualified this resolution, i.e. the $\mathtt{this}$ function defined below, to map a definition-site mixin back to the inheritance-site paths that incorporate it:
(2) $\mathtt{supers}(p)=\bigl\{\;(\mathtt{init}(p_{\mathrm{base}}),\;p_{\mathrm{override}})\;\big|\;p_{\mathrm{base}}\in\mathtt{bases}^{*}(p),\;p_{\mathrm{override}}\in\mathtt{overrides}(p_{\mathrm{base}})\;\bigr\}$
Here $\mathtt{bases}^{*}$ denotes the reflexive-transitive closure of $\mathtt{bases}$. The $\mathtt{supers}$ formula depends on two functions: $\mathtt{overrides}$ and the one-hop reference targets $\mathtt{bases}$. We define $\mathtt{overrides}$ first.
Overrides
Because inheritance-calculus has no separate notion of “method,” a mixin simultaneously plays the role of a class and the role of a method in object-oriented languages. When a mixin is viewed as a method, a child scope should be able to find it through the parent’s label even if the child does not redefine that label—this is method override. Concretely, inheritance can introduce multiple definitions of the same label at the same scope level; $\mathtt{overrides}(p)$ collects all such paths that share the same identity as $p$, so that their definitions merge rather than shadow each other, enabling deep merge. The overrides of $p$ include $p$ itself and $p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)$ for every branch $p_{\mathrm{branch}}$ of $\mathtt{init}(p)$ that also defines $\mathtt{last}(p)$:
(3) $\mathtt{overrides}(p)=\begin{cases}\{p\}&\text{if }p=()\\[6.0pt] \{p\}\;\cup\;\left\{\;p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)\;\left|\;\begin{aligned} &(\_,\;p_{\mathrm{branch}})\in\mathtt{supers}(\mathtt{init}(p)),\\ &\text{s.t.}\;\mathtt{last}(p)\in\mathtt{defines}(p_{\mathrm{branch}})\end{aligned}\right.\right\}&\text{if }p\neq()\end{cases}$
It remains to define $\mathtt{bases}$, the other dependency of $\mathtt{supers}$.
Bases
Intuitively, $\mathtt{bases}(p)$ are the paths that $p$ directly inherits from via references, analogous to the direct base classes in object-oriented languages. Concretely, $\mathtt{bases}$ resolves every reference in $p$’s $\mathtt{overrides}$ one step:
[…]
The $\mathtt{bases}$ formula calls the reference resolution function $\mathtt{resolve}$, which we define next.
Reference resolution
Intuitively, $\mathtt{resolve}$ turns a syntactic reference into the set of paths it points to in the fully inherited tree. It takes an inheritance-site path $p_{\mathrm{site}}$, a definition-site path $p_{\mathrm{def}}$, a de Bruijn index $n$, and downward projections $\ell_{\mathrm{down},*}$. Resolution proceeds in two phases: $\mathtt{this}$ performs $n$ upward steps starting from the enclosing scope $\mathtt{init}(p_{\mathrm{def}})$, mapping the definition-site path to inheritance-site paths; then the downward projections $\ell_{\mathrm{down},*}$ are appended:
(5) $\mathtt{resolve}(p_{\mathrm{site}},\;p_{\mathrm{def}},\;n,\;\ell_{\mathrm{down},*})=\\ \bigl\{\;p_{\mathrm{current}}\mathbin{\triangleright}\ell_{\mathrm{down},1}\mathbin{\triangleright}\cdots\mathbin{\triangleright}\ell_{\mathrm{down},k}\;\big|\;p_{\mathrm{current}}\in\mathtt{this}(\{p_{\mathrm{site}}\},\;\mathtt{init}(p_{\mathrm{def}}),\;n)\;\bigr\}$
When multiple routes exist due to multi-path inheritance, they may yield different target paths; all are collected. We now define $\mathtt{this}$.
Qualified this resolution
$\mathtt{this}$ answers the question: “in the fully inherited tree, where does the definition-site scope $p_{\mathrm{def}}$ actually live?” Each step finds, among the supers of every path in the frontier $S$, those whose override component matches $p_{\mathrm{def}}$, and collects the corresponding inheritance-site paths as the new frontier. After $n$ steps the frontier contains the answer:
(6) $\mathtt{this}(S,\;p_{\mathrm{def}},\;n)=\begin{cases}S&\text{if }n=0\\[6.0pt] \mathtt{this}\!\left(\left\{\;p_{\mathrm{site}}\;\left|\;\begin{aligned} &p_{\mathrm{current}}\in S,\\ &(p_{\mathrm{site}},\;p_{\mathrm{override}})\in\mathtt{supers}(p_{\mathrm{current}}),\\ &\text{s.t.}\;p_{\mathrm{override}}=p_{\mathrm{def}}\end{aligned}\right.\right\},\;\mathtt{init}(p_{\mathrm{def}}),\;n-1\right)&\text{if }n>0\end{cases}$
At each step $\mathtt{init}$ shortens $p_{\mathrm{def}}$ by one label and $n$ decreases by one. Since $n$ is a non-negative integer, the recursion terminates.
This completes the chain of definitions needed to compute $\mathtt{properties}(p)$.
Observational semantics without reduction
The definitions above constitute the complete semantics of inheritance-calculus. There are no reduction rules, no evaluation contexts, and no small-step or big-step judgments. In the $\lambda$-calculus, a term such as $(\lambda x.\,x)\;y$ must be reduced to $y$ before its meaning is apparent; the operational semantics specifies how and in what order reductions occur. In inheritance-calculus, an inheritance tree such as $\{a\mapsto\{\},\;b\mapsto\{\}\}$ does not reduce to anything: it is the value. The functions $\mathtt{properties}$, $\mathtt{supers}$, etc. determine what is observable at each path, and that is all. Computation arises not from rewriting terms but from querying an inductively deeper path in a lazily constructed tree: the observer drives the computation by choosing which path to inspect, and the mutually recursive definitions unfold on demand. Formally, the mutually recursive equations (1)–(6) define set membership inductively: $\ell\in\mathtt{properties}(p)$ holds if and only if it can be established by a finite chain of applications of these equations. The corresponding operator on the product lattice of path sets is monotone, and its least fixed point exists by the Knaster–Tarski theorem (Tarski, 1955). The detailed well-definedness argument, which identifies the lattice, verifies positivity of each equation, and handles the transitive closure $\mathtt{bases}^{*}$, is given in Appendix C. When no finite derivation exists for a given query, the membership is not in this least fixed point and the semantics is undefined at that point, analogous to divergence in the $\lambda$-calculus, and the expected consequence of Turing completeness. Identity is determined by observable behavior rather than by internal structure.
4. Embedding the $\lambda$-Calculus
The following translation maps the $\lambda$-calculus in A-normal form (ANF) (Flanagan et al., 1993) to inheritance-calculus. In ANF, every intermediate result is bound to a name and each expression contains at most one application:
[…]
Every $\lambda$-term can be mechanically converted to ANF by naming intermediate results, and every ANF term is a $\lambda$-term, so ANF is neither a restriction nor an extension of the $\lambda$-calculus. The ANF structure matches inheritance-calculus naturally: each $\mathbf{let}$-binding becomes a named record property, and each application is wrapped and projected via $\mathrm{result}.$
Let $\mathcal{T}$ denote the translation function. Each $\lambda$-abstraction introduces one scope level in the translated mixin tree.
$x$ ($\lambda$-bound, index $n$)   $\longrightarrow$ $\uparrow^{\!n}.\mathrm{argument}$
$x$ (let-bound)   $\longrightarrow$ $x.\mathrm{result}$
$\lambda x.\,M$   $\longrightarrow$ $\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
$\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M$   $\longrightarrow$ $\{x\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
$V_{1}\;V_{2}$ (tail call)   $\longrightarrow$ $\{\_\mathrm{tail}\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\_\mathrm{tail}.\mathrm{result}\}$
[…]
An abstraction $\lambda x.\,M$ translates to a record with own properties $\{\mathrm{argument},\mathrm{result}\}$; we call this the abstraction shape. In the $\mathbf{let}$-binding rule, the name $x$ binds to the inherited record $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$, and $\mathcal{T}(M)$ may reference $x.\mathrm{result}$ to obtain the value of the application. In the tail-call rule, the fresh label $\mathrm{tail}$ serves the same purpose, and $\mathrm{result}$ projects the answer. In both cases the inheritance that performs the application is encapsulated behind a named property, so its internal structure (the $\mathrm{argument}$ and $\mathrm{result}$ labels of the callee) does not leak to the enclosing scope.
The five rules above are a complete translation from the $\lambda$-calculus (via ANF) to inheritance-calculus: every ANF term has an image, and the construction is compositional. Since the $\lambda$-calculus is Turing complete, so is inheritance-calculus.
An elementary semantics of the $\lambda$-calculus
Composing the translation $\mathcal{T}$ with the semantic functions of Section 3 induces a semantics on the $\lambda$-calculus itself. Given a $\lambda$-term $M$, one translates it to $\mathcal{T}(M)$ and then determines convergence behavior by querying $\ell\in\mathtt{properties}(p)$ via equations (1)–(6), following the $\mathrm{result}$ chain until the abstraction shape appears or the fixed point yields the empty set.
This induced semantics is more elementary than traditional approaches in the mathematical machinery it requires:
No function spaces:
Scott’s denotational semantics (Scott, 1976) constructs a reflexive domain $D$ satisfying $D\cong[D\to D]$ via an inverse limit in a category of continuous lattices. Here, the semantic domain is a product of powerset lattices, with no function spaces, no continuity conditions, and no inverse limits. The only fixed-point construction is Knaster–Tarski on powerset lattices (Appendix C), which requires only set-theoretic monotonicity, which is strictly weaker than Scott continuity.
No substitution:
Operational semantics defines $\beta$-reduction via capture-avoiding substitution, a mechanism whose subtleties motivated de Bruijn indices (de Bruijn, 1972) and explicit substitution calculi. Here, variable references are resolved by the $\mathtt{this}$ function walking the inheritance tree; no substitution occurs at any point.
No reduction rules:
There are no $\beta$-reduction steps, no evaluation contexts, and no choice between call-by-name and call-by-value (Plotkin, 1975). The semantics is observational: one queries path membership, and the mutually recursive equations unfold on demand.
The key step is that function application is encoded as record inheritance, and the resulting record is observed through path queries rather than applied as a function; no function spaces are required. Section 4.1 shows that the convergence behavior of the translated mixin tree corresponds exactly to the Böhm tree (Barendregt, 1984) of the original term: the translation is adequate, meaning a term converges iff it has a head normal form, and fully abstract with respect to $\lambda$-definable contexts, meaning two $\lambda$-terms are equivalent under the translation iff they have the same Böhm tree.
Combining these two facts, elementary machinery and full abstraction, yields a noteworthy consequence: the lazy $\lambda$-calculus admits a fully abstract semantics (w.r.t. $\lambda$-definable contexts) that lives entirely in naive set theory. The only mathematical prerequisites are powerset lattices, set-theoretic monotonicity, and the Knaster–Tarski fixed-point theorem. No continuous lattices, no inverse limits, no Scott topology, no game-theoretic machinery. Scott’s $D_{\infty}$ model (Scott, 1976) uses heavier machinery yet is only adequate, not fully abstract (Plotkin, 1975); Abramsky and McCusker’s game semantics (Abramsky and McCusker, 1995) achieves full abstraction but at the cost of even greater mathematical complexity. The mixin tree semantics achieves full abstraction with strictly simpler foundations than either.
4.1. Böhm Tree Correspondence
The translation $\mathcal{T}$ induces a semantics on the $\lambda$-calculus via the mixin tree equations (1)–(6). We show that the induced semantics corresponds precisely to the Böhm tree (Barendregt, 1984) of the original $\lambda$-term.
A mixin tree $T$ converges, written $T{\Downarrow}$, if following finitely many $\mathrm{result}$ projections from the root reaches a node with both $\mathrm{argument}$ and $\mathrm{result}$ among its properties, that is, the abstraction shape produced by $\mathcal{T}(\lambda x.\,M)$. We write $T{\Uparrow}$ if no such depth exists.
Definition 4.1 (Inheritance-convergence).
A mixin tree $T$ converges, written $T{\Downarrow}$, if there exists $n\geq 0$ such that:
$\{\mathrm{argument},\,\mathrm{result}\}\;\subseteq\;\mathtt{properties}\!\bigl(\,\underbrace{\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}\cdots\mathbin{\triangleright}\mathrm{result}}_{n}\,\bigr)$
in the least fixed point of Section 3.
Theorem 4.2 (Adequacy).
A closed ANF $\lambda$-term $M$ has a head normal form if and only if $\mathcal{T}(M){\Downarrow}$.
Theorem 4.3 (Full abstraction w.r.t. $\lambda$-definable contexts).
For closed $\lambda$-terms $M$ and $N$, define $\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)$ iff for every closing $\lambda$-calculus context $C[\cdot]$, $\mathcal{T}(C[M]){\Downarrow}\Leftrightarrow\mathcal{T}(C[N]){\Downarrow}$. Then:
$\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)\quad\Longleftrightarrow\quad\mathrm{BT}(M)=\mathrm{BT}(N)$
Here $\approx_{\lambda}$ quantifies over $\lambda$-calculus contexts only; inheritance-calculus contexts are strictly more discriminating (Section 4.2). Together, Adequacy and Full Abstraction establish that the translation $\mathcal{T}$ is a fully faithful embedding of the lazy $\lambda$-calculus into inheritance-calculus with respect to Böhm tree equivalence (Barendregt, 1984), using only powerset lattices and the Knaster–Tarski fixed-point theorem, without the topological infrastructure of domain theory. This is stronger than Scott’s denotational semantics (Scott, 1976; Wadsworth, 1976), which is adequate but not fully abstract, as it identifies terms that observational equivalence distinguishes (Plotkin, 1975). The full proof, including the Single-Path Lemma, Substitution Lemma, and Convergence Preservation, is given in Appendix D.
The relationship between $\approx_{\lambda}$ and full inheritance-contextual equivalence is clarified by the expressive asymmetry of the next section: inheritance-calculus contexts can separate $\lambda$-terms that are Böhm-tree equivalent, so ${\approx_{\lambda}}\subsetneq{(\approx_{\mathrm{inh}}|_{\mathrm{im}(\mathcal{T})})}$. This gap is precisely the source of inheritance-calculus’s additional expressiveness (Theorem A.3).
4.2. Expressive Asymmetry
Two $\lambda$-terms that are Böhm-tree equivalent—and therefore indistinguishable in the $\lambda$-calculus—can be separated in inheritance-calculus by inheriting new observable projections onto an existing definition via open extension. By Felleisen’s expressiveness criterion (Felleisen, 1991), inheritance-calculus is strictly more expressive than the $\lambda$-calculus: the forward embedding is macro-expressible (Theorem A.2), but the reverse is not (Theorem A.3). The formal definitions, constructions, and proofs are given in Appendix A; the separating construction is itself a miniature instance of the Expression Problem, adding a new observable projection to an existing definition without modifying it.
The underlying mechanism is open extension: mixin definitions compose via open recursion, so existing definitions gain new observable properties without modification. In the $\lambda$-calculus, a function’s result is closed—there is no way to add new observable projections to an already-computed value without modifying the function’s definition.
5. Case Study: Nat Arithmetic
We implemented boolean logic and natural number arithmetic, both unary and binary, as MIXINv2 files (Yang, 2026). This section traces the unary Nat case in detail. The implementation follows a declarative object-oriented style: each concern is one file; UpperCamelCase names are constructors and operations; lowerCamelCase names are fields and parameters; and since there are no methods, all names are nouns or adjectives: $\mathrm{Plus}$ is a structure, not an action. Operations expose their result through a projection field (e.g., $\mathrm{sum}$).
NatData
Following the Church-encoding convention (Yang, 2026), we define a factory containing the abstract product type and its two constructors. A $\mathrm{Zero}$ value inherits $\mathrm{Product}$; a $\mathrm{Successor}$ value inherits $\mathrm{Product}$ and exposes a $\mathrm{predecessor}$ field. The file name $\mathrm{NatData}$ is itself the top-level scope:
$\mathrm{NatData}\mapsto\{\;$ $\mathrm{NatFactory}\mapsto\{\mathrm{Product}\mapsto\{\},$
$\quad\mathrm{Zero}\mapsto\{\mathrm{Product}\},$
$\quad\mathrm{Successor}\mapsto\{\mathrm{Product},\;\mathrm{predecessor}\mapsto\{\mathrm{Product}\}\}\},$
$\mathrm{Nat}\mapsto\{\mathrm{NatFactory},\;\mathrm{NatFactory}.\mathrm{Product}\}\}$
NatPlus
Addition is a separate MIXINv2 file that inherits $\mathrm{NatData}$ and extends each constructor with a $\mathrm{Plus}$ operation exposing a $\mathrm{sum}$ field. The base case is $0+m=m$; the recursive case reduces $\mathrm{S}(n_{0})+m$ to $n_{0}+\mathrm{S}(m)$ by delegating to $n_{0}$’s $\mathrm{Plus}$ with an incremented addend:
$\mathrm{NatPlus}\mapsto\{\;$ $\mathrm{NatData},$
$\mathrm{NatFactory}\mapsto\{$
$\quad\mathrm{Product}\mapsto\{\mathrm{Plus}\mapsto\{\mathrm{sum}\mapsto\{\mathrm{Product}\}\}\},$
$\quad\mathrm{Zero}\mapsto\{\mathrm{Plus}\mapsto\{\mathrm{addend}\mapsto\{\mathrm{Product}\},\;\mathrm{sum}\mapsto\{\mathrm{addend}\}\}\},$
$\quad\mathrm{Successor}\mapsto\{\mathrm{Plus}\mapsto\{\mathrm{addend}\mapsto\{\mathrm{Product}\},$
$\qquad\mathrm{\_increasedAddend}\mapsto\{\mathrm{Successor},\;\mathrm{predecessor}\mapsto\{\mathrm{addend}\}\},$
$\qquad\mathrm{\_recursiveAddition}\mapsto\{\mathrm{Successor}.\mathbf{this}.\mathrm{predecessor}.\mathrm{Plus},$
$\qquad\quad\mathrm{addend}\mapsto\{\mathrm{\_increasedAddend}\}\},$
$\qquad\mathrm{sum}\mapsto\{\mathrm{\_recursiveAddition}.\mathrm{sum}\}\}\}\}\}$
NatEquality
To test that $2+3=5$, we needed equality on Nat values. Equality requires case analysis on the constructor; we use the Visitor pattern. $\mathrm{NatVisitor}$ adds a $\mathrm{Visitor}$ operation to each constructor, with a $\mathrm{Visit}$ field that delegates to the appropriate $\mathrm{VisitZero}$ or $\mathrm{VisitSuccessor}$ branch:
$\mathrm{NatVisitor}\mapsto\{\;$ $\mathrm{NatData},$
$\mathrm{NatFactory}\mapsto\{$
$\quad\mathrm{Product}\mapsto\{\mathrm{Visitor}\mapsto\{\mathrm{Visit}\mapsto\{\}\}\},$
$\quad\mathrm{Zero}\mapsto\{\mathrm{Visitor}\mapsto\{\mathrm{VisitZero}\mapsto\{\},\;\mathrm{Visit}\mapsto\{\mathrm{VisitZero}\}\}\},$
$\quad\mathrm{Successor}\mapsto\{\mathrm{Visitor}\mapsto\{\mathrm{VisitSuccessor}\mapsto\{\},\;\mathrm{Visit}\mapsto\{\mathrm{VisitSuccessor}\}\}\}\}\}$
$\mathrm{BooleanData}$ defines the output sort:
$\mathrm{BooleanData}\mapsto\{\;$ $\mathrm{BooleanFactory}\mapsto\{\mathrm{Product}\mapsto\{\},\;\mathrm{True}\mapsto\{\mathrm{Product}\},\;\mathrm{False}\mapsto\{\mathrm{Product}\}\},$
$\mathrm{Boolean}\mapsto\{\mathrm{BooleanFactory},\;\mathrm{BooleanFactory}.\mathrm{Product}\}\}$
$\mathrm{NatEquality}$ inherits $\mathrm{NatVisitor}$ and $\mathrm{BooleanData}$, and adds an $\mathrm{Equal}$ operation to each constructor. $\mathrm{Zero}$ is equal only to another $\mathrm{Zero}$; $\mathrm{Successor}$ is equal only to another $\mathrm{Successor}$ with an equal predecessor. Case analysis on $\mathrm{other}$ is done via its $\mathrm{Visitor}$:
$\mathrm{NatEquality}\mapsto\{\;$ $\mathrm{NatVisitor},\;\mathrm{BooleanData},$
$\mathrm{NatFactory}\mapsto\{$
$\qquad\mathrm{Product}\mapsto\{\mathrm{Equal}\mapsto\{\mathrm{other}\mapsto\{\mathrm{Product}\},$
$\hskip 18.49988pt\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{Boolean}\}\}\},$
$\qquad\mathrm{Zero}\mapsto\{\mathrm{Equal}\mapsto\{\mathrm{other}\mapsto\{\mathrm{Product}\},$
$\hskip 18.49988pt\mathrm{OtherVisitor}\mapsto\{\mathrm{other}.\mathrm{Visitor},$
$\hskip 18.49988pt\;\mathrm{VisitZero}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{True}\}\},$
$\hskip 18.49988pt\;\mathrm{VisitSuccessor}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{False}\}\},$
$\hskip 18.49988pt\;\mathrm{Visit}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{Boolean}\}\}\},$
$\hskip 18.49988pt\mathrm{equal}\mapsto\{\mathrm{OtherVisitor}.\mathrm{Visit}.\mathrm{equal}\}\}\},$
$\qquad\mathrm{Successor}\mapsto\{\mathrm{Equal}\mapsto\{\mathrm{other}\mapsto\{\mathrm{Product},\;\mathrm{predecessor}\mapsto\{\mathrm{Product}\}\},$
$\hskip 18.49988pt\mathrm{RecursiveEquality}\mapsto\{\mathrm{Successor}.\mathbf{this}.\mathrm{predecessor}.\mathrm{Equal},$
$\hskip 18.49988pt\;\mathrm{other}\mapsto\{\mathrm{Equal}.\mathbf{this}.\mathrm{other}.\mathrm{predecessor}\}\},$
$\hskip 18.49988pt\mathrm{OtherVisitor}\mapsto\{\mathrm{other}.\mathrm{Visitor},$
$\hskip 18.49988pt\;\mathrm{VisitZero}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{False}\}\},$
$\hskip 18.49988pt\;\mathrm{VisitSuccessor}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{RecursiveEquality}.\mathrm{equal}\}\},$
$\hskip 18.49988pt\;\mathrm{Visit}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{Boolean}\}\}\},$
$\hskip 18.49988pt\mathrm{equal}\mapsto\{\mathrm{OtherVisitor}.\mathrm{Visit}.\mathrm{equal}\}\}\}\}\}$
Composing $\mathrm{NatPlus}$ with $\mathrm{NatEquality}$ by inheriting both, with no modifications to either file, gives the resulting Nat values both $\mathrm{Plus}$ and $\mathrm{Equal}$ automatically. The concrete numerals live in a shared file that inherits only $\mathrm{NatData}$:
$\mathrm{NatConstants}\mapsto\{\;$ $\mathrm{NatData},$
$\mathrm{One}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Successor},$
$\quad\quad\mathrm{predecessor}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Zero}\}\},$
$\mathrm{Two}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Successor},$
$\quad\quad\mathrm{predecessor}\mapsto\{\mathrm{One}\}\},$
$\mathrm{Three}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Successor},$
$\quad\quad\mathrm{predecessor}\mapsto\{\mathrm{Two}\}\},$
$\mathrm{Four}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Successor},$
$\quad\quad\mathrm{predecessor}\mapsto\{\mathrm{Three}\}\},$
$\mathrm{Five}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Successor},$
$\quad\quad\mathrm{predecessor}\mapsto\{\mathrm{Four}\}\}\}$
The arithmetic test inherits $\mathrm{NatConstants}$ together with the operations it needs:
$\mathrm{Test}\mapsto\{\;$ $\mathrm{NatConstants},\;\mathrm{NatPlus},\;\mathrm{NatEquality},$
$\mathrm{Addition}\mapsto\{\mathrm{Test}.\mathbf{this}.\mathrm{Two}.\mathrm{Plus},\;\mathrm{addend}\mapsto\{\mathrm{Test}.\mathbf{this}.\mathrm{Three}\}\},$
$\mathrm{Test2plus3}\mapsto\{\mathrm{Test}.\mathbf{this}.\mathrm{Five}.\mathrm{Equal},\;\mathrm{other}\mapsto\{\mathrm{Addition}.\mathrm{sum}\}\}\}$
Querying $\mathrm{Test}.\mathrm{Test2plus3}.\mathrm{equal}$ via the host-language API returns $\mathrm{True}$. No existing file was modified. $\mathrm{NatEquality}$, a new operation, and $\mathrm{Boolean}$, a new data type, each required only a new file: extensibility in both dimensions of the Expression Problem (Wadler, 1998).
Tree-level inheritance and return types
Each operation is an independent mixin that extends a factory’s subtree. Inheritance recursively merges subtrees that share a label, so every entity in the factory acquires all operations from all inherited mixins. Crucially, this applies to return types: $\mathrm{NatPlus}$ returns values from $\mathrm{NatFactory}$, and $\mathrm{NatEquality}$ independently adds $\mathrm{Equal}$ to the same factory. Inheriting both causes the values returned by $\mathrm{Plus}$ to carry $\mathrm{Equal}$ automatically, without modifying either file. In object algebras (d. S. Oliveira and Cook, 2012) and finally tagless interpreters (Carette et al., 2009), independently defined operations do not automatically enrich each other’s return types; additional boilerplate or type-class machinery is required.
Multi-path qualified this (equation 6) is essential. In the test file above, $\mathrm{NatPlus}.\mathrm{Successor}$ and $\mathrm{NatEquality}.\mathrm{Successor}$ both inherit $\mathrm{NatData}.\mathrm{Successor}$. When the two operations are composed, the resulting $\mathrm{Successor}$ inherits the $\mathrm{NatData}$ schema through two independent routes. The reference $\mathrm{Successor}.\mathbf{this}.\mathrm{predecessor}$ inside each operation resolves through both routes, and since inheritance is idempotent, all routes contribute the same properties. In Scala, the same pattern, where two independently defined inner classes extend the same outer class’s trait, triggers a “conflicting base types” rejection (Appendix B), because DOT’s self variable is single-valued and cannot resolve through multiple inheritance routes simultaneously. The Expression Problem can still be solved in Scala using object algebras or similar patterns, but composing independently defined operations inherently produces multi-path inheritance to shared schemas, which DOT’s single-valued self variable cannot accommodate.
Church-encoded Nats are tries
A Church-encoded Nat is a record whose structure mirrors the constructor path used to build it: $\mathrm{Zero}$ is a flat record; $\mathrm{S}(\mathrm{Zero})$ is a record with a $\mathrm{predecessor}$ field pointing to a $\mathrm{Zero}$ record; and so on. This is exactly the shape of a trie. A scope that simultaneously inherits two Nat values is a trie union. The following test applies $\mathrm{Plus}$ to a trie union of $\{1,2\}$ with addend a trie union of $\{3,4\}$, then compares the result to itself with $\mathrm{Equal}$:
$\mathrm{CartesianTest}\mapsto\{\;$ $\mathrm{NatConstants},\;\mathrm{NatPlus},\;\mathrm{NatEquality},$
$\mathrm{OneOrTwo}\mapsto\{\mathrm{CartesianTest}.\mathbf{this}.\mathrm{One},\;\mathrm{CartesianTest}.\mathbf{this}.\mathrm{Two}\},$
$\mathrm{ThreeOrFour}\mapsto\{\mathrm{CartesianTest}.\mathbf{this}.\mathrm{Three},\;\mathrm{CartesianTest}.\mathbf{this}.\mathrm{Four}\},$
$\mathrm{Result}\mapsto\{\mathrm{OneOrTwo}.\mathrm{Plus},\;\mathrm{addend}\mapsto\{\mathrm{ThreeOrFour}\}\},$
$\mathrm{Check}\mapsto\{\mathrm{Result}.\mathrm{sum}.\mathrm{Equal},\;\mathrm{other}\mapsto\{\mathrm{Result}.\mathrm{sum}\}\}\}$
$\mathrm{Result}.\mathrm{sum}$ evaluates to the trie $\{4,5,6\}$, the set of all pairwise sums. $\mathrm{Equal}$ then applies to every pair drawn from $\{4,5,6\}\times\{4,5,6\}$; querying $\mathrm{CartesianTest}.\mathrm{Check}.\mathrm{equal}$ via the host-language API returns $\{\mathrm{True},\mathrm{False}\}$. The Cartesian product arises directly from the semantic equations. $\mathrm{OneOrTwo}$ inherits both $\mathrm{One}$ and $\mathrm{Two}$, so $\mathtt{bases}(\mathrm{OneOrTwo})$ contains two paths. When $\mathrm{Result}$ inherits $\mathrm{OneOrTwo}.\mathrm{Plus}$, the $\mathtt{bases}$ of $\mathrm{Result}$ include the $\mathrm{Plus}$ of both $\mathrm{One}$ and $\mathrm{Two}$. Each $\mathrm{Plus}$ resolves its $\mathrm{addend}$ via $\mathtt{this}$, reaching $\mathrm{ThreeOrFour}$, which is itself a trie union of $\mathrm{Three}$ and $\mathrm{Four}$: the $\mathtt{supers}$ of the $\mathrm{addend}$ path contain both values. Since $\mathrm{Successor}.\mathrm{Plus}$ recursively delegates to $\mathrm{predecessor}.\mathrm{Plus}$ with an incremented addend, and $\mathrm{predecessor}$ also resolves through the trie union, the recursion branches at every trie node. The result $\mathrm{sum}$ therefore collects all paths reachable through any combination of the two operands, which is the Cartesian product.
The same $\mathrm{Plus}$ and $\mathrm{Equal}$ files that operate on single values thus produce, without modification, the relational semantics of logic programming: $\mathtt{overrides}$ and $\mathtt{supers}$ collect all inheritance paths, and every operation distributes over them.
6. Discussion
6.1. Emergent Phenomena
Several features of inheritance-calculus are primitive yet have no direct counterpart in the $\lambda$-calculus. Encoding any one of them requires pervasive, non-local changes to program structure.
Self-referential records with lazy evaluation:
A mixin’s properties can reference other properties and enclosing scopes through qualified this without explicit binding. In the $\lambda$-calculus, encoding such self-referential structures requires explicit allocation of mutable or lazy references (Boudol, 2004) (e.g., Haskell-style thunks or ML-style ref cells), obscuring the declarative intent.
Open recursion with symmetric inheritance:
Inheritance merges definitions from independent sources, all sharing the same inherited mixin tree. Encoding this in the $\lambda$-calculus requires maintaining an extensible dictionary of methods that can be merged from multiple directions. Cook’s denotational semantics of inheritance (Cook, 1989) devoted an entire dissertation to formalizing this pattern in a functional setting.
Commutative, idempotent merge:
The inheritance operation is symmetric: a record’s elements form a set, so reordering them produces the same observable result. In the $\lambda$-calculus, function composition $(f\circ g)$ is neither commutative nor idempotent, and simulating these properties requires additional machinery (Harper and Pierce, 1991; Cardelli, 1994).
Relational semantics:
Every inheritance-calculus operation is defined by recursion on mixin trees, not on individual values. When a mixin tree is viewed as a trie, the correspondence with relational algebra becomes precise: a path from the root to a leaf encodes a single value (e.g., $\mathrm{Successor}\to\mathrm{predecessor}\to\mathrm{Zero}$ encodes the natural number $1$); a trie union encodes a set of such values, i.e., a unary relation. An operation that takes two trie-union operands traverses every combination of their paths through $\mathtt{bases}$ (equation 4), which is a cross-join; the transitive closure $\mathtt{bases}^{*}$ in $\mathtt{supers}$ (equation 2) computes the union of all reachable paths; and $\mathtt{properties}$ (equation 1) projects the observable labels. The carrier is a trie rather than a flat relation, and observation is coinductive—the observer drives computation by querying progressively deeper paths—rather than Datalog’s bottom-up materialization, but the relational structure of the computation is the same. In the Case Study (Section 5), the $\mathrm{Plus}$ and $\mathrm{Equal}$ operations are defined on single Nat values. When applied to set-valued mixins $\{1,2\}$ and $\{3,4\}$, the same definitions, without any modification, produce the set of all pairwise sums $\{4,5,6\}$. This is relational semantics in action: each operation distributes over every combination of its inputs. The behavior is native: the semantic equations of Section 3 resolve every property to a set of paths, so single-valued and set-valued computation are the same mechanism. No code change is needed to move from one to the other.
Function color blindness:
In a web application (Yang, 2026) initially targeting Python’s synchronous sqlite3 and http.server modules, the business logic MIXINv2 file contained no async-specific code. When the FFI layer was replaced with one built on aiosqlite and starlette, the business logic MIXINv2 file did not change. This is an instance of function color blindness (Nystrom, 2015): async and sync functions are distinct types in the host language, yet the MIXINv2 file is agnostic to the distinction because its empty property slots are filled at assembly time, with coroutine handlers under asyncio, or with direct implementations under synchronous execution. More generally, empty slots can be filled with effect handlers (Plotkin and Power, 2003; Pretnar, 2015) under a CPS interpretation, or with direct implementations under a direct-style interpretation, without explicit monad machinery.
Immunity to nonextensibility:
Programming languages typically provide a mix of extensible and nonextensible constructs. In such languages, extensibility is opt-in: frameworks must encode nonextensible constructs into extensible constructs using sophisticated mechanisms (Liang et al., 1995; Lämmel and Jones, 2003; Löh and Hinze, 2006; Swierstra, 2008; Carette et al., 2009; d. S. Oliveira and Cook, 2012; Kiselyov et al., 2013; Wu et al., 2014; Kiselyov and Ishii, 2015; Wang and d. S. Oliveira, 2016; Leijen, 2017; Poulsen and van der Rest, 2023), and code that uses nonextensible constructs cannot be extended unless rewritten in the framework. In contrast, every construct of inheritance-calculus (Section 2) is extensible via deep merge (equation 3). There is no nonextensible construct, so it is impossible to write a nonextensible program. In the Case Study (Section 5), new operations and data types are added without modifying existing files. Tree-level inheritance (Section 5) and multi-path qualified this (equation 6; Appendix B) account for this. The Expressive Asymmetry (Section 4.2; Theorem A.3) provides the formal counterpart: the mixin trees in its separating construction are mechanical translations of $\lambda$-terms via $\mathcal{T}$ (Section 4), yet a new observable projection is added to them without modification. This is the extensibility that the Expression Problem (Wadler, 1998) demands. [^3] As an untyped calculus, inheritance-calculus cannot violate the extensibility requirement of the Expression Problem but does not encounter its type-safety requirement. A typed language built on inheritance-calculus (Yang, 2026) inherits the immunity to nonextensibility and need only ensure that its type system does not introduce nonextensible constructs.
Each of these features is primitive in inheritance-calculus. In the $\lambda$-calculus, encoding them requires pervasive changes to program structure: every scope must thread a mutable store for lazy allocation, every method invocation must consult a dynamically constructed dictionary for open recursion, and every composition must be wrapped in order-insensitive infrastructure for commutativity. These are global, cascading transformations, not local rewrites of individual terms. In common sense, inheritance-calculus is strictly more expressive than the $\lambda$-calculus.
6.2. Semantic Alternatives
The preceding item established that inheritance-calculus is immune to nonextensibility. We now ask whether an alternative design could share this property by surveying every candidate known to us.
Degrees of freedom
Every language whose programs are tree-structured ASTs with references can be characterized by two finite sets: a set $\mathcal{N}$ of node types, each determining how a subtree organizes its children, and a set $\mathcal{R}$ of reference types, each determining how one position in the tree refers to another. The syntax is fixed by the choice of $\mathcal{N}$ and $\mathcal{R}$; the semantics is determined by how each node type structures composition and how each reference type resolves its target.
Language Node types $\mathcal{N}$ Reference types $\mathcal{R}$
Java (OOP) class, interface, method, block, conditional, loop, throw, lambda, … variable, field access, method call, extends/implements, import, …
Haskell (FP) $\lambda$, application, let, case, data, type class, instance, … variable, constructor, type class dispatch, import, …
Scala (multi-paradigm) class, trait, object, method, block, conditional, match, while, throw, … variable, member access, extends, with, implicits, import, …
$\lambda$-calculus abstraction, application variable
Inheritance-calculus record inheritance
Immunity to nonextensibility
A path $p$ in a program $P$ is extensible if there exists a program $Q$ such that composing $P$ with $Q$ through a reference makes new properties observable at $p$ without modifying $P$. A language is immune to nonextensibility if every path in every well-formed program is extensible.
We now classify the design space to explain why we chose the semantics of Section 3. Two axes remain free: the composition mechanism—how the single reference type incorporates definitions; and the reference resolution—how the single reference type finds its target.
The composition mechanism
Every composition mechanism determines how the definitions of one abstraction are incorporated into another. The following table classifies the known candidates by whether they satisfy commutativity (C), idempotence (I), and associativity (A), and by their extensibility model.
Candidate C I A Extensibility Representative
Deep merge $\checkmark$ $\checkmark$ $\checkmark$ universal this paper
Shallow merge $\times$ $\checkmark$ $\checkmark$ top-level only JS spread
Conflict rejection $\checkmark$ $\checkmark$ $\checkmark$ none on overlap Harper–Pierce (Harper and Pierce, 1991)
Priority merge $\times$ $\times$ $\checkmark$ universal Jsonnet +, Nix //
$\beta$-reduction $\times$ $\times$ $\times$ opt-in Java, Haskell, Scala, …
Conditional breaks monotonicity
Deep merge:
Merging same-label definitions recurses into their subtrees ($\mathtt{overrides}$, equation 3). Every label is an extension point, whether or not the original author anticipated extension.
Shallow merge:
Nested labels cannot be independently extended.
Conflict rejection:
Same-label definitions are rejected as errors. All three identities hold—but vacuously: the identities are satisfied because the compositions that would test them are refused. Conflict rejection prevents extending existing labels (Harper and Pierce, 1991)—the composability that the Expression Problem (Wadler, 1998) requires.
Priority merge:
$A+B\neq B+A$: one source takes precedence (Cunningham, 2014; Dolstra et al., 2010), introducing an ordering dependency.
$\beta$-reduction:
Only the parameters placed by the author are extension points; the function body is closed. To make a new aspect extensible, the original function must be rewritten to accept an additional parameter. This applies to every language whose composition mechanism is function application—functional, object-oriented, and multi-paradigm alike—because method calls, virtual dispatch, and type-class resolution all reduce to $\beta$-reduction. The Expression Problem (Wadler, 1998) is hard in all of them for the same reason: extensibility is opt-in. The frameworks that simulate universal extensibility (visitor pattern, object algebras (d. S. Oliveira and Cook, 2012), finally tagless interpreters (Carette et al., 2009)) are the machinery needed to work around this limitation.
Conditional incorporation:
Removing or conditionally including definitions breaks monotonicity, so the semantic equations (Appendix C) may have no fixed point (Tarski, 1955).
Among the candidates above, deep merge is the only one we found that is commutative, idempotent, associative, and universally extensible. The other candidates each introduce a vulnerability to nonextensibility or break monotonicity.
The reference resolution
Once deep merge over records is fixed as the composition mechanism, four of the six semantic equations are determined by the syntax: $\mathtt{properties}$ (1) reads the labels of a record, $\mathtt{bases}$ (4) reads its inheritance sources, $\mathtt{supers}$ (2) takes the transitive closure, and $\mathtt{overrides}$ (3) implements the merge. The only remaining degree of freedom is how references resolve: the $\mathtt{resolve}$ (5) and $\mathtt{this}$ (6) equations. We examine the three known candidates.
Early binding:
With early binding, references resolve at definition time to a fixed path, independently of how the record is later inherited. CUE (van Lohuizen, 2019) exemplifies this: field references are statically resolved, and a nested field cannot refer to properties of its enclosing context contributed by later composition. This prevents the mixin from serving as a call stack for the $\lambda$-calculus: $\beta$-reduction requires the callee to observe the argument supplied at the call site, which is a later composition. With early binding, references are frozen before this composition occurs. In the equations, $\mathtt{resolve}$ (5) calls $\mathtt{this}$ (6), which walks the inherited tree structure; late binding is what allows $\mathtt{this}$ to see properties contributed by later-composed mixins. Without a Turing-complete mixin, functions must be reintroduced as the composition mechanism, re-exposing the opt-in extensibility problem above.
Dynamic scope:
With dynamic scope, references resolve at the point of use rather than the definition site. Jsonnet (Cunningham, 2014) follows this approach: self always refers to the final merged object, not the object at the definition site. Union file systems exhibit the same property at the file-system level: a symbolic link’s relative path resolves against the mount context at dereference time, not against the layer that defined the link. Dynamic scope breaks the $\lambda$-calculus embedding (Section 4). The Substitution Lemma (Lemma D.2) relies on the base case $M=y$ ($y\neq x$): the de Bruijn index $m$ of $y$ walks up from the definition-site path $\mathtt{init}(p_{\mathrm{def}})$ to a scope level different from $x$’s binder, so the substitution $\mathrm{argument}\mapsto\mathcal{T}(V)$ at $x$’s scope has no effect. With dynamic scope this separation fails. [^4] Counterexample: $(\lambda x.\,\lambda y.\,x)\;V_{1}\;V_{2}$. The reference to $x$ inside $\lambda y.\,x$ translates to $\uparrow^{\!1}.\mathrm{argument}$ (de Bruijn index $1$). With late binding, $\mathtt{this}$ starts from the definition-site scope (the inner $\lambda$) and walks up one level to the outer $\lambda$, reaching the slot $\mathrm{argument}\mapsto\mathcal{T}(V_{1})$. With dynamic scope, the reference resolves from the use-site context, where the innermost enclosing scope is the application $\{C_{1}.\mathrm{result},\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$; walking up one level reaches this scope’s $\mathrm{argument}$ slot, returning $\mathcal{T}(V_{2})$ instead of $\mathcal{T}(V_{1})$. This is the classical variable-capture failure of dynamic scoping. As with early binding, the mixin cannot serve as a complete replacement for functions.
Single-target $\mathtt{this}$:
Single-target $\mathtt{this}$ is a promising candidate: the Single-Path Lemma (D.1) shows that for translated $\lambda$-terms, the frontier set $S$ in $\mathtt{this}$ (6) always contains exactly one path, so the $\lambda$-calculus embedding does not require multi-target resolution. However, single-target $\mathtt{this}$ is vulnerable to nonextensibility for a different reason. When multiple inheritance routes reach the same scope, single-target resolution must reject the situation or select one route. The NixOS module system (NixOS Contributors, 2024) raises a duplicate-declaration error (Appendix B). This rejects nested extensions where two independently authored modules define the same option—a composition that arises when solving the Expression Problem (Wadler, 1998) (Section 5).
Our first three implementations of MIXINv2 used standard techniques for self-reference resolution: closure-based fixed points following Cook (Cook, 1989), and stack-based environment lookup using de Bruijn indices (de Bruijn, 1972). All three produced subtle bugs when inheritance introduced multiple routes to the same enclosing scope. The difficulty was structural: a closure captures a single environment, so a fixed-point combinator solves for a single self-reference; a stack-based environment is a linear chain where each scope level has one parent. Both data structures embed a single-valued assumption incompatible with the multi-path situation that multiple inheritance naturally produces (Appendix B). Equation (6) emerged from abandoning this assumption: it tracks a frontier set $S$ rather than a single current scope.
The following table summarizes the vulnerabilities of all alternatives at both levels.
Candidate Vulnerability Representative
composition Shallow merge nested extensions not composable JS spread
Conflict rejection same-label extensions rejected Harper–Pierce (Harper and Pierce, 1991)
Priority merge ordering dependency Jsonnet + (Cunningham, 2014), Nix // (Dolstra et al., 2010)
$\beta$-reduction opt-in extensibility Java, Haskell, Scala, …
Conditional breaks monotonicity
resolution Early binding mixin cannot replace function CUE (van Lohuizen, 2019)
Dynamic scope mixin cannot replace function Jsonnet (Cunningham, 2014), union FS
Single-target $\mathtt{this}$ nested extensions rejected NixOS modules (NixOS Contributors, 2024)
No alternative we examined at either level is immune to nonextensibility.
The semantics of inheritance-calculus was not chosen from a menu. We arrived at it by trying each alternative in this section and finding that it failed: shallow merge could not compose nested extensions; priority merge introduced ordering dependencies; early binding froze references before composition could supply them; dynamic scope broke the substitution that the $\lambda$-calculus embedding relies on; single-target $\mathtt{this}$ rejected the multi-path compositions that the Expression Problem demands (three successive implementations confirmed this—see above). Deep merge with multi-target late-binding $\mathtt{this}$ is what remained.
7. Related Work
Minimal Turing-complete models
Schönfinkel (Schönfinkel, 1924) and Curry (Curry and Feys, 1958) showed that the $\lambda$-calculus reduces to three combinators $S$, $K$, II (II is redundant: $I=SKK$). The bracket abstraction that compiles $\lambda$-terms to SKI is mechanical but incurs exponential blow-up in term size; no new computational patterns emerge beyond those already present in the $\lambda$-calculus. Dolan (Dolan, 2013) showed that the x86 mov instruction alone is Turing complete by exploiting addressing modes as implicit arithmetic on mutable memory; the resulting programs are correct but provide no abstraction mechanisms beyond those of the RAM Machine. The Turing Machine itself, while universal, lacks random access: reading a cell at distance $d$ requires $d$ sequential tape movements. Each of these models achieves Turing completeness by removing structure from an existing model: fewer abstractions, worse ergonomics, no new expressiveness. None of these reductions exhibits emergent phenomena absent from the model it reduces.
Semantic descriptions of the $\lambda$-calculus
The semantics of the $\lambda$-calculus can be defined in several fundamentally different ways. Operational semantics takes $\beta$-reduction $(\lambda x.\,M)\,N\to M[N/x]$ as the primitive notion of computation (Church, 1936, 1941; Barendregt, 1984). This requires capture-avoiding substitution, a mechanism whose subtleties motivated de Bruijn indices (de Bruijn, 1972) and explicit substitution calculi (Abadi et al., 1991), and a choice of evaluation strategy: call-by-name versus call-by-value (Plotkin, 1975) or the lazy strategy of Abramsky (Abramsky, 1990). Denotational semantics (Scott, 1970, 1976) interprets terms in a mathematical domain, but giving meaning to $\beta$-reduction requires a reflexive domain $D\cong[D\to D]$, which has no set-theoretic solution by Cantor’s theorem; Scott’s $D_{\infty}$ resolves this by constructing $D$ as an inverse limit of continuous lattices (Scott, 1972). The resulting semantics is adequate but not fully abstract (Plotkin, 1977; Milner, 1977): it identifies terms that observational equivalence distinguishes (Plotkin, 1975). Game semantics (Abramsky et al., 2000; Hyland and Ong, 2000; Nickau, 1994) achieves full abstraction for PCF and for the lazy $\lambda$-calculus (Abramsky and Ong, 1993; Abramsky and McCusker, 1995) at the cost of considerably more intricate mathematical machinery (arenas, strategies, innocence conditions). In each of these approaches, the function space $D\to D]$ is the central obstacle: operational semantics needs it implicitly (via substitution), denotational semantics constructs it explicitly, and game semantics refines it via strategies. None of these approaches yields a semantics of the lazy $\lambda$-calculus that simultaneously lives in naive set theory and is fully abstract: Scott’s $D_{\infty}$ is adequate but not fully abstract, and game semantics is fully abstract but requires machinery (arenas, strategies, innocence conditions) far beyond set theory. None of them eliminates the function space from the semantic domain; none avoids capture-avoiding substitution; and none dispenses with reduction rules or evaluation strategies.
Set-theoretic models of the $\lambda$-calculus
Several families of $\lambda$-calculus models avoid the inverse-limit construction of Scott’s $D_{\infty}$ by working in set-theoretic rather than domain-theoretic settings. Graph models (Engeler, 1981; Plotkin, 1993) take a web $A$ equipped with an injection $\iota:\mathcal{P}_{\mathrm{fin}}(A)\times A\hookrightarrow A$ and set $D=\mathcal{P}(A)$. Application is $f\cdot x=\{\,b\mid\exists\,S\subseteq x,\;\iota(S,b)\in f\,\}$: the function space $D\to D]$ is not constructed explicitly but encoded inside $D$ via $\iota$. The construction uses only powerset lattices, no Scott topology or continuity conditions. Graph models are sensible (they equate all unsolvable terms) but not fully abstract: Bucciarelli and Salibra (Bucciarelli and Salibra, 2008) proved that the greatest sensible graph theory equals the Böhm tree theory $\mathcal{B}$, which is strictly coarser than observational equivalence for the lazy $\lambda$-calculus. Filter models (Barendregt et al., 1983; Coppo and Dezani-Ciancaglini, 1978) interpret a term as the filter (upward-closed, intersection-closed set) of all intersection types derivable for it. The collection of all such filters forms an $\omega$-algebraic lattice. Barendregt, Coppo, and Dezani-Ciancaglini proved that type assignment is sound and complete with respect to the filter $\lambda$-model. Dezani-Ciancaglini et al. (Dezani-Ciancaglini et al., 1998) showed that the $\lambda$-theory of the BCD filter model coincides exactly with Böhm tree equality: two terms receive the same set of types if and only if they have the same Böhm tree. Filter models are more discriminating than graph models but are still not fully abstract for the lazy $\lambda$-calculus, and require the additional machinery of intersection type theory and algebraic lattices. Relational models (Ehrhard, 2005; Bucciarelli and Ehrhard, 2001) arise from the relational semantics of linear logic: the exponential modality !! is interpreted via finite multisets $\mathcal{M}_{\mathrm{fin}}(A)$, giving a co-Kleisli category whose morphisms are relations from $\mathcal{M}_{\mathrm{fin}}(A)$ to $B$. The multiset structure corresponds to non-idempotent intersection types (de Carvalho, 2018), which carry quantitative information (bounds on reduction length) that idempotent systems lack. Ehrhard (Ehrhard, 2005) showed that in the finiteness-space refinement of this model, standard fixed-point operators are not finitary, and a standard reflexive object for the pure $\lambda$-calculus cannot exist in this category. For the lazy $\lambda$-calculus, full abstraction has been achieved only through game semantics (Abramsky and Ong, 1993; Abramsky and McCusker, 1995); none of the three set-theoretic families attains it. All three families are semantic models of the $\lambda$-calculus: they interpret $\lambda$-terms in an external mathematical structure and validate $\beta$-reduction as an equation in that structure. None of them is a self-contained calculus with its own primitives into which the $\lambda$-calculus embeds. Moreover, all three identify terms more coarsely than observational equivalence—graph and filter models coincide with Böhm tree equality, while relational models cannot even host a standard reflexive object (Ehrhard, 2005)—and none provides a fully abstract semantics for the lazy $\lambda$-calculus without resorting to game-theoretic machinery.
Denotational semantics of inheritance
Cook (Cook, 1989) gave the first denotational semantics of inheritance, modeling objects as recursive records and inheritance as composition of generators, which are functions from self to complete object, and wrappers, which are functions that modify generators, solved by a fixed-point construction. Cook’s key insight, that inheritance is a general mechanism applicable to any form of recursive definition and not only object-oriented methods, is one of the starting points of the present work. The model has four primitives—records, functions, generators, and wrappers—and requires a fixed-point combinator; composition is asymmetric, as a wrapper modifies a generator but not vice versa.
Mixin-based inheritance
Bracha and Cook (Bracha and Cook, 1990) formalized mixins as abstract subclasses, that is, functions from a superclass parameter to a subclass, unifying the inheritance mechanisms of Smalltalk, Beta, and CLOS under a single model. Because mixin application is function composition, it is neither commutative nor idempotent: applying the same mixin twice may produce different results, and applying two mixins in different orders may yield different linearizations. The C3 linearization algorithm (Barrett et al., 1996) was later developed to impose a deterministic order; whether linearization is inherent to inheritance or an artifact of non-commutativity had not been established. Mixin composition operates at the method level: applying a mixin replaces or extends individual methods in a flat namespace. There is no recursive merging of nested structure: composing two mixins that define the same method name results in an override, not deep merge. Finally, mixins presuppose the $\lambda$-calculus as their computational foundation: method bodies are functions, and mixin application is function composition. Mixins organize and compose functions but cannot replace them; without the underlying $\lambda$-calculus, there is nothing to compose. The mixin mechanism itself is not Turing complete.
Traits
Schärli et al. (Schärli et al., 2003) introduced traits as composable units of behavior. Trait composition is symmetric and commutative: the sum of two traits is order-independent, which is a departure from mixin ordering. However, when two traits provide methods with the same name, a conflict arises and must be explicitly resolved by the programmer through exclusion or aliasing. Ducasse et al. (Ducasse et al., 2006) extended the model with a flattening property: the semantics of a class is independent of whether its methods originate from traits or are defined directly. Like mixins, trait composition is flat: it operates on individual method names, not on nested structures, so same-name definitions are conflicts rather than recursive merges. And like mixins, traits presuppose the $\lambda$-calculus: trait methods are functions, and the computational power resides in those functions, not in the trait mechanism itself.
Family polymorphism and virtual classes
Ernst (Ernst, 2001) introduced family polymorphism, where families of mutually dependent classes can be refined together in subclasses. Ernst, Ostermann, and Cook (Ernst et al., 2006) formalized this in the virtual class calculus, where classes are class-valued attributes of objects, accessed via path expressions of the form this.out.C. The path this.out navigates to an enclosing scope, analogous to qualified this ($\ell_{\mathrm{up}}.\mathbf{this}$) in this paper. A notable restriction is that each such path resolves to a single class in a single enclosing object. When multiple inheritance routes lead to the same scope, a common situation under inheritance, single-valued resolution must choose one route and discard the others, losing information. The framework provides no mechanism for set-valued resolution.
Object calculi
Abadi and Cardelli (Abadi and Cardelli, 1996) developed a comprehensive theory of objects, treating objects rather than classes as the primitive notion. Their calculi include method update and self-referential objects, but methods are fundamentally functions that take self as a parameter, and object extension is asymmetric since method override replaces the previous definition. Boudol (Boudol, 2004) studied the recursive record semantics of objects in a call-by-value setting, showing that self-referential records require an unsafe fixed-point operator. Whether self-reference can be handled without a fixed-point operator in the object language, for instance through coalgebraic observation rather than reduction, was not explored in these calculi.
Record calculi
Harper and Pierce (Harper and Pierce, 1991) presented a record calculus based on symmetric concatenation with a type system that tracks the absence of fields. When two records define the same label, concatenation is undefined and raises a type error; the calculus leaves same-label composition undefined. Cardelli (Cardelli, 1994) studied extensible records with subtyping; Rémy (Rémy, 1989) gave a type system for records and variants in ML. These calculi operate on flat records with no nesting and treat records as data, that is, initial algebras: a record is a finite map from labels to values, fully determined at construction time. Nested records, self-reference, and lazy observation, which are the properties needed for records to serve as a computational substrate, are outside their scope.
DOT calculus
Amin et al. (Amin et al., 2016) formalized path-dependent types in the DOT calculus, the theoretical foundation of Scala’s type system. DOT paths navigate through objects to reach nested type members, providing a rich dependent type structure. DOT’s self variable in the object constructor $\nu(x:T)\,d$ binds $x$ to a single object; when multiple inheritance routes lead to the same enclosing scope, the resulting intersection type is unrealizable and Scala rejects it at compile time (Appendix B). DOT stratifies its syntax into terms and types, with types further divided into abstract type members bounded by $L..U$ and concrete definitions; each stratum requires its own subtyping rules. Whether these three roles could be unified under a single construct remains an open question in the DOT line of work. DOT includes the $\lambda$-calculus as a sub-language; removing it while retaining path-dependent types would yield a calculus closer to pure mixin inheritance (see Section 8).
Configuration languages
CUE (van Lohuizen, 2019) unifies types and values in a single lattice where unification (&) is commutative, associative, and idempotent. CUE’s design was motivated by Google’s Borg Configuration Language (GCL), which used graph unification. CUE’s lattice includes scalar types with a conflict semantics where unifying incompatible scalars yields $\bot$; whether scalars are necessary for a configuration language, or whether tree structure alone suffices, is an interesting design question. Jsonnet (Cunningham, 2014) provides object inheritance via the + operator with mixin semantics: composition is not commutative since the right-hand side wins on scalar conflicts, and deep merging requires explicit opt-in via the +: syntax) on a per-field basis, so deep merge is available but not the default. Dhall (Gonzalez, 2017) takes a functional approach: it is a typed $\lambda$-calculus with records, where composition is a record merge operator, not an inheritance mechanism.
Module systems and deep merge
The NixOS module system (Dolstra et al., 2010; NixOS Contributors, 2024) composes modules by recursively merging nested attribute sets, not by linearization or method-level override, which is the same deep-merge semantics that distinguishes tree-level inheritance from method-level composition in prior mixin and trait calculi. When two modules define the same nested path, their subtrees are merged rather than overridden. The NixOS module system incorporates deep merge, recursive self-reference, deferred evaluation of modules, and a rich type-checking layer on top. The same mechanism powers NixOS system configuration, Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), flake-parts (Hensing, 2022), and dozens of other ecosystems, collectively managing configurations of arbitrary complexity. The one aspect that appears not to have been fully addressed is the semantics of self-reference under multiple inheritance. When two independent modules that inherit from the same deferred module are composed, each import introduces its own declaration of the shared options; the module system’s mergeOptionDecls rejects duplicate declarations with a static error (Appendix B). The module system cannot express that both routes contribute equally to the same option, that is, the multi-path situation that recursive merging naturally gives rise to, which motivated the observational semantics of Section 3.
Modular software architectures
Plugin systems, component frameworks, and dependency injection frameworks enable software extensibility through declarative inheritance. The symmetric, associative nature of inheritance (Section 3) explains why such systems work: components can be inherited in any order without changing behavior. These systems implement inheritance-calculus patterns ad hoc.
Union file systems
Union file systems (UnionFS, OverlayFS, AUFS) layer multiple directory hierarchies to present a unified view. Their semantics resembles inheritance-calculus in that later layers override earlier layers and files from all layers remain accessible, but without a formal foundation these systems resort to ad-hoc choices: asymmetric inheritance, non-commutative conflict resolution for scalar files, and no late-binding or dynamic dispatch across layers. We implemented inheritance-calculus directly as a union file system (Yang, 2025), obtaining late-binding semantics and dynamic dispatch for file lookups with no additional mechanism beyond the three primitives. In package managers, build systems, and OS distributions, external toolchains transform configuration into file trees; here, the file system serves as configuration to itself, and the transformation is inheritance: a file-system-as-compiler in which the directory tree is both the source program and the compilation target.
8. Future Work
MIXINv2 (Yang, 2026) is an executable implementation of inheritance-calculus that already goes beyond the untyped calculus presented here: it includes compile-time checking that all references resolve to valid paths in the mixin tree, and a foreign-function interface (FFI) that introduces scalar values from the host language. The future work described below concerns the distance between this implementation and a fully practical language, spanning both theoretical foundations and library-level encodings.
8.1. Type System
MIXINv2’s existing compile-time checks verify that every reference path resolves to a property that exists in the inherited result. Formalizing the soundness of these checks, that is, proving that well-typed programs do not produce dangling references at runtime, is one direction of future work. Such a formalization would relate to the DOT calculus (Amin et al., 2016) in a manner analogous to how inheritance-calculus relates to the $\lambda$-calculus: a Typed inheritance-calculus can be viewed as DOT without $\lambda$, retaining path-dependent types while replacing functions with mixin inheritance. A formal type system would be amenable to mechanization in proof assistants (Coq, Agda) for verifying type safety, progress, and preservation properties.
Beyond soundness of the existing checks, additional type system features would strengthen the language. A key example is totality checking: verifying that an inheritance provides implementations for all required slots, not merely that references resolve. In untyped inheritance-calculus, writing $\mathrm{magnitude}\mapsto\{\}$ is purely documentary: it defines a structural slot but imposes no constraint on what must fill it. Totality checking would enforce such declarations statically, rejecting inheritances that leave required slots unfilled.
8.2. Standard Library
The remaining items concern the standard library: they require no extensions to the calculus or type system, but involve nontrivial encodings within the existing framework.
Inheritance-calculus is open by default: inheritance can freely merge any two mixins, and a value may simultaneously inhabit multiple constructors (e.g., $\{\mathrm{Zero},\;\mathrm{Odd},\;\mathrm{half}\mapsto\mathrm{Zero}\}$). This is not a defect; it is the natural trie semantics of the calculus, and the basis for solving the expression problem (Section 5). However, many operations (e.g., equality testing) assume that values are linear, inhabiting exactly one constructor. The observer pattern (Section 5) does not enforce this: applied to a multi-constructor value, all callbacks fire and their results are inherited.
Schema validation can be implemented at the library level using existing primitives. An observer can test whether a particular constructor is present, returning a Boolean; a second Boolean dispatch then branches on the result. By chaining such tests, a factory can validate that a value matches exactly one constructor and carries the required fields. This does not restrict inheritance itself, which remains unconstrained, but provides a way to detect invariant violations before they propagate.
The same technique, an observer returning Boolean followed by Boolean dispatch, yields closed pattern matching. The observer pattern of Section 5 is inherently open: new constructor cases can be added through inheritance. Closed matching, where exactly one branch is taken, can be encoded as a visitor chain: a linked list of if-then-else nodes, each testing one constructor and falling through on mismatch. This is analogous to GHC.Generics’ $\mathrm{(:+:)}$ sum representation (Magalhães et al., 2010), where each link corresponds to one summand. An important consequence is that such chains do not commute, as they have a fixed order, and therefore cannot be extended through open inheritance, which is the correct semantics for closed dispatch.
9. Conclusion
Immunity to nonextensibility guided the design of inheritance-calculus. The emergent phenomena exhibit an expressive asymmetry between inheritance-calculus and the $\lambda$-calculus. Inheritance is not a mechanism for organizing functions; functions are syntactic sugar for inheritance.
Appendix A Expressive Asymmetry: Proofs
Definition A.1 (Macro-expressibility, after Felleisen (Felleisen, 1991)).
A construct $C$ of language $\mathscr{L}_{1}$ is macro-expressible in language $\mathscr{L}_{0}$ if there exists a translation $\mathcal{E}$ such that for every program $P$ containing occurrences of $C$, the translation $\mathcal{E}(P)$ is obtained by replacing each occurrence of $C$ with an $\mathscr{L}_{0}$ expression that depends only on $C$’s subexpressions, leaving all other constructs of $P$ unchanged.
Theorem A.2 (Forward macro-expressibility).
The translation $\mathcal{T}$ (Section 4) is a macro-expressible embedding of the $\lambda$-calculus into inheritance-calculus.
Proof.
Each of the five rules of $\mathcal{T}$ replaces one $\lambda$-calculus construct with an inheritance-calculus expression that depends only on the construct’s immediate subexpressions. No rule inspects the surrounding context. ∎
Theorem A.3 (Non-expressibility of inheritance-calculus).
There is no macro-expressible translation from inheritance-calculus to the $\lambda$-calculus.
Proof.
Let $\mathscr{L}_{1}$ be the inheritance-calculus and let $\mathscr{L}_{0}$ be the sublanguage consisting of mixin trees in the image of the translation $\mathcal{T}$ (Section 4), so that $\mathscr{L}_{0}\subset\mathscr{L}_{1}$. Define the operational equivalence $\cong_{0}$ on $\mathscr{L}_{0}$ via Böhm trees: $\mathcal{T}(M)\cong_{0}\mathcal{T}(N)$ iff $\mathrm{BT}(M)=\mathrm{BT}(N)$. Define $\cong_{1}$ on $\mathscr{L}_{1}$ as inheritance-contextual equivalence: $S\cong_{1}S^{\prime}$ iff for every inheritance-calculus context $C[\cdot]$, $C[S]{\Downarrow}\Leftrightarrow C[S^{\prime}]{\Downarrow}$.
By the Böhm Tree Correspondence (Theorems 4.2 and 4.3), $\mathscr{L}_{1}$ conservatively extends $\mathscr{L}_{0}$: for closed $\lambda$-terms $M$, $\mathcal{T}(M){\Downarrow}$ if and only if $M$ has a head normal form.
Define Church booleans via the translation $\mathcal{T}$ (Section 4): $\mathrm{true}=\lambda t.\,\lambda f.\,t$ and $\mathrm{false}=\lambda t.\,\lambda f.\,f$. In de Bruijn notation:
$\mathrm{true}$ $\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\uparrow^{\!1}.\mathrm{argument}\}\}$
$\mathrm{false}$ $\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\uparrow^{\!0}.\mathrm{argument}\}\}$
Define Church boolean equality $\mathrm{eq}=\lambda a.\,\lambda b.\,(a\;b)\;(b\;\mathrm{false}\;\mathrm{true})$. In ANF:
$\mathrm{eq}=\lambda a.\,\lambda b.\,\mathbf{let}\;x_{1}=b\;\mathrm{false}\;\mathbf{in}\;\mathbf{let}\;x_{2}=x_{1}\;\mathrm{true}\;\mathbf{in}\;\mathbf{let}\;x_{3}=a\;b\;\mathbf{in}\;x_{3}\;x_{2}$
Applying $\mathcal{T}$ (Section 4):
$\mathrm{eq}$ $\mapsto\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\{\mathrm{argument}\mapsto\{\},$
$\quad x_{1}\mapsto\{\uparrow^{\!0}.\mathrm{argument},\;\mathrm{argument}\mapsto\mathrm{false}\},$
$\quad x_{2}\mapsto\{x_{1}.\mathrm{result},\;\mathrm{argument}\mapsto\mathrm{true}\},$
$\quad x_{3}\mapsto\{\uparrow^{\!1}.\mathrm{argument},\;\mathrm{argument}\mapsto\uparrow^{\!0}.\mathrm{argument}\},$
$\quad\_\mathrm{tail}\mapsto\{x_{3}.\mathrm{result},\;\mathrm{argument}\mapsto x_{2}.\mathrm{result}\},$
$\quad\mathrm{result}\mapsto\_\mathrm{tail}.\mathrm{result}\}\}$
Here $\uparrow^{\!1}.\mathrm{argument}$ in $x_{3}$ reaches the outer $\lambda a$ scope (one $\lambda$ up from $\lambda b$), and $\uparrow^{\!0}.\mathrm{argument}$ reaches $\lambda b$ itself.
In inheritance-calculus, applying a function record to an argument is: inherit the function, override its $\mathrm{argument}$ slot, and project $\mathrm{result}$. Define mixin trees $e_{1}$ and $e_{2}$ that apply $\mathrm{eq}$ to its two arguments via this mechanism:
$e_{1}\;=\;\{$ $\mathrm{eqFalse}\mapsto\{\mathrm{eq},\;\mathrm{argument}\mapsto\mathrm{false}\},$
$\mathrm{applied}\mapsto\{\mathrm{eqFalse}.\mathrm{result},\;\mathrm{argument}\mapsto\mathrm{false}\},$
$\mathrm{result}\mapsto\mathrm{applied}.\mathrm{result}\;\}$
$e_{2}\;=\;\{$ $\mathrm{eqTrue}\mapsto\{\mathrm{eq},\;\mathrm{argument}\mapsto\mathrm{true}\},$
$\mathrm{applied}\mapsto\{\mathrm{eqTrue}.\mathrm{result},\;\mathrm{argument}\mapsto\mathrm{true}\},$
$\mathrm{result}\mapsto\mathrm{applied}.\mathrm{result}\;\}$
That is, $e_{1}=\mathcal{T}(\mathrm{ANF}(\mathrm{eq}\;\mathrm{false}\;\mathrm{false}))$ and $e_{2}=\mathcal{T}(\mathrm{ANF}(\mathrm{eq}\;\mathrm{true}\;\mathrm{true}))$, where $\mathrm{ANF}$ converts a $\lambda$-term to A-normal form by naming every intermediate application (Section 4). Since both $\mathrm{ANF}$ and $\mathcal{T}$ are macro-expressible, their composition is too, so $e_{1},e_{2}\in\mathscr{L}_{0}$. Both $e_{1}.\mathrm{result}$ and $e_{2}.\mathrm{result}$ evaluate to Church $\mathrm{true}$. By Full Abstraction (Theorem 4.3), $\mathrm{BT}(e_{1})=\mathrm{BT}(e_{2})$, so $e_{1}\cong_{0}e_{2}$.
[…]
$C={}$ $\{\mathrm{eq}\mapsto\{\mathrm{result}\mapsto\{\mathrm{repr}\mapsto\{$
$\qquad\mathrm{firstOperand}\mapsto\uparrow^{\!2}.\mathrm{argument},\;\mathrm{secondOperand}\mapsto\uparrow^{\!1}.\mathrm{argument}\}\}\}\}$
The de Bruijn indices count scope levels from $\mathrm{repr}$: $\uparrow^{\!0}$ is $\mathrm{repr}$ itself, $\uparrow^{\!1}$ is the $\lambda b$ scope, and $\uparrow^{\!2}$ is the $\lambda a$ scope. Thus $\mathrm{firstOperand}$ projects $\lambda a$’s argument and $\mathrm{secondOperand}$ projects $\lambda b$’s argument.
In the composite scope $\{S,\,C\}$, the definition of $\mathrm{eq}$ gains $\mathrm{repr}$ as a new sibling of $x_{1},x_{2},x_{3}$ inside the $\lambda b$ scope. By open recursion, the de Bruijn references $\uparrow^{\!2}.\mathrm{argument}$ and $\uparrow^{\!1}.\mathrm{argument}$ resolve through the same scope chain as the existing let-bindings. Since $e_{1}$ and $e_{2}$ inherit $\mathrm{eq}$, they also inherit the newly added $\mathrm{repr}$: $e_{1}.\mathrm{repr}.\mathrm{firstOperand}=\mathrm{false}$ and $e_{2}.\mathrm{repr}.\mathrm{firstOperand}=\mathrm{true}$. These produce structurally distinct mixin trees (one converges to an abstraction selecting its first argument, the other to an abstraction selecting its second), so $e_{1}\not\cong_{1}e_{2}$.
Since $e_{1}\cong_{0}e_{2}$ but $e_{1}\not\cong_{1}e_{2}$, we have ${\cong_{0}}\neq{(\cong_{1}|_{\mathscr{L}_{0}})}$. By Theorem 3.14(i) of Felleisen (Felleisen, 1991), $\mathscr{L}_{0}$ cannot macro-express the inheritance facilities of $\mathscr{L}_{1}$. ∎
Corollary A.4.
Inheritance-calculus is strictly more expressive than the $\lambda$-calculus in the sense of Felleisen (Felleisen, 1991): the forward embedding is macro-expressible (Theorem A.2), but the reverse is not (Theorem A.3).
Appendix B Multi-Path Self-Reference in Other Systems
This appendix demonstrates that two representative systems, Scala 3 and the NixOS module system, reject the multi-path self-reference pattern that inheritance-calculus handles naturally.
Scala 3
Consider two objects that independently extend an outer class, each providing its own copy of an inner trait:
class MyOuter:
trait MyInner:
def outer = MyOuter.this
object Object1 extends MyOuter
object Object2 extends MyOuter
object HasMultipleOuters extends Object1.MyInner
with Object2.MyInner
Scala 3 rejects HasMultipleOuters with the error:
trait MyInner is extended twice
object HasMultipleOuters cannot be instantiated since
it has conflicting base types
Object1.MyInner and Object2.MyInner
The rejection is not a surface-level restriction but a consequence of DOT’s type-theoretic foundations (Amin et al., 2016). In DOT, an object is constructed as $\nu(x:T)\,d$, where the self variable $x$ binds to a single object. Object1.MyInner and Object2.MyInner are distinct path-dependent types, each constraining the outer self-reference to a different object. Their intersection requires outer to simultaneously return Object1 and Object2, but DOT’s self variable is single-valued, making this intersection unrealizable.
The equivalent inheritance-calculus definition is:
$\{\quad\mathrm{MyOuter}$ $\mapsto\{\mathrm{MyInner}\mapsto\{\mathrm{outer}\mapsto\mathrm{MyOuter}.\mathbf{this}\}\},$
$\mathrm{Object1}$ $\mapsto\{\mathrm{MyOuter}\},$
$\mathrm{Object2}$ $\mapsto\{\mathrm{MyOuter}\},$
$\mathrm{HasMultipleOuters}$ $\mapsto\{\mathrm{Object1}.\mathrm{MyInner},\;\mathrm{Object2}.\mathrm{MyInner}\}\quad\}$
This is well-defined in inheritance-calculus. The reference $\mathrm{MyOuter}.\mathbf{this}$ inside $\mathrm{outer}$ has de Bruijn index $n=1$: starting from $\mathrm{outer}$’s enclosing scope $\mathrm{MyInner}$, one $\mathtt{this}$ step reaches $\mathrm{MyOuter}$. When $\mathrm{HasMultipleOuters}$ inherits from both $\mathrm{Object1}.\mathrm{MyInner}$ and $\mathrm{Object2}.\mathrm{MyInner}$, the $\mathtt{this}$ function (equation 6) resolves $\mathrm{MyOuter}.\mathbf{this}$ by searching through $\mathtt{supers}(\mathrm{HasMultipleOuters})$ for override paths matching $\mathrm{MyInner}$’s definition site. It finds two inheritance-site paths, one through $\mathrm{Object1}$ and one through $\mathrm{Object2}$, and returns both. Both paths lead to records that inherit from the same $\mathrm{MyOuter}$, so querying $\mathrm{HasMultipleOuters}.\mathrm{outer}$ yields the $\mathtt{properties}$ of $\mathrm{MyOuter}$ regardless of which route is taken. Since inheritance is idempotent, the two routes contribute the same properties in this semantics, so no route-priority disambiguation rule is introduced.
NixOS Module System
The NixOS module system rejects the same pattern with a static error. The translation uses the module system’s own abstractions: deferredModule for traits, which are unevaluated module values imported into other fixpoints, and submoduleWith for objects, which are evaluated in their own fixpoints.
let
lib = (import <nixpkgs> { }).lib;
result = lib.evalModules {
modules = [
(toplevel@{ config, ... }: {
# class MyOuter { trait MyInner {
# def outer = MyOuter.this } }
options.MyOuter = lib.mkOption {
default = { };
type = lib.types.deferredModuleWith {
staticModules = [
(MyOuter: {
options.MyInner = lib.mkOption {
default = { };
type = lib.types.deferredModuleWith {
staticModules = [
(MyInner: {
options.outer = lib.mkOption {
default = { };
type =
lib.types.deferredModuleWith {
staticModules =
[ toplevel.config.MyOuter ];
# object Object1 extends MyOuter
options.Object1 = lib.mkOption {
default = { };
type = lib.types.submoduleWith {
modules = [ toplevel.config.MyOuter ];
# object Object2 extends MyOuter
options.Object2 = lib.mkOption {
default = { };
type = lib.types.submoduleWith {
modules = [ toplevel.config.MyOuter ];
# object HasMultipleOuters extends
# Object1.MyInner with Object2.MyInner
options.HasMultipleOuters = lib.mkOption {
default = { };
type = lib.types.submoduleWith {
modules = [
toplevel.config.Object1.MyInner
toplevel.config.Object2.MyInner
in
builtins.attrNames result.config.HasMultipleOuters.outer
The NixOS module system rejects this with:
error: The option HasMultipleOuters.outer
in ‘<unknown-file>’
is already declared
in ‘<unknown-file>’.
The rejection occurs because Object1.MyInner and Object2.MyInner each carry the full staticModules of the MyInner deferred module, including the declaration options.outer. When MyObjectA imports both, the module system’s mergeOptionDecls encounters two declarations of the same option and raises a static error; it cannot express that both declarations originate from the same trait and should be unified.
The rejection is analogous to Scala’s “conflicting base types” error above: both systems assume that each option or type member has a single declaration site, and reject the multi-path situation where two inheritance routes introduce the same declaration independently. In inheritance-calculus, the same pattern is well-defined because $\mathtt{overrides}$ (equation 3) recognizes that both routes lead to the same definition, and $\mathtt{supers}$ (equation 2) collects both inheritance-site contexts without duplication.
Appendix C Well-Definedness of the Semantic Functions
This appendix proves that the mutually recursive equations (1)–(6) have a unique least fixed point, and therefore the semantics of Section 3 is well-defined.
C.1. The Product Lattice
Fix an AST with its primitive functions $\mathtt{defines}$ and $\mathtt{inherits}$. Let $\mathcal{L}$ denote the set of all labels and $\mathcal{P}$ the set of all paths (finite sequences of labels). Define the following domains, one for each semantic function:
$D_{\mathtt{properties}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{L})$ (sets of labels)
$D_{\mathtt{supers}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P}\times\mathcal{P})$ (sets of site–override pairs)
$D_{\mathtt{overrides}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P})$ (sets of paths)
$D_{\mathtt{bases}}$ $=\mathcal{P}\to\mathcal{P}(\mathcal{P})$ (sets of paths)
$D_{\mathtt{resolve}}$ $=(\mathcal{P}\times\mathcal{P}\times\mathbb{N}\times\mathcal{L}^{*})\to\mathcal{P}(\mathcal{P})$ (sets of paths)
$D_{\mathtt{this}}$ $=(\mathcal{P}(\mathcal{P})\times\mathcal{P}\times\mathbb{N})\to\mathcal{P}(\mathcal{P})$ (sets of paths)
Each $D_{i}$ is a complete lattice under the pointwise subset ordering: $f\sqsubseteq g$ iff $f(x)\subseteq g(x)$ for all $x$. The product $D=D_{\mathtt{properties}}\times D_{\mathtt{supers}}\times D_{\mathtt{overrides}}\times D_{\mathtt{bases}}\times D_{\mathtt{resolve}}\times D_{\mathtt{this}}$ is a complete lattice under the componentwise ordering, with bottom element $\bot=(\lambda x.\,\varnothing,\ldots,\lambda x.\,\varnothing)$.
C.2. The Operator
Equations (1)–(6) define a function $F:D\to D$ that takes a tuple of candidate interpretations $(\mathtt{properties},\mathtt{supers},\mathtt{overrides},\mathtt{bases},\mathtt{resolve},\mathtt{this})$ and produces a new tuple by evaluating the right-hand side of each equation. The transitive closure $\mathtt{bases}^{*}$ appearing in equation (2) is derived from $\mathtt{bases}$: for any fixed $\mathtt{bases}$, the set $\mathtt{bases}^{*}(p)$ is defined as the least fixed point of the monotone operator $X\mapsto\{p\}\cup\bigcup_{q\in X}\mathtt{bases}(q)$ on $\mathcal{P}(\mathcal{P})$. If $\mathtt{bases}\sqsubseteq\mathtt{bases}^{\prime}$ (pointwise), then $\mathtt{bases}^{*}\sqsubseteq\mathtt{bases}^{\prime*}$ (pointwise), since enlarging the one-step relation can only enlarge the transitive closure.
The $\mathtt{this}$ function involves a recursion on the de Bruijn index $n$. Since $n$ decreases by one at each step and is a non-negative integer, this recursion terminates for any fixed interpretation of $\mathtt{supers}$. The result of $\mathtt{this}(S,p_{\mathrm{def}},n)$ for all valid $(S,p_{\mathrm{def}},n)$ is therefore well-defined as a function of $\mathtt{supers}$, and it remains to show that it is monotone in $\mathtt{supers}$.
C.3. Monotonicity
Proposition C.1.
The operator $F$ is monotone: if $\vec{d}\sqsubseteq\vec{d}^{\prime}$ in $D$, then $F(\vec{d})\sqsubseteq F(\vec{d}^{\prime})$.
Proof.
It suffices to show that each component of $F$ is monotone in every argument, i.e., that enlarging any of the input interpretations (in the $\subseteq$ ordering) can only enlarge the output. We verify this by inspecting the right-hand side of each equation and confirming that every occurrence of a mutually defined function appears in a positive position—as the domain of an existential quantifier in a set comprehension (Aczel, 1977).
Equation (1): $\mathtt{properties}$
$\mathtt{properties}(p)=\bigl\{\,\ell\bigm|(\_,p_{\mathrm{override}})\in\mathtt{supers}(p),\;\ell\in\mathtt{defines}(p_{\mathrm{override}})\,\bigr\}.$
The function $\mathtt{supers}$ appears as the set being iterated over ($\in\mathtt{supers}(p)$). The predicate $\ell\in\mathtt{defines}(\cdot)$ involves only the AST primitive $\mathtt{defines}$, not any mutually defined function. Enlarging $\mathtt{supers}(p)$ can only add pairs, hence can only add labels to the result. Positive in $\mathtt{supers}$.
Equation (2): $\mathtt{supers}$
$\mathtt{supers}(p)=\bigl\{\,(\mathtt{init}(p_{\mathrm{base}}),p_{\mathrm{override}})\bigm|p_{\mathrm{base}}\in\mathtt{bases}^{*}(p),\;p_{\mathrm{override}}\in\mathtt{overrides}(p_{\mathrm{base}})\,\bigr\}$. Both $\mathtt{bases}^{*}$ and $\mathtt{overrides}$ appear as sets being iterated over. As noted above, $\mathtt{bases}^{*}$ is monotone in $\mathtt{bases}$. Enlarging either $\mathtt{bases}$ or $\mathtt{overrides}$ can only add elements to the result. Positive in $\mathtt{bases}$ and $\mathtt{overrides}$.
Equation (3): $\mathtt{overrides}$
For $p\neq()$:
$\mathtt{overrides}(p)=\{p\}\cup\bigl\{\,p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)\bigm|{}\\ (\_,p_{\mathrm{branch}})\in\mathtt{supers}(\mathtt{init}(p)),\;\mathtt{last}(p)\in\mathtt{defines}(p_{\mathrm{branch}})\,\bigr\}.$
The function $\mathtt{supers}$ appears as the set being iterated over. The filtering condition $\mathtt{last}(p)\in\mathtt{defines}(\cdot)$ involves only the AST primitive. Enlarging $\mathtt{supers}$ can only add branches, hence can only add paths to the result. Positive in $\mathtt{supers}$.
Equation (4): $\mathtt{bases}$
$\mathtt{bases}(p)=\bigl\{\,p_{\mathrm{target}}\bigm|p_{\mathrm{override}}\in\mathtt{overrides}(p),\\ (n,\ell_{\mathrm{down},*})\in\mathtt{inherits}(p_{\mathrm{override}}),\;p_{\mathrm{target}}\in\mathtt{resolve}(\ldots)\,\bigr\}.$
Both $\mathtt{overrides}$ and $\mathtt{resolve}$ appear as sets being iterated over. The function $\mathtt{inherits}$ is an AST primitive. Enlarging either $\mathtt{overrides}$ or $\mathtt{resolve}$ can only add targets to the result. Positive in $\mathtt{overrides}$ and $\mathtt{resolve}$.
Equation (5): $\mathtt{resolve}$
$\mathtt{resolve}(\ldots)=\bigl\{\,p_{\mathrm{current}}\mathbin{\triangleright}\ell_{\mathrm{down},1}\mathbin{\triangleright}\cdots\bigm|p_{\mathrm{current}}\in\mathtt{this}(\{p_{\mathrm{site}}\},\mathtt{init}(p_{\mathrm{def}}),n)\,\bigr\}.$
The function $\mathtt{this}$ appears as the set being iterated over. Enlarging $\mathtt{this}$ can only add paths to the result. Positive in $\mathtt{this}$.
Equation (6): $\mathtt{this}$
For $n>0$: $\mathtt{this}(S,p_{\mathrm{def}},n)=\mathtt{this}(S^{\prime},\mathtt{init}(p_{\mathrm{def}}),n-1)$ where $S^{\prime}=\bigl\{\,p_{\mathrm{site}}\bigm|p_{\mathrm{current}}\in S,\;(p_{\mathrm{site}},p_{\mathrm{override}})\in\mathtt{supers}(p_{\mathrm{current}}),\;p_{\mathrm{override}}=p_{\mathrm{def}}\,\bigr\}$. The function $\mathtt{supers}$ appears as the set being iterated over. The filtering condition $p_{\mathrm{override}}=p_{\mathrm{def}}$ is an equality test on path data, not involving any mutually defined function. Enlarging $\mathtt{supers}$ can only enlarge $S^{\prime}$, and by induction on $n$ (the base case $n=0$ returns $S$ unchanged), the recursive call is monotone in $S^{\prime}$. Positive in $\mathtt{supers}$.
In summary, every mutually defined function appears on the right-hand side of every equation only inside the pattern $x\in f(\ldots)$ within a set comprehension, i.e., as the domain of an existential quantifier. No complement, set difference, or negated membership test appears in any equation. Therefore $F$ is a monotone operator on the complete lattice $D$. ∎
C.4. Existence of the Least Fixed Point
Theorem C.2.
The semantic functions $\mathtt{properties}$, $\mathtt{supers}$, $\mathtt{overrides}$, $\mathtt{bases}$, $\mathtt{resolve}$, $\mathtt{this}$ are well-defined as the least fixed point of $F$.
Proof.
By Proposition C.1, $F$ is monotone on the complete lattice $D$. By the Knaster–Tarski theorem (Tarski, 1955), $F$ has a least fixed point $\vec{d}^{*}=\bigsqcup_{n\geq 0}F^{n}(\bot)$, where $\bot$ assigns the empty set to every input. The components of $\vec{d}^{*}$ are the semantic functions of Section 3.
Concretely, the least fixed point is the union of the finite approximation chain: $\bot\sqsubseteq F(\bot)\sqsubseteq F^{2}(\bot)\sqsubseteq\cdots$. A membership $\ell\in\mathtt{properties}(p)$ holds in $\vec{d}^{*}$ if and only if it appears at some finite stage $F^{k}(\bot)$, which corresponds to a finite derivation tree using the equations. When no finite derivation exists, the membership is absent from the least fixed point and the query is undefined, the analogue of divergence. ∎
Appendix D Böhm Tree Correspondence: Proofs
This appendix contains the full proofs of Adequacy (Theorem 4.2) and Full Abstraction (Theorem 4.3) stated in Section 4.1.
D.1. Böhm Trees
We briefly recall the definition of Böhm trees (Barendregt, 1984). A head reduction $M\to_{h}M^{\prime}$ contracts the outermost $\beta$-redex only: if $M$ has the form $(\lambda x.\,e)\;v\;M_{1}\cdots M_{k}$, then $M\to_{h}e[v/x]\;M_{1}\cdots M_{k}$. A term $M$ is in head normal form (HNF) if it has no head redex, i.e., $M=\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}$ where $y$ is a variable.
The Böhm tree $\mathrm{BT}(M)$ of a $\lambda$-term $M$ is an infinite labeled tree defined by:
$\mathrm{BT}(M)=\begin{cases}\bot&\text{if }M\text{ has no HNF}\\[6.0pt] \lambda x_{1}\ldots x_{m}.\;y\bigl(\mathrm{BT}(M_{1}),\;\ldots,\;\mathrm{BT}(M_{k})\bigr)&\begin{aligned} &\text{if }M\to^{*}_{h}\\[-2.0pt] &\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}\end{aligned}\end{cases}$
Böhm tree equivalence $\mathrm{BT}(M)=\mathrm{BT}(N)$ is the standard observational equivalence of the lazy $\lambda$-calculus (Barendregt, 1984; Abramsky and Ong, 1993).
D.2. Path Encoding
We define a correspondence between positions in the Böhm tree and paths in the mixin tree. When the translation $\mathcal{T}$ is applied to an ANF $\lambda$-term, the resulting mixin tree has a specific shape:
  • An abstraction $\lambda x.\,M$ translates to a record with own properties $\{\mathrm{argument},\mathrm{result}\}$ (the abstraction shape).
  • A $\mathbf{let}$-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M$ translates to a record with own properties $\{x,\mathrm{result}\}$, where $x$ holds the encapsulated application $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$.
  • A tail call $V_{1}\;V_{2}$ translates to a record with own properties $\{\_\mathrm{tail},\mathrm{result}\}$, where $\mathrm{tail}$ holds the encapsulated application and $\mathrm{result}$ projects $\mathrm{tail}.\mathrm{result}$.
  • A variable $x$ with de Bruijn index $n$ translates to an indexed reference $\uparrow^{\!n}.\mathrm{argument}$.
Each Böhm tree position is a sequence of navigation steps from the root. In the Böhm tree, the possible steps at a node $\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}$ are: entering the body (peeling off one abstraction layer) and entering the $i$-th argument $M_{i}$ of the head variable. In the mixin tree, these correspond to following the $\mathrm{result}$ label (for entering the body) and the $\mathrm{argument}$ label after $\mathtt{this}$ resolution (for entering an argument position).
Rather than comparing absolute paths across different mixin trees (which would be sensitive to internal structure), we observe only the convergence behavior accessible via the $\mathrm{result}$ projection. Recall the definition of inheritance-convergence (Definition 4.1). Each $\mathrm{result}$-following step corresponds to one head reduction step in the $\lambda$-calculus: a tail call’s $\mathrm{result}$ projects through the encapsulated application, and a $\mathbf{let}$-binding’s $\mathrm{result}$ enters the continuation. The abstraction shape at the end signals that a weak head normal form has been reached.
D.3. Single-Path Lemma
The $\mathtt{this}$ function (equation 6) is designed for the general case of multi-path mixin inheritance. For translated $\lambda$-terms, we show it degenerates to single-path resolution.
Lemma D.1 (Single Path).
For any closed ANF $\lambda$-term $M$, at every invocation of $\mathtt{this}(S,\;p_{\mathrm{def}},\;n)$ during evaluation of $\mathcal{T}(M)$, the frontier set $S$ contains exactly one path.
Proof.
By induction on the structure of the ANF term.
Case $\mathcal{T}(\lambda x.\,M)=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
This is a record literal with $\mathtt{defines}=\{\mathrm{argument},\mathrm{result}\}$ and no inheritance sources, so $\mathtt{inherits}=\varnothing$. At the root path $p$, we have $\mathtt{overrides}(p)=\{p\}$ (a single path) since there are no inheritance sources to introduce additional branches. Inside $\mathcal{T}(M)$, any reference to $x$ is $\uparrow^{\!n}.\mathrm{argument}$ where $n$ is the de Bruijn index pointing to this scope level. Since $\mathcal{T}(\lambda x.\,M)$ introduces exactly one scope level, there is exactly one mixin at that level, hence $\mathtt{this}$ finds exactly one matching pair $(p_{\mathrm{site}},\;p_{\mathrm{override}})$ in $\mathtt{supers}$.
Case $\mathcal{T}(\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M)=\{x\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
The outer record has own properties $\{x,\mathrm{result}\}$ and no inheritance sources at the outer level ($\mathtt{inherits}=\varnothing$), so $\mathtt{overrides}(\text{root})=\{\text{root}\}$. Inside the $x$ subtree, the inheritance $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$ has exactly one inheritance source ($\mathcal{T}(V_{1})$), creating exactly one $\mathtt{bases}$ entry. By induction on $V_{1}$, the single-path property holds inside $\mathcal{T}(V_{1})$. Inside $\mathcal{T}(M)$, the induction hypothesis applies directly.
Case $\mathcal{T}(V_{1}\;V_{2})=\{\_\mathrm{tail}\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\_\mathrm{tail}.\mathrm{result}\}$
Identical to the $\mathbf{let}$-binding case: the outer record has own properties $\{\_\mathrm{tail},\mathrm{result}\}$ with no inheritance sources at the outer level, and the $\mathrm{tail}$ subtree contains exactly one inheritance source.
Case $\mathcal{T}(x)=\uparrow^{\!n}.\mathrm{argument}$
This is a reference, not a record. The $\mathtt{this}$ function walks up $n$ steps from the enclosing scope. By the inductive hypothesis on the enclosing translations, each step of $\mathtt{this}$ encounters exactly one matching pair in $\mathtt{supers}$.
The key structural invariant is that the translation $\mathcal{T}$ never produces two independent inheritance sources at the same scope level that both define the same label. In every case, application inheritances are encapsulated inside a named property ($x$ or $\mathrm{tail}$), and each such encapsulated inheritance has exactly one inheritance source. This prevents diamond inheritance and ensures single-path resolution throughout. ∎
D.4. Substitution Lemma
The core mechanism of the translation is that inheritance with $\mathrm{argument}\mapsto\mathcal{T}(V)$ plays the role of substitution. We make this precise.
Lemma D.2 (Substitution).
Let $M$ be an ANF term in which $x$ may occur free, and let $V$ be a closed value. Define the inherited tree $C=\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$. Then for every path $p$ under $C\mathbin{\triangleright}\mathrm{result}$ and every label $\ell$:
$\ell\in\mathtt{properties}(p)\text{ in }C\mathbin{\triangleright}\mathrm{result}\quad\Longleftrightarrow\quad\ell\in\mathtt{properties}(p)\text{ in }\mathcal{T}(M[V/x])$
where $M[V/x]$ is the usual capture-avoiding substitution.
Proof.
By structural induction on $M$.
Base case: $M=x$
Then $\mathcal{T}(x)=\uparrow^{\!n}.\mathrm{argument}$ (where $n$ is the de Bruijn index of $x$), and $C\mathbin{\triangleright}\mathrm{result}$ contains this reference. In the inherited tree $C$, $\mathtt{resolve}$ and $\mathtt{this}$ resolve the reference by navigating from the reference’s enclosing scope up $n$ steps to the binding $\lambda$, where the inheritance $\{\mathcal{T}(\lambda x.\,x),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$ inherits into the $\mathrm{argument}$ slot, providing $\mathcal{T}(V)$. Since $C$ is the $\mathrm{tail}$ subtree of $\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)$, which is a closed ANF term, Lemma D.1 applies and $\mathtt{this}$ finds exactly one path, so $\mathtt{resolve}$ returns $\mathcal{T}(V)$’s subtree. Since $M[V/x]=V$, we have $\mathcal{T}(M[V/x])=\mathcal{T}(V)$, and the properties coincide.
Base case: $M=y$ where $y\neq x$
Then $\uparrow^{\!m}.\mathrm{argument}$ (where $m$ is the de Bruijn index of $y$, $m\neq n$) walks up to a different scope level, so the inheritance with $\mathrm{argument}\mapsto\mathcal{T}(V)$ at the binding scope of $x$ has no effect. Since $M[V/x]=y$, the properties are identical.
Inductive case: $M=\lambda y.\,M^{\prime}$
Then $\mathcal{T}(M)=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$. The subtree at $C\mathbin{\triangleright}\mathrm{result}$ is another abstraction-shaped record. References to $x$ inside $\mathcal{T}(M^{\prime})$ resolve through $\mathtt{this}$ in exactly the same way (with one additional scope level from the inner $\lambda$), and by induction on $M^{\prime}$, the properties under $C\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}\mathrm{result}$ coincide with those of $\mathcal{T}(M^{\prime}[V/x])$. Since $(\lambda y.\,M^{\prime})[V/x]=\lambda y.\,(M^{\prime}[V/x])$ (assuming $y$ is fresh), the result follows.
Inductive case: $M=\mathbf{let}\;z=V_{1}\;V_{2}$ $\mathbf{in}\;M^{\prime}$
Then
$\mathcal{T}(M)=\{z\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}.$
The substitution distributes: $M[V/x]=\mathbf{let}\;z=V_{1}[V/x]\;V_{2}[V/x]\;\mathbf{in}\;M^{\prime}[V/x]$. By induction on $V_{1}$, $V_{2}$, and $M^{\prime}$, the properties under each subtree ($z$ and $\mathrm{result}$) coincide, since references to $x$ inside each subtree resolve to $\mathcal{T}(V)$ via the same $\mathtt{this}$ mechanism.
Inductive case: $M=V_{1}\;V_{2}$ (tail call)
Analogous to the $\mathbf{let}$-binding case with $\mathrm{tail}$ in place of $z$. ∎
D.5. Convergence Preservation
The convergence criterion (Definition 4.1) follows the $\mathrm{result}$ chain from the root. Each step in this chain corresponds to one head reduction step in the $\lambda$-calculus. We make this precise.
Lemma D.3 (Result-step).
Let $(\lambda x.\,M)\;V$ be a tail call (a $\beta$-redex). Then for every path $p$ and label $\ell$:
$\ell\in\mathtt{properties}\!\bigl(\,\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p\,\bigr)\text{ in }\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)\;\Longleftrightarrow\;\ell\in\mathtt{properties}\!\bigl(\,\text{root}\mathbin{\triangleright}p\,\bigr)\text{ in }\mathcal{T}(M[V/x])$
That is, following one $\mathrm{result}$ projection in the tail call’s mixin tree yields the same properties as the root of the reduct’s mixin tree.
Proof.
The translation gives:
$\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)=\{\_\mathrm{tail}\mapsto\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\},\;\mathrm{result}\mapsto\_\mathrm{tail}.\mathrm{result}\}$
The $\mathrm{result}$ label at the root is defined as $\mathrm{tail}.\mathrm{result}$, so $\text{root}\mathbin{\triangleright}\mathrm{result}$ resolves to the $\mathrm{result}$ path inside the inheritance $\mathrm{tail}=\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$. This inheritance is exactly the inherited tree $C$ of Lemma D.2, and $\mathrm{tail}\mathbin{\triangleright}\mathrm{result}$ is $C\mathbin{\triangleright}\mathrm{result}$. By Lemma D.2, the properties under $C\mathbin{\triangleright}\mathrm{result}$ coincide with those of $\mathcal{T}(M[V/x])$.
Let-binding variant. An analogous result holds for $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}$ where $V_{1}=\lambda y.\,M^{\prime\prime}$. The translation gives:
$\mathcal{T}\bigl(\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}\bigr)=\{x\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$
Since $\mathrm{result}\mapsto\mathcal{T}(M^{\prime})$ is a property definition, the subtree at $\text{root}\mathbin{\triangleright}\mathrm{result}$ is $\mathcal{T}(M^{\prime})$. Inside $\mathcal{T}(M^{\prime})$, each reference to $x$ is $x.\mathrm{result}$, which resolves to the $\mathrm{result}$ path inside the $x$ subtree $\{\mathcal{T}(\lambda y.\,M^{\prime\prime}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$. This is the inherited tree $C$ of Lemma D.2, so $x.\mathrm{result}$ has the same properties as $\mathcal{T}(M^{\prime\prime}[V_{2}/y])$. In the $\lambda$-calculus, $M=(\lambda x.\,M^{\prime})(V_{1}\;V_{2})$, and the head reduct is $N=M^{\prime}[(V_{1}\;V_{2})/x]$. In $\mathcal{T}(N)$, each occurrence of $x$ is replaced by $(V_{1}\;V_{2})$, whose result, by the tail-call variant above, also yields $\mathcal{T}(M^{\prime\prime}[V_{2}/y])$. Therefore the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p$ in $\mathcal{T}(M)$ coincide with those at $\text{root}\mathbin{\triangleright}p$ in $\mathcal{T}(N)$ for all $p$. ∎
Theorem D.4 (Convergence preservation).
If $M\to_{h}N$ (head reduction), then $\mathcal{T}(M){\Downarrow}$ if and only if $\mathcal{T}(N){\Downarrow}$.
Proof.
A head reduction step in ANF takes one of two forms: a tail call $(\lambda x.\,M^{\prime})\;V\to_{h}M^{\prime}[V/x]$, or a $\mathbf{let}$-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}\to_{h}M^{\prime}[(V_{1}\;V_{2})/x]$ (since $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}=(\lambda x.\,M^{\prime})(V_{1}\;V_{2})$ in the $\lambda$-calculus). In both cases, Lemma D.3 (tail-call and let-binding variants respectively) shows that the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}^{n}$ in $\mathcal{T}(M)$ coincide with the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}^{n-1}$ in $\mathcal{T}(N)$ for all $n\geq 1$, where $N$ is the head reduct. Hence the abstraction shape appears at depth $n$ in the pre-reduct if and only if it appears at depth n1n-1 in the post-reduct. ∎
D.6. Adequacy
A closed ANF term at the top level is either an abstraction (a value) or a computation (a $\mathbf{let}$-binding or tail call). An abstraction translates to a record whose root immediately has the abstraction shape; a computation’s root has the form $\{x,\mathrm{result}\}$ or $\{\_\mathrm{tail},\mathrm{result}\}$, and one must follow the $\mathrm{result}$ chain to find the eventual value.
Proof of Theorem 4.2.
($\Rightarrow$) Suppose $M\to^{*}_{h}\lambda x.\,M^{\prime}$ in $k$ head reduction steps. We show $\mathcal{T}(M)$ converges at depth $\leq k$ by induction on $k$.
Base case ($k=0$): $M$ is already an abstraction $\lambda x.\,M^{\prime}$. Then $\mathcal{T}(M)=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$, whose root has $\mathtt{defines}=\{\mathrm{argument},\mathrm{result}\}$. These labels appear at stage $F^{1}(\bot)$ of the approximation chain, so $\mathcal{T}(M)$ converges at depth $n=0$.
Inductive step ($k\geq 1$): $M$ is not an abstraction, so $M$ is either a tail call or a $\mathbf{let}$-binding, both of which are $\beta$-redexes. Let $N$ be the head reduct, so $M\to_{h}N\to^{k-1}_{h}\lambda x.\,M^{\prime}$. By Lemma D.3 (tail-call or let-binding variant), the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p$ in $\mathcal{T}(M)$ coincide with those at $\text{root}\mathbin{\triangleright}p$ in $\mathcal{T}(N)$. By the induction hypothesis, $\mathcal{T}(N)$ converges at depth $\leq k-1$, so $\mathcal{T}(M)$ converges at depth $\leq k$.
($\Leftarrow$) Suppose $\mathcal{T}(M){\Downarrow}$ at depth $n$. We show $M$ has a head normal form by induction on $n$.
Base case ($n=0$): the root of $\mathcal{T}(M)$ has abstraction shape $\{\mathrm{argument},\mathrm{result}\}$. Only the abstraction rule $\mathcal{T}(\lambda x.\,M^{\prime})$ produces a root with own property $\mathrm{argument}$. (For tail calls and $\mathbf{let}$-bindings, the root has $\mathtt{defines}=\{\_\mathrm{tail},\mathrm{result}\}$ or $\{x,\mathrm{result}\}$, and $\mathrm{argument}$ cannot appear via $\mathtt{supers}$ because $\mathtt{inherits}=\varnothing$ at the root, giving $\mathtt{bases}(\text{root})=\varnothing$.) Therefore $M$ is an abstraction, hence already in head normal form.
Inductive step ($n\geq 1$): the abstraction shape appears at depth $n$ but not at depth 0. By the base case argument, $M$ is not an abstraction, so $M$ is either a tail call $V_{1}\;V_{2}$ or a $\mathbf{let}$-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}$. In both cases, $M$ is a $\beta$-redex (since $V_{1}=\lambda x.\,M^{\prime}$ or $V_{1}=\lambda y.\,M^{\prime\prime}$ respectively, the only closed ANF value). By Lemma D.3 (tail-call or let-binding variant respectively), the properties at depth n1n-1 in $\mathcal{T}(N)$ match those at depth $n$ in $\mathcal{T}(M)$, where $N$ is the head reduct. So $\mathcal{T}(N)$ converges at depth n1n-1. By the induction hypothesis, $N$ has a head normal form, and since $M\to_{h}N$, so does $M$. ∎
D.7. Full Abstraction
Adequacy relates a single term’s head normal form to its inheritance-convergence. Full abstraction lifts this to an equivalence between terms: the translation $\mathcal{T}$ preserves and reflects observational equivalence.
Definition D.5 ($\lambda$-contextual equivalence under $\mathcal{T}$).
For closed $\lambda$-terms $M$ and $N$, write $\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)$ if for every closing $\lambda$-calculus context $C[\cdot]$:
$\mathcal{T}(C[M]){\Downarrow}\;\Longleftrightarrow\;\mathcal{T}(C[N]){\Downarrow}$
The subscript $\lambda$ emphasizes that the quantification ranges over $\lambda$-calculus contexts, not over all inheritance-calculus contexts. This definition observes only convergence (a single bit), but quantifying over all $\lambda$-contexts makes it a fine-grained equivalence: different contexts can supply different arguments and project different results, probing every aspect of the term’s behavior within the $\lambda$-definable fragment.
Proof of Theorem 4.3.
The translation $\mathcal{T}$ is compositional: for every $\lambda$-calculus context $C[\cdot]$, the mixin tree $\mathcal{T}(C[M])$ is determined by the mixin tree $\mathcal{T}(M)$ and the translation of the context. This means $\mathcal{T}$ preserves the context structure needed for the argument.
($\Leftarrow$) Suppose $\mathrm{BT}(M)=\mathrm{BT}(N)$. Böhm tree equivalence is a congruence (Barendregt, 1984), so for every context $C[\cdot]$, $\mathrm{BT}(C[M])=\mathrm{BT}(C[N])$. In particular, $C[M]$ has a head normal form iff $C[N]$ does. By Theorem 4.2, $\mathcal{T}(C[M]){\Downarrow}$ iff $\mathcal{T}(C[N]){\Downarrow}$. Hence $\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)$.
($\Rightarrow$) Suppose $\mathrm{BT}(M)\neq\mathrm{BT}(N)$. Böhm tree equivalence coincides with observational equivalence for the lazy $\lambda$-calculus (Barendregt, 1984; Abramsky and Ong, 1993): there exists a context $C[\cdot]$ such that $C[M]$ converges and $C[N]$ diverges (or vice versa). By Theorem 4.2, $\mathcal{T}(C[M]){\Downarrow}$ and $\mathcal{T}(C[N]){\Uparrow}$. Hence $\mathcal{T}(M)\not\approx_{\lambda}\mathcal{T}(N)$. ∎
The proof rests on two pillars: Adequacy (Theorem 4.2), which is our contribution, and the classical result that Böhm tree equivalence equals observational equivalence for the lazy $\lambda$-calculus (Barendregt, 1984; Abramsky and Ong, 1993). The mixin tree semantics inherits full abstraction from the $\lambda$-calculus without needing to define or compare absolute paths inside mixin trees.