### v7

# A Calculus of Inheritance

Abstract.
Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) as the foundation of functional programming, inheritance-calculus uses three primitives (record, definition, inheritance) as the foundation of declarative programming. By unifying modules, classes, objects, methods, fields, and locals under a single record abstraction, the calculus models inheritance simply as set union. Consequently, composition is inherently commutative, idempotent, and associative, structurally eliminating the multiple-inheritance linearization problem. Its semantics is first-order (van Emden and Kowalski, 1976; Reynolds, 1972; Aczel, 1977), denotational, and computable by tabling (Tamaki and Sato, 1986), even for cyclic inheritance hierarchies. These three properties extend to the $\lambda$-calculus, since Böhm tree equivalence (Barendregt, 1984) is fully abstract for the first-iteration approximation of a sublanguage of inheritance-calculus. As a corollary, this establishes a convergence hierarchy $\text{eager}\subsetneq\text{lazy}\subsetneq\text{fixpoint}$ among $\lambda$-calculi sharing the same $\lambda$-syntax.
Inheritance-calculus is distilled from MIXINv2, a practical implementation in which the same code acts as different function colors (Nystrom, 2015); ordinary arithmetic yields the relational semantics of logic programming (van Emden and Kowalski, 1976); $\mathtt{this}$ resolves to multiple targets; and programs are immune to nonextensibility in the sense of the Expression Problem (Wadler, 1998). This makes inheritance-calculus strictly more expressive than the $\lambda$-calculus in both common sense and Felleisen’s sense (Felleisen, 1991).
declarative programming, inheritance, fixpoint semantics, self-referential records, expression problem, first-order semantics, $\lambda$-calculus, Böhm trees, configuration languages
This work was conducted independently prior to the author’s employment at Figure AI.
$c$s: Theory of computation Program semanticsccs: Theory of computation Object oriented constructsccs: Theory of computation Functional constructsccs: Theory of computation Program schemesccs: Software and its engineering Formal language definitionsccs: Software and its engineering Extensible languagesccs: Software and its engineering Object oriented languages
1. Introduction
Declarative and configuration languages are ubiquitous in modern software engineering. Systems such as NixOS modules (Dolstra et al., 2010; NixOS Contributors, ), Jsonnet (Cunningham, 2014), Hydra (Yadan and Droppo, 2023), CUE (van Lohuizen, 2019), Dhall (Gonzalez, 2017), Kustomize (Kubernetes SIG CLI, 2018), and JSON Patch (Bryan et al., 2013) all provide mechanisms for composing structured data through inheritance or merging. Among these, the NixOS module system stands out: its recursive attribute set merging with fixpoint semantics and deferred modules (NixOS Contributors, ) allows independently authored modules to extend the system configuration without modifying existing modules; in practice, the ecosystem has not been troubled by the multi-dimensional extension difficulty analogous to the Expression Problem (Wadler, 1998). [^1] As a rough proxy for ecosystem scale: according to Repology (Repology Contributors, 2025), nixpkgs contains over 113,000 packages, nearly three times the count of Debian or Ubuntu, with a comparable number of contributors. Not every package exercises the module system’s extensibility, but the same caveat applies to other distributions. The mechanism has been adopted well beyond NixOS itself: Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), disko (Lassulus and nix-community Contributors, 2022), flake-parts (Hensing, 2022), dream2nix (Hauer, 2021), devenv (Kožar, 2022), KubeNix (Xtruder and Hall, 2022), and nixidy (Ingólfsson, 2024) use it to manage user environments, macOS configuration, disk partitioning, flake structure, multi-language packaging, developer environments, and Kubernetes clusters respectively. Yet no computational theory captures the extensibility that this mechanism exhibits.
The $\lambda$-calculus serves as the foundational computational model for functional programming. No analogous calculus exists for declarative programming. This gap matters because the two paradigms differ in fundamental ways. Configuration languages are inherently declarative: their values are immutable and they have no first-class functions. Note that a function body is closed: its internal structure cannot be extended without modifying the definition, whereas the configuration languages above all compose by extending existing structure from outside. At first glance, Turing completeness appears incompatible with these constraints. The three classical models of computation each violate at least one of them:
Model Turing complete Immutable No $\lambda$
Turing Machine $\checkmark$ $\times$ $\checkmark$
$\lambda$-calculus $\checkmark$ $\checkmark$ $\times$
RAM Machine $\checkmark$ $\times$ $\checkmark$
Table 1. Classical computational models and the declarative gap.
The Turing Machine and RAM Machine require mutable state; the $\lambda$-calculus requires first-class functions. A computational model that is Turing complete, immutable, and function-free would sit in the empty cell of Table 1. The NixOS module system already suggests that such a model exists: its inheritance-based composition over recursive records, without explicit functions, is expressive enough to configure entire operating systems.
Most configuration languages operate on finite, well-founded structures (initial algebras in the sense of universal algebra). Nix is an exception: although designed specifically as a configuration language for NixOS (Dolstra, 2006), it is a complete programming language whose values are lazily observable and possibly infinite. The nixpkgs package collection (over 100,000 packages) behaves as a lazily observed structure, allowing any single package to be evaluated without materializing the entire set. Guided by this observation, we set out to reduce the NixOS module system to a minimal set of primitives. The reduction produced MIXIN, an early implementation in Nix, and subsequently MIXINv2 (Yang, 2026), now implemented in Python: a declarative programming language with only three constructs (record literals, definitions, and inheritance) and no functions or let-bindings. [^2] The executable MIXINv2 implementation (Yang, 2026) contains every example in this paper.
In developing real programs in MIXINv2, including boolean logic, natural number arithmetic, and a web application, several phenomena arose from off-label use of the adopted design patterns. Section 3 shows how the three constructs give rise to a single abstraction that simultaneously serves as module, class, object, method, field, and local binding. The linearization problem of mixin-based systems (Bracha and Cook, 1990; Barrett et al., 1996) did not arise: inheritance is inherently commutative, idempotent, and associative, and $\mathtt{this}$ naturally resolves to multiple targets rather than one, where prior systems such as Scala and the NixOS module system reject the multi-target situation. Section 5 traces how boolean logic and arithmetic, implemented as separate files, extend in both dimensions of the Expression Problem (Wadler, 1998) without any dedicated mechanism, and how ordinary arithmetic yields the relational semantics of logic programming (van Emden and Kowalski, 1976). Section 6.1 discusses these and further emergent phenomena, including function color blindness (Nystrom, 2015).
This paper distills these observations into inheritance-calculus, a minimal computational model for declarative programming. Section 3 defines the semantics via six mutually recursive first-order functions on paths and sets of paths, with no function spaces, closures, or higher-order constructs; the functions are computable by tabling (Tamaki and Sato, 1986). Section 4 shows that these functions extend to the $\lambda$-calculus, embedding it. Remark B.5 derives a convergence hierarchy among eager, lazy, and fixpoint $\lambda$-calculi sharing the same syntax as a corollary. Section 4.2 establishes that the lazy $\lambda$-calculus cannot macro-express the inheritance constructors, so inheritance-calculus is strictly more expressive in Felleisen’s sense (Felleisen, 1991).
2. Syntax
Let $e$ denote an expression, $\ell$ a label serving as a property name, and $k$ a nonnegative integer serving as the length of a downward projection.
$e\quad::=\quad$ $\{c_{1},\;\ldots,\;c_{m}\}$ $\text{(}m\geq 0\text{)}$
$\mid\quad$ $c$ $\text{(sugar for }\{c\}\text{)}$
$c\quad::=\quad$ $\ell\mapsto e$ (definition)
$\mid\quad$ $(\ell_{\mathrm{up}}.\mathbf{this}\mid\uparrow^{\!n}).\;\ell_{\mathrm{down},1}.\ell_{\mathrm{down},2}\ldots\ell_{\mathrm{down},k}$ (inheritance: qualifier + downward path)
$\mid\quad$ $\ell_{\mathrm{up}}.\mathbf{this}\mid\uparrow^{\!n}$ (inheritance: qualifier only)
$\mid\quad$ $\ell_{\mathrm{down},1}.\ell_{\mathrm{down},2}\ldots\ell_{\mathrm{down},k}$ (inheritance: downward path only)
An expression is a set of elements enclosed in braces. Each element is either a definition $\ell\mapsto e$ or an inheritance source. Because expressions are sets, element order is irrelevant and duplicates have no effect; inheritance is inherently commutative, idempotent, and associative.
The $\mapsto$ in $\ell\mapsto e$ defines a property named $\ell$ with value $e$. MIXINv2 also provides scalar types through a foreign-function interface (FFI); inheritance-calculus is what remains after removing the FFI. None of the examples in this paper uses the FFI; no scalar values are needed in the examples.
Inheritance
An inheritance is written as a reference that navigates the scope hierarchy upward and then projects properties downward. The upward navigation admits two equivalent notations:
Named form:
$\ell_{\mathrm{up}}.\mathbf{this}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$, analogous to Java’s Outer.this.field. Each $\{\ldots\}$ record has a label in its enclosing scope; $\ell_{\mathrm{up}}$ names the target enclosing scope.
Indexed form:
$\uparrow^{\!n}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$, where $n\geq 0$ is a de Bruijn index (de Bruijn, 1972): $n=0$ refers to the reference’s enclosing scope, and each increment of $n$ moves one scope level further out.
Both forms retrieve the target scope, that is, the fully inherited record after all inheritance has been applied, and then project properties $\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$ from it. The named form desugars to the indexed form during parsing (Section 3).
A new scope level is created at each $\{\ldots\}$ record literal. [^3] We use record for the syntactic construct, scope for the name-resolution context it creates, and mixin for its role as a composable unit; in inheritance-calculus these are three views of the same entity. The scope contains all properties of that mixin, including those inherited via inheritance sources. For example, $\{a\mapsto\{\},\;b\mapsto\{\}\}$ is a single mixin whose scope contains both $a$ and $b$; so is $\{r,\;a\mapsto\{\}\}$ where $r$ is a reference to a record containing $b$. Inheritance sources do not create additional scope levels.
  • $\mathrm{Foo}.\mathbf{this}$ (equivalently, $\uparrow^{\!n}$ where $\mathrm{Foo}$ is $n$ steps above the enclosing scope) retrieves the record named $\mathrm{Foo}$.
  • $\mathrm{Foo}.\mathbf{this}.\ell$, or equivalently, $\uparrow^{\!n}.\ell$, projects property $\ell$ from $\mathrm{Foo}$.
  • When $k=0$, the reference inherits the entire enclosing scope, which may produce an infinitely deep tree.
When the qualifier is unambiguous, it may be omitted: the shorthand $\ell_{\mathrm{down},1}.\ell_{\mathrm{down},2}\ldots\ell_{\mathrm{down},k}$ ($k\geq 1$) resolves $\ell_{\mathrm{down},1}$ against the innermost enclosing scope whose $\mathtt{defines}$ contains it, and desugars to the corresponding indexed form. Since this uses $\mathtt{defines}$, the shorthand can only access properties defined in record literals, not those inherited through inheritance; one of the two qualified forms is required for inherited properties or to bypass variable shadowing.
3. Mixin Trees
In object-oriented programming, a mixin (Bracha and Cook, 1990) is a fragment of behavior that can be composed onto a class via inheritance. Inheritance-calculus has a single abstraction, the deep-mergeable mixin, [^4] Where no ambiguity arises, we write simply “mixin.” The distinction from the Bracha–Cook mixin (Bracha and Cook, 1990) is detailed in Section 7. which can simultaneously serve as module, class, object, method, field, and local binding, because there are no opaque methods: every value is a transparent record, and inheritance always performs a deep merge of subtrees. We call the resulting structure a mixin tree. Its semantics is purely observational: given an inheritance-calculus program, one queries whether a label $\ell$ is a property at a given position $p$.
This section defines the semantics via six mutually recursive first-order [^5] First-order in three senses simultaneously: the semantic domain contains only paths (finite sequences of labels) and sets of paths, with no function spaces, closures, or strategies (van Emden and Kowalski, 1976); the six equations are first-order functions whose inputs and outputs are paths and sets of paths, never functions (Reynolds, 1972); and the judgments are first-order inductive definitions (Aczel, 1977) over paths and labels, where quantifiers range over paths and labels, never over functions or predicates. functions on paths. The six functions map object-oriented concepts to set-theoretic operations: $\mathtt{properties}$ and $\mathtt{supers}$ generalize member lookup and inheritance chains, $\mathtt{overrides}$ and $\mathtt{bases}$ capture method override and base classes, and $\mathtt{resolve}$ and $\mathtt{this}$ handle name resolution and qualified $\mathtt{this}$ resolution.
3.1. Definitions
Path
A $\mathtt{Path}$ is a sequence of labels $(\ell_{1},\ell_{2},\ldots,\ell_{n})$ that identifies a position in the tree. We write $p$ for a path. The root path is the empty sequence $()$. Given a path $p$ and a label $\ell$, $p\mathbin{\triangleright}\ell$ is the sequence $p$ extended with $\ell$. For a nonroot path, the parent path $\mathtt{init}(p)$ is $p$ with its last element removed, and the final label $\mathtt{last}(p)$ is the last element of $p$. For convenience, we sometimes write $\ell$ for $\mathtt{last}(p)$ when the path is clear from context. With paths in hand, we can state what the AST provides at each path.
AST
An expression $e$ (Section 2) is parsed into an AST. The AST provides two primitive functions at each path $p$:
  • $\mathtt{defines}(p)$: the set of labels $\ell$ for which $p$ contains a definition $\ell\mapsto e$.
  • $\mathtt{inherits}(p)$: the set of reference pairs $(n,\;\ell_{\mathrm{down},*})$, where $n$ is the de Bruijn index and $\ell_{\mathrm{down},*}$ is the list of downward projections.
During parsing, all three syntactic forms are resolved to de Bruijn index pairs $(n,\;\ell_{\mathrm{down},*})$. The indexed form $\uparrow^{\!n}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$ already carries the de Bruijn index $n$ directly; no resolution is needed. A named reference $\ell_{\mathrm{up}}.\mathbf{this}.\ell_{\mathrm{down},1}\ldots\ell_{\mathrm{down},k}$ at path $p$ is resolved by finding the last occurrence of $\ell_{\mathrm{up}}$ among the labels of $p$. Let $p_{\mathrm{target}}$ be the prefix of $p$ up to and including that occurrence. The de Bruijn index is $n=|p|-|p_{\mathrm{target}}|-1$ and the projections are $\ell_{\mathrm{down},*}=(\ell_{\mathrm{down},1},\ldots,\ell_{\mathrm{down},k})$. A lexical reference $\ell_{\mathrm{down},1}.\ell_{\mathrm{down},2}\ldots\ell_{\mathrm{down},k}$ at path $p$ is resolved by finding the nearest prefix $p^{\prime}$ of $p$ such that $\ell_{\mathrm{down},1}\in\mathtt{defines}(p^{\prime})$. The de Bruijn index is $n=|p|-|p^{\prime}|-1$ and the projections are $\ell_{\mathrm{down},*}=(\ell_{\mathrm{down},1},\ldots,\ell_{\mathrm{down},k})$. All three forms produce the same representation; the semantic functions below operate only on de Bruijn index pairs.
Parsing populates $\mathtt{defines}$ and $\mathtt{inherits}$ from each element of a record. A definition $\ell\mapsto e$ contributes $\ell$ to $\mathtt{defines}(p)$. A reference contributes $(n,\;\ell_{\mathrm{down},*})$ to $\mathtt{inherits}(p)$. Both $\mathtt{defines}$ and $\mathtt{inherits}$ are pure data; they are not functions of runtime state. Given these two primitives, we can now answer the question posed at the beginning of this section.
Properties
The properties of a path are its own properties together with those inherited from all supers:
(1) $\mathtt{properties}(p)=\bigl\{\;\ell\;\big|\;(\_,\;p_{\mathrm{override}})\in\mathtt{supers}(p),\;\ell\in\mathtt{defines}(p_{\mathrm{override}})\;\bigr\}$
The remainder of this section defines $\mathtt{supers}$, the transitive inheritance closure, and its dependencies.
Supers
Intuitively, $\mathtt{supers}(p)$ collects every path that $p$ inherits from: the $\mathtt{overrides}$ of $p$ itself, plus the $\mathtt{overrides}$ of each direct base $\mathtt{bases}$ of $p$, and so on transitively. Each result is paired with the inheritance-site context $\mathtt{init}(p_{\mathrm{base}})$ through which it is reached. This provenance is needed by qualified $\mathtt{this}$ resolution, that is, the $\mathtt{this}$ function defined below, to map a definition-site mixin back to the inheritance-site paths that incorporate it:
(2) $\mathtt{supers}(p)=\bigl\{\;(\mathtt{init}(p_{\mathrm{base}}),\;p_{\mathrm{override}})\;\big|\;p_{\mathrm{base}}\in\mathtt{bases}^{*}(p),\;p_{\mathrm{override}}\in\mathtt{overrides}(p_{\mathrm{base}})\;\bigr\}$
Here $\mathtt{bases}^{*}$ denotes the reflexive-transitive closure of $\mathtt{bases}$. The $\mathtt{supers}$ formula depends on two functions: $\mathtt{overrides}$ and the one-hop reference targets $\mathtt{bases}$. We define $\mathtt{overrides}$ first.
Overrides
Since a mixin can play the role of a method, a child scope should be able to find it through the parent’s label even if the child does not redefine that label; this is method override. Concretely, inheritance can introduce multiple definitions of the same label at the same scope level; $\mathtt{overrides}(p)$ collects all such paths that share the same identity as $p$, so that their definitions merge rather than shadow each other, enabling deep merge. The overrides of $p$ include $p$ itself and $p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)$ for every branch $p_{\mathrm{branch}}$ of $\mathtt{init}(p)$ that also defines $\mathtt{last}(p)$:
(3) $\mathtt{overrides}(p)=\begin{cases}\{p\}&\text{if }p=()\\[6.0pt] \{p\}\;\cup\;\left\{\;p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)\;\left|\;\begin{aligned} &(\_,\;p_{\mathrm{branch}})\in\mathtt{supers}(\mathtt{init}(p)),\\ &\text{s.t.}\;\mathtt{last}(p)\in\mathtt{defines}(p_{\mathrm{branch}})\end{aligned}\right.\right\}&\text{if }p\neq()\end{cases}$
It remains to define $\mathtt{bases}$, the other dependency of $\mathtt{supers}$.
Bases
Intuitively, $\mathtt{bases}(p)$ are the paths that $p$ directly inherits from via references, analogous to the direct base classes in object-oriented languages. Concretely, $\mathtt{bases}$ resolves every reference in $p$’s $\mathtt{overrides}$ one step:
[…]
The $\mathtt{bases}$ formula calls the reference resolution function $\mathtt{resolve}$, which we define next.
Reference resolution
Intuitively, $\mathtt{resolve}$ turns a syntactic reference into the set of paths it points to in the fully inherited tree. It takes an inheritance-site path $p_{\mathrm{site}}$, a definition-site path $p_{\mathrm{def}}$, a de Bruijn index $n$, and downward projections $\ell_{\mathrm{down},*}$. Resolution proceeds in two phases: $\mathtt{this}$ performs $n$ upward steps starting from the enclosing scope $\mathtt{init}(p_{\mathrm{def}})$, mapping the definition-site path to inheritance-site paths; then the downward projections $\ell_{\mathrm{down},*}$ are appended:
(5) $\mathtt{resolve}(p_{\mathrm{site}},\;p_{\mathrm{def}},\;n,\;\ell_{\mathrm{down},*})=\\ \bigl\{\;p_{\mathrm{current}}\mathbin{\triangleright}\ell_{\mathrm{down},1}\mathbin{\triangleright}\cdots\mathbin{\triangleright}\ell_{\mathrm{down},k}\;\big|\;p_{\mathrm{current}}\in\mathtt{this}(\{p_{\mathrm{site}}\},\;\mathtt{init}(p_{\mathrm{def}}),\;n)\;\bigr\}$
When multiple routes exist due to multi-target $\mathtt{this}$ resolution, they may yield different target paths; all are collected. We now define $\mathtt{this}$.
Qualified $\mathtt{this}$ resolution
$\mathtt{this}$ answers the question: “in the fully inherited tree, where does the definition-site scope $p_{\mathrm{def}}$ actually live?” Each step finds, among the supers of every path in the frontier $S$, those whose override component matches $p_{\mathrm{def}}$, and collects the corresponding inheritance-site paths as the new frontier. After $n$ steps the frontier contains the answer:
(6) $\mathtt{this}(S,\;p_{\mathrm{def}},\;n)=\begin{cases}S&\text{if }n=0\\[6.0pt] \mathtt{this}\!\left(\left\{\;p_{\mathrm{site}}\;\left|\;\begin{aligned} &p_{\mathrm{current}}\in S,\\ &(p_{\mathrm{site}},\;p_{\mathrm{override}})\in\mathtt{supers}(p_{\mathrm{current}}),\\ &\text{s.t.}\;p_{\mathrm{override}}=p_{\mathrm{def}}\end{aligned}\right.\right\},\;\mathtt{init}(p_{\mathrm{def}}),\;n-1\right)&\text{if }n>0\end{cases}$
At each step $\mathtt{init}$ shortens $p_{\mathrm{def}}$ by one label and $n$ decreases by one. Since $n$ is a nonnegative integer, the recursion terminates. [^6] If no matching pair $(p_{\mathrm{site}},\;p_{\mathrm{override}})$ exists at some step, the frontier becomes the empty set, and $\mathtt{resolve}$ returns $\varnothing$: the reference has no target; no properties are inherited. This is the most self-consistent semantics for the untyped calculus. MIXINv2 (Yang, 2026) is stricter: its resolver rejects references whose frontier becomes empty, reporting a resolution error at compile time.
This completes the chain of definitions needed to compute $\mathtt{properties}(p)$.
3.2. Recursive Evaluation
The six equations (1)–(6) define $\mathtt{properties}(p)$ by mutual recursion: $\ell\in\mathtt{properties}(p)$ holds if and only if it can be established by a finite chain of applications of these functions. The functions are the interpreter: the observer drives the computation by choosing which path to inspect, and each equation unfolds on demand. Because the functions are pure and order-independent, the reflexive-transitive closure $\mathtt{bases}^{*}$ in (2) may be explored breadth-first or depth-first without affecting the result; paths may be interned so that structural equality reduces to pointer equality; and the six functions form a memoizable dynamic-programming recurrence keyed by paths.
Operationally, evaluation uses tabled evaluation (Tamaki and Sato, 1986): each query is memoised on first encounter, and re-entrant calls, which correspond to cycles in the dependency graph, return the current accumulator rather than diverging. Denotationally, the six equations define a monotone immediate consequence operator $T_{P}$ (van Emden and Kowalski, 1976) whose right-hand sides use only $\cup$, $\in$, equality, and positive set comprehension, with no negation, set difference, or complement. By the Knaster–Tarski theorem (Tarski, 1955), $T_{P}$ has a unique least fixed point $\mathrm{lfp}(T_{P})$, which serves as the denotational semantics. This denotational characterization formalizes the observational semantics introduced above: the observer-driven queries are computed by the same fixed-point equations. Whether evaluation terminates depends on the program. Recall that $\mathtt{supers}$ (equation 2) takes the reflexive-transitive closure of $\mathtt{bases}$ (equation 4); the resulting graph is the inheritance hierarchy. Since each inheritance source occupies a distinct node in the mixin tree, this hierarchy is directly readable from the tree structure. Programs whose inheritance hierarchy is acyclic terminate in one pass. When the hierarchy contains cycles, evaluation terminates when each query has a finite answer domain: the ascending chain condition on the finite lattice guarantees convergence (Bancilhon and Ramakrishnan, 1986), as in Datalog (Appendix F). When the hierarchy is infinite, evaluation may diverge; this is inherent to Turing completeness. Appendix A proves that tabled evaluation never returns a wrong answer and formalizes these termination conditions.
4. Embedding the $\lambda$-Calculus
The following translation maps the $\lambda$-calculus in A-normal form (ANF) (Flanagan et al., 1993) to inheritance-calculus. In ANF, every intermediate result is bound to a name and each expression contains at most one application. We extend the standard ANF grammar with a value binding $\mathbf{let}\;x=V\;\mathbf{in}\;M$, which gives a name to a value that is not an application result: [^7] Standard ANF allows values inline as operands; the value-binding production names them so that every subexpression of the translated program is accessible via a lexical reference (Appendix C).
[…]
Every $\lambda$-term can be mechanically converted to ANF by naming intermediate results. [^8] This can optionally include naming values as well; since every ANF term is also a $\lambda$-term, ANF is neither a restriction nor an extension of the $\lambda$-calculus. The ANF structure matches inheritance-calculus naturally: each $\mathbf{let}$-binding becomes a named record property, and each application is wrapped and projected via $\mathrm{result}.$
Let $\mathcal{T}$ denote the translation function. Each $\lambda$-abstraction introduces one scope level in the translated mixin tree.
$x$ ($\lambda$-bound, index $n^{\prime}$)    $\longrightarrow$ $\uparrow^{\!n^{\prime}}.\mathrm{argument}$
$x$ (let-bound)    $\longrightarrow$ $x.\mathrm{result}$
$\lambda x.\,M$    $\longrightarrow$ $\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
$\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M$    $\longrightarrow$ $\{x\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},$
   $\longrightarrow$ $\mathrm{result}\mapsto\mathcal{T}(M)\}$
$\mathbf{let}\;x=V\;\mathbf{in}\;M$    $\longrightarrow$ $\{x\mapsto\{\mathrm{result}\mapsto\mathcal{T}(V)\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
$V_{1}\;V_{2}$ (tail call)    $\longrightarrow$ $\{\mathrm{tailCall}\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},$
   $\longrightarrow$ $\mathrm{result}\mapsto\mathrm{tailCall}.\mathrm{result}\}$
[…]
An abstraction $\lambda x.\,M$ translates to a record with own properties $\{\mathrm{argument},\mathrm{result}\}$; we call this the abstraction shape. In the application $\mathbf{let}$-binding rule, the name $x$ binds to the inherited record $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$, and $\mathcal{T}(M)$ may reference $x.\mathrm{result}$ to obtain the value of the application. In the value $\mathbf{let}$-binding rule, $x$ wraps $\mathcal{T}(V)$ under a $\mathrm{result}$ child so that the uniform accessor $x.\mathrm{result}$ yields $\mathcal{T}(V)$, the translated value itself. This extra wrapper is semantically inert but proof-relevant: it lets the same $\mathrm{result}$-following argument used for applications extend uniformly to value $\mathbf{let}$-bindings in the adequacy proof (Appendix B). In the tail-call rule, the fresh label $\mathrm{tailCall}$ serves the same purpose, and $\mathrm{result}$ projects the answer. In the application and tail-call cases, the inheritance that performs the application is encapsulated behind a named property, so its internal structure (the $\mathrm{argument}$ and $\mathrm{result}$ labels of the callee) does not leak to the enclosing scope.
The six rules above are a complete translation from the $\lambda$-calculus (via ANF) to inheritance-calculus: every ANF term has an image, and the construction is compositional. Since the $\lambda$-calculus is Turing complete, so is inheritance-calculus. [^10] MIXINv2 separates symbol resolution (which resolves lexical references and qualified $\mathtt{this}$ references) from runtime mixin evaluation, but symbol resolution is performed just-in-time per symbol rather than ahead of time, so it does not provide static type checking in the traditional sense. Nevertheless, MIXINv2’s resolution phase is stricter than the untyped inheritance-calculus. All examples in this paper are ported from MIXINv2 and conform to both semantics. Elements added to satisfy the resolver are harmless for the algorithms presented here and may aid readability.
4.1. Böhm Tree Correspondence
The translation $\mathcal{T}$ embeds the $\lambda$-calculus into a sublanguage of inheritance-calculus. Intuitively, a mixin tree augments the AST with additional semantic structure; for translated $\lambda$-terms, the projection of this additional structure is isomorphic to the Böhm tree. A mixin tree $T$ converges, written $T{\Downarrow}$, if following finitely many $\mathrm{result}$ projections from the root reaches a node with both $\mathrm{argument}$ and $\mathrm{result}$ among its properties, that is, the abstraction shape produced by $\mathcal{T}(\lambda x.\,M)$. The translation preserves and reflects convergence (adequacy). Write $\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)$ when $\mathcal{T}(C[M]){\Downarrow}\Leftrightarrow\mathcal{T}(C[N]){\Downarrow}$ for every closing $\lambda$-calculus context $C[\cdot]$; then $\approx_{\lambda}$ coincides with Böhm tree equivalence (Barendregt, 1984). Formal statements and proofs are in Appendix B. [^11] Appendix E applies the same first-order technique directly on ANF to give the $\lambda$-calculus a self-contained fixpoint semantics, without passing through inheritance-calculus. Note that $\approx_{\lambda}$ quantifies over $\lambda$-calculus contexts only; inheritance-calculus contexts are strictly more discriminating (Section 4.2). Böhm tree equivalence is fully abstract for the first-iteration approximation $T_{P}\!\uparrow\!1$ of the equations. The full $\mathrm{lfp}(T_{P})$ is strictly more defined. For example, $\mathbf{let}\;x=x\;\mathbf{in}\;x$ diverges under $\beta$-reduction, but its translation $\{x\mapsto\{\mathrm{result}\mapsto x.\mathrm{result}\},\;\mathrm{result}\mapsto x.\mathrm{result}\}$ stabilizes under $\mathrm{lfp}(T_{P})$: [^12] This simple cycle could also be resolved by a fixed-point combinator such as $Y$. The $\mathrm{lfp}(T_{P})$ semantics is more general: it computes fixed points at the container level, e.g. finding a set $S$ such that $S=F(S)$. Recursive Datalog on cyclic graphs (Appendix F, Section F.8) requires exactly this, and neither $Y$ nor Haskell’s mfix (RecursiveDo) can express it. tabled evaluation detects the cycle and returns $\mathtt{properties}(\mathrm{root})=\{x,\mathrm{result}\}$ with no abstraction shape, reflecting an undefined value rather than diverging (Remark B.5). [^13] We call the sublanguage of inheritance-calculus corresponding to $\lambda$-terms, evaluated under $\mathrm{lfp}(T_{P})$ rather than $T_{P}\!\uparrow\!1$, the fixpoint $\lambda$-calculus; Remark B.5 elaborates.
The relationship between $\approx_{\lambda}$ and full inheritance-contextual equivalence is clarified by the expressive asymmetry of the next section: contexts in the language universe $\mathscr{U}$ (Definition C.3) that mix $\lambda$-calculus and inheritance-calculus constructors can separate $\lambda$-terms that are Böhm-tree equivalent, so ${\approx_{\lambda}}\subsetneq{(\approx_{\mathscr{U}}|_{L_{0}})}$ where $L_{0}$ is the lazy $\lambda$-calculus. This gap is precisely the source of inheritance-calculus’s additional expressiveness (Theorem C.8).
4.2. Expressive Asymmetry
Two $\lambda$-terms that are Böhm-tree equivalent, and therefore indistinguishable in the $\lambda$-calculus, can be separated by a context that mixes $\lambda$-calculus and inheritance-calculus constructors, inheriting new observable projections onto an existing definition via open extension. By Felleisen’s expressiveness criterion (Felleisen, 1991), inheritance-calculus is strictly more expressive than the lazy $\lambda$-calculus (Theorem C.9): the $\lambda$-calculus can be macro-expressed in inheritance-calculus via ANF conversion and $\mathcal{T}$ (Lemma C.6, Theorem C.7), but the lazy $\lambda$-calculus cannot macro-express inheritance (Theorem C.8). The formal definitions, constructions, and proofs are given in Appendix C; the separating construction is itself a miniature instance of the Expression Problem, adding a new observable projection to an existing definition without modifying it.
The underlying mechanism is open extension: mixin definitions compose via open extension, so existing definitions gain new observable properties without modification. In the $\lambda$-calculus, a function’s result is closed: there is no way to add new observable projections to an already-computed value without modifying the function’s definition.
5. Case Study: Nat Arithmetic
We implemented boolean logic and natural number arithmetic, both unary and binary, as MIXINv2 files (Yang, 2026). [^14] The examples in this section are rendered in inheritance-calculus notation; the executable MIXINv2 sources are available as supplementary material. This section traces the unary Nat case in detail. The implementation follows a declarative object-oriented style: each concern is one top-level mixin; UpperCamelCase names serve as data constructors and operations; lowerCamelCase names serve as fields and parameters; and since there are no methods, all names are nouns or adjectives: $\mathrm{Plus}$ is a structure, not an action. Operations expose their result through a projection field (e.g., $\mathrm{sum}$).
NatData
The data representation (Zero/Successor) is a Peano encoding, an initial algebra of the functor $F(X)=1+X$. The visitor pattern (NatVisitor, below) adds Church-encoding behavior: a fold (catamorphism) over this algebra. These two concerns are defined in separate files and composed via inheritance (Yang, 2026), which is itself an instance of the Expression Problem. Following this convention, we define a factory containing the abstract product type and two labels that serve as data constructors. A $\mathrm{Zero}$ value inherits $\mathrm{Product}$; a $\mathrm{Successor}$ value inherits $\mathrm{Product}$ and exposes a $\mathrm{predecessor}$ field. The name $\mathrm{NatData}$ is the top-level scope:
$\mathrm{NatData}\mapsto\{\;$ $\mathrm{NatFactory}\mapsto\{\mathrm{Product}\mapsto\{\},$
$\quad\mathrm{Zero}\mapsto\{\mathrm{Product}\},$
$\quad\mathrm{Successor}\mapsto\{\mathrm{Product},\;\mathrm{predecessor}\mapsto\{\mathrm{Product}\}\}\},$
$\mathrm{Nat}\mapsto\{\mathrm{NatFactory},\;\mathrm{NatFactory}.\mathrm{Product}\}\}$
NatPlus
Addition is a separate mixin that inherits $\mathrm{NatData}$ and extends each data constructor with a $\mathrm{Plus}$ operation exposing a $\mathrm{sum}$ field. The base case is $0+m=m$; the recursive case reduces $\mathrm{S}(n_{0})+m$ to $n_{0}+\mathrm{S}(m)$ by delegating to $n_{0}$’s $\mathrm{Plus}$ with an incremented addend:
$\mathrm{NatPlus}\mapsto\{\;$ $\mathrm{NatData},$
$\mathrm{NatFactory}\mapsto\{$
$\quad\mathrm{Product}\mapsto\{\mathrm{Plus}\mapsto\{\mathrm{sum}\mapsto\{\mathrm{Product}\}\}\},$
$\quad\mathrm{Zero}\mapsto\{\mathrm{Plus}\mapsto\{\mathrm{addend}\mapsto\{\mathrm{Product}\},\;\mathrm{sum}\mapsto\{\mathrm{addend}\}\}\},$
$\quad\mathrm{Successor}\mapsto\{\mathrm{Plus}\mapsto\{\mathrm{addend}\mapsto\{\mathrm{Product}\},$
$\qquad\mathrm{\_increasedAddend}\mapsto\{\mathrm{Successor},\;\mathrm{predecessor}\mapsto\{\mathrm{addend}\}\},$
$\qquad\mathrm{\_recursiveAddition}\mapsto\{\mathrm{Successor}.\mathbf{this}.\mathrm{predecessor}.\mathrm{Plus},$
$\qquad\quad\mathrm{addend}\mapsto\{\mathrm{\_increasedAddend}\}\},$
$\qquad\mathrm{sum}\mapsto\{\mathrm{\_recursiveAddition}.\mathrm{sum}\}\}\}\}\}$
NatVisitor
Equality requires case analysis on the data constructor; we use the visitor pattern. $\mathrm{NatVisitor}$ adds an $\mathrm{Acceptance}$ operation to each data constructor, containing a $\mathrm{VisitorMap}$ with the appropriate $\mathrm{ZeroVisitor}$ or $\mathrm{SuccessorVisitor}$ branch, and an $\mathrm{Accepted}$ field that delegates to the matching branch:
$\mathrm{NatVisitor}\mapsto\{\;$ $\mathrm{NatData},$
$\mathrm{NatFactory}\mapsto\{$
$\quad\mathrm{Product}\mapsto\{\mathrm{Acceptance}\mapsto\{\mathrm{Accepted}\mapsto\{\}\}\},$
$\quad\mathrm{Zero}\mapsto\{\mathrm{Acceptance}\mapsto\{\mathrm{VisitorMap}\mapsto\{\mathrm{ZeroVisitor}\mapsto\{\}\},$
$\qquad\mathrm{Accepted}\mapsto\{\mathrm{VisitorMap}.\mathrm{ZeroVisitor}\}\}\},$
$\quad\mathrm{Successor}\mapsto\{\mathrm{Acceptance}\mapsto\{\mathrm{VisitorMap}\mapsto\{\mathrm{SuccessorVisitor}\mapsto\{\}\},$
$\qquad\mathrm{Accepted}\mapsto\{\mathrm{VisitorMap}.\mathrm{SuccessorVisitor}\}\}\}\}\}$
$\mathrm{BooleanData}$ defines the output sort:
$\mathrm{BooleanData}\mapsto\{\;$ $\mathrm{BooleanFactory}\mapsto\{\mathrm{Product}\mapsto\{\},\;\mathrm{True}\mapsto\{\mathrm{Product}\},\;\mathrm{False}\mapsto\{\mathrm{Product}\}\},$
$\mathrm{Boolean}\mapsto\{\mathrm{BooleanFactory},\;\mathrm{BooleanFactory}.\mathrm{Product}\}\}$
NatEquality
To test that $2+3=5$, we needed equality on Nat values. The Cartesian product test (page 5) will apply this same $\mathrm{Equal}$ to a trie union of multiple Nat values without modification. $\mathrm{NatEquality}$ inherits $\mathrm{NatVisitor}$ and $\mathrm{BooleanData}$, and adds an $\mathrm{Equal}$ operation to each data constructor. $\mathrm{Zero}$ is equal only to another $\mathrm{Zero}$; $\mathrm{Successor}$ is equal only to another $\mathrm{Successor}$ with an equal predecessor. Case analysis on $\mathrm{other}$ is done via its $\mathrm{Acceptance}$:
$\mathrm{NatEquality}\mapsto\{\;$ $\mathrm{NatVisitor},\;\mathrm{BooleanData},$
$\mathrm{NatFactory}\mapsto\{$
$\qquad\mathrm{Product}\mapsto\{\mathrm{Equal}\mapsto\{\mathrm{other}\mapsto\{\mathrm{Product}\},$
$\hskip 18.49988pt\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{Boolean}\}\}\},$
$\qquad\mathrm{Zero}\mapsto\{\mathrm{Equal}\mapsto\{\mathrm{other}\mapsto\{\mathrm{Product}\},$
$\hskip 18.49988pt\mathrm{OtherAcceptance}\mapsto\{\mathrm{other}.\mathrm{Acceptance},$
$\hskip 18.49988pt\;\mathrm{VisitorMap}\mapsto\{\mathrm{ZeroVisitor}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{True}\}\},$
$\hskip 18.49988pt\;\qquad\mathrm{SuccessorVisitor}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{False}\}\}\},$
$\hskip 18.49988pt\;\mathrm{Accepted}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{Boolean}\}\}\},$
$\hskip 18.49988pt\mathrm{equal}\mapsto\{\mathrm{OtherAcceptance}.\mathrm{Accepted}.\mathrm{equal}\}\}\},$
$\qquad\mathrm{Successor}\mapsto\{\mathrm{Equal}\mapsto\{\mathrm{other}\mapsto\{\mathrm{Product},\;\mathrm{predecessor}\mapsto\{\mathrm{Product}\}\},$
$\hskip 18.49988pt\mathrm{RecursiveEquality}\mapsto\{\mathrm{Successor}.\mathbf{this}.\mathrm{predecessor}.\mathrm{Equal},$
$\hskip 18.49988pt\;\mathrm{other}\mapsto\{\mathrm{Equal}.\mathbf{this}.\mathrm{other}.\mathrm{predecessor}\}\},$
$\hskip 18.49988pt\mathrm{OtherAcceptance}\mapsto\{\mathrm{other}.\mathrm{Acceptance},$
$\hskip 18.49988pt\;\mathrm{VisitorMap}\mapsto\{\mathrm{ZeroVisitor}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{False}\}\},$
$\hskip 18.49988pt\;\qquad\mathrm{SuccessorVisitor}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{RecursiveEquality}.\mathrm{equal}\}\}\},$
$\hskip 18.49988pt\;\mathrm{Accepted}\mapsto\{\mathrm{equal}\mapsto\{\mathrm{NatEquality}.\mathbf{this}.\mathrm{Boolean}\}\}\},$
$\hskip 18.49988pt\mathrm{equal}\mapsto\{\mathrm{OtherAcceptance}.\mathrm{Accepted}.\mathrm{equal}\}\}\}\}\}$
Composing $\mathrm{NatPlus}$ with $\mathrm{NatEquality}$ by inheriting both, with no modifications to either mixin, gives the resulting Nat values both $\mathrm{Plus}$ and $\mathrm{Equal}$ automatically. The concrete numerals live in a shared mixin that inherits only $\mathrm{NatData}$:
$\mathrm{NatConstants}\mapsto\{\;$ $\mathrm{NatData},$
$\mathrm{One}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Successor},$
$\quad\quad\mathrm{predecessor}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Zero}\}\},$
$\mathrm{Two}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Successor},$
$\quad\quad\mathrm{predecessor}\mapsto\{\mathrm{One}\}\},$
$\mathrm{Three}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Successor},$
$\quad\quad\mathrm{predecessor}\mapsto\{\mathrm{Two}\}\},$
$\mathrm{Four}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Successor},$
$\quad\quad\mathrm{predecessor}\mapsto\{\mathrm{Three}\}\},$
$\mathrm{Five}\mapsto\{\mathrm{NatConstants}.\mathbf{this}.\mathrm{NatFactory}.\mathrm{Successor},$
$\quad\quad\mathrm{predecessor}\mapsto\{\mathrm{Four}\}\}\}$
The arithmetic test inherits $\mathrm{NatConstants}$ together with the operations it needs:
$\mathrm{Test}\mapsto\{\;$ $\mathrm{NatConstants},\;\mathrm{NatPlus},\;\mathrm{NatEquality},$
$\mathrm{Addition}\mapsto\{\mathrm{Test}.\mathbf{this}.\mathrm{Two}.\mathrm{Plus},\;\mathrm{addend}\mapsto\{\mathrm{Test}.\mathbf{this}.\mathrm{Three}\}\},$
$\mathrm{Test2plus3}\mapsto\{\mathrm{Test}.\mathbf{this}.\mathrm{Five}.\mathrm{Equal},\;\mathrm{other}\mapsto\{\mathrm{Addition}.\mathrm{sum}\}\}\}$
The path $\mathrm{Test}.\mathrm{Test2plus3}.\mathrm{equal}$ inherits $\mathrm{True}$. No existing mixin was modified. $\mathrm{NatEquality}$, a new operation, and $\mathrm{Boolean}$, a new data type, each required only a new mixin: extensibility in both dimensions of the Expression Problem (Wadler, 1998).
Tree-level inheritance and return types
Each operation is an independent mixin that extends a factory’s subtree. Inheritance recursively merges subtrees that share a label, so every label in the factory acquires all operations from all inherited mixins. Crucially, this applies to return types: $\mathrm{NatPlus}$ returns values from $\mathrm{NatFactory}$, and $\mathrm{NatEquality}$ independently adds $\mathrm{Equal}$ to the same factory. Inheriting both causes the values returned by $\mathrm{Plus}$ to carry $\mathrm{Equal}$ automatically, without modifying either mixin. In object algebras (d. S. Oliveira and Cook, 2012) and finally tagless interpreters (Carette et al., 2009), independently defined operations do not automatically enrich each other’s return types; additional boilerplate or type-class machinery is required.
Multi-target qualified $\mathtt{this}$ (equation 6) is essential. In the test mixin above, $\mathrm{NatPlus}.\mathrm{Successor}$ and $\mathrm{NatEquality}.\mathrm{Successor}$ both inherit $\mathrm{NatData}.\mathrm{Successor}$. When the two operations are composed, the resulting $\mathrm{Successor}$ inherits the $\mathrm{NatData}$ schema through two independent routes. The reference $\mathrm{Successor}.\mathbf{this}.\mathrm{predecessor}$ inside each operation resolves through both routes, and since inheritance is idempotent, all routes contribute the same properties. In Scala, the same pattern, where two independently defined inner classes extend the same outer class’s trait, triggers a “conflicting base types” rejection (Appendix D), because DOT’s self variable is single-valued and cannot resolve through multiple inheritance routes simultaneously. The Expression Problem can still be solved in Scala using object algebras or similar patterns, but composing independently defined operations inherently produces multi-target $\mathtt{this}$ resolution to shared schemas, which DOT’s single-valued self variable cannot accommodate.
Church-encoded Nats are tries
A Church-encoded Nat is a record whose structure mirrors the data-constructor path used to build it: $\mathrm{Zero}$ is a flat record; $\mathrm{S}(\mathrm{Zero})$ is a record with a $\mathrm{predecessor}$ field pointing to a $\mathrm{Zero}$ record; and so on. This is exactly the shape of a trie. [^15] The unary Nat trie can be viewed as a sparse linked list whose depth equals the number it represents: in a trie union, some $\mathrm{Successor}$ nodes along the list point to $\mathrm{Zero}$ (leaf nodes), so not every position in the list is occupied. The MIXINv2 source (Yang, 2026) includes a binary natural number (BinNat) arithmetic whose trie depth is logarithmic. A scope that simultaneously inherits two Nat values is a trie union. The following test applies $\mathrm{Plus}$ to a trie union of $\{1,2\}$ with addend a trie union of $\{3,4\}$, then compares the result to itself with $\mathrm{Equal}$:
$\mathrm{CartesianTest}\mapsto\{\;$ $\mathrm{NatConstants},\;\mathrm{NatPlus},\;\mathrm{NatEquality},$
$\mathrm{OneOrTwo}\mapsto\{\mathrm{CartesianTest}.\mathbf{this}.\mathrm{One},\;\mathrm{CartesianTest}.\mathbf{this}.\mathrm{Two}\},$
$\mathrm{ThreeOrFour}\mapsto\{\mathrm{CartesianTest}.\mathbf{this}.\mathrm{Three},\;\mathrm{CartesianTest}.\mathbf{this}.\mathrm{Four}\},$
$\mathrm{Result}\mapsto\{\mathrm{OneOrTwo}.\mathrm{Plus},\;\mathrm{addend}\mapsto\{\mathrm{ThreeOrFour}\}\},$
$\mathrm{Check}\mapsto\{\mathrm{Result}.\mathrm{sum}.\mathrm{Equal},\;\mathrm{other}\mapsto\{\mathrm{Result}.\mathrm{sum}\}\}\}$
$\mathrm{Result}.\mathrm{sum}$ evaluates to the trie $\{4,5,6\}$, the set of all pairwise sums. $\mathrm{Equal}$ then applies to every pair drawn from $\{4,5,6\}\times\{4,5,6\}$; the path $\mathrm{CartesianTest}.\mathrm{Check}.\mathrm{equal}$ inherits both $\mathrm{True}$ and $\mathrm{False}$. The Cartesian product arises directly from the semantic equations. $\mathrm{OneOrTwo}$ inherits both $\mathrm{One}$ and $\mathrm{Two}$, so $\mathtt{bases}(\mathrm{OneOrTwo})$ contains two paths. When $\mathrm{Result}$ inherits $\mathrm{OneOrTwo}.\mathrm{Plus}$, the $\mathtt{bases}$ of $\mathrm{Result}$ include the $\mathrm{Plus}$ of both $\mathrm{One}$ and $\mathrm{Two}$. Each $\mathrm{Plus}$ resolves its $\mathrm{addend}$ via $\mathtt{this}$, reaching $\mathrm{ThreeOrFour}$, which is itself a trie union of $\mathrm{Three}$ and $\mathrm{Four}$: the $\mathtt{supers}$ of the $\mathrm{addend}$ path contain both values. Since $\mathrm{Successor}.\mathrm{Plus}$ recursively delegates to $\mathrm{predecessor}.\mathrm{Plus}$ with an incremented addend, and $\mathrm{predecessor}$ also resolves through the trie union, the recursion branches at every trie node. The result $\mathrm{sum}$ therefore collects all paths reachable through any combination of the two operands, which is the Cartesian product.
The same $\mathrm{Plus}$ and $\mathrm{Equal}$ files that operate on single values thus produce, without modification, the relational semantics of logic programming: [^16] Appendix F gives the systematic encoding of Datalog into inheritance-calculus. $\mathtt{overrides}$ and $\mathtt{supers}$ collect all inheritance paths, and every operation distributes over them.
6. Discussion
6.1. Emergent Phenomena
In developing real programs in MIXINv2, three phenomena arose from off-label use of the adopted design patterns. [^17] MIXINv2 programs adopt two standard patterns: data types follow a factory pattern isomorphic to Scala’s sealed trait with case classes; pattern matching uses Scott encoding. Both are named after Gang-of-Four terminology (Gamma et al., 1994) (factory, visitor, acceptance), with verbs replaced by nouns since the calculus has no methods. They can be seen in the source code (Yang, 2026) and the Case Study (Section 5). Below we explain why they arise from the three primitives.
Relational semantics:
Single-valued computation and set-valued computation are the same mechanism. The semantic equations of Section 3 resolve every property to a set of paths, so operations defined on single values automatically distribute over all combinations of set-valued inputs (Section 5). When a mixin tree is viewed as a trie, the correspondence with relational algebra becomes precise. The carrier is a trie rather than a flat relation, and observation is coinductive: the observer drives computation by querying progressively deeper paths. This contrasts with Datalog’s bottom-up materialization, but the relational structure of the computation is the same. Appendix F makes this correspondence precise with a systematic encoding of Datalog into inheritance-calculus, demonstrating that variable joins, constant matches, multi-body rules, and recursive rules are all expressible using only the three primitives of Section 2.
Immunity to nonextensibility:
It is structurally impossible to write a nonextensible program. Every construct of inheritance-calculus (Section 2) is extensible via deep merge (equation 3). The Case Study (Section 5) and the Expressive Asymmetry (Section 4.2; Theorem C.8) demonstrate and formalize this: new operations and data types are added, and new observable projections are introduced. This is the extensibility that the Expression Problem (Wadler, 1998) demands. [^18] As an untyped calculus, inheritance-calculus cannot violate the extensibility requirement of the Expression Problem but does not encounter its type-safety requirement. A typed language built on inheritance-calculus (Yang, 2026) inherits the immunity to nonextensibility and need only ensure that its type system does not introduce nonextensible constructs. Programming languages typically provide a mix of extensible and nonextensible constructs, where extensibility is opt-in: frameworks must encode nonextensible constructs into extensible constructs using sophisticated mechanisms (Liang et al., 1995; Lämmel and Jones, 2003; Löh and Hinze, 2006; Swierstra, 2008; Carette et al., 2009; d. S. Oliveira and Cook, 2012; Kiselyov et al., 2013; Wu et al., 2014; Kiselyov and Ishii, 2015; Wang and d. S. Oliveira, 2016; Leijen, 2017; Poulsen and van der Rest, 2023), and code that uses nonextensible constructs cannot be extended unless rewritten in the framework. Section 6.2 analyzes why deep merge is the only composition mechanism that achieves this.
Function color blindness:
Business logic in MIXINv2 declares abstract property slots; different assemblies fill them with synchronous or asynchronous implementations via inheritance. Because the slot’s color varies at assembly time without affecting the code that uses it, the same MIXINv2 file runs under both synchronous and asynchronous runtimes (Yang, 2026), an instance of function color blindness (Nystrom, 2015). The mechanism is dependency injection via inheritance: business logic declares empty slots, and the assembly inherits an FFI module that fills them, so the same business logic accepts implementations of any color. [^19] Function color (Nystrom, 2015) arises when effectful functions have signatures $a\to f\,b$ (Kleisli arrows), requiring monadic bind to compose and making the effect $f$ visible at every call site. Because inheritance-calculus composes program structure rather than executing effects, this type distinction does not arise at the calculus level. The color is determined at assembly time by the choice of inherited implementation, and effect execution is delegated to the host language. Unstructured control flow, such as capturing and resuming continuations as in call/cc, cannot be abstracted this way, as it requires a global CPS transformation that inheritance-calculus does not express.
6.2. Semantic Alternatives
The preceding item established that inheritance-calculus is immune to nonextensibility. We adopt this as a design goal and ask whether an alternative design could share this property by surveying every candidate known to us.
Degrees of freedom
The analysis below assumes that the AST is a named tree:every edge from parent to child carries a label, so that every node is uniquely identified by the path from the root. For configuration languages this is naturally the case: every key in a JSON, YAML, or Nix attribute set is a label, and nesting produces paths. [^20] Nix is a functional language with first-class functions; its attribute sets alone do not form a named-tree DSL. The NixOS module system (NixOS Contributors, ), however, restricts the user-facing interface to nested attribute sets composed by recursive merging, and it is this DSL-level structure that naturally forms a named tree. For mainstream programming languages the AST is not directly a named tree, since anonymous expression positions and implicit scopes lack labels. After ANF transformation (Flanagan et al., 1993), however, every intermediate result is bound to a name, and the remaining anonymous positions [^21] For example, the body of a $\mathbf{let}$. can be assigned synthetic fresh labels such as $\mathrm{result}$ and $\mathrm{tailCall}$ in Section 4, yielding a named tree.
Under this premise, every language whose programs are such named trees with references can be characterized by two finite sets: a set $\mathcal{N}$ of node types, each determining how a subtree organizes its children, and a set $\mathcal{R}$ of reference types, each determining how one position in the tree refers to another. The syntax is fixed by the choice of $\mathcal{N}$ and $\mathcal{R}$; the semantics is determined by how each node type structures composition and how each reference type resolves its target.
Language Node types $\mathcal{N}$ Reference types $\mathcal{R}$
Java (OOP) class, interface, method, block, conditional, loop, throw, lambda, … variable, field access, method call, extends/implements, import, …
Haskell (FP) $\lambda$, application, let, case, data, type class, instance, … variable, constructor, type class dispatch, import, …
Scala [^22] Scala is a multi-paradigm language. class, trait, object, method, block, conditional, match, while, throw, … variable, member access, extends, with, implicits, import, …
$\lambda$-calculus abstraction, application variable
Inheritance-calculus record inheritance
Immunity to nonextensibility
A path $p$ in a program $P$ is extensible if there exists a program $Q$ such that composing $P$ with $Q$ through a reference makes new properties observable at $p$ without modifying $P$. A language is immune to nonextensibility if every path in every well-formed program is extensible.
We now classify the design space to explain why we chose the semantics of Section 3. Two axes remain free: the composition mechanism, which determines how the single reference type incorporates definitions; and the reference resolution, which determines how the single reference type finds its target.
The composition mechanism
Every composition mechanism determines how the definitions of one abstraction are incorporated into another. The following table classifies the known candidates by whether they satisfy commutativity (C), idempotence (I), and associativity (A), and by their extensibility model.
Candidate C I A Extensibility Representative
Deep merge $\checkmark$ $\checkmark$ $\checkmark$ universal this paper
Shallow merge $\times$ $\checkmark$ $\checkmark$ top-level only JS spread
Conflict rejection $\checkmark$ $\checkmark$ $\checkmark$ none on overlap Harper–Pierce (Harper and Pierce, 1991)
Priority merge $\times$ $\times$ $\checkmark$ universal Jsonnet +, Nix //
$\beta$-reduction $\times$ $\times$ $\times$ opt-in Java, Haskell, Scala, …
Conditional causes divergence
Deep merge:
Merging same-label definitions recurses into their subtrees ($\mathtt{overrides}$, equation 3). Every label is an extension point, whether or not the original author anticipated extension.
Shallow merge:
Nested labels cannot be independently extended.
Conflict rejection:
Same-label definitions are rejected as errors. All three identities hold, but vacuously: the identities are satisfied because the compositions that would test them are refused. Conflict rejection prevents extending existing labels (Harper and Pierce, 1991), precisely the composability that the Expression Problem (Wadler, 1998) requires.
Priority merge:
One source takes precedence (Cunningham, 2014; Dolstra et al., 2010), so $A+B\neq B+A$, introducing an ordering dependency.
$\beta$-reduction:
Only the parameters placed by the author are extension points; the function body is closed. To make a new aspect extensible, the original function must be rewritten to accept an additional parameter. Any composition mechanism based on function application inherits this limitation, because $\beta$-reduction closes the function body. The Expression Problem (Wadler, 1998) is hard in such languages for the same reason: extensibility is opt-in. The frameworks that simulate immunity to nonextensibility (visitor pattern, object algebras (d. S. Oliveira and Cook, 2012), finally tagless interpreters (Carette et al., 2009)) are the machinery needed to work around this limitation.
Conditional incorporation:
Removing or conditionally including definitions based on the presence of other definitions creates a dependency cycle: whether a definition exists depends on whether another definition exists, which in turn depends on the first. The recursive evaluation (Appendix A) cannot reach a fixed point when addition and removal alternate, causing divergence. This differs from the divergence inherent to Turing completeness: Turing-complete programs diverge because computation is unbounded, whereas conditional incorporation diverges because the immediate consequence operator $T_{P}$ oscillates rather than ascending monotonically.
Among the candidates above, deep merge is the only one we found that is commutative, idempotent, associative, and immune to nonextensibility. The other candidates each introduce a vulnerability to nonextensibility or cause divergence.
The reference resolution
Once deep merge over records is fixed as the composition mechanism, four of the six semantic equations are determined by the syntax: $\mathtt{properties}$ (1) reads the labels of a record, $\mathtt{bases}$ (4) reads its inheritance sources, $\mathtt{supers}$ (2) takes the transitive closure, and $\mathtt{overrides}$ (3) implements the merge. The only remaining degree of freedom is how references resolve: the $\mathtt{resolve}$ (5) and $\mathtt{this}$ (6) equations. We examine the three known candidates.
Early binding:
With early binding, references resolve at definition time to a fixed path, independently of how the record is later inherited. CUE (van Lohuizen, 2019) exemplifies this: field references are statically resolved, and a nested field cannot refer to properties of its enclosing context contributed by later composition. This prevents the mixin from serving as a call stack for the $\lambda$-calculus: $\beta$-reduction requires the callee to observe the argument supplied at the call site, which is a later composition. With early binding, references are frozen before this composition occurs. In the equations, $\mathtt{resolve}$ (5) calls $\mathtt{this}$ (6), which walks the inherited tree structure; late binding is what allows $\mathtt{this}$ to see properties contributed by later-composed mixins. Without a Turing-complete mixin, functions must be reintroduced as the composition mechanism, re-exposing the opt-in extensibility problem above.
Dynamic scope:
With dynamic scope, references resolve at the point of use rather than the definition site. Jsonnet (Cunningham, 2014) follows this approach: self always refers to the final merged object, not the object at the definition site. Union file systems exhibit the same property at the file-system level: a symbolic link’s relative path resolves against the mount context at dereference time, not against the layer that defined the link. Dynamic scope breaks the $\lambda$-calculus embedding (Section 4). The Substitution Lemma (Lemma B.2) relies on the base case $M=y$ ($y\neq x$): the de Bruijn index $m$ of $y$ walks up from the definition-site path $\mathtt{init}(p_{\mathrm{def}})$ to a scope level different from $x$’s binder, so the substitution $\mathrm{argument}\mapsto\mathcal{T}(V)$ at $x$’s scope has no effect. With dynamic scope this separation fails. [^23] Counterexample: $(\lambda x.\,\lambda y.\,x)\;V_{1}\;V_{2}$. The reference to $x$ inside $\lambda y.\,x$ translates to $\uparrow^{\!1}.\mathrm{argument}$ (de Bruijn index $1$). With late binding, $\mathtt{this}$ starts from the definition-site scope (the inner $\lambda$) and walks up one level to the outer $\lambda$, reaching the slot $\mathrm{argument}\mapsto\mathcal{T}(V_{1})$. With dynamic scope, the reference resolves from the use-site context, where the innermost enclosing scope is the application $\{C_{1}.\mathrm{result},\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$; walking up one level reaches this scope’s $\mathrm{argument}$ slot, returning $\mathcal{T}(V_{2})$ instead of $\mathcal{T}(V_{1})$. This is the classical variable-capture failure of dynamic scoping. As with early binding, the mixin cannot serve as a complete replacement for functions.
Single-target $\mathtt{this}$:
Single-target $\mathtt{this}$ is a promising candidate: the Single-Path Lemma (B.1) shows that for translated $\lambda$-terms, the frontier set $S$ in $\mathtt{this}$ (6) always contains exactly one path, so the $\lambda$-calculus embedding does not require multi-target resolution. However, single-target $\mathtt{this}$ is vulnerable to nonextensibility for a different reason. When multiple inheritance routes reach the same scope, single-target resolution must reject the situation or select one route. The NixOS module system (NixOS Contributors, ) raises a duplicate-declaration error (Appendix D). This rejects nested extensions where two independently authored modules define the same option, a composition that arises when solving the Expression Problem (Wadler, 1998) (Section 5).
Our first three implementations of MIXINv2 used standard techniques for $\mathtt{this}$ resolution: closure-based fixed points following Cook (Cook, 1989), and stack-based environment lookup using de Bruijn indices (de Bruijn, 1972). All three produced subtle bugs when inheritance introduced multiple routes to the same enclosing scope. The difficulty was structural. A closure captures a single environment, so a fixed-point combinator solves for a single $\mathtt{this}$ target. A stack-based environment is a linear chain where each scope level has one parent. Both data structures embed a single-valued assumption incompatible with the multi-target situation that multiple inheritance naturally produces (Appendix D). Equation (6) emerged from abandoning this assumption: it tracks a frontier set $S$ rather than a single current scope.
The following table summarizes the vulnerabilities of all alternatives at both levels.
Candidate Vulnerability Representative
composition Shallow merge nested extensions not composable JS spread
Conflict rejection same-label extensions rejected Harper–Pierce (Harper and Pierce, 1991)
Priority merge ordering dependency Jsonnet + (Cunningham, 2014), Nix // (Dolstra et al., 2010)
$\beta$-reduction opt-in extensibility Java, Haskell, Scala, …
Conditional causes divergence
resolution Early binding mixin cannot replace function CUE (van Lohuizen, 2019)
Dynamic scope mixin cannot replace function Jsonnet (Cunningham, 2014), union FS
Single-target $\mathtt{this}$ nested extensions rejected NixOS modules (NixOS Contributors, )
No alternative we examined at either level is immune to nonextensibility.
We arrived at this semantics by trying each alternative in this section and finding that it failed: shallow merge could not compose nested extensions; priority merge introduced ordering dependencies; early binding froze references before composition could supply them; dynamic scope broke the substitution that the $\lambda$-calculus embedding relies on; single-target $\mathtt{this}$ rejected the multi-target compositions that the Expression Problem demands. Deep merge with multi-target late-binding $\mathtt{this}$ is what remained.
7. Related Work
7.1. Computational Models and $\lambda$-Calculus Semantics
Minimal Turing-complete models
Schönfinkel (Schönfinkel, 1924) and Curry (Curry and Feys, 1958) showed that the $\lambda$-calculus reduces to three combinators $S$, $K$, II  [^24] The $I$ combinator is redundant since $I=SKK$. . The bracket abstraction that compiles $\lambda$-terms to SKI is mechanical but incurs exponential blow-up in term size; no new computational patterns emerge beyond those already present in the $\lambda$-calculus. Semantically, the two share the same models: Meyer (Meyer, 1982) showed that combinatory algebras satisfying five additional conditions are exactly lambda algebras. No independent semantic domain for SKI has ever been exhibited from which $\lambda$-calculus semantics could be derived. The direction has always been the reverse: Scott’s $D_{\infty}$ was built for the $\lambda$-calculus, and SKI was interpreted inside it afterwards. Dolan (Dolan, 2013) showed that the x86 mov instruction alone is Turing complete by exploiting addressing modes as implicit arithmetic on mutable memory; the resulting programs are correct but provide no abstraction mechanisms beyond those of the RAM Machine. The Turing Machine itself, while universal, lacks random access: reading a cell at distance $d$ requires $d$ sequential tape movements. Each of these models achieves Turing completeness within the semantic framework of an existing model, and no independent semantic domain has been exhibited for any of them.
Semantic descriptions of the $\lambda$-calculus
The semantics of the $\lambda$-calculus can be defined in several fundamentally different ways. Operational semantics takes $\beta$-reduction $(\lambda x.\,M)\,N\to M[N/x]$ as the primitive notion of computation (Church, 1936, 1941; Barendregt, 1984). This requires capture-avoiding substitution, a mechanism whose subtleties motivated de Bruijn indices (de Bruijn, 1972) and explicit substitution calculi (Abadi et al., 1991). It also requires a choice of evaluation strategy: call-by-name versus call-by-value (Plotkin, 1975) or the lazy strategy of Abramsky (Abramsky, 1990). Denotational semantics (Scott, 1970, 1976) interprets terms in a mathematical domain, but giving meaning to $\beta$-reduction requires a reflexive domain $D\cong[D\to D]$, which has no set-theoretic solution by Cantor’s theorem. Scott’s $D_{\infty}$ resolves this by constructing $D$ as an inverse limit of continuous lattices (Scott, 1972). The resulting semantics is adequate but not fully abstract (Plotkin, 1977; Milner, 1977): it identifies terms that observational equivalence distinguishes. Game semantics (Abramsky et al., 2000; Hyland and Ong, 2000; Nickau, 1994) achieves full abstraction for PCF and for the lazy $\lambda$-calculus (Abramsky and Ong, 1993; Abramsky and McCusker, 1995), requiring arenas, strategies, and innocence conditions. In each of these approaches, the function space $D\to D]$ is the central obstacle: operational semantics needs it implicitly via substitution. In contrast, denotational semantics constructs it explicitly, and game semantics refines it via strategies. Each of these approaches involves higher-order structure [^25] Function spaces in denotational semantics, strategies in game semantics, and substitution in operational semantics. and retains the function space in its semantic domain. The denotational and operational accounts remain separate formalisms in each case. All three inherit $\beta$-reduction’s divergence on recursive programs: a naïve transitive closure on a cyclic graph, for instance, diverges in every framework.
Set-theoretic models of the $\lambda$-calculus
Several families of $\lambda$-calculus models avoid the inverse-limit construction of Scott’s $D_{\infty}$ by working in set-theoretic rather than domain-theoretic settings. These models eliminate Scott topology but not the function space itself: they remain higher-order, encoding $D\to D]$ inside $D$ via injections, intersection types, or multiset relations. Graph models (Engeler, 1981; Plotkin, 1993) encode functions via a web injection on powerset lattices; Bucciarelli and Salibra (Bucciarelli and Salibra, 2008) proved that the greatest sensible graph theory equals the Böhm tree theory $\mathcal{B}$. Filter models (Barendregt et al., 1983; Coppo and Dezani-Ciancaglini, 1978) interpret a term as the set of intersection types derivable for it; Dezani-Ciancaglini et al. (Dezani-Ciancaglini et al., 1998) showed that the BCD filter model’s $\lambda$-theory coincides exactly with Böhm tree equality. Relational models (Ehrhard, 2005; Bucciarelli and Ehrhard, 2001) arise from the relational semantics of linear logic via finite multisets; Ehrhard (Ehrhard, 2005) showed that a standard reflexive object for the pure $\lambda$-calculus cannot exist in this category. All three families are semantic models: they interpret $\lambda$-terms in an external mathematical structure and validate $\beta$-reduction as an equation in that structure, rather than defining an independent calculus with its own primitives. Each encodes the function space rather than eliminating it, and each requires higher-order machinery to obtain full abstraction for the lazy $\lambda$-calculus.
Recursive equations and fixpoint semantics
Van Emden and Kowalski (van Emden and Kowalski, 1976) showed that the semantics of predicate logic as a programming language admits three equivalent characterizations: operational [^26] Typically SLD resolution. , model-theoretic [^27] Based on least Herbrand models. , and fixed-point [^28] Namely, the least fixed point of the immediate consequence operator $T_{P}$ on the powerset of ground atoms. . Aczel (Aczel, 1977) gave a systematic treatment of inductive definitions via monotone operators on powerset lattices, providing the logical foundations that Datalog-style semantics rests on. Leroy and Grall (Leroy and Grall, 2006) treated divergence in big-step semantics via coinductive definitions, requiring a separate coinductive judgment to account for non-termination alongside the inductive evaluation judgment. Van Emden and Kowalski’s $T_{P}$ computes $\mathrm{lfp}$ for recursive Datalog programs over a first-order domain. Conventional $\lambda$-calculus semantics requires function spaces in the semantic domain, and recursive $\lambda$-terms ordinarily diverge under $\beta$-reduction; using $T_{P}$ over a first-order domain to give the $\lambda$-calculus a fixpoint semantics has not been explored.
7.2. Inheritance, Objects, and Records
Denotational semantics of inheritance
Cook (Cook, 1989) gave the first denotational semantics of inheritance, modeling objects as recursive records. Inheritance is composition of generators  [^29] These are functions from self to complete object. and wrappers [^30] Wrappers are functions that modify generators. , solved by a fixed-point construction. Cook’s key insight is that inheritance is a general mechanism applicable to any form of recursive definition, not only object-oriented methods. This insight is one of the starting points of the present work. The model has four primitives (records, functions, generators, and wrappers) and requires a fixed-point combinator. Composition is asymmetric: a wrapper modifies a generator but not vice versa. The fixed-point construction relies on Scott continuity in a domain of functions. The resulting denotations are not directly executable, so a separate operational semantics is needed to define method dispatch and object construction at runtime.
Mixins and traits
Bracha and Cook (Bracha and Cook, 1990) formalized mixins as abstract subclasses, that is, functions from a superclass parameter to a subclass. Because mixin application is function composition, it is neither commutative nor idempotent; the C3 linearization algorithm (Barrett et al., 1996) was later developed to impose a deterministic order. Schärli et al. (Schärli et al., 2003) introduced traits, whose composition is commutative but rejects same-name conflicts (Ducasse et al., 2006). Both mechanisms operate at the method level: composing two definitions of the same name results in an override or a conflict, not a recursive merge of nested structure. Both presuppose the $\lambda$-calculus: method bodies are functions, and the composition mechanism itself is not Turing complete. Neither mixins nor traits are deep-mergeable.
Object and record calculi
Abadi and Cardelli (Abadi and Cardelli, 1996) treat objects as the primitive notion, but methods are functions that take self as a parameter and object extension is asymmetric. Boudol (Boudol, 2004) showed that self-referential records require an unsafe fixed-point operator whose well-definedness depends on the evaluation strategy. Harper and Pierce (Harper and Pierce, 1991) gave a record calculus with symmetric concatenation that rejects same-label composition; Cardelli (Cardelli, 1994) studied extensible records with subtyping; Rémy (Rémy, 1989) introduced row polymorphism for records and variants in ML. All of these calculi operate on flat records: there is no recursive merging of nested structure, no lazy observation, and $\mathtt{this}$ requires an explicit fixed-point combinator.
Family polymorphism and DOT
Ernst (Ernst, 2001) introduced family polymorphism; Ernst, Ostermann, and Cook (Ernst et al., 2006) formalized it in the virtual class calculus, where classes are accessed via path expressions this.out.C. Amin et al. (Amin et al., 2016) formalized path-dependent types in the DOT calculus, the theoretical foundation of Scala. In both frameworks, each path resolves to a single class in a single enclosing object. When multiple inheritance routes lead to the same scope, single-valued resolution must choose one route and discard the others; DOT’s self variable $\nu(x:T)\,d$ binds $x$ to a single object, and Scala rejects the resulting intersection type at compile time (Appendix D). Neither framework provides a mechanism for set-valued resolution.
7.3. Configuration Languages and Module Systems
Configuration languages
CUE (van Lohuizen, 2019) unifies types and values in a single lattice where unification, denoted by &, is commutative, associative, and idempotent. CUE’s design was motivated by Google’s Borg Configuration Language (GCL), which used graph unification. CUE’s lattice includes scalar types with a conflict semantics where unifying incompatible scalars yields $\bot$; whether scalars are necessary for a configuration language, or whether tree structure alone suffices, is an interesting design question. Jsonnet (Cunningham, 2014) provides object inheritance via the + operator with mixin semantics: composition is not commutative since the right-hand side wins on scalar conflicts, and deep merging requires explicit opt-in via the +: syntax on a per-field basis, so deep merge is available but not the default. Dhall (Gonzalez, 2017) takes a functional approach: it is a typed $\lambda$-calculus with records, where composition is a record merge operator, not an inheritance mechanism.
Module systems and deep merge
The NixOS module system (Dolstra et al., 2010; NixOS Contributors, ) composes modules by recursively merging nested attribute sets, not by linearization or method-level override, which is the same deep-merge semantics that distinguishes tree-level inheritance from method-level composition in prior mixin and trait calculi. When two modules define the same nested path, their subtrees are merged rather than overridden. The NixOS module system incorporates deep merge, recursive $\mathtt{this}$, deferred evaluation of modules, and a rich type-checking layer on top. The same mechanism powers NixOS system configuration, Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), flake-parts (Hensing, 2022), and dozens of other ecosystems, collectively managing configurations of arbitrary complexity. The one aspect that appears not to have been fully addressed is the semantics of $\mathtt{this}$ under multiple inheritance. When two independent modules that inherit from the same deferred module are composed, each import introduces its own declaration of the shared options; the module system’s mergeOptionDecls rejects duplicate declarations with a static error (Appendix D). The module system cannot express that both routes contribute equally to the same option, that is, the multi-target situation that recursive merging naturally gives rise to, which motivated the observational semantics of Section 3.
7.4. Our Companion Work
Modular software architectures
Plugin systems, component frameworks, and dependency injection frameworks enable software extensibility through declarative inheritance. The symmetric, associative nature of inheritance (Section 3) explains why such systems work: components can be inherited in any order without changing behavior. MIXINv2 (Yang, 2026) is itself a dependency injection framework: each scope declares its dependencies as named slots, and composition resolves them by name across scope boundaries, making the inheritance-calculus patterns explicit rather than ad hoc.
Union file systems
Union file systems, such as UnionFS, OverlayFS, and AUFS, layer multiple directory hierarchies to present a unified view. Their semantics resembles inheritance-calculus in that later layers override earlier layers and files from all layers remain accessible, but their designs differ in specific ways: inheritance is asymmetric, conflict resolution for scalar files is noncommutative, and there is no late-binding or dynamic dispatch across layers. We implemented inheritance-calculus directly as a union file system (Yang, 2025), obtaining late-binding semantics and dynamic dispatch for file lookups with no additional mechanism beyond the three primitives. In package managers, build systems, and OS distributions, external toolchains transform configuration into file trees; here, the file system serves as configuration to itself, and the transformation is inheritance: a file-system-as-compiler in which the directory tree is both the source program and the compilation target.
8. Future Work
MIXINv2 (Yang, 2026) is an executable implementation of inheritance-calculus that already goes beyond the untyped calculus presented here: it includes compile-time checking that all references resolve to valid paths in the mixin tree, and a foreign-function interface (FFI) that introduces scalar values from the host language. The future work described below concerns the distance between this implementation and a fully practical language, spanning both theoretical foundations and library-level encodings.
8.1. Type System
MIXINv2’s existing compile-time checks verify that every reference path resolves to a property that exists in the inherited result. Formalizing the soundness of these checks, that is, proving that well-typed programs do not produce dangling references at runtime, is one direction of future work. Such a formalization would relate to the DOT calculus (Amin et al., 2016) in a manner analogous to how inheritance-calculus relates to the $\lambda$-calculus: a Typed inheritance-calculus can be viewed as DOT without $\lambda$, retaining path-dependent types while replacing functions with mixin inheritance. A formal type system would be amenable to mechanization in proof assistants, such as Coq and Agda, for verifying type safety, progress, and preservation properties.
Beyond soundness of the existing checks, additional type system features would strengthen the language. A key example is totality checking: verifying that an inheritance provides implementations for all required slots, not merely that references resolve. In untyped inheritance-calculus, writing $\mathrm{magnitude}\mapsto\{\}$ is purely documentary: it defines a structural slot but imposes no constraint on what must fill it. Totality checking would enforce such declarations statically, rejecting inheritances that leave required slots unfilled.
8.2. Standard Library
The remaining items concern the standard library: they require no extensions to the calculus or type system, but involve nontrivial encodings within the existing framework.
Inheritance-calculus is open by default: inheritance can freely merge any two mixins, and a value may simultaneously inhabit multiple data constructors (e.g., $\{\mathrm{Zero},\;\mathrm{Odd},\;\mathrm{half}\mapsto\mathrm{Zero}\}$). This is the natural trie semantics of the calculus, and the basis for solving the Expression Problem (Section 5). However, many operations, such as equality testing, assume that values are linear, inhabiting exactly one data constructor. The observer pattern (Section 5) does not enforce this: applied to a multi-data-constructor value, all callbacks fire and their results are inherited.
Schema validation can be implemented at the library level using existing primitives. An observer can test whether a particular data constructor is present, returning a Boolean; a second Boolean dispatch then branches on the result. By chaining such tests, a factory can validate that a value matches exactly one data constructor and carries the required fields. This does not restrict inheritance itself, which remains unconstrained, but provides a way to detect invariant violations before they propagate.
The same technique, an observer returning Boolean followed by Boolean dispatch, yields closed pattern matching. The observer pattern of Section 5 is inherently open: new data-constructor cases can be added through inheritance. Closed matching, where exactly one branch is taken, can be encoded as a visitor chain: a linked list of if-then-else nodes, each testing one data constructor and falling through on mismatch. This is analogous to GHC.Generics’ $\mathrm{(:+:)}$ sum representation (Magalhães et al., 2010), where each link corresponds to one summand. An important consequence is that such chains do not commute, as they have a fixed order, and therefore cannot be extended through open inheritance, which is the correct semantics for closed dispatch.
9. Conclusion
Immunity to nonextensibility guided the design of inheritance-calculus. Section 6.2 eliminated every alternative composition mechanism and reference resolution, leaving deep merge with multi-target late-binding qualified $\mathtt{this}$ as the unique survivor. The mixin served as the single abstraction unifying module, class, object, method, field, and local binding; deep merge of mixin trees made inheritance commutative, idempotent, and associative, so the linearization problem of multiple inheritance did not arise. The six first-order equations of Section 3, denotational and computable by tabling (Tamaki and Sato, 1986), suffice to define the semantics; the observer drives computation by querying paths in a lazily constructed tree. These equations are a minimal departure from standard object-oriented semantics: the only change is the $\mathtt{overrides}$ equation (3), which replaces method-level shadowing with deep merge of subtrees. Object-oriented programming is traditionally classified as imperative; removing mutable state and function bodies yields pure object-oriented programming,313131“Pure” carries a double meaning: not impure, since values are immutable and composition has no side effects; and not mixed, since composition relies entirely on inheritance, without functional abstractions such as $\lambda$ or $\beta$-reduction. a minimal computational model for declarative programming that fills the empty cell of Table 1.
Since $\mathtt{overrides}$ merges subtrees rather than method bodies, functions are no longer primitive. The $\lambda$-calculus macro-expressed into inheritance-calculus in six translation rules in Section 4. The Böhm tree correspondence of Section 4.1 extended this first-order semantics to the $\lambda$-calculus: the prior higher-order semantic frameworks proved faithful to the first-iteration approximation of a sublanguage of inheritance-calculus’s first-order equations; Appendix E reformulated this semantics directly on ANF. A convergence hierarchy among the eager, lazy, and fixpoint $\lambda$-calculi sharing the same $\lambda$-syntax followed as a corollary, detailed in Remark B.5.
Inheritance-calculus further gains three phenomena that the $\lambda$-calculus cannot express; indeed, inheritance-calculus is strictly more expressive than the lazy $\lambda$-calculus in Felleisen’s sense (Felleisen, 1991) (Section 4.2). Section 5 established immunity to nonextensibility in the sense of the Expression Problem (Wadler, 1998). The case study of Section 5 showed that ordinary arithmetic yields the relational semantics of logic programming: cyclic inheritance hierarchies converge by the same ascending chain condition that governs recursive Datalog, and Appendix F gave a systematic encoding. Section Function color blindness: exhibited function color blindness (Nystrom, 2015).
Inheritance is not a mechanism for organizing functions. Functions are syntactic sugar for inheritance.
Appendix A Well-Definedness of the Semantic Functions
The six semantic equations (1)–(6) are first-order: their inputs and outputs are paths and sets of paths, with no function spaces, closures, or higher-order constructs. This section establishes well-definedness by showing that the six equations define a monotone operator whose least fixed point is computed soundly by tabled evaluation (Tamaki and Sato, 1986).
Definition A.1 (Dependency graph).
Fix an AST with its primitive functions $\mathtt{defines}$ and $\mathtt{inherits}$. The dependency graph $G$ has as vertices all semantic-function applications ($\mathtt{properties}(p)$, $\mathtt{supers}(p)$, $\mathtt{overrides}(p)$, $\mathtt{bases}(p)$, $\mathtt{resolve}(\ldots)$, $\mathtt{this}(\ldots)$) that arise from the AST. There is a directed edge from vertex $u$ to vertex $v$ whenever computing $u$ via equations (1)–(6) requires the value of $v$.
Definition A.2 (Immediate consequence operator $T_{P}$).
Let $\mathcal{Q}$ be the set of all semantic-function queries (vertices of $G$). For each query $q\in\mathcal{Q}$, let $\mathcal{V}_{q}$ be the set of answer values of the appropriate shape for $q$ (paths, sets of paths, or sets of pairs of paths, depending on the equation defining $q$). A query–answer map is an element of the pointwise product $\prod_{q\in\mathcal{Q}}\mathcal{P}(\mathcal{V}_{q})$; equivalently, it assigns to each query $q$ a set $M(q)\subseteq\mathcal{V}_{q}$ of candidate answers. The immediate consequence operator (van Emden and Kowalski, 1976) $T_{P}$ maps a query–answer map $M$ to a new query–answer map $T_{P}(M)$ by evaluating, for each query $q$, the right-hand side of the equation that defines $q$, interpreting every sub-query $q^{\prime}$ by the current approximation $M(q^{\prime})$.
Lemma A.3 (Monotonicity).
The operator $T_{P}$ is monotone on the complete lattice $\Bigl(\prod_{q\in\mathcal{Q}}\mathcal{P}(\mathcal{V}_{q}),\;\subseteq\Bigr)$.
Proof.
Each equation’s right-hand side is built from $\cup$, $\in$, equality tests, and positive set comprehension $\{\,x\mid x\in S,\ldots\,\}$. None of the six equations uses negation, set difference, or complement. Therefore, if $M\subseteq M^{\prime}$ (pointwise), then $T_{P}(M)\subseteq T_{P}(M^{\prime})$. ∎
Theorem A.4 (Existence and uniqueness of $\mathrm{lfp}(T_{P})$).
The operator $T_{P}$ has a unique least fixed point $\mathrm{lfp}(T_{P})$.
Proof.
Each powerset lattice $(\mathcal{P}(\mathcal{V}_{q}),\subseteq)$ is complete, so the pointwise product lattice $\Bigl(\prod_{q\in\mathcal{Q}}\mathcal{P}(\mathcal{V}_{q}),\subseteq\Bigr)$ is also complete. By the Knaster–Tarski theorem (Tarski, 1955), every monotone function on a complete lattice has a unique least fixed point. ∎
Theorem A.5 (Well-definedness).
The semantic functions $\mathtt{properties}$, $\mathtt{supers}$, $\mathtt{overrides}$, $\mathtt{bases}$, $\mathtt{resolve}$, $\mathtt{this}$ are well-defined: their denotational semantics is $\mathrm{lfp}(T_{P})$, the unique least fixed point of the immediate consequence operator.
Theorem A.6 (Tabling soundness).
Tabled evaluation of equations (1)–(6) never returns a wrong answer, even when the dependency graph $G$ is infinite. Specifically:
  1. (1)
    Soundness (no false positives). Every fact returned by tabling was produced by evaluating some equation’s right-hand side on its inputs. By induction on the evaluation trace, every returned fact belongs to $\mathrm{lfp}(T_{P})$.
  2. (2)
    Monotone accumulation. The tabling accumulator only grows between iterations via set union; no fact is ever removed.
  3. (3)
    Correct fixpoint detection. If no sub-query was re-entered during evaluation, all sub-queries were fully computed with no approximation, and the result is exact. If re-entry occurred but the accumulator is stable, meaning no new facts were added, the accumulator equals $\mathrm{lfp}(T_{P})$ restricted to the reachable queries, because it is a pre-fixed point reached by ascending from $\bot$.
  4. (4)
    Sound failure. When tabling cannot determine the answer, whether because the iteration budget is exhausted or because of an infinite non-cyclic chain, it raises an error rather than returning a wrong answer.
Proof.
For (1), each returned fact is the output of some equation’s right-hand side applied to previously returned facts; by induction, all inputs are in $\mathrm{lfp}(T_{P})$, so the output is too (by monotonicity and the fixed-point property).
For (2), the implementation accumulates via $A\leftarrow A\cup T_{P}(A)$, which is monotone in $A$.
For (3), when no reentry occurs, the evaluation is ordinary memoised recursion on a finite acyclic reachable subgraph, so every sub-query is computed exactly and no approximation is involved. When reentry occurs, let $R\subseteq\mathcal{Q}$ be the set of queries reachable from the initial query, and let $T_{P}^{R}$ denote the restriction of $T_{P}$ to maps on $R$. If the accumulator is stable, that is, $A=A\cup T_{P}^{R}(A)$, then $T_{P}^{R}(A)\subseteq A$, so $A$ is a pre-fixed point of $T_{P}^{R}$.
By (1), every fact ever added to the accumulator was produced by some right-hand side evaluation and therefore belongs to $\mathrm{lfp}(T_{P}^{R})$. Hence $A\subseteq\mathrm{lfp}(T_{P}^{R})$.
Conversely, by the Knaster–Tarski characterization of the least fixed point as the least pre-fixed point, $\mathrm{lfp}(T_{P}^{R})$ is contained in every pre-fixed point of $T_{P}^{R}$, in particular in $A$. Thus $\mathrm{lfp}(T_{P}^{R})\subseteq A$.
Therefore $A=\mathrm{lfp}(T_{P}^{R})$, that is, the stable accumulator equals the exact denotation on the reachable queries.
For (4), budget exhaustion and stack overflow both produce exceptions, never silent wrong answers. ∎
Termination
Whether evaluation terminates depends on the program:
  • Acyclic, finite reachable subgraph. One pass suffices; evaluation terminates.
  • Cyclic, finite per-query domain. Multiple passes converge by the ascending chain condition on a finite lattice (Bancilhon and Ramakrishnan, 1986). This is the Datalog case (Section F.8).
  • Infinite reachable subgraph. Evaluation may diverge; this is inherent to Turing completeness and not a defect of the evaluation strategy.
Appendix B Böhm Tree Correspondence: Proofs
This appendix proves two results about the sublanguage of inheritance-calculus corresponding to the $\lambda$-calculus (Section 4.1). First, the first-iteration approximation $T_{P}\!\uparrow\!1$ is fully abstract with respect to Böhm tree equivalence (Theorems B.7 and B.9), establishing that the first iteration coincides with the lazy $\lambda$-calculus. Second, the full $\mathrm{lfp}(T_{P})$ semantics, the fixpoint $\lambda$-calculus, is strictly more defined: $\lambda$-terms that diverge under $\beta$-reduction may converge (Remark B.5), establishing a convergence hierarchy among the eager, lazy, and fixpoint $\lambda$-calculi.
B.1. Böhm Trees
We briefly recall the definition of Böhm trees (Barendregt, 1984). A head reduction $M\to_{h}M^{\prime}$ contracts the outermost $\beta$-redex only: if $M$ has the form $(\lambda x.\,e)\;v\;M_{1}\cdots M_{k}$, then $M\to_{h}e[v/x]\;M_{1}\cdots M_{k}$. A term $M$ is in head normal form (HNF) if it has no head redex, that is, $M=\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}$ where $y$ is a variable.
The Böhm tree $\mathrm{BT}(M)$ of a $\lambda$-term $M$ is an infinite labeled tree defined by:
$\mathrm{BT}(M)=\begin{cases}\bot&\text{if }M\text{ has no HNF}\\[6.0pt] \lambda x_{1}\ldots x_{m}.\;y\bigl(\mathrm{BT}(M_{1}),\;\ldots,\;\mathrm{BT}(M_{k})\bigr)&\begin{aligned} &\text{if }M\to^{*}_{h}\\[-2.0pt] &\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}\end{aligned}\end{cases}$
Böhm tree equivalence $\mathrm{BT}(M)=\mathrm{BT}(N)$ is the standard observational equivalence of the lazy $\lambda$-calculus (Abramsky and Ong, 1993; Abramsky and McCusker, 1995).
B.2. Path Encoding
We define a correspondence between positions in the Böhm tree and paths in the mixin tree. When the translation $\mathcal{T}$ is applied to an ANF $\lambda$-term, the resulting mixin tree has a specific shape:
  • An abstraction $\lambda x.\,M$ translates to a record with own properties $\{\mathrm{argument},\mathrm{result}\}$, which is referred to as the abstraction shape.
  • A $\mathbf{let}$-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M$ translates to a record with own properties $\{x,\mathrm{result}\}$, where $x$ holds the encapsulated application $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$.
  • A tail call $V_{1}\;V_{2}$ translates to a record with own properties $\{\mathrm{tailCall},\mathrm{result}\}$, where $\mathrm{tailCall}$ holds the encapsulated application and $\mathrm{result}$ projects $\mathrm{tailCall}.\mathrm{result}$.
  • A variable $x$ with de Bruijn index $n$ translates to an indexed reference $\uparrow^{\!n}.\mathrm{argument}$.
Each Böhm tree position is a sequence of navigation steps from the root. In the Böhm tree, the possible steps at a node $\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}$ are: entering the body by peeling off one abstraction layer, and entering the $i$-th argument $M_{i}$ of the head variable. In the mixin tree, these correspond to following the $\mathrm{result}$ label (for entering the body) and the $\mathrm{argument}$ label after $\mathtt{this}$ resolution (for entering an argument position).
Rather than comparing absolute paths across different mixin trees, which would be sensitive to internal structure; hence, we observe only the convergence behavior accessible via the $\mathrm{result}$ projection. Recall the definition of inheritance-convergence (Definition B.6). Each $\mathrm{result}$-following step corresponds to one head reduction step in the $\lambda$-calculus: a tail call’s $\mathrm{result}$ projects through the encapsulated application, and a $\mathbf{let}$-binding’s $\mathrm{result}$ enters the continuation. The abstraction shape at the end signals that a head normal form has been reached.
B.3. Single-Path Lemma
The $\mathtt{this}$ function (equation 6) is designed for the general case of multi-target $\mathtt{this}$ resolution. For translated $\lambda$-terms, we show it degenerates to single-target resolution.
Lemma B.1 (Single-Path).
For any closed ANF $\lambda$-term $M$, at every invocation of $\mathtt{this}(S,\;p_{\mathrm{def}},\;n)$ during evaluation of $\mathcal{T}(M)$, the frontier set $S$ contains exactly one path.
Proof.
By case analysis on the translation rules of $\mathcal{T}$. Since $\mathcal{T}$ is compositional, the structural invariant established for each rule propagates to all inherited subtrees: every subtree appearing in the mixin tree is itself the image of some sub-term under $\mathcal{T}$, so the invariant holds throughout the runtime composition.
Case $\mathcal{T}(\lambda x.\,M)=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
This is a record literal with $\mathtt{defines}=\{\mathrm{argument},\mathrm{result}\}$ and no inheritance sources, so $\mathtt{inherits}=\varnothing$. At the root path $p$, we have $\mathtt{overrides}(p)=\{p\}$ (a single path) since there are no inheritance sources to introduce additional branches. Inside $\mathcal{T}(M)$, any reference to $x$ is $\uparrow^{\!n}.\mathrm{argument}$ where $n$ is the de Bruijn index pointing to this scope level. Since $\mathcal{T}(\lambda x.\,M)$ introduces exactly one scope level, there is exactly one mixin at that level, hence $\mathtt{this}$ finds exactly one matching pair $(p_{\mathrm{site}},\;p_{\mathrm{override}})$ in $\mathtt{supers}$.
Case $\mathcal{T}(\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M)=\{x\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
The outer record has own properties $\{x,\mathrm{result}\}$ and no inheritance sources at the outer level ($\mathtt{inherits}=\varnothing$), so $\mathtt{overrides}(\text{root})=\{\text{root}\}$. Inside the $x$ subtree, the inheritance $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$ has exactly one inheritance source ($\mathcal{T}(V_{1})$), creating exactly one $\mathtt{bases}$ entry. By induction on $V_{1}$, the single-path property holds inside $\mathcal{T}(V_{1})$. Inside $\mathcal{T}(M)$, the induction hypothesis applies directly.
Case $\mathcal{T}(\mathbf{let}\;x=V\;\mathbf{in}\;M)=\{x\mapsto\{\mathrm{result}\mapsto\mathcal{T}(V)\},\;\mathrm{result}\mapsto\mathcal{T}(M)\}$
The outer record has own properties $\{x,\mathrm{result}\}$ and no inheritance sources at the outer level, so $\mathtt{overrides}(\text{root})=\{\text{root}\}$. The $x$ subtree is itself a record literal with own property $\{\mathrm{result}\}$ and no inheritance sources. Thus neither entering the bound value nor entering the body introduces branching in $\mathtt{supers}$. By induction, the single-path property holds inside both $\mathcal{T}(V)$ and $\mathcal{T}(M)$.
Case $\mathcal{T}(V_{1}\;V_{2})=\{\mathrm{tailCall}\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathrm{tailCall}.\mathrm{result}\}$
Identical to the $\mathbf{let}$-binding case: the outer record has own properties $\{\mathrm{tailCall},\mathrm{result}\}$ with no inheritance sources at the outer level, and the $\mathrm{tailCall}$ subtree contains exactly one inheritance source.
Case $\mathcal{T}(x)=\uparrow^{\!n}.\mathrm{argument}$
This is a reference, not a record. The $\mathtt{this}$ function walks up $n$ steps from the enclosing scope. By the compositionality argument above, every scope level encountered during the walk, including inherited subtrees reached through reference resolution, is the image of some sub-term under $\mathcal{T}$, so each step of $\mathtt{this}$ encounters exactly one matching pair in $\mathtt{supers}$.
The key structural invariant is the following. For every subtree generated by $\mathcal{T}$, its root has either no inheritance source or exactly one inheritance source. Moreover, the only roots with an inheritance source are the application wrappers introduced by the application $\mathbf{let}$-binding and tail-call rules, and in both cases that source is unique and encapsulated under a distinguished child ($x$ or $\mathrm{tailCall}$), never at the enclosing scope level. Therefore every scope level traversed by $\mathtt{this}$ contributes at most one matching pair in $\mathtt{supers}$, so the frontier set remains a singleton throughout evaluation. ∎
B.4. Substitution Lemma
The core mechanism of the translation is that inheritance with $\mathrm{argument}\mapsto\mathcal{T}(V)$ plays the role of substitution. We make this precise.
Lemma B.2 (Substitution).
Let $M$ be an ANF term in which $x$ may occur free, and let $V$ be a closed value. [^32] The proof uses the Single-Path Lemma (Lemma B.1), which requires the inherited tree to occur within a closed term. This holds whenever $\lambda x.\,M$ applied to $V$ is a subterm of a closed ANF program. Define the inherited tree $C=\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$. Then for every path $p$ under $C\mathbin{\triangleright}\mathrm{result}$ and every label $\ell$:
$\ell\in\mathtt{properties}(p)\text{ in }C\mathbin{\triangleright}\mathrm{result}\quad\Longleftrightarrow\quad\ell\in\mathtt{properties}(p)\text{ in }\mathcal{T}(M[V/x])$
where $M[V/x]$ is the usual capture-avoiding substitution.
Proof.
By structural induction on $M$.
Base case: $M=x$
Then $\mathcal{T}(x)=\uparrow^{\!n}.\mathrm{argument}$ (where $n$ is the de Bruijn index of $x$), and $C\mathbin{\triangleright}\mathrm{result}$ contains this reference. In the inherited tree $C$, $\mathtt{resolve}$ and $\mathtt{this}$ resolve the reference by navigating from the reference’s enclosing scope up $n$ steps to the binding $\lambda$, where the inheritance $\{\mathcal{T}(\lambda x.\,x),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$ inherits into the $\mathrm{argument}$ slot, providing $\mathcal{T}(V)$. Since $C$ is the $\mathrm{tailCall}$ subtree of $\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)$, which is a closed ANF term, Lemma B.1 applies and $\mathtt{this}$ finds exactly one path, so $\mathtt{resolve}$ returns $\mathcal{T}(V)$’s subtree. Since $M[V/x]=V$, we have $\mathcal{T}(M[V/x])=\mathcal{T}(V)$, and the properties coincide.
Base case: $M=y$ where $y\neq x$ and $y$ is $\lambda$-bound
Then $\mathcal{T}(y)=\uparrow^{\!m}.\mathrm{argument}$ where $m$ is the de Bruijn index of $y$. This walks up to the scope that binds $y$, which is different from the scope that binds $x$, so the inheritance with $\mathrm{argument}\mapsto\mathcal{T}(V)$ at $x$’s binding scope has no effect. Since $M[V/x]=y$, the properties are identical.
Base case: $M=y$ where $y\neq x$ and $y$ is let-bound
Then $\mathcal{T}(y)=y.\mathrm{result}$. This is a lexical reference to the sibling property $y$ in the enclosing record, so the inherited tree $C$ does not alter its resolution. Since $M[V/x]=y$, the properties are identical.
Inductive case: $M=\lambda y.\,M^{\prime}$
Then $\mathcal{T}(M)=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$. The subtree at $C\mathbin{\triangleright}\mathrm{result}$ is another abstraction-shaped record. References to $x$ inside $\mathcal{T}(M^{\prime})$ resolve through $\mathtt{this}$ in exactly the same way (with one additional scope level from the inner $\lambda$), and by induction on $M^{\prime}$, the properties under $C\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}\mathrm{result}$ coincide with those of $\mathcal{T}(M^{\prime}[V/x])$. Since $(\lambda y.\,M^{\prime})[V/x]=\lambda y.\,(M^{\prime}[V/x])$ (assuming $y$ is fresh), the result follows.
Inductive case: $M=\mathbf{let}\;z=V_{1}\;V_{2}$ $\mathbf{in}\;M^{\prime}$
Then
$\mathcal{T}(M)=\{z\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}.$
The substitution distributes: $M[V/x]=\mathbf{let}\;z=V_{1}[V/x]\;V_{2}[V/x]\;\mathbf{in}\;M^{\prime}[V/x]$. By induction on $V_{1}$, $V_{2}$, and $M^{\prime}$, the properties under each subtree ($z$ and $\mathrm{result}$) coincide, since references to $x$ inside each subtree resolve to $\mathcal{T}(V)$ via the same $\mathtt{this}$ mechanism.
Inductive case: $M=\mathbf{let}\;z=V^{\prime}\;\mathbf{in}\;M^{\prime}$
Then
$\mathcal{T}(M)=\{z\mapsto\{\mathrm{result}\mapsto\mathcal{T}(V^{\prime})\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}.$
The substitution distributes: $M[V/x]=\mathbf{let}\;z=V^{\prime}[V/x]\;\mathbf{in}\;M^{\prime}[V/x]$. By induction on $V^{\prime}$ and $M^{\prime}$, the properties under both subtrees, specifically $z$ and $\mathrm{result}$, coincide. In particular, occurrences of $z$ in $M^{\prime}$ are translated as the lexical reference $z.\mathrm{result}$ on both sides, so the wrapper introduced by the value $\mathbf{let}$-binding is preserved by the substitution.
Inductive case: $M=V_{1}\;V_{2}$ (tail call)
Analogous to the $\mathbf{let}$-binding case with $\mathrm{tailCall}$ in place of $z$. ∎
B.5. Convergence Preservation
The convergence criterion (Definition B.6) follows the $\mathrm{result}$ chain from the root. Each step in this chain corresponds to one head reduction step in the $\lambda$-calculus. We make this precise.
Lemma B.3 (Result-step).
Let $(\lambda x.\,M)\;V$ be a tail call, which is a $\beta$-redex. Then for every path $p$ and label $\ell$:
$\ell\in\mathtt{properties}\!\bigl(\,\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p\,\bigr)\text{ in }\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)\;\Longleftrightarrow\;\ell\in\mathtt{properties}\!\bigl(\,\text{root}\mathbin{\triangleright}p\,\bigr)\text{ in }\mathcal{T}(M[V/x])$
That is, following one $\mathrm{result}$ projection in the tail call’s mixin tree yields the same properties as the root of the reduct’s mixin tree.
Proof.
The translation gives:
$\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)=\{\mathrm{tailCall}\mapsto\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\},\;\mathrm{result}\mapsto\mathrm{tailCall}.\mathrm{result}\}$
The $\mathrm{result}$ label at the root is defined as $\mathrm{tailCall}.\mathrm{result}$, so $\text{root}\mathbin{\triangleright}\mathrm{result}$ resolves to the $\mathrm{result}$ path inside the inheritance $\mathrm{tailCall}=\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$. This inheritance is exactly the inherited tree $C$ of Lemma B.2, and $\mathrm{tailCall}\mathbin{\triangleright}\mathrm{result}$ is $C\mathbin{\triangleright}\mathrm{result}$. By Lemma B.2, the properties under $C\mathbin{\triangleright}\mathrm{result}$ coincide with those of $\mathcal{T}(M[V/x])$.
Let-binding variant. An analogous result holds for $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}$ where $V_{1}=\lambda y.\,M^{\prime\prime}$. The translation gives:
$\mathcal{T}\bigl(\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}\bigr)=\{x\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$
Since $\mathrm{result}\mapsto\mathcal{T}(M^{\prime})$ is a property definition, the subtree at $\text{root}\mathbin{\triangleright}\mathrm{result}$ is $\mathcal{T}(M^{\prime})$. Inside $\mathcal{T}(M^{\prime})$, each reference to $x$ is $x.\mathrm{result}$, which resolves to the $\mathrm{result}$ path inside the $x$ subtree $\{\mathcal{T}(\lambda y.\,M^{\prime\prime}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$. This is the inherited tree $C$ of Lemma B.2, so $x.\mathrm{result}$ has the same properties as $\mathcal{T}(M^{\prime\prime}[V_{2}/y])$. In the $\lambda$-calculus, $M=(\lambda x.\,M^{\prime})(V_{1}\;V_{2})$, and the head reduct is $N=M^{\prime}[(V_{1}\;V_{2})/x]$. In $\mathcal{T}(N)$, each occurrence of $x$ is replaced by $(V_{1}\;V_{2})$, whose result, by the tail-call variant above, also yields $\mathcal{T}(M^{\prime\prime}[V_{2}/y])$. Therefore the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p$ in $\mathcal{T}(M)$ coincide with those at $\text{root}\mathbin{\triangleright}p$ in $\mathcal{T}(N)$ for all $p$.
Value let-binding variant. An analogous result holds for $\mathbf{let}\;x=V\;\mathbf{in}\;M^{\prime}$. The translation gives:
$\mathcal{T}\bigl(\mathbf{let}\;x=V\;\mathbf{in}\;M^{\prime}\bigr)=\{x\mapsto\{\mathrm{result}\mapsto\mathcal{T}(V)\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}.$
Since $\mathrm{result}\mapsto\mathcal{T}(M^{\prime})$ is a property definition, the subtree at $\text{root}\mathbin{\triangleright}\mathrm{result}$ is $\mathcal{T}(M^{\prime})$. Inside $\mathcal{T}(M^{\prime})$, each occurrence of $x$ is translated as the lexical reference $x.\mathrm{result}$, which resolves to the subtree $\mathcal{T}(V)$. This is exactly the translation of the reduct $M^{\prime}[V/x]$. Therefore the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p$ in $\mathcal{T}(\mathbf{let}\;x=V\;\mathbf{in}\;M^{\prime})$ coincide with those at $\text{root}\mathbin{\triangleright}p$ in $\mathcal{T}(M^{\prime}[V/x])$ for all $p$. ∎
Theorem B.4 (Convergence preservation).
If $M\to_{h}N$ (head reduction), then $\mathcal{T}(M){\Downarrow}$ if and only if $\mathcal{T}(N){\Downarrow}$.
Proof.
A head reduction step in ANF takes one of three forms: a tail call $(\lambda x.\,M^{\prime})\;V\to_{h}M^{\prime}[V/x]$, an application $\mathbf{let}$-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}\to_{h}M^{\prime}[(V_{1}\;V_{2})/x]$, or a value $\mathbf{let}$-binding $\mathbf{let}\;x=V\;\mathbf{in}\;M^{\prime}\to_{h}M^{\prime}[V/x]$. The second and third forms are simply the ANF presentation of one $\beta$-step in the underlying $\lambda$-term. In all three cases, Lemma B.3 (tail-call, application let-binding, and value let-binding variants) shows that the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}^{n}$ in $\mathcal{T}(M)$ coincide with the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}^{n-1}$ in $\mathcal{T}(N)$ for all $n\geq 1$, where $N$ is the head reduct. Hence the abstraction shape appears at depth $n$ in the pre-reduct if and only if it appears at depth n1n-1 in the post-reduct. ∎
Remark B.5 (Convergence hierarchy).
Section 4.1 showed that the Böhm tree correspondence holds at the first iteration $T_{P}\!\uparrow\!1$ and that $\mathrm{lfp}(T_{P})$ is strictly more defined. Here we make the connection precise. Tabled evaluation (Tamaki and Sato, 1986) computes $\mathrm{lfp}(T_{P})$ by iterating: $T_{P}\!\uparrow\!0=\bot$, $T_{P}\!\uparrow\!1=T_{P}(\bot)$, $T_{P}\!\uparrow\!2=T_{P}(T_{P}(\bot))$, … until stabilisation (Theorem A.6).
$T_{P}\!\uparrow\!1$ corresponds to $\beta$-reduction.:
$T_{P}\!\uparrow\!1$ evaluates each equation once, with re-entrant calls returning $\bot$ (the value from $T_{P}\!\uparrow\!0$). The Result-Step Lemma (Lemma B.3) establishes a structural isomorphism between the equation systems of $\mathcal{T}(M)$ and $\mathcal{T}(N)$, shifted by one $\mathrm{result}$ step. This isomorphism is purely structural and holds for any number of iterations. Theorems B.7 and B.9 are proved at this level.
$\mathrm{lfp}(T_{P})$ defines the fixpoint $\lambda$-calculus.:
We call the sublanguage with $\mathrm{lfp}(T_{P})$ semantics the fixpoint $\lambda$-calculus; it shares the same syntax as the lazy $\lambda$-calculus but has strictly greater computational power: $M{\Downarrow}_{\lambda}$ implies $\mathcal{T}(M){\Downarrow}_{\mathrm{lfp}}$, but not conversely.
B.6. Adequacy
Definition B.6 (Inheritance-convergence).
A mixin tree $T$ converges, written $T{\Downarrow}$, if there exists $n\geq 0$ such that:
$\{\mathrm{argument},\,\mathrm{result}\}\;\subseteq\;\mathtt{properties}\!\bigl(\,\underbrace{\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}\cdots\mathbin{\triangleright}\mathrm{result}}_{n}\,\bigr)$
in the semantics of Section 3.
Theorem B.7 (Adequacy).
A closed ANF $\lambda$-term $M$ has a head normal form if and only if $\mathcal{T}(M){\Downarrow}$.
A closed ANF term at the top level is either an abstraction (a value) or a computation (an application $\mathbf{let}$-binding, a value $\mathbf{let}$-binding, or a tail call). An abstraction translates to a record whose root immediately has the abstraction shape; a computation’s root has the form $\{x,\mathrm{result}\}$ or $\{\mathrm{tailCall},\mathrm{result}\}$, and one must follow the $\mathrm{result}$ chain to find the eventual value.
Proof of Theorem B.7.
($\Rightarrow$) Suppose $M\to^{*}_{h}\lambda x.\,M^{\prime}$ in $k$ head reduction steps. We show $\mathcal{T}(M)$ converges at depth $\leq k$ by induction on $k$.
Base case ($k=0$)
$M$ is already an abstraction $\lambda x.\,M^{\prime}$. Then $\mathcal{T}(M)=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$, whose root has $\mathtt{defines}=\{\mathrm{argument},\mathrm{result}\}$. These labels are directly in $\mathtt{defines}$ of the root, so the recursive evaluation returns them without further recursion, and $\mathcal{T}(M)$ converges at depth $n=0$.
Inductive step ($k\geq 1$)
$M$ is not an abstraction, so $M$ is either a tail call, an application $\mathbf{let}$-binding, or a value $\mathbf{let}$-binding. Each form corresponds to one head $\beta$-step in the underlying $\lambda$-term. Let $N$ be the head reduct, so $M\to_{h}N\to^{k-1}_{h}\lambda x.\,M^{\prime}$. Note that $N$ is closed: head reduction substitutes a closed value into a term whose only free variable is the substitution target, preserving closedness. By Lemma B.3 (tail-call, application let-binding, or value let-binding variant), the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p$ in $\mathcal{T}(M)$ coincide with those at $\text{root}\mathbin{\triangleright}p$ in $\mathcal{T}(N)$. By the induction hypothesis, $\mathcal{T}(N)$ converges at depth $\leq k-1$, so $\mathcal{T}(M)$ converges at depth $\leq k$.
($\Leftarrow$) Suppose $\mathcal{T}(M){\Downarrow}$ at depth $n$. We show $M$ has a head normal form by induction on $n$.
Base case ($n=0$)
The root of $\mathcal{T}(M)$ has abstraction shape $\{\mathrm{argument},\mathrm{result}\}$. Only the abstraction rule $\mathcal{T}(\lambda x.\,M^{\prime})$ produces a root with own property $\mathrm{argument}$. (For tail calls and $\mathbf{let}$-bindings, the root has $\mathtt{defines}=\{\mathrm{tailCall},\mathrm{result}\}$ or $\{x,\mathrm{result}\}$, and $\mathrm{argument}$ cannot appear via $\mathtt{supers}$ because $\mathtt{inherits}=\varnothing$ at the root, giving $\mathtt{bases}(\text{root})=\varnothing$.) Therefore $M$ is an abstraction, hence already in head normal form.
Inductive step ($n\geq 1$)
The abstraction shape appears at depth $n$ but not at depth 0. By the base case argument, $M$ is not an abstraction, so at the head position of this closed top-level ANF term, $M$ is either a tail call, an application $\mathbf{let}$-binding, or a value $\mathbf{let}$-binding. In each case, the ANF grammar requires the function position to be a value $V_{1}$ ($V::=x\mid\lambda x.\,M$). Since $M$ is closed, $V_{1}$ cannot be a free variable, so $V_{1}$ must be a $\lambda$-abstraction; the head form is therefore a $\beta$-redex. By Lemma B.3 (tail-call, application let-binding, or value let-binding variant), the properties at depth n1n-1 in $\mathcal{T}(N)$ match those at depth $n$ in $\mathcal{T}(M)$, where $N$ is the head reduct. So $\mathcal{T}(N)$ converges at depth n1n-1. Since $M$ is closed and head reduction preserves closedness , which substitutes a closed value for the only free variable, $N$ is a closed ANF term. By the induction hypothesis, $N$ has a head normal form, and since $M\to_{h}N$, so does $M$. ∎
B.7. First-Order Semantics of the Lazy $\lambda$-Calculus
Adequacy relates a single term’s head normal form to its inheritance-convergence. The following theorem lifts this to an equivalence between terms, establishing that the translation $\mathcal{T}$ gives the lazy $\lambda$-calculus a first-order semantics that preserves and reflects Böhm tree equivalence.
Definition B.8 ($\lambda$-contextual equivalence under $\mathcal{T}$).
For closed $\lambda$-terms $M$ and $N$, write $\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)$ if for every closing $\lambda$-calculus context $C[\cdot]$:
$\mathcal{T}(C[M]){\Downarrow}\;\Longleftrightarrow\;\mathcal{T}(C[N]){\Downarrow}$
The subscript $\lambda$ emphasizes that the quantification ranges over $\lambda$-calculus contexts, not over all inheritance-calculus contexts. This definition observes only convergence [^33] Convergence provides exactly a single bit of information. , but quantifying over all $\lambda$-contexts makes it a fine-grained equivalence: different contexts can supply different arguments and project different results, probing every aspect of the term’s behavior within the $\lambda$-definable fragment.
Theorem B.9 (First-order semantics of the lazy $\lambda$-calculus).
For closed $\lambda$-terms $M$ and $N$:
$\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)\quad\Longleftrightarrow\quad\mathrm{BT}(M)=\mathrm{BT}(N)$
Proof of Theorem B.9.
The translation $\mathcal{T}$ is compositional: for every $\lambda$-calculus context $C[\cdot]$, the mixin tree $\mathcal{T}(C[M])$ is determined by the mixin tree $\mathcal{T}(M)$ and the translation of the context. This means $\mathcal{T}$ preserves the context structure needed for the argument.
($\Leftarrow$) Suppose $\mathrm{BT}(M)=\mathrm{BT}(N)$. Böhm tree equivalence is a congruence (Barendregt, 1984), so for every context $C[\cdot]$, $\mathrm{BT}(C[M])=\mathrm{BT}(C[N])$. In particular, $C[M]$ has a head normal form iff $C[N]$ does. By Theorem B.7, $\mathcal{T}(C[M]){\Downarrow}$ iff $\mathcal{T}(C[N]){\Downarrow}$. Hence $\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)$.
($\Rightarrow$) Suppose $\mathrm{BT}(M)\neq\mathrm{BT}(N)$. Böhm tree equivalence coincides with observational equivalence for the lazy $\lambda$-calculus (Abramsky and Ong, 1993; Abramsky and McCusker, 1995): there exists a context $C[\cdot]$ such that $C[M]$ converges and $C[N]$ diverges, or vice versa. By Theorem B.7, $\mathcal{T}(C[M]){\Downarrow}$ and $\mathcal{T}(C[N]){\Uparrow}$. Hence $\mathcal{T}(M)\not\approx_{\lambda}\mathcal{T}(N)$. ∎
The proof rests on two pillars: Adequacy (Theorem B.7), which is our contribution, and the classical result that Böhm tree equivalence equals observational equivalence for the lazy $\lambda$-calculus (Abramsky and Ong, 1993; Abramsky and McCusker, 1995). Neither pillar requires defining or comparing absolute paths inside mixin trees.
Appendix C Expressive Asymmetry: Proofs
Definition C.1 (Macro-expressibility, after Felleisen (Felleisen, 1991)).
A construct $C$ of language $\mathscr{L}_{1}$ is macro-expressible in language $\mathscr{L}_{0}$ if there exists a translation $\mathcal{E}$ such that for every program $P$ containing occurrences of $C$, the translation $\mathcal{E}(P)$ is obtained by replacing each occurrence of $C$ with an $\mathscr{L}_{0}$ expression that depends only on $C$’s subexpressions, leaving all other constructs of $P$ unchanged.
Lemma C.2 (Composition of macro-expressible translations).
If $\mathcal{E}_{1}$ macro-expresses the constructs of $\mathscr{L}_{1}$ in $\mathscr{L}_{0}$ and $\mathcal{E}_{2}$ macro-expresses the constructs of $\mathscr{L}_{2}$ in $\mathscr{L}_{1}$, then the composition $\mathcal{E}_{1}\circ\mathcal{E}_{2}$ macro-expresses the constructs of $\mathscr{L}_{2}$ in $\mathscr{L}_{0}$.
Proof.
Each construct $F$ of $\mathscr{L}_{2}$ has a fixed syntactic abstraction $A_{F}^{(2)}$ over $\mathscr{L}_{1}$. Each $\mathscr{L}_{1}$-construct $G$ appearing in $A_{F}^{(2)}$ has a fixed syntactic abstraction $A_{G}^{(1)}$ over $\mathscr{L}_{0}$. Substituting $A_{G}^{(1)}$ for every $G$ in $A_{F}^{(2)}$ yields a fixed syntactic abstraction over $\mathscr{L}_{0}$ for $F$, satisfying E4. ∎
Definition C.3 (Language universe).
Let $\mathscr{U}$ be the language whose constructors are the union of the constructors of inheritance-calculus, the constructors of the lazy $\lambda$-calculus , including $\lambda$-abstraction, application, and variable, and $\mathbf{let}$-binding. The semantics of $\mathscr{U}$ is given by translating every $\mathscr{U}$-program to inheritance-calculus in three steps:
  1. (1)
    Convert every maximal $\lambda$-calculus subexpression (that is, every subtree whose root and all internal nodes use only $\lambda$-calculus constructors) to A-normal form (Flanagan et al., 1993).
  2. (2)
    Apply the translation $\mathcal{T}$ (Section 4) to every ANF $\lambda$-calculus subexpression.
  3. (3)
    Evaluate the resulting inheritance-calculus program by the rules of Section 3.
When a $\mathscr{U}$-expression interleaves $\lambda$-calculus and inheritance-calculus constructors (e.g. a $\mathbf{let}$-binding whose bound value is an inheritance-calculus expression), the translation is applied compositionally from leaves to root: each $\lambda$-calculus construct is replaced by its inheritance-calculus equivalent via $\mathcal{T}$, treating inheritance-calculus subexpressions as opaque values. Since each rule of $\mathcal{T}$ (Section 4) is a fixed syntactic abstraction depending only on the translated subexpressions, this bottom-up application is well-defined and yields a pure inheritance-calculus program for step (3).
A $\mathscr{U}$-program $P$ converges, $\mathrm{eval}_{\mathscr{U}}(P)$, iff its translation converges.
Remark C.4.
The lazy $\lambda$-calculus and inheritance-calculus are conservative restrictions of $\mathscr{U}$: each is obtained by removing the other’s constructors. For a pure inheritance-calculus program, steps (1) and (2) are vacuous, so $\mathrm{eval}_{\mathscr{U}}$ agrees with the inheritance-calculus semantics. For a pure $\lambda$-calculus program, the composition of steps (1)–(3) agrees with lazy evaluation by Adequacy (Theorem B.7).
The three-step translation is effective (ANF conversion and $\mathcal{T}$ are syntactic transformations) and the recursive evaluation of inheritance-calculus (Section 3) is a semi-decision procedure, so $\mathrm{eval}_{\mathscr{U}}$ is recursively enumerable, as required by Felleisen’s framework (Felleisen, 1991).
Lemma C.5 ($\mathbf{let}$-binding is macro-expressible in the lazy $\lambda$-calculus).
The translation $\mathbf{let}\;x=e_{1}\;\mathbf{in}\;e_{2}\;\mapsto\;(\lambda x.\,e_{2})\;e_{1}$ is a macro-expressible elimination of $\mathbf{let}$ in the lazy $\lambda$-calculus.
Proof.
The translation replaces each $\mathbf{let}$-binding with a $\lambda$-expression that depends only on the two subexpressions $e_{1}$ and $e_{2}$. This is Proposition 3.7 of Felleisen (Felleisen, 1991). ∎
Lemma C.6 (Equal expressiveness of the lazy $\lambda$-calculus and its ANF fragment).
In $\mathscr{U}$, the lazy $\lambda$-calculus and the ANF lazy $\lambda$-calculus (Section 4) are equally macro-expressive.
Proof.
Every ANF term is a $\lambda$-term, so the embedding of the ANF fragment into the full $\lambda$-calculus is the identity, which is trivially macro-expressible. For the reverse direction, the ANF transformation (Flanagan et al., 1993) consists of two steps, each macro-expressible:
  1. (1)
    Application flattening. Replace each nested application $(e_{1}\;e_{2})\;e_{3}$ with $\mathbf{let}\;x=e_{1}\;e_{2}\;\mathbf{in}\;x\;e_{3}$. The syntactic abstraction is $A(\cdot_{1},\cdot_{2},\cdot_{3})=\mathbf{let}\;x=\cdot_{1}\;\cdot_{2}\;\mathbf{in}\;x\;\cdot_{3}$, depending only on the three subexpressions.
  2. (2)
    Value lifting. Replace each free value $V$ used in an application position with $\mathbf{let}\;x=V\;\mathbf{in}\;x$. The syntactic abstraction is $A(\cdot_{1})=\mathbf{let}\;x=\cdot_{1}\;\mathbf{in}\;x$.
By Lemma C.5, each resulting $\mathbf{let}$-binding is itself macro-expressible in the $\lambda$-calculus. By Lemma C.2, composing these macro-expressible translations yields a macro-expressible translation from the full $\lambda$-calculus to its ANF fragment. ∎
Theorem C.7 (Forward macro-expressibility).
The translation $\mathcal{T}$ (Section 4) is a macro-expressible embedding of the $\lambda$-calculus into inheritance-calculus.
Proof.
We verify condition E4 of Felleisen (Felleisen, 1991, Definition 3.11): for each $\lambda$-calculus construct $F$ of arity $a$, there must exist a fixed $a$-ary syntactic abstraction $A_{F}(\cdot_{1},\ldots,\cdot_{a})$ over inheritance-calculus such that $\mathcal{T}(F(e_{1},\ldots,e_{a}))=A_{F}(\mathcal{T}(e_{1}),\ldots,\mathcal{T}(e_{a}))$. We check each rule of $\mathcal{T}$ (Section 4):
  1. (1)
    $\lambda$-bound variable ($x$ with de Bruijn index $n^{\prime}$). This is a nullary construct and therefore has no subexpressions. The syntactic abstraction is the constant $A=\uparrow^{\!n^{\prime}}.\mathrm{argument}$.
  2. (2)
    Let-bound variable ($x$). Nullary. $A=x.\mathrm{result}$.
  3. (3)
    $\lambda$-abstraction, denoted $\lambda x.\,M$. Unary in the body $M$. $A(\cdot_{1})=\{\mathrm{argument}\mapsto\{\},\;\mathrm{result}\mapsto\cdot_{1}\}$.
  4. (4)
    Application let-binding ($\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M$). Ternary in $V_{1}$, $V_{2}$, and $M$. $A(\cdot_{1},\cdot_{2},\cdot_{3})=\{x\mapsto\{\cdot_{1},\;\mathrm{argument}\mapsto\cdot_{2}\},\;\mathrm{result}\mapsto\cdot_{3}\}$.
  5. (5)
    Value let-binding ($\mathbf{let}\;x=V\;\mathbf{in}\;M$). Binary in $V$ and $M$. $A(\cdot_{1},\cdot_{2})=\{x\mapsto\{\mathrm{result}\mapsto\cdot_{1}\},\;\mathrm{result}\mapsto\cdot_{2}\}$.
  6. (6)
    Tail call ($V_{1}\;V_{2}$). Binary.
    $A(\cdot_{1},\cdot_{2})=\{\mathrm{tailCall}\mapsto\{\cdot_{1},\;\mathrm{argument}\mapsto\cdot_{2}\},\;\mathrm{result}\mapsto\mathrm{tailCall}.\mathrm{result}\}$.
In each case the syntactic abstraction $A_{F}$ is a fixed inheritance-calculus context whose holes are filled by the recursive translations of the subexpressions, satisfying E4. ∎
Theorem C.8 (Nonexpressibility of inheritance).
The lazy $\lambda$-calculus cannot macro-express the inheritance constructors of $\mathscr{U}$ (Definition C.3).
Proof.
We apply Theorem 3.14(i) of Felleisen (Felleisen, 1991) to the pair $L_{0}=\text{lazy $\lambda$-calculus}$, $L_{1}=\mathscr{U}$. By Definition C.3, $\mathscr{U}=L_{0}+\{$mixin-tree constructors}\} is a conservative extension of $L_{0}$ (Definition 3.2 of Felleisen (Felleisen, 1991)). We verify the four conditions:
  1. (i)
    $L_{0}\subseteq L_{1}$: The syntax of $\mathscr{U}$ includes all $L_{0}$-phrases.
  2. (ii)
    No new $L_{0}$-phrases: The mixin-tree constructors do not introduce new phrases in $L_{0}$.
  3. (iii)
    Disjoint constructors: The mixin-tree constructors are disjoint from the $L_{0}$-constructors.
  4. (iv)
    Semantics preserved: For pure $\lambda$-terms, steps (1) and (2) of Definition C.3 are vacuous and step (3) yields lazy evaluation by Adequacy (Theorem B.7), so $\mathrm{eval}_{\mathscr{U}}$ agrees with $\mathrm{eval}_{L_{0}}$ on $L_{0}$-phrases.
The theorem requires us to exhibit two $L_{0}$-terms $M$ and $N$ with $M\cong_{0}N$ but $M\not\cong_{1}N$, where $\cong_{0}$ is operational equivalence in $L_{0}$ and $\cong_{1}$ is operational equivalence in $\mathscr{U}$.
Operational equivalence in $L_{0}$
By Böhm tree equivalence (Abramsky and Ong, 1993; Abramsky and McCusker, 1995) (Theorem B.9), for closed $\lambda$-terms $M$ and $N$: $M\cong_{0}N$ iff $\mathrm{BT}(M)=\mathrm{BT}(N)$.
Separating pair
Define Church booleans and Church boolean equality: $\mathrm{true}=\lambda t.\,\lambda f.\,t$, $\mathrm{false}=\lambda t.\,\lambda f.\,f$, and $\mathrm{eq}=\lambda a.\,\lambda b.\,(a\;b)\;(b\;\mathrm{false}\;\mathrm{true})$. Let
$M=\mathrm{eq}\;\mathrm{false}\;\mathrm{false}\qquad N=\mathrm{eq}\;\mathrm{true}\;\mathrm{true}.$
Both reduce to Church $\mathrm{true}$ in the lazy $\lambda$-calculus, so $\mathrm{BT}(M)=\mathrm{BT}(N)$ and hence $M\cong_{0}N$.
Distinguishing $\mathscr{U}$-context
We now construct a $\mathscr{U}$-context that separates $M$ and $N$. $M$ and $N$ are syntactically closed $\lambda$-terms; by Definition C.3, their semantics in $\mathscr{U}$ is given by translating them to inheritance-calculus via ANF conversion followed by $\mathcal{T}$. Under the extended ANF convention (Section 4), every free value used in an application is first bound to a name:
$\mathrm{ANF}(M)={}$ $\mathbf{let}\;\mathrm{eq}=\lambda a.\,\lambda b.\,\mathbf{let}\;x_{1}=a\;b\;\mathbf{in}\;$
$\qquad\mathbf{let}\;f_{1}=\mathrm{false}\;\mathbf{in}\;\mathbf{let}\;x_{2}=b\;f_{1}\;\mathbf{in}\;$
$\qquad\mathbf{let}\;t_{1}=\mathrm{true}\;\mathbf{in}\;\mathbf{let}\;x_{3}=x_{2}\;t_{1}\;\mathbf{in}\;x_{1}\;x_{3}$
$\mathbf{in}\;\;\mathbf{let}\;f_{1}=\mathrm{false}\;\mathbf{in}\;\mathbf{let}\;\mathrm{eqFalse}=\mathrm{eq}\;f_{1}$
$\quad\;\mathbf{in}\;\mathbf{let}\;f_{2}=\mathrm{false}\;\mathbf{in}\;\mathrm{eqFalse}\;f_{2}$
Applying $\mathcal{T}$:
$\mathcal{T}(\mathrm{ANF}(M))\;=\;\{$ $\mathrm{eq}\mapsto\{\mathrm{result}\mapsto\mathcal{T}(\lambda a.\,\lambda b.\,\cdots)\},$
$f_{1}\mapsto\{\mathrm{result}\mapsto\mathcal{T}(\mathrm{false})\},$
$\mathrm{eqFalse}\mapsto\{\mathrm{eq}.\mathrm{result},\;\mathrm{argument}\mapsto f_{1}.\mathrm{result}\},$
$f_{2}\mapsto\{\mathrm{result}\mapsto\mathcal{T}(\mathrm{false})\},$
$\mathrm{tailCall}\mapsto\{\mathrm{eqFalse}.\mathrm{result},\;\mathrm{argument}\mapsto f_{2}.\mathrm{result}\},$
$\mathrm{result}\mapsto\mathrm{tailCall}.\mathrm{result}\;\}$
and symmetrically for $N$ with $\mathrm{true}$ in place of $\mathrm{false}$. Note that $\mathrm{eq}$ is a named child of $\mathcal{T}(\mathrm{ANF}(M))$, bound via a value let-binding. Inside the translation of $\mathrm{eq}$, the inner $\lambda b$ scope contains the let-bindings for $x_{1}$, $f_{1}$, $x_{2}$, $t_{1}$, $x_{3}$, and a tail call. In particular, $\uparrow^{\!1}.\mathrm{argument}$ reaches the outer $\lambda a$ scope and $\uparrow^{\!0}.\mathrm{argument}$ reaches $\lambda b$ itself. [^34] These indices use the condensed convention, counting only $\lambda$-scopes. As noted in the translation footnote (Section 4), each let-binding and tail call introduces an additional scope level in inheritance-calculus; the formal index is correspondingly larger, but the target scope is the same.
The $\mathscr{U}$-context uses inheritance to add a $\mathrm{repr}$ record inside the $\lambda b$ scope of $\mathrm{eq}$ that projects the first operand. Define the context $C[\cdot]$ as:
$C[\cdot]={}$ $\mathbf{let}\;\mathrm{extended}=\{[\cdot],\;\mathrm{eq}\mapsto\{\mathrm{result}\mapsto\{$
$\quad\mathrm{result}\mapsto\{\mathrm{repr}\mapsto\{\mathrm{firstOperand}\mapsto\uparrow^{\!2}.\mathrm{argument}\}\}\}\}\}$
$\mathbf{in}\;\mathrm{extended}.\mathrm{tailCall}.\mathrm{repr}.\mathrm{firstOperand}$
Here $\cdot$] is the hole. Note the two levels of $\mathrm{result}$ nesting in the mixin: the outer $\mathrm{result}$ enters the value let-binding wrapper for $\mathrm{eq}$, and the inner $\mathrm{result}$ enters the $\lambda a$ scope. The second $\mathrm{result}$ then contains the $\lambda b$ scope, so $\mathrm{repr}$ is placed at the $\lambda b$ scope level. The mixin $\{\mathrm{eq}\mapsto\{\mathrm{result}\mapsto\{\mathrm{result}\mapsto\{\mathrm{repr}\mapsto\cdots\}\}\}\}$ is an inheritance-calculus expression: it uses only the mixin-tree constructors of $\mathscr{U}$. The outer $\mathbf{let}$-binding and projection are $\lambda$-calculus constructs (Lemma C.5). Thus $C[\cdot]$ is a valid $\mathscr{U}$-context.
The de Bruijn index $\uparrow^{\!2}$ in $\mathrm{repr}$ counts scope levels (in the condensed convention): $\uparrow^{\!0}$ is $\mathrm{repr}$ itself, $\uparrow^{\!1}$ is the $\lambda b$ scope, and $\uparrow^{\!2}$ is the $\lambda a$ scope. So $\mathrm{firstOperand}$ projects $\lambda a$’s argument.
Since $\mathrm{eqFalse}$ (respectively $\mathrm{eqTrue}$) inherits $\mathrm{eq}.\mathrm{result}$, and $\mathrm{tailCall}$ inherits $\mathrm{eqFalse}.\mathrm{result}$ (the $\lambda b$ scope), the $\mathrm{repr}$ record placed by the mixin at the $\lambda b$ scope level propagates through the inheritance chain to $\mathrm{tailCall}$. In $C[M]$: the $\mathrm{eqFalse}$ binding supplies $\mathrm{argument}\mapsto f_{1}.\mathrm{result}$ (that is, $\mathrm{false}$) to the $\lambda a$ scope, so $\mathrm{firstOperand}$ evaluates to $\mathrm{false}=\lambda t.\,\lambda f.\,f$. In $C[N]$: the $\mathrm{eqTrue}$ binding supplies $\mathrm{argument}\mapsto\mathrm{true}$, so $\mathrm{firstOperand}$ evaluates to $\mathrm{true}=\lambda t.\,\lambda f.\,t$.
Define $D[\cdot]=(\lambda x.\,x\;\Omega\;I)\;C[\cdot]$. Then $\mathrm{eval}_{\mathscr{U}}$ holds for $D[M]$ (since $\mathrm{false}\;\Omega\;I\to I$, which converges) but not for $D[N]$ (since $\mathrm{true}\;\Omega\;I\to\Omega$, which diverges). Therefore $M\not\cong_{1}N$.
Conclusion
We have $M\cong_{0}N$ but $M\not\cong_{1}N$, so ${\cong_{0}}\neq{(\cong_{1}|_{L_{0}})}$. By Theorem 3.14(i) of Felleisen (Felleisen, 1991), the lazy $\lambda$-calculus cannot macro-express the inheritance constructors of $\mathscr{U}$. ∎
Theorem C.9 (Expressive asymmetry).
In the language universe $\mathscr{U}$ (Definition C.3), inheritance-calculus is strictly more expressive than the lazy $\lambda$-calculus: it can macro-express every $\mathscr{U}$-construct that the $\lambda$-calculus can, but not conversely (Definition 3.17 of Felleisen (Felleisen, 1991)).
Proof.
We must show: (a) inheritance-calculus can macro-express every $\mathscr{U}$-construct that the lazy $\lambda$-calculus can macro-express, and (b) the converse fails.
Part (b): the lazy $\lambda$-calculus $\not\geq$ inheritance-calculus
The $\lambda$-calculus contains its own constructors ($\lambda$, application, variable) and can macro-express $\mathbf{let}$-binding (Lemma C.5), but cannot macro-express the inheritance constructors of $\mathscr{U}$ (Theorem C.8). Inheritance-calculus contains those constructors. Therefore the lazy $\lambda$-calculus is not at least as expressive as inheritance-calculus with respect to $\mathscr{U}$.
Part (a): inheritance-calculus $\geq$ the lazy $\lambda$-calculus
We show that inheritance-calculus can macro-express every $\mathscr{U}$-construct that the $\lambda$-calculus contains or can macro-express. The $\lambda$-calculus contains $\lambda$-abstraction, application, and variable; it can macro-express $\mathbf{let}$-binding (Lemma C.5). We must show that inheritance-calculus can macro-express all four of these constructs. The translation proceeds in two stages:
  1. (1)
    ANF conversion. The A-normal form transformation (Flanagan et al., 1993) replaces nested applications with $\mathbf{let}$-bindings; it is macro-expressible in the $\lambda$-calculus by Lemma C.6. In the reverse direction, $\mathbf{let}$-bindings are macro-expressible in the $\lambda$-calculus by Proposition 3.7 of Felleisen (Felleisen, 1991) (restated as Lemma C.5). Thus ANF conversion is a macro-expressible equivalence between the $\lambda$-calculus and its ANF fragment.
  2. (2)
    The translation $\mathcal{T}$. Each ANF $\lambda$-calculus construct maps to a fixed syntactic abstraction over inheritance-calculus (Theorem C.7), satisfying condition E4 of Felleisen (Felleisen, 1991, Definition 3.11).
By Lemma C.2, composing the two macro-expressible translations gives a macro-expressible translation from the $\lambda$-calculus (with $\mathbf{let}$) into inheritance-calculus. Since inheritance-calculus trivially contains its own constructors, it can macro-express every $\mathscr{U}$-construct that the $\lambda$-calculus contains or can macro-express. ∎
Appendix D Multi-Target $\mathtt{this}$ Resolution in Other Systems
This appendix demonstrates that two representative systems, Scala 3 and the NixOS module system, reject the multi-target $\mathtt{this}$ resolution pattern that inheritance-calculus handles naturally.
Scala 3
Consider two objects that independently extend an outer class, each providing its own copy of an inner trait:
class MyOuter:
trait MyInner:
def outer = MyOuter.this
object Object1 extends MyOuter
object Object2 extends MyOuter
object HasMultipleOuters extends Object1.MyInner
with Object2.MyInner
Scala 3 rejects HasMultipleOuters with the error:
trait MyInner is extended twice
object HasMultipleOuters cannot be instantiated since
it has conflicting base types
Object1.MyInner and Object2.MyInner
The rejection is not a surface-level restriction but a consequence of DOT’s type-theoretic foundations (Amin et al., 2016). In DOT, an object is constructed as $\nu(x:T)\,d$, where the self variable $x$ binds to a single object. Object1.MyInner and Object2.MyInner are distinct path-dependent types, each constraining the outer $\mathtt{this}$ to a different object. Their intersection requires outer to simultaneously return Object1 and Object2, but DOT’s self variable is single-valued, making this intersection unrealizable.
The equivalent inheritance-calculus definition is:
$\{\quad\mathrm{MyOuter}$ $\mapsto\{\mathrm{MyInner}\mapsto\{\mathrm{outer}\mapsto\mathrm{MyOuter}.\mathbf{this}\}\},$
$\mathrm{Object1}$ $\mapsto\{\mathrm{MyOuter}\},$
$\mathrm{Object2}$ $\mapsto\{\mathrm{MyOuter}\},$
$\mathrm{HasMultipleOuters}$ $\mapsto\{\mathrm{Object1}.\mathrm{MyInner},\;\mathrm{Object2}.\mathrm{MyInner}\}\quad\}$
This is well-defined in inheritance-calculus. The reference $\mathrm{MyOuter}.\mathbf{this}$ inside $\mathrm{outer}$ has de Bruijn index $n=1$: starting from $\mathrm{outer}$’s enclosing scope $\mathrm{MyInner}$, one $\mathtt{this}$ step reaches $\mathrm{MyOuter}$. When $\mathrm{HasMultipleOuters}$ inherits from both $\mathrm{Object1}.\mathrm{MyInner}$ and $\mathrm{Object2}.\mathrm{MyInner}$, the $\mathtt{this}$ function (equation 6) resolves $\mathrm{MyOuter}.\mathbf{this}$ by searching through $\mathtt{supers}(\mathrm{HasMultipleOuters})$ for override paths matching $\mathrm{MyInner}$’s definition site. It finds two inheritance-site paths, one through $\mathrm{Object1}$ and one through $\mathrm{Object2}$, and returns both. Both paths lead to records that inherit from the same $\mathrm{MyOuter}$, so querying $\mathrm{HasMultipleOuters}.\mathrm{outer}$ yields the $\mathtt{properties}$ of $\mathrm{MyOuter}$ regardless of which route is taken. Because attribute resolution aggregates results via set union (which is an intrinsically idempotent operation), identical property definitions from the two routes merge without conflict. Thus, the calculus eliminates the need for any artificial route-priority or linearization rules.
NixOS Module System
The NixOS module system rejects the same pattern with a static error. The translation uses the module system’s own abstractions: deferredModule for traits, which are unevaluated module values imported into other fixpoints, and submoduleWith for objects, which are evaluated in their own fixpoints.
let
lib = (import <nixpkgs> { }).lib;
result = lib.evalModules {
modules = [
(toplevel@{ config, ... }: {
# class MyOuter { trait MyInner {
# def outer = MyOuter.this } }
options.MyOuter = lib.mkOption {
default = { };
type = lib.types.deferredModuleWith {
staticModules = [
(MyOuter: {
options.MyInner = lib.mkOption {
default = { };
type = lib.types.deferredModuleWith {
staticModules = [
(MyInner: {
options.outer = lib.mkOption {
default = { };
type =
lib.types.deferredModuleWith {
staticModules =
[ toplevel.config.MyOuter ];
# object Object1 extends MyOuter
options.Object1 = lib.mkOption {
default = { };
type = lib.types.submoduleWith {
modules = [ toplevel.config.MyOuter ];
# object Object2 extends MyOuter
options.Object2 = lib.mkOption {
default = { };
type = lib.types.submoduleWith {
modules = [ toplevel.config.MyOuter ];
# object HasMultipleOuters extends
# Object1.MyInner with Object2.MyInner
options.HasMultipleOuters = lib.mkOption {
default = { };
type = lib.types.submoduleWith {
modules = [
toplevel.config.Object1.MyInner
toplevel.config.Object2.MyInner
in
builtins.attrNames result.config.HasMultipleOuters.outer
The NixOS module system rejects this with:
error: The option HasMultipleOuters.outer
in ‘<unknown-file>’
is already declared
in ‘<unknown-file>’.
The rejection occurs because Object1.MyInner and Object2.MyInner each carry the full staticModules of the MyInner deferred module, including the declaration options.outer. When HasMultipleOuters imports both, the module system’s mergeOptionDecls encounters two declarations of the same option and raises a static error; it cannot express that both declarations originate from the same trait and should be unified.
The rejection is analogous to Scala’s “conflicting base types” error above: both systems assume that each option or type member has a single declaration site, and reject the multi-target situation where two inheritance routes introduce the same declaration independently. In inheritance-calculus, the same pattern is well-defined because $\mathtt{overrides}$ (equation 3) recognizes that both routes lead to the same definition, and $\mathtt{supers}$ (equation 2) collects both inheritance-site contexts without duplication.
Appendix E First-Order Semantics of the ANF Fixpoint $\lambda$-Calculus
This appendix reformulates the semantics of Section 3, specialized to $\lambda$-terms via the Single-Path Lemma (Lemma B.1), directly on ANF syntax. The first-iteration approximation gives the lazy $\lambda$-calculus semantics.
The Böhm tree has exactly two node types—scope and application—with separate path types and separate first-order functions for each. All dynamic functions are single-valued: the Single-Path Lemma (Lemma B.1) guarantees that no multi-inheritance arises for translated $\lambda$-terms.
Syntax
Let $M$ denote a computation and $V$ a value.
[…]
The grammar includes application let-bindings ($\mathbf{let}\;x=V_{1}\;V_{2}$), value let-bindings ($\mathbf{let}\;x=V$), tail calls ($V_{1}\;V_{2}$), and tail values ($V$). Variables use de Bruijn indices: $\uparrow^{\!n}$ counts only $\lambda$-abstractions between the reference and its binder, following the standard convention. Let-bound variables are referenced by their binding name $x$.
Two node types
Each position in the Böhm tree is either a scope node or an application node. These have distinct path types $\mathcal{S}$ and $\mathcal{A}$ and distinct first-order functions.
A scope node (path type $\mathcal{S}$) represents a body: a chain of let-bindings ending in a tail expression. A scope may or may not be a $\lambda$-abstraction (recorded by $\mathrm{is\_abstraction}$). Its edges are the names of application let-bindings $\{x_{1},\ldots,x_{k}\}$; value let-bindings do not create edges (see AST below). Navigation from a scope to a binding: $s\mathbin{\triangleright}x_{i}$ gives an application path $a\in\mathcal{A}$.
An application node (path type $\mathcal{A}$) represents a function application $V_{1}\;V_{2}$ bound by a $\mathbf{let}$. In the AST, $\mathtt{head\_ref}$ and $\mathtt{arg\_ref}$ are references (de Bruijn index plus projection, or an inline $\lambda$-literal); in the evaluated Böhm tree, $\mathtt{head}$ and $\mathtt{argument}$ lazily resolve these references to scope nodes. Navigation from an application to a scope: $\mathtt{head}(a)\in\mathcal{S}$ and $\mathtt{argument}(a)\in\mathcal{S}$.
The path types alternate strictly:
$\mathcal{S}\xrightarrow{\;\mathbin{\triangleright}\,x_{i}\;}\mathcal{A}\xrightarrow{\;\mathtt{head}/\mathtt{argument}\;}\mathcal{S}\xrightarrow{\;\mathbin{\triangleright}\,y_{j}\;}\mathcal{A}\;\cdots$
There is no scope-to-scope or application-to-application direct transition.
Appendix F Encoding Datalog in Inheritance-Calculus
Section 6.1 observed that inheritance-calculus natively produces the relational semantics of logic programming. This appendix makes that observation precise by giving a systematic encoding of Datalog into inheritance-calculus. The encoding uses no extensions to the calculus; it is expressible in the three primitives of Section 2. All examples in this appendix have executable MIXINv2 implementations (Yang, 2026). There are three ingredients to keep separate. First, relations are stored as tries, so tuple lookup is indexed by prefixes rather than by scanning a flat set of facts. Second, the control skeleton is written in the visitor-pattern style already used in Section 5; from a functional-programming viewpoint, this is the same role played by Scott-encoded case analysis. Third, inheritance union and tabled evaluation give the resulting program its relational, least-fixed-point semantics. The appendix proceeds in that order: data representation first, then control flow, then recursion.
F.1. Relations as Tries
A binary relation over a finite domain $D=\{a,b,c,\ldots\}$ is represented as a trie: a mixin tree whose paths encode tuples. The trie is built from three base scopes: $\mathrm{Relation}$ (abstract base), $\mathrm{Cons}$ (non-empty node), and $\mathrm{Nil}$ (terminator). Each constant $d\in D$ inherits from $\mathrm{Cons}$ and carries a $\mathrm{TailMap}$ that provides per-constant sub-tries:
$\mathrm{Relation}\mapsto\{\},\quad\mathrm{Cons}\mapsto\{\mathrm{Relation},\;\mathrm{Tail}\mapsto\{\mathrm{Relation}\}\},\quad\mathrm{Nil}\mapsto\{\mathrm{Relation}\}$
$d\mapsto\{\mathrm{Cons},\;\mathrm{Tail}\mapsto\{\},\;\mathrm{TailMap}\mapsto\{d\mapsto\{\mathrm{Tail}\}\}\}$
A fact $R(d_{1},d_{2})$ is encoded as a trie path of depth 2, with a $\mathrm{Nil}$ terminator. The per-constant sub-tries are accessed via $\mathrm{TailMap}.d$:
$R(a,b)\;\leadsto\quad R\mapsto\{a,\;\mathrm{TailMap}\mapsto\{a\mapsto\{b,\;\mathrm{TailMap}\mapsto\{b\mapsto\{\mathrm{Nil}\}\}\}\}\}$
Multiple facts for the same relation are accumulated by inheritance (trie union). Adding $R(b,c)$:
$R\mapsto\{a,\;b,\;\mathrm{TailMap}\mapsto\{$ $a\mapsto\{b,\;\mathrm{TailMap}\mapsto\{b\mapsto\{\mathrm{Nil}\}\}\},$
$b\mapsto\{c,\;\mathrm{TailMap}\mapsto\{c\mapsto\{\mathrm{Nil}\}\}\}\}\}$
The point of the trie representation is not merely that it is a tree-shaped encoding. It also provides the indexing structure used by the subsequent joins: once a constant $d$ has been selected, the encoding moves directly to $\mathrm{TailMap}.d$ and continues from the matching sub-trie. This is why the appendix speaks about tries rather than about an arbitrary encoding of finite sets of tuples.
F.2. Infrastructure
Each constant $d$ carries four groups of boilerplate definitions used by the encoding. These play a role analogous to derived type-class instances in Haskell: each constant must provide a fixed set of properties that the encoding’s generic patterns rely on. Inheritance-calculus has no built-in derivation mechanism, so the boilerplate is written out manually; however, the amount of boilerplate per constant is $O(1)$ and does not grow with the domain size $|D|$, so adding a new constant does not require modifying any existing scope, preserving the expression-problem property.
Acceptance:
The visitor-pattern infrastructure from Section 5. Operationally, this is the object-oriented spelling of a case split on the current constant. Each constant $d$ defines a $\mathrm{VisitorMap}$ containing a branch keyed by $d$, a $\mathrm{Visitor}$ with a $\mathrm{Visited}$ slot for per-constant results, and an $\mathrm{Accepted}$ property that delegates to the matching branch’s $\mathrm{Visited}$ value. This is the dispatch mechanism: given a scope that may contain multiple constants, $\mathrm{Acceptance}$ selects the branch corresponding to each constant present and collects the per-constant $\mathrm{Visited}$ results.
WithVisitorTail:
Maps each constant $d$ to its sub-trie entry $\mathrm{TailMap}.d$, labeled as $\mathrm{VisitorTail}$. This provides the right-hand side of a join [^36] In the context of the visitor pattern, this acts as the visitor. : when two relations’ $\mathrm{VisitorMap}$ entries are composed, $\mathrm{VisitorTail}$ carries the joined relation’s row data for the matched constant.
WithVisiteeTail:
Maps each constant $d$ to its sub-trie entry $\mathrm{TailMap}.d$, labeled as $\mathrm{VisiteeTail}$. This provides the left-hand side [^37] This acts as the visitee. : the relation being iterated. Together, $\mathrm{WithVisitorTail}$ and $\mathrm{WithVisiteeTail}$ enable value-level joins where only constants present in both relations produce output.
Replacement:
Given a $\mathrm{NewTail}$ value, produces a $\mathrm{Replaced}$ scope that inherits the label $d$ but substitutes $\mathrm{TailMap}.d$ with $\mathrm{NewTail}$. This is the output construction mechanism: it builds output tuples by substituting new sub-tries into constant wrappers. Crucially, $\mathrm{Replacement}$ is defined on the per-constant $\mathrm{TailMap}$ entry (not on the merged scope), ensuring that output construction operates on individual constants rather than their union.
As a running example, we use the executable two-hop encoding $\mathrm{twoHop}(X,Z)\mathrel{{:}\!{-}}\mathrm{edge}(X,Y),\;\mathrm{edge}(Y,Z)$ from RelationalTwoHopRepeatedVisitor.mixin.yaml (Yang, 2026). Its top-level scopes are named after the stages of the encoding: _XAcceptance iterates the first column, _YAcceptance performs the shared-variable join, _NilAcceptance0 confirms the end of the first fact, _ZAcceptance iterates the output column, and _NilAcceptance1 triggers the final Replacement chain. The schematic encodings below should be read as abstractions of that executable MIXINv2 file. If one temporarily ignores relational accumulation and reads each dispatch as an ordinary case split, the control flow is: choose $X$ from the first relation, choose a matching $Y$, confirm that the first tuple is complete, choose $Z$ from the second relation’s matching row, confirm that the second tuple is complete, and emit $(X,Z)$. The rest of the appendix explains how that familiar single-valued control skeleton is realised by visitor dispatch over tries and then lifted to relational semantics by inheritance union.
F.3. Encoding Rules
A Datalog rule $H(\bar{X})\mathrel{{:}\!{-}}B_{1}(\bar{Y}_{1}),\ldots,B_{n}(\bar{Y}_{n})$ is encoded by nesting three mechanisms inside that control skeleton:
  1. (1)
    Column iteration. For each column position, an $\mathrm{Acceptance}$ dispatch iterates over the constants present. The $\mathrm{VisitorMap}$ is populated from $\mathrm{WithVisiteeTail}$, so each matched constant’s $\mathrm{Visitor}$ receives $\mathrm{VisiteeTail}$: the per-constant sub-trie at that position. The result of processing each constant is placed in $\mathrm{Visited}$, and $\mathrm{Accepted}$ collects all per-constant results. (Details in Sections F.4, F.5, and F.6.)
  2. (2)
    Join. For each shared variable between body atoms, the inner $\mathrm{Acceptance}$’s $\mathrm{VisitorMap}$ inherits from both $\mathrm{WithVisitorTail}$ (from the joined relation) and $\mathrm{WithVisiteeTail}$ (from the iterated relation). A constant $d$ produces a $\mathrm{Visitor}$ entry only if $d$ appears in both relations, achieving a value-level join. The $\mathrm{Visitor}$ then receives both $\mathrm{VisiteeTail}$ and $\mathrm{VisitorTail}$, providing access to both relations’ data for the matched constant.
  3. (3)
    Output construction. Inside the innermost $\mathrm{Nil}$ confirmation (verifying that each body atom’s tuple is complete), chain $\mathrm{Replacement}$ operations from the innermost output column outward, each setting $\mathrm{NewTail}$ to the previous $\mathrm{Replaced}$ value, terminated by $\mathrm{Nil}$. The final $\mathrm{Replaced}$ value is assigned to $\mathrm{Visited}$.
F.4. Variable Join
Consider the simplest join: $\mathrm{composed}(X,Z)\mathrel{{:}\!{-}}\alpha(X,Y),\;\beta(Y,Z)$. The executable self-join instance of this pattern is RelationalTwoHopRepeatedVisitor.mixin.yaml (Yang, 2026), whose scope names are shown in parentheses below. This is the right place to see the division of labour from the section introduction: visitor dispatch provides the control flow, the trie provides direct access to the matching tails, and $\mathtt{supers}$ accumulates the outputs produced by each successful match.
The shared variable $Y$ requires matching constants between $\alpha$’s second column and $\beta$’s first column. The encoding nests five $\mathrm{Acceptance}$ dispatches: iterate $X$ from $\alpha$, join on $Y$ between $\alpha$ and $\beta$, confirm $\mathrm{Nil}$ after $Y$ in $\alpha$, iterate $Z$ from $\beta$’s matched row, and confirm $\mathrm{Nil}$ after $Z$. We present each level as a named sub-scope.
Level 1: iterate $X$ from $\alpha$
(Executable scope: _XAcceptance.)
$\mathrm{xAcc}\mapsto\{\alpha.\mathrm{Acceptance},\;\mathrm{VisitorMap}\mapsto\{\alpha.\mathrm{WithVisiteeTail}\},$
$\qquad\mathrm{Visitor}\mapsto\{\_X\mapsto\{\mathrm{VisiteeTail}\},\;\ldots,\;\mathrm{Visited}\mapsto\{\mathrm{yAcc}.\mathrm{Accepted}\}\}\}$
Level 2: join $\alpha$’s $Y$-column with $\beta$’s first column
(Executable scope: _YAcceptance.)
The $\mathrm{VisitorMap}$ inherits from both $\beta.\mathrm{WithVisitorTail}$ and $X.\mathrm{WithVisiteeTail}$, producing an entry for constant $d$ only if $d$ appears in both $\alpha$’s second column (via _X\_X) and $\beta$’s first column.
$\mathrm{yAcc}\mapsto\{\_X.\mathrm{Acceptance},$
$\qquad\mathrm{VisitorMap}\mapsto\{\beta.\mathrm{WithVisitorTail},\;\_X.\mathrm{WithVisiteeTail}\},$
$\qquad\mathrm{Visitor}\mapsto\{\_Y_{0}\mapsto\{\mathrm{VisiteeTail}\},\;\mathrm{VisitorTail}\mapsto\{\beta.\mathrm{Tail}\},$
$\hskip 18.49988pt\ldots,\;\mathrm{Visited}\mapsto\{\mathrm{nilAcc}_{0}.\mathrm{Accepted}\}\}\}$
Level 3: confirm $\mathrm{Nil}$ after $Y$ in $\alpha$
(Executable scope: _NilAcceptance0.)
This checks that $\alpha(X,Y)$ is a complete two-column fact (its trie path terminates at $\mathrm{Nil}$ after $Y$). Inside the $\mathrm{Nil}$ branch, the encoding proceeds to iterate $\beta$’s second column via $\mathrm{VisitorTail}$:
$\mathrm{nilAcc}_{0}\mapsto\{\_Y_{0}.\mathrm{Acceptance},\;\mathrm{VisitorMap}\mapsto\{\mathrm{Nil}\mapsto\{\ldots,\;\mathrm{Visited}\mapsto\{\mathrm{zAcc}.\mathrm{Accepted}\}\}\}\}$
Level 4: iterate $Z$ from $\beta$’s matched row
(Executable scope: _ZAcceptance.)
$\mathrm{zAcc}\mapsto\{\mathrm{VisitorTail}.\mathrm{Acceptance},\;\mathrm{VisitorMap}\mapsto\{\mathrm{VisitorTail}.\mathrm{WithVisiteeTail}\},$
$\qquad\mathrm{Visitor}\mapsto\{\_Z\mapsto\{\mathrm{VisiteeTail}\},\;\ldots,\;\mathrm{Visited}\mapsto\{\mathrm{nilAcc}_{1}.\mathrm{Accepted}\}\}\}$
Level 5: confirm $\mathrm{Nil}$ after $Z$ and construct output
(Executable scope: _NilAcceptance1, with _ZReplacement and _XReplacement.)
$\mathrm{nilAcc}_{1}\mapsto\{\mathrm{VisiteeTail}.\mathrm{Acceptance},\;\mathrm{VisitorMap}\mapsto\{\mathrm{Nil}\mapsto\{$
$\qquad\mathrm{zR}\mapsto\{\_Z.\mathrm{Replacement},\;\mathrm{NewTail}\mapsto\{\mathrm{Nil}\}\},$
$\qquad\mathrm{xR}\mapsto\{\_X.\mathrm{Replacement},\;\mathrm{NewTail}\mapsto\{\mathrm{zR}.\mathrm{Replaced}\}\},$
$\qquad\mathrm{Visited}\mapsto\{\mathrm{xR}.\mathrm{Replaced}\}\}\}\}$
Final result
$\mathrm{composed}\mapsto\{\mathrm{xAcc}.\mathrm{Accepted}\}$
Semantics
At Level 2, the combined $\mathrm{VisitorMap}$ ensures that a constant $d$ produces a $\mathrm{Visitor}$ only when $d$ appears in both $\alpha$’s second column and $\beta$’s first column. The $\mathrm{Visitor}$ for each matched constant $d$ independently receives $\mathrm{VisiteeTail}$ ($\alpha$’s data for $d$) and $\mathrm{VisitorTail}$ ($\beta$’s row for $d$). Output construction (Level 5) operates inside this per-constant $\mathrm{Visitor}$, so $\mathrm{Replacement}$ acts on the individual constant’s data, not on the merged scope. Each matched constant $d$ produces its own output tuples in $\mathrm{Visited}$, and these are collected by $\mathrm{Accepted}$ via trie union ($\mathtt{supers}$). This yields standard Datalog tuple-level semantics: only constants that actually match between both relations contribute to the output. For the concrete instance $\mathrm{edge}=\{(a,b),\;(b,c),\;(c,a)\}$, as encoded in RelationalTwoHopRepeatedVisitor.mixin.yaml (Yang, 2026), Level 1 selects $a$, Level 2 matches $b$, Level 3 confirms that $\mathrm{edge}(a,b)$ ends in $\mathrm{Nil}$, Level 4 iterates $\mathrm{edge}.\mathrm{TailMap}.b$ and selects $c$, and Level 5 confirms $\mathrm{Nil}$ again before the two $\mathrm{Replacement}$ operations reconstruct $(a,c)$.
F.5. Constant Match and Navigation
Constants in Datalog atoms use two mechanisms, neither of which is a join. They are simpler than variable joins because no value from a second relation needs to be matched against the currently chosen constant:
Type existence match
An atom $R(X,b)$ with a constant in the second position requires checking that constant $b$ exists in $R$’s second column at the matched row. This is encoded as an $\mathrm{Acceptance}$ with a hand-crafted $\mathrm{VisitorMap}$ containing only the $b$ branch (instead of composing from $\mathrm{WithVisitorTail}$, which would create branches for all constants):
$\mathrm{VisitorMap}\mapsto\{b\mapsto\{\ldots\}\}$
If $b$ exists in the dispatching scope, the $b$ branch fires; otherwise, $\mathrm{Accepted}$ receives no matching branch.
Direct navigation
An atom $R(b,Z)$ with a constant in the first position navigates directly to $b$’s sub-trie:
$\mathrm{edgeB}_{Z}\mapsto\{R.\mathrm{TailMap}.b\}$
This bypasses the visitor mechanism entirely, accessing the per-constant $\mathrm{TailMap}$ entry to restrict the scope to only $b$’s data.
Example
$\mathrm{viaB}(X,Z)\mathrel{{:}\!{-}}\mathrm{edge}(X,b),\;\mathrm{edge}(b,Z)$ combines both mechanisms. The executable MIXINv2 encoding is RelationalConstant.mixin.yaml (Yang, 2026). The encoding iterates $\mathrm{edge}$ for $X$ (Level 1), then checks for constant $b$ in $X$’s row with a hand-crafted $\mathrm{VisitorMap}$ (Level 2), navigates $\mathrm{edge}.\mathrm{TailMap}.b$ for $Z$ (Level 3), confirms $\mathrm{Nil}$ (Level 4), and constructs $(X,Z)$ via $\mathrm{Replacement}$.
F.6. Multi-Body Rules
Rules with $n>2$ body atoms chain additional join levels. Each shared variable introduces one combined $\mathrm{WithVisitorTail}$$\mathrm{WithVisiteeTail}$ join inside the previous level’s $\mathrm{Visitor}$, and the output construction is placed inside the innermost $\mathrm{Nil}$ confirmation. So the binary-rule encoding is the whole story structurally: larger bodies only add more nested copies of the same pattern.
Three-body rule
$\mathrm{chain}(X,W)\mathrel{{:}\!{-}}\alpha(X,Y),\;\beta(Y,Z),\;\gamma(Z,W)$ requires two joins: on $Y$ (between $\alpha$ and $\beta$) and on $Z$ (between $\beta$ and $\gamma$). The structure is: iterate $X$ from $\alpha$, join on $Y$ (getting both $\alpha$’s and $\beta$’s data), confirm $\mathrm{Nil}$ after $Y$ in $\alpha$, then inside the $\mathrm{Nil}$ branch join $\beta$’s $Z$-column with $\gamma$’s first column, iterate $W$ from $\gamma$’s matched row, confirm $\mathrm{Nil}$, and construct $(X,W)$. The $Z$-join is nested inside the $Y$-join’s $\mathrm{Nil}$ confirmation, ensuring that $\beta$’s data for the matched $Y$ constant (available via $\mathrm{VisitorTail}$) is used to iterate $\beta$’s second column.
Mixed constants and variables
$\mathrm{constJoin}(X,W)\mathrel{{:}\!{-}}\mathrm{edge}(X,Y),\;\mathrm{edge}(Y,b),\;\mathrm{edge}(b,W)$ combines a variable join on $Y$, a constant match on $b$ (hand-crafted $\mathrm{VisitorMap}$ with only the $b$ entry), and direct navigation via $\mathrm{edge}.\mathrm{TailMap}.b$ for $W$. The corresponding executable MIXINv2 file is RelationalConstantJoin.mixin.yaml (Yang, 2026).
F.7. Multi-Variable Join
When a rule shares multiple variables between two body atoms, each shared variable requires its own $\mathrm{Acceptance}$ check. The checks are chained: the outer $\mathrm{Acceptance}$ joins on one column pair using combined $\mathrm{WithVisitorTail}$$\mathrm{WithVisiteeTail}$, and its $\mathrm{Visitor}$ contains an inner $\mathrm{Acceptance}$ joining on another column pair.
Example
$\mathrm{multiJoin}(X,Y)\mathrel{{:}\!{-}}r(X,Y),\;s(X,Y)$ shares both $X$ and $Y$. The executable MIXINv2 encoding is RelationalMultiVariableJoin.mixin.yaml (Yang, 2026). The outer $\mathrm{Acceptance}$ joins $r$’s and $s$’s first columns (variable $X$) using combined $\mathrm{WithVisitorTail}$$\mathrm{WithVisiteeTail}$. Inside the outer $\mathrm{Visitor}$, an inner $\mathrm{Acceptance}$ joins $r$’s and $s$’s second columns (variable $Y$) similarly. Both joins must match for the output to be produced, and the per-constant $\mathrm{Visited}$ pattern ensures that only constants present in both relations contribute.
F.8. Recursive Rules
Recursive Datalog rules, such as transitive closure $\mathrm{path}(X,Y)\mathrel{{:}\!{-}}\mathrm{edge}(X,Z),\;\mathrm{path}(Z,Y)$, are encoded identically to non-recursive rules: the head relation $\mathrm{path}$ appears both as a defined scope and as a reference in the body (providing $\mathrm{WithVisitorTail}$ for the $Z$-join). This should not be read as ordinary $\mathtt{this}$ of a single-valued lazy program. Operationally, tabled evaluation (Section 3, Appendix A) computes $\mathrm{lfp}(T_{P})$: each query into $\mathrm{path}$ triggers further evaluation through the recursive body, and the trie union ($\mathtt{supers}$) accumulates all reachable tuples. This corresponds exactly to Naive evaluation (Bancilhon and Ramakrishnan, 1986) of the $T_{P}$ operator (van Emden and Kowalski, 1976): each iteration applies every rule to the current fact set and adds newly derived facts. Because Datalog-encoded relations have finitely many constants per query, the per-query answer domain is finite, and convergence is guaranteed by the ascending chain condition on a finite lattice. Semi-Naive evaluation (Bancilhon and Ramakrishnan, 1986), which restricts each iteration to rules that use at least one newly derived fact, is a natural optimisation left for future work.
The recursive rule as monadic pseudocode
The recursive rule above roughly corresponds to the following pseudocode (Haskell-like syntax with RebindableSyntax and OverloadedLists semantics, where do-notation desugars to a Set-based bind and list literals denote sets):
{-# LANGUAGE RebindableSyntax, OverloadedLists #-}
s >>= f = unions [f x | x <- s]
return x = [x]
guard True = [()]
guard False = []
edge = [("a","b"), ("b","c"), ("c","a")]
path = edge <> do
(x, y0) <- edge
(y1, z) <- path
guard (y0 == y1)
return (x, z)-}
Each line of the pseudocode has a direct counterpart in the inheritance-calculus encoding (Section F.4):
edge:
corresponds to the base-fact trie;
path = edge <> do …:
corresponds to $\mathrm{path}$ inheriting $\mathrm{edge}$ (base facts) and the recursive body (derived facts via $\mathtt{supers}$);
(x, y0) <- edge:
corresponds to Level 1 ($\mathrm{xAcc}$), iterating $X$ from $\mathrm{edge}$;
(y1, z) <- path:
corresponds to Level 2 ($\mathrm{yAcc}$), joining on $Y$ with $\mathrm{path}$, and Level 4 ($\mathrm{zAcc}$), iterating $Z$;
guard (y0 == y1):
corresponds to the $\mathrm{WithVisitorTail}\cap\mathrm{WithVisiteeTail}$ intersection in Level 2: a constant produces a $\mathrm{Visitor}$ entry only if it appears in both relations;
return (x, z):
corresponds to the $\mathrm{Replacement}$ chain in Level 5, constructing the output tuple $(X,Z)$.
The only difference is the evaluation strategy: Haskell evaluates path lazily (diverges), whereas inheritance-calculus evaluates it via tabled $T_{P}$ iteration (converges to $\mathrm{lfp}$).
However, this monadic program does not converge because path is defined recursively. [^38] Enabling the RecursiveDo extension does not help: mfix ties a knot at the element level via laziness, whereas the transitive closure equation requires a fixed point at the container level, that is, finding a set $S$ such that $S=F(S)$. Inheritance-calculus solves this through tabled evaluation (Section 3), which computes the least fixed point of the $T_{P}$ operator by iterating over the lattice of mixin trees, precisely the container-level fixed point that the pseudocode above cannot express.
F.9. Semantic Correspondence with Standard Datalog
The encoding produces standard Datalog tuple-level semantics. Putting the previous subsections together: tries provide indexed tuple storage, visitor dispatch supplies the control skeleton, and per-constant $\mathrm{Replacement}$ keeps output construction local to each witness. The key mechanism is the per-constant $\mathrm{Visited}$ pattern: output construction ($\mathrm{Replacement}$) occurs inside each matched constant’s $\mathrm{Visitor}$ scope, operating on the individual constant’s data rather than on the merged scope. The $\mathrm{Accepted}$ field then collects all per-constant $\mathrm{Visited}$ results via trie union ($\mathtt{supers}$).
For example, given $\mathrm{edge}=\{(a,b),\;(b,c),\;(c,a)\}$ (a 3-cycle), the two-hop self-join $\mathrm{twoHop}(X,Z)\mathrel{{:}\!{-}}\mathrm{edge}(X,Y),\;\mathrm{edge}(Y,Z)$ produces exactly $3$ pairs: $\{(a,c),\;(b,a),\;(c,b)\}$, matching standard Datalog. This is the behavior exhibited by RelationalTwoHopRepeatedVisitor.mixin.yaml (Yang, 2026). When the $Y$-join matches constant $b$ (present in both $\mathrm{edge}$’s second column from $(a,b)$ and $\mathrm{edge}$’s first column), the $\mathrm{Visitor}$ for $b$ receives $\mathrm{VisitorTail}=\mathrm{edge}.\mathrm{TailMap}.b$ (containing only $c$), and $\mathrm{Replacement}$ constructs a single output tuple $(a,c)$, rather than all of $\{a,b,c\}$.
This value-level precision arises because each constant’s $\mathrm{Replacement}$ operates on $\mathrm{TailMap}.d$ (the per-constant sub-trie) rather than on the merged trie. The same point is visible in the repeated-visitor self-join regression example (Yang, 2026), where repeated visitor dispatch prevents witnesses from different constants from being merged spuriously. Although $\mathtt{bases}$ (equation 4) still computes cross-joins at the primitive level, the encoding routes computation through per-constant $\mathrm{TailMap}$ entries before any merging occurs, achieving the same effect as standard Datalog’s tuple-level evaluation.
Acknowledgements.
Claude Code, GitHub Copilot, Gemini Code Assist, and Codex were utilized to generate sections of this Work, including text, tables, graphs, code, formulas, citations, etc.